A Complete Bibliography of Publications in Information Processing Letters: 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

25 April 2024 Version 1.92

Title word cross-reference

# [SW13]. #P [Mil19]. #P-completeness [Mil19].

(2, 0, 0) [WX13]. (2, ℓ) [GLY11]. (2p + 1) [LLL^{+}11]. (2P_{2}, H) [DP18]. (3 + ϵ)k [LX19]. (a, b, k) [ZJ11, ZXS19]. (Δ + 1) [WLW11, ZW10]. (g, f, n) [Liu10]. (k, l) [DFdFT16]. (k, m) [ZSY13]. (k, s) [PS19a]. (X , S) [YWC11]. (n, k) [CLS13, WXZ^{+}12, YLG10]. (P_{5}, H) [DP18]. (→, ∃) [SeF14]. (s, t) [DL13, FLP^{+}18]. (t, n) [QD16]. 0/1 [OBT12]. 1 [AIR17, AMRR11, Cha19, CZZ^{+}12, Cza13, Did13, EKN11, LTK^{+}19, SSK12, Wu14, XL15, ZW11, ZL12, ZLLZ18]. 1.5 [EKN11].

10 [CL15, KSV15]. 11 [CL15, Hoc11]. 121 [KTUY17]. 13 [Blo15]. 16 [Mil15]. 2 [AMT12, ACL18, CLZ18, Che16c, CLS13, DFMS10, EAA^{+}16, EKN11, FY14, Jan12b, KR16, KM13, LZ10a, LH11, LT13c, LM15, NN17, OM17, Och17, Sch13b, Shi12, Tak16, TXQ11, TZ11, Wid17, XGX11, YBMK15, vGLM12]. 2/3 [Vid13]. 22 [MNP12]. 24 [Ehl17]. 2^{k} [Sun16]. 2^{n} [GKM14]. 2^{w} [Bae13]. 2k − c log k [Lam11]. 2p [EP16]. 2p^{m} [KZ12]. 2π/5 [BF19]. 2\sqrt{2} [GKW15]. 3 [AD18, BS11, DS15, Dur13, FWS13, FM19, GH14, JSZC15, Kat16, Lei17, Mil15, NS11, NT14, PI19, Xia10, YL11a, YYYZ12, YYZH13, Zha10b, ZHXS10]. 3 × 3 [GSA^{+}19, OKM13]. 4 [AD18, LS15, ZLSX16]. 4n [CCT14]. 5 [CQ12, Kar13, Moo11, WLW11, Zha10b]. 6 [Hoc11, WLW11, Zha10b, ZW10]. 7 [DN12, WW11]. 7 − [3] [SWZ12]. 7/6 [Tak16]. 8 [Mou19]. 9 [Zha10b]. [2] [BFRV15]. [a, b] [YH19]. ^{∗} [BKS16, WH16]. R [ZZZ15]. α [KMMN15, KP14b]. b [Fuj16]. B^{+} [YBM^{+}10]. β [Dur12]. c [vBCH^{+}15]. C_{7} [PS19b]. PS^{−} [Pas15a, Pas15b]. d [Ezr10, LT13b, LT15b, QD16, RV13]. D^{2} [JKY15]. Δ [ZW10]. Δ = 9 [CL15]. Δ ≥ 6 [WLW11]. d × d [KA17]. e [MNV11]. E = I + T [BV11]. ϵ [Aka14, Ezr10]. f [ZZ18]. g [WH16]. γ [Dur12]. GF(2) [VSP18]. GF(2)[x] [SF12]. GF(2^{m}) [NCP18, nXlCL14]. GF(3^{m}) [PCH13]. GF(m^{2}) [CLY11]. GF (p) [FLMQ10]. GF (q) [LPdS10]. H [Szy12, GLPR19, GHvtHP13]. H_{1} [SR11]. {k} [ALT15, HK16, LT13b, LT15a, LT15b, SGM13, Tur12, ALT15, ALT18, BK10, BYK14, BT16a, BIR16, BN10b, BJMC17, CLZ18, CCH^{+}14a, CP10, DM13, DW10a, DG14, DT10, DK14, EFKR10, FLPS15, FGKU15, GLPR19, GGG^{+}14, GGF13, Gra15, HLY11, HL12, Hil14, JKY15, Joh14a, KSK13, KZP10, KP14b, KM16b, Kun18, Lau18b, LLT14, Lee10, LSW17, Li10, ML10b, Mor11, NR17, Nut18, Pra12, Pue11, SW13, SSS15, SSKA17, Suk19, Tsu13, WZL12, Xav12, YL16b, Zal11, ZJN10, ZXZL11, ZLS^{+}17]. k = 2 [SSS15]. K_{4} [LSK17, WWW18]. kP + IQ [LZ12]. L [LM16]. L(2, 1) [DL13, JSR11, PG12]. L(δ_{1}, δ_{2}, 1) [Cal13]. L(j, k) [WL12]. L_{1} [DGKS14]. λ′ [CLM12]. λ_{2} [LWXZ14]. m [Kok17, LX13c, Nut18]. M(k) [WWZ15]. F_{2} [AH17]. F_{2m} [GSA^{+}19]. F_{3^{m}} [CCH14b]. Z [GPSD17]. GF(2^{m}) [LlChL11]. m ≤ 3 [Hav10]. n [BN10b, GH14, Kok17, LWL11, LL10c, MPRC19, RARM12, SD13, WH16, YYYZ12, YYZH13]. O [MPRC19]. O(1) [YBMK15]. O(1.587^{n}) [MTA10]. O(d13^{d}) [KA17]. O(n) [CG10b]. O(n^{2}) [BCNPL14, CPW11, Pol18]. O^{∗}(1.84^{k}) [CCF14]. ω [FS14, IL12]. P [Lan11, Tel19, NHK15]. p^{2} [WXCK19]. P_{2} [WNFC10]. P_{3} [MS18b, TY13, Tu15, TZ11]. P_{4} [BDNVP15, BKN^{+}12]. P_{5} [BL14]. P_{6} [Mos13a]. π [vG18]. π/3 [BF17]. Q [CB18, MS16, CA17, Sal12]. r [AK10, BGJ^{+}16, BWZ12, KdER^{+}14, LZ10a]. r = m − 1 [LP10]. S_{n}(132) [Vaj18]. Σ [BR10b]. st [SS19a]. TF2|ν = 1, c = 2|C_{max} [LHW^{+}16]. W [1] [MS18b]. X [TX11].

-acyclicity [Dur12]. -Approximation [ACL18, EKN11, GKW15, Tak16, YBMK15]. -arboricity [NN17]. -ary [GH14, YYYZ12, YYZH13, Bae13]. -Attribute-Anonymity [SSS15]. -automata [IL12]. -bent [GPSD17]. -bottleneck [YWC11]. -calculus [vG18]. -choosability [Zha10b]. -clauses [SR11]. -CNFs [BT16a]. -code [TX11]. -colorable [Hoc11, WX13, ZLSX16]. -coloring [AD18, NS11]. -colorings [Szy12, ZZ18]. -coloured [vGLM12]. -colouring [Lei17]. -connected [CLZ18, DFMS10, GLY11, LZ10a, Nut18, Wid17]. -Connection [LM15]. -connectivity [HL12]. -covered [YH19]. -critical [Liu10, ZJ11, ZXS19]. -cube [GH14, LWL11, YYYZ12, YYZH13]. -Curry [SeF14]. -cycle [WLW11]. -cycles [Zha10b]. -deleted [ZSY13]. -dimensional [Jan12b, LL10c, QD16, RARM12, RV13, WH16, YBMK15, ZHXS10]. -discrepancy [Aka14]. -disjoint [Che16c, LL10c, PI19]. -distance [Kun18, Suk19]. -DNF [Lau18b]. -dominating [BYK14, Nut18]. -domination [DW10a, ALT15]. -dominations [ALT18]. -edge [GLPR19, NT14, Tak16]. -edge- [SS19a]. -edge-colored [OM17]. -edge-connected [Mou19]. -edge-connectivity [Sch13b]. -edge-critical [FM19]. -element [Zal11]. -entropy [CA17]. -extra [GH14]. -factors [GGG^{+}14]. -fault [AMRR11, SSK12]. -flows [DS15, YL11a]. -free [LX13c, Mos13a, BL14, DP18, GLPR19, GHvtHP13, LSK17]. -Gabriel [KSV15]. -gap [LWXZ14].

-good-neighbor [LTK^{+}19, WH16]. -gram [Sal12]. -hard [MS18b]. -hardness [Lan11]. -heavy [CLZ18]. -hop [KM16b, LZ10a]. -hulls [KP14b]. -intervals [vBCH^{+}15]. -isogenies [Moo11]. -labeling [KSK13, Cal13, JSR11, PG12]. -labelings [DL13, WL12]. -laden [BKN^{+}12]. -languages [FS14]. -length [DG14]. -letter [Dur13]. -line [BS11]. -manifold [LT13c]. -matching [Fuj16]. -means [JKY15, LSW17]. -median [Cha19, Wu14]. -minor [WWW18]. -mismatch [CP10]. -mismatches [GGF13, Gra15]. -move [CCT14]. -nearest [Lee10]. -nets [Ezr10]. -NN [KZP10]. -numberings [SS19a]. -Optimality [CLM12]. -ordered [Li10]. -orientations [LLL^{+}11]. -packing [Shi12, WNFC10]. -paging [KMMN15]. -partite [BWZ12, ML10b]. -path [FLP^{+}18, Kat16, ZLS^{+}17]. -periodic [AMT12]. -planar [Cza13, Did13, ZW11, ZL12, ZLLZ18]. -quasi-Horn [Pue11]. -query [KR16]. -rankings [ZJN10]. -regular [BJMC17]. -relaxed [DL13]. -resolution [CB18, MS16]. -restricted [MNV11, WZL12, YL16b]. -rollback [CZZ^{+}12]. -round [Blo15, Kar13]. -sampling [JKY15]. -SAT [PS19a, SW13, XGX11]. -search [ZXZL11]. -server [EFKR10, Hil14]. -set [BIR16, CCH^{+}14a, Mor11, AK10]. -shapes [KP14b]. -sink [XL15]. -sorter [Ehl17]. -space [Ezr10, Jan12b]. -sparse [BDNVP15]. -stage [LH11]. -star [CLS13, NHK15, YLG10, WXZ^{+}12]. -stars [BGJ^{+}16]. -subcoloring [Och17]. -subgraph [BK10]. -Subset [Xav12]. -sumset [BFRV15]. -terminal [LT13b, LT15a, LT15b]. -total-colorability [WLW11, ZW10]. -trapezoid [LT13b, LT15b]. -tree [YBM^{+}10]. -trees [CLS13]. -trivial [Tur12]. -tuple [ALT15, ALT18, Pra12]. -utilities [OBT12]. -valued [KM13]. -vertex [LX19]. -vertex- [Sch13b]. -vertex-colored [OM17]. -visibility [LM16]. -way [Xia10]. -wheel [KdER^{+}14].

/poly [Tel19].

1 [Tho18]. 1-median [Cha13]. 104 [LPWZ14]. 110 [BB11, Zha13]. 111 [KMS19, KT12]. 113 [GR14]. 115 [Pas15a, SSW16]. 118 [KMN18]. 128 [MGPI12]. 132 [KMS19]. 192 [Blo15].

2 [CCRS14, MB14]. 2-FTP [MB14]. 2-K-reliability [CCRS14]. 2019 [Ano19a, Ano19o]. 256 [AKY13]. 2D [EFMA10].

3 [GSY15]. 3-choosability [Zha13].

42-step [AKY13].

5- [Zha13]. 59 [HAM11]. 5D [KTUY17].

6- [Zha13].

802.11e [YWcW^{+}14].

9-cycles [Zha13].

= [MST16, Tel19].

ABE [TY16a]. Abelian [CIK^{+}13]. above [CGJ10, GY10, Vid13]. absent [WJS10]. Abstract [vGGS11, DLMV10, GCNR10, NAD19]. abstraction [AK13]. accelerate [VLV15]. acceleration [Bab17]. acceptor [HI12]. Access [MGNAB11, AC10, CPTZ13, Csi12, Ria17, Ria19]. accordance [MSV14]. achieving [ML10b]. Ack [Pér17]. Ack-complete [Pér17]. acquaintances [CS16b]. acquisitions [Gou15]. activated [ZDQ^{+}17]. activation [CC14, XZBX16]. active [SD19]. activities [JC10]. Activity [SSZW16]. ACTL [XZ17b]. Acyclic [Fie11, FH12, HRW11, AKMR19, DX10, FH10, GKPP16, GKW15, HS19, MU15, MMS15, PT11, RT13, SM17, SY15b, Xu11]. Acyclically [ZLSX16, Hoc11]. acyclicity [Dur12]. ad [HKW18, Kor12]. ad-hoc [HKW18]. adaptive [AZ14, CY18, DG16, GKCK11, Nic19, Oan11, SaBG17, TWW17, YYK17]. adaptively [Elm15]. adaptivity [ZZJ11]. Addendum [BG15, BMW13, HAM11]. addition [DGR15, JKSS19, Laz10]. additional [Shi18]. additive [CX18, EFKR10, Har18, MMS15, OBT12]. Adjacent [CL15, Zha10b, ZW10, ZL12, Zha13, vZBSY16]. Adjacent-vertex-distinguishing [CL15]. adjusting [MMKK18]. adjustment [TWW17]. admissibility [NAD19]. Advanced [SKK10]. Advantages [Sto16]. adversary [LXLY12]. advice [Miy14]. AES [DK10, GLMS18, eSKAI10]. Affine [Agg15, KR16]. Affine-evasive [Agg15]. after [YL16a]. after-the-fact [YL16a]. against [Bee16, DBB12, FS13b, GSY15, MNP12, SB19, SY15a, Suz18, TC11, Yam14]. agent [DP12]. agents [BDPP18, YMSA14]. aggregate [SGM13]. aggregation [CKY15]. aggregations [BKZ15]. agreeable [MY13]. Agreement [Shi15, BIR16, CDPB^{+}10, MT18, Nos11, Nos14]. Aiding [RM18]. AKF [KDH15]. al [PLPW13]. algebra [BNRC10, Str16, GN10]. algebraic [GNG11, KSBT13]. Algorithm [HZSL05, WZQH16, AR18, AK10, Ada11, Aku10, AR13a, Ale13, AR13b, Ara10, Ars18, AGW13, ACL18, Bae13, BGI^{+}12, BL12, BMB16, BP11, Bir11, Bol14, BES17, CCF14, CC13, CJ11, CdA13, CR18, CCL10, CZD14, CCXW12, CCT14, CNPS15, CP10, Deo12, DJZ^{+}15, Dur13, EN17, EGK13, EKN11, FBD18, FZ13, FCLCR17, FKR^{+}16, FGG^{+}10, FLP^{+}18, FV13, FB10, FS12, GL11, Ger12, GM13a, Ghi14, Gio18, GM15, GK10, GM13b, GPT16, GJ19, HCCG15, HHJ^{+}12, HWJ96, HMS16, HAM11, Hou10, HHTL10, IS10, Ili13, IH18, JAH16, JSO11, JSR11, Kak15, Kat16, KC11, KT16, KM16a, KN13c, KKR^{+}13, KM16b, Kun18, Kut12, Lee10, LlCyChL10, LTWS11, LFZJ14, LSS15, LT15b, LFW^{+}18, LLWH13, LXX14, M.15, MK16, Man10, MNV11, ML10a, Mat15, MKI11, Min11, Möm15, MOW17, MN14, NB12]. algorithm [NG10, NHK15, NT14, iP13, PP13, PC18, PI19, Pea16, PR17, RR16, RS12b, Rza14, SS18, Shi12, Sor10, SWZ12, SW19, Svi12, TLL16, TNN11, Tak16, TJ18, TCXT10, TFY11, TVB15, Tsu19b, TZ11, Tu15, Vyg11, WS10, WCZZ12, WWBB15, Waw14, XN12, XTTH12, nXlCL14, Yam19, YZZX12, YLZZ18a, YBMK15, YW12, YL17, Zeh16, ZZH10a, ZGLY12, ZXJ^{+}11, EFKR10, LREIMBMV16]. algorithm-based [GPT16]. Algorithmic [CCKP13, Pra12, BFRV15, Car19, DS16]. Algorithms [BFP18, Gav11, AMMY10, Bab17, BK10, BLYL17, BYK14, BDH^{+}11, BRFGL10, BI14, BFLM15, CC12, CCH^{+}14a, CZCD13, CP13, Dam17, DDK^{+}15, DDER18, DG14, Doe13, EL10, FLRS11, FY14, Fuj16, GKM^{+}15, GG13, GJ14, Jan12b, KMK18, KV10, Kim10, KK15b, KMMN15, Kra12, Las17, LZ10a, LXDX12, LZY12, LLG14, LS11, LLL18b, LMS14, LV15, LC17, LFXH17, LM11, MO12, MLW11, Mor11, NS11, Nut18, OKM13, OG10, OZL16, PYHA10, PPN^{+}17, Sal12, Tur13, YBHK13, YC11, ZJN10, ZXZL11]. aliasing [LR10]. aligned [Shi13]. alignment [DM10]. all-one [nXlCL14]. all-pairs [Ars18, MN19, OG10]. alliance [YBHK13]. Allocation [YWcW^{+}14, FP18, Jia11]. almost [BKS16, Ehl17, Elm10, FS13a, Tel19, Zho16, LFW^{+}18]. almost-increasing [Elm10]. almost-universal [BKS16]. alphabet [AdFEGRI11, CK11a, CD19, Kos16]. Alphabetic [Bae10]. alphabets [HN10, LNP11, LGT17, NII^{+}15]. alternate [SW12]. alternating [Den19, KDH15, Tsa11, Tsa15, ZXY10, ZX10]. Alternative [CG15, AAC^{+}10, GLO12, ST18]. always [BKS16, BvdZ19]. always-connected [BvdZ19]. Amdahl [YBC11]. among [Mis15, SS17]. Amortized [ALT16]. Analogues [BHLM13]. analysed [XZ10]. Analysis [AR13b, BS11, Cho12, HKK12, JF15, Ada11, ABPS15, AT18, Bol10, CRJ10, FOSC18, FS12, Gra18, JKY15, JB16, KRV16, LR10, LLG10, MS13, NZX19, SB19, SL19, WWS12, WZQH16, Waw14, Wit14, WDH16, XZBX16, YWC11, YL16a, YK15, ZW14, ZZ19, Zun18, LSLY11]. analytic [Mat12]. Analyzing [FG18a]. anarchy [DE19, YSGY10]. ancestor [FH10]. ancestors [GBH10]. AND-circuits [Mor11]. and/or [SD13, MTA10]. Android [CC14]. Angle [BF17, AGÁRCS^{+}19, BF19, DKL^{+}12]. Angle-constrained [BF17]. angular [FMHL11]. anisotropic [MG16]. Annotating [CV12]. annotation [DWQ10, KC17, NVB15]. annulus [AAB^{+}19]. anonymity [HLR11, SSS15]. anonymous [BEFP11]. answering [KP12, Wij10]. answers [Ama10]. Anti [ZJ16, BFP18, WZS^{+}18]. Anti-Collusion [ZJ16, WZS^{+}18]. anti-powers [BFP18]. Antimagic [Bar10, Slí12]. antisymmetric [SY15b]. any [CVV10, Tys13]. App [LCC17, HPP17]. applicability [Tho12]. Application [LQL^{+}17, BDNVP15, CFJ12, CA17, HH15, Kim10, RK15, SL19, XZ17b, ZLM17, ZD18]. applications [BT16a, BT16b, CW13, LLWH13, VB19, WD11]. applicative [SSSM11]. apply [Bol14]. Applying [MN19]. approach [AR13a, AC10, AMN^{+}10, CLR13, CPHS18, CNKS15, DJRB15, GPT16, LH10, PCK10, YWWW14, YTYZ15, ZMZ19]. approval [BT16b]. Approximability [MM19a, vBCH^{+}15, HKT17, MS15]. Approximate [Dat15, DFRS13, GGF13, AGW13, DTS15, Fil18, RS12b, Sim16, ZA17, ZXJ^{+}11]. approximately [HKW14]. Approximating [BV19, Gou15, JT10, KKZ13, LMV17, SLC15, ZLB11, BTW15, Wu14]. Approximation [ACL18, CZCD13, DDK^{+}15, GKM^{+}15, GHRT17, GM15, KMK18, KV10, LLG14, LS11, LLL18b, LV15, YC11, BK10, BU17, BL12, BLYL17, BDH^{+}11, BC13, BI14, BK18, CR18, Çiv13, DGK^{+}17, Doe13, DJZ^{+}15, EKN11, Fuj16, GS10a, Gen14, GKW15, HM10, KS11, LXDX12, LM11, M.15, MKI11, Möm15, Mor11, NPR17, Nut18, SW13, Tak16, Tat19, TZ11, YBMK15]. approximations [Cha13, Cha19]. APX [Lee17]. APX-hardness [Lee17]. arbitrarily [BPW12]. Arbitrary [Bra15a, BKS16, BDPP18, GNV14, KKO10]. Arboral [DGK^{+}17]. arborescences [KK15a]. arboricity [BU17, CQ12, GZM15, NN17, WWLC14]. arc [CGJY12, GKM^{+}15, LX13c, LJL18, XZW15]. arc-chromatic [XZW15]. architectural [CV12]. architecture [GCH^{+}10]. architectures [Pie15]. arcs [GLLX12]. Area [CY17, PP14]. Area-universal [CY17]. areas [BNRC10]. argument [Fre14]. argumentation [DW10b, NAD19]. arithmetic [ACK11, IM17, Jan12a, JS18, SM17]. Armstrong [KL10]. arrangement [EH18, RV13, ZX11]. arrangements [AD18, CW13, KPR18]. array [Liu15, LGT17, RK15, TNN11, TX11, YYYZ12, YYZH13, ZDQ^{+}17]. array-based [YYYZ12]. arrays [GRSS17, GS10c, KM10, KTV13]. art [IM12]. Arthur [Ned17]. ARTICLE [Zha10b]. artificial [GL11]. ary [Bae13, GH14, YYYZ12, YYZH13]. aspect [CCKP13]. aspects [Pra12]. Assignment [LREIMBMV16, ADFM13, BTW15, CK11b, GM13a, KM13, Liu15, WS16, YYYZ12, YYZH13]. assignments [DHW11]. associative [BKZ15]. assortative [SW18]. assumption [JSO11]. asymmetric [BGL10, RV17]. asymptotic [TD14]. asymptotically [Ama11]. asynchronous [BR10b, vG18]. AT-free [BFLM15]. atomic [BR10b, Che18b]. atoms [Ivá16, KP12, SM17]. attack [ABPS15, AT18, ASA13, Bee16, Blo15, GSY15, KA17, LR18, LC13, LYHH14, Men12, MNP12, OPS14, TC11]. attacking [GJ19]. Attacks [KN13a, ACD18, Dra16, GLMS18, KM10, Kim10, LJP19, RS15, TS16, TY16b, WS13, WWBC14]. attestation [CWW10]. attractor [MTA10]. attractors [AMT12]. Attribute [XWLJ16, XWS17, SSS15]. Attribute-Based [XWLJ16, XWS17]. auctions [Ger19]. augmentation [IZ10, ZZ14]. augmented [CN18, Fu10]. augmenting [EKN11]. authenticated [LWS10, LHH11, Nos11, Nos14]. authentication [ASA13, Jia16]. autocorrelation [EP16]. automata [ADF13, AY12a, AL18, BN10a, DK14, HS19, HHK17, IL12, Kos18, RS12a, VB15, Vrg15, ZZH16]. automated [CNKS15]. Automatic [Ghi14, Li15, JKSS19, SSZW16, YEMR19]. automaton [EGKL11, Fre10, FG14]. Automorphism [DZ12]. automorphisms [YTN10]. autoregressive [CZD14]. auxiliary [GGG^{+}14]. Average [Sal12, Cho12, EO13, GFG11, KPSZ11, Li12, LL16, MWZ12, YL11b, CGJ10]. averaging [CG15]. AVL [ALT16]. avoiding [BV10, CH12]. aware [AZ14, AHLS18, VB19, ZGLY12]. axes [AAJ15]. axiomatization [AGI15]. axioms [GNV14, KL10]. axis [Ezr10, MM19a]. axis-parallel [Ezr10, MM19a]. backbone [JT15, ZXJ^{+}11]. backbones [JT15]. Background [ZD18]. backward [Sal12]. bacteria [LTWS11]. bag [ADG10, KL10]. bag-set [ADG10]. bakery [Ara10]. Balanced [FP18, AADB^{+}18, AMO19, Bog10, CGLS16, DM16, FLMQ10, Kam19, LW19, SWF18, Til19, YZ19]. Balancedness [MS18a]. balancing [AY12b, DHW11]. ball [ZLS^{+}17]. Bandwidth [STU12, LL10b, WCW11]. banner [Mos13a]. Base [HM13, LST11]. Based [XWLJ16, ASTD14, AdFEGRI11, ABPS15, AT18, Aku10, ASM17, AC10, AMN^{+}10, ABS12, Bab17, Bol10, CPHS18, CYQ13, CZD14, CTHP13, CZZ^{+}10, CDM^{+}11, DWQ10, Dra16, jDX11, EH18, EZ15, Fay16, Gal13, GWJ11, GKCK11, GPT16, GS19, GW16, HF14, HLR11, HHTL10, JKY15, yJxW16, JK18, JCC11, KSBT13, KM10, KC17, LXLY12, LK14, LCC17, LZJX10, LFZJ14, Li15, LPdS10, LH10, LMCG16, MG16, MHFMSa11, MMZ12, MM13, Mes15, MGPI12, MS13, MHIU18, NS11, PLPW13, PYYC16, QYWX16, RPL19, SZC^{+}17, Sun16, TNN11, TPL16, TWW17, Tia15, VN17, WWYY11, XPC^{+}10, XW12, XWS17, XZ17b, YCL11, YHLC12, YYYZ12, YZH14, YL11b, ZpH15, ZMZ19, ZXJ^{+}11, ZZ13, LZJX10, MGNAB11]. bases [FY14, Lag14, LL10d]. basis [CLY11, CCH14b, FS12, Ghi14, LLP^{+}18, LlCyChL10]. Batch [MM12, FL16, FTYL14, GY15, LY11, LYF11, LZY12, LZLY12, LZ14, MZC11, SP18]. batching [FCNY10, Oro11, TFY11, Zhu12]. Bayesian [DJRB15]. BC [LJP19]. BCCSP [AdFEGRI11]. BDD [LWXZ14, YKD^{+}12, ZZ19]. be [AIR17, Att17, CLZ18, DF11, MN15, WZL12, YL16b]. Beam [VAC13]. beat [JS19, Pol18]. Bee [WZQH16, GL11]. behavior [Lee10]. behaviors [ZZ11]. behavioural [vB12]. Bell [QD16]. Benczúr [Ber17]. bent [BPRMS14, DZQF13, GPSD17, GNG11, PZ12, Pas15a, Pas15b, PGZB19]. Bernstein [DZ11]. best [FT15]. best-response [FT15]. Beth [CTHP13]. better [BCNPL14, CSX16]. between [Bar13, BDNVP15, DD14, Feg19, HLR11, Jan12a, yJxW16, Kle13, KK11, Mat19, MHIU18, Mos13b, MOW17, NB12, SC12, Sim16, Suk19]. bi [CGLS10]. bi-enhancement [CGLS10]. Biased [LREIMBMV16, LLP^{+}18]. bichromatic [AABCC19]. Biclique [KDH13, Gav11, Sie19]. biclique-free [Sie19]. bicliques [BRFGL10, CK12a, Dam14]. Bicolored [CK12a, DBFMPLV17, GGG^{+}14]. biconnected [CY17]. bicyclic [MBW19]. bidding [EL10]. bidirectional [Vay13]. bijections [Bar13]. bilinear [OKM13, ZY17]. bin [CP15, DE19, HPY^{+}10, Jan12b, LMU15]. bin-packing [CP15]. Binary [BFKL13, AG19, BC15, DCH12, DHR13, EP16, GG13, GY19, JS18, KYC13, KM12, Luc10, LEP10, Maß15, Mil15, MR10, WXCK19, YKD^{+}12, YEMR19, ZST13]. binomial [CG10a, Pas15a, Pas15b]. bins [AY12b, BTW15]. biometrics [LXLY12, SS17]. biometrics-based [LXLY12]. bipanconnectivity [Che10b]. bipartite [ALT18, BK10, Che10c, CL15, CGLS10, DP17, Dam14, DEL10, FS13a, KO16, Kut12, LC17, Miy14, Mor16, MT10, NN17, PP10, Pul16, STU12, Tak16, XPC^{+}10, ZZ18, CLM12, GM13b, KV10]. bipolar [GS17]. Birthday [ACD18, GJ19]. birthday-type [GJ19]. bisection [Aku10, GY10]. bisector [EH18]. bisimilarity [AIS10, AGI15, Kie13]. bit [ASM17, BG11, JS18, LlChL11, LMCG16, RH10, VN17, nXlCL14, YA13]. bit-parallel [LlChL11, nXlCL14]. bit-parallelism [ASM17]. bit-rate [YA13]. bit-vector [JS18]. Bitcoin [Bee16, DSPSHJNA18]. bitwise [CdA13]. bivariate [AH17]. Black [JP11, Sto16, DW12, HBL14, Shp13]. Black-box [Sto16, DW12]. black-boxes [Shp13]. Black-peg [JP11]. BLAKE [VNP10]. BLAKE2 [NA19]. blind [BBB^{+}17]. Block [EFMA10, HCCG15, Ama11, HHTL10, LM91, LC13, LYHH14, MNP12, SB19, Sar11, SKN11, SKK10, Vir11, WB12, WWBC14]. block-interchanges [HHTL10]. Block-wise [EFMA10]. blocking [MÖ15]. blocks [MS12]. Bloom [PRM14, CRJ10, Gra18, PRM16]. blue [CCL10]. Board [Ano18g, Ano18h, Ano18i, Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano10m, Ano10n, Ano10o, Ano10p, Ano10q, Ano10r, Ano10s, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano11n, Ano11o, Ano11p, Ano11q, Ano11r, Ano11s, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano12r, Ano12s, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano13n, Ano14a, Ano14b, Ano14c, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano16a]. Board [Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18j, Ano18k, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m]. body [KHKS16]. body-hinge [KHKS16]. Bohr [Aka14]. Boolean [AMT12, CKK13, DKKY10, GJ11, KS18, Lag14, MTA10, MGPI12, Sim16, SV19, Vir11, WT13]. Bootstrap [KdER^{+}14]. Borderedness [FWH14]. Borderedness-preserving [FWH14]. borders [RSwW11]. both [AIR17, OG11]. Bottleneck [AABCC19, BLC10, YWC11]. bound [AS18, ADFM13, Ber11, BGL10, BCKM15, BGSM10, Dar15, DE19, FZ13, Ger12, GD13, GNG11, GY10, HYC12, KM11, KTUY17, MCS12, PP14, Pod12, PS19b, Pud12, Shr19, VK18, XGX11, Zha10a, Zho15, vZBSY16]. Boundary [LP13a]. Bounded [MZC11, BEFP11, BU17, BDF^{+}18, BM19, BN10b, DDER18, DTS15, EL10, Feg19, Fie11, FGvL11, GKM14, GJ15, HMS16, IS10, JT15, Jan12b, JSR10, Kor12, LLL18b, MSZ11, MRZ10, Nis92, Pod12, Rac10, RR16, SLdAMP17, TCXT10, XZW15, vE17, CFJ12]. bounded-degree [JT15]. Boundedness [Oan11]. Bounding [DMS12, Fre14, JF15]. Bounds [GWL18, KR16, LST11, Qi17, XS15, ASTD14, ABH19, BU17, BT16a, BS17, Cha12, CA17, CP16, DKKY10, DEH^{+}19, Fil11, GW16, KS11, Kos17, LEP10, MO12, Par11, SI18, VSP18, WXZ^{+}12, WLLS08, Wit14]. box [DW12, Sto16]. boxes [BCNPL14, Ezr10, KSBT13, Shp13, eSKAI10]. Boyce [KL10]. Boyer [AR13b]. BPA [Kie13]. BPP [Mos11a]. branching [EGK13, Laz10, PPN^{+}17]. Breadth [CG10b]. Breadth-First [CG10b]. break [YBC11]. Breaking [TPL16]. bribery [BT16b]. brief [Val12]. Broadcast [TVB15, IS10, MPRC19, VLV15, WWYY11]. broadcasting [WCW11, YWY12]. broken [LTK^{+}19]. broken-down [LTK^{+}19]. Brook [KS12]. Bruijn [BH11, GZJ^{+}12]. Bryant [Bol14]. bubble [zSfNbL11, YWL15]. bubble-sort [zSfNbL11, YWL15]. budget [Gou15, HKT17, TL12, Vou19]. budget-constrained [HKT17]. Buffer [ABS13, CXD^{+}13, SCB13, SF13]. Building [BBR^{+}18, DM13, Fre10, VB19]. bundles [LZ15]. burnt [IK10]. Burrows [MN19, DGG^{+}19, SKN11]. burst [RV17, SCB13]. business [BBR^{+}18]. byte [LKR17, RV17]. byte-oriented [LKR17]. Byzantine [BIR16, DP12, IS10, TVB15].

C1P [MCS12]. cache [LMSN16]. caching [YBM^{+}10]. cactus [Dat15]. cactuses [AH11]. cake [Brâ15b]. calculi [Ind15, SSSM11]. calculus [YHLC12, vG18]. call [DJRB15, SSSM11, ZCT^{+}12]. call-by-need [SSSM11]. Camellia [Blo15]. Camellia-192 [Blo15]. camera [JSZC15]. can [Att17, Bon13, DF11, FIPS11, JS19]. candidate [CY18]. cannot [AIR17]. canonical [Bar10]. capability [SL19]. Capacitated [CPW11, RX17]. Capacity [WDH16, PS19b]. capped [Gou15]. CAPTCHA [KC17]. CAPTCHA-based [KC17]. captures [AT15]. cardinality [BES17, CKY15, Pet12]. cares [CP10, NR17, ZZH10a]. Carlo [Lof14]. carry [GK10]. carry-less [GK10]. Cartesian [SSW16, Che11, DM13, LJL18, LZCM10, SSW15, TWZ17, XS15, ZLCM10]. cascades [Cha11]. cascading [NLZX14]. Case [RK15, ARdSP15, KN13a, KM13, SS17]. cases [AAB^{+}19, SI18, WS16]. cash [BB15]. causal [CZZ^{+}10, MPRC19]. Cayley [CLS13, LM15, YLM10, YL11a]. CBC [ZWWL12]. CBC-MAC [ZWWL12]. CCA2 [Gal13, ZY17]. CCA2-secure [ZY17]. CCM [SKK10]. cell [YWC11]. cells [KPR18]. cellular [ADF13, ZCT^{+}12]. Center [ABS12, GHL18]. Center-based [ABS12]. centerpoints [AG15]. centers [SLL13]. centralities [LLT14]. certain [Ama10, CZ13, Tur12]. Certificateless [YW14, ZY17]. certification [CRTU15]. certified [TVB15]. ChaCha [NA19]. chaff [NTD16]. Chain [EAA^{+}16, LP13b, DKZ18, FHvtH^{+}15, GWL18, KO16, MO12, RDX13]. chains [AAOW15, CHK13, GXZZ13, GP17, Lu15]. chair [BG12, BG15]. Chaitin [BHLM13]. challenges [BVF12, JF15]. changing [BL10, Rac10]. Channel [CK11b, ABPS15, JF15]. channels [DDPBT11, TK15]. Characterization [DRdSS12, GLY11, Pel15, AAC^{+}10, BGL13, DEL10, LN11, LX17, MS12]. characterizations [Dur12]. characterized [Koc18]. Characterizing [KHKS16]. chasing [WYLP17]. Chebyshev [LPdS10, LMCG16]. Checking [GQ17, CHK13, DLMV10, Dem18, DBB12, GXZZ13, JUY15, NAD19, SC12, YYK17]. Checkpoint [XMLL11]. checkpointing [CZZ^{+}12]. Chen [CL10]. Chinese [GWL18, HF14, LCC17, PYW^{+}13]. chip [BGP12]. choice [LLP^{+}18, MT18, WMLN10]. Choosability [GHvtHP13, CHR10, Zha10b, Zha13]. choosable [Maß15]. chord [CG10a]. chordal [AKH16, BP11, BFLM15, CGLS10, KP13, LN11, Lei17, PP10, PG12, WLW11]. chordal-bipartite [CGLS10]. chordality [DEG^{+}12]. chromatic [AH11, BS17, DGŚN15, DN12, DS14, FH12, HRW11, JSR10, Mat19, WWW18, XZW15]. cipher [KDH15, LC13, LYHH14, MNP12, SB19, Sar11, WWBC14]. ciphers [KM10, SKK10, WB12]. Ciphertext [KA17, XWLJ16, XWS17]. Ciphertext-only [KA17]. Ciphertext-Policy [XWLJ16, XWS17]. circle [HMS16, NG10]. Circuit [Rud17, XWS17, XWLJ16, DKKY10, IM17, Jan12a, VSP18, ZDQ^{+}17]. circuits [ASTD14, Ale13, Ber11, GJ11, HP18, Mor11, Pod12]. Circulant [vzGS15, CT16a, Den14, GSR^{+}14]. Circular [AAB^{+}19, PS19b]. circumference [FL13b]. clairvoyant [Sun19]. class [ABS13, Che10c, EAE21, GSR^{+}14, GNG11, KP13, MY18, Pas15a, Pas15b, SWF18]. classes [AH11, Dar15, IL12, Joh14b, KZ12, LV15, Mos13b]. classical [Bar13]. Classification [MHIU18, SLdAMP17, SaBG17, MS15]. classifier [GE12, PR17, STAR15]. Classifying [GLPR19]. clauses [SR11]. claw [CZ16, Hua14, LS15, LSK17]. claw-free [LS15]. claw-heavy [CZ16, Hua14]. CLEFIA [TS16, WB12]. CLEFIA-type [WB12]. Clique [Cou14, Juk12, Wid17, Iba17, LS11, LS15, LSK17, LV15, Val10, YL17]. clique-covering [Val10]. Clique-heavy [Wid17]. clique-independent [LS11]. clique-separator [Iba17]. clique-transversal [LS11, LS15, LSK17]. Clique-width [Cou14]. Cliques [Lau18a, BP11, BKN^{+}12, Kut12, VML18]. CLL [ZZZ15]. clones [Lag14]. closed [BK13, CGG14, Dar15, MDB14]. closest [DGKS14]. closure [HS15, MDB14, vB12]. Closures [SSS10, Vág18]. Cloud [XWLJ16, ZJ16, AT18, PCK10, WZS^{+}18]. clouds [Jia11]. Cluster [Man10, BD11, BDNVP15, LSS15, TZF16, WZQH16, XPC^{+}10]. clustered [BL12, BLYL17]. clustering [ABS12, DD14, JKY15, LFZJ14, LLWH13, SD19, ZD18, ZXJ^{+}11]. clusterings [QK15]. clusters [BLM10, Dam16, SLL13, ZZJ11]. CMAC [SKK10]. CNFs [BT16a]. co [BG12, BG15, Lag14, Mos13a]. co-banner [Mos13a]. co-chair-free [BG12, BG15]. co-clones [Lag14]. coarse [LH10]. coarse-grained [LH10]. coarseness [DBFMPLV17]. cocomparability [DM19, KM16a]. Codd [KL10]. code [CWW10, GMS19, MNV11, Mil15, TX11, ZLM17, ŽGH10]. coded [ZST13]. codes [AAB^{+}16, BB11, GWL18, GM12b, Har18, HHL15, Jia18, Kle13, LZ15, Mol18, RV10, RV17, SQLS17, VV11, Vid13, WS10, KR16]. coding [Bae10, KK15b, LLP^{+}18, Maß15, SPdR13]. coefficient [SWW16]. coefficients [AH17, LT18]. cographs [JPB19, PP10]. coincide [AIS10]. collaborative [GWJ11]. collapsing [SO10]. collection [FC14]. collective [SZC^{+}17]. collision [AKY13, WS13]. Collisions [VNP10, LT13a]. Collusion [ZJ16, WZS^{+}18]. colony [GL11]. color [BKPP18, FX11, KMK18]. color-spanning [FX11]. colorability [Koc18, WLW11, ZW10]. colorable [Hoc11, San11, WX13, XLZ16, ZLSX16]. Colored [SGM13, AADB^{+}18, OM17, Tsu18, WD11]. Colorful [PT12]. Coloring [BGP12, Rom11, AD18, BJH15, BGR13, BDNVP15, DS16, DX10, FKL^{+}11, FS13a, FL12, FL13a, FM11, FA17, Gol18, GJ14, Hal10, JT15, KSK13, KLMP18, LRS11, LSZX15, MSM14, MKI11, NS11, RZ10, SSW15, SSW16, Shi18, SWZ12, WW11, XXZ14, Xu11, YeCM14, LM11]. colorings [CL15, Cza13, LZSX17, Szy12, YYR16, ZL11, ZW11, ZL12, ZZ18]. colors [FA17]. coloured [AC12, vGLM12]. colouring [AH11, DP18, GLPR19, HOV13, Lei17, Qi17, SK12]. colourings [Feg19, Fie11, Rac10]. combination [SWLX15]. Combinatorial [SO19, ADF13, FMHL11, KKSŠ11, Lib10]. combinatorially [Ivá16]. combinatorics [KTV13]. Combined [HYYZ15, MK14]. Combining [MZQL14, YLLL16]. comment [CWYP14, Jha15, THS12]. Comments [BB11, KMS19, LMU15, Par11, WZS^{+}18, XWS17]. commitment [ZC12]. Common [FKR^{+}16, Pol18, SW18, AK14, BBDS12, BVDP10, DG14, Dur13, FH10, FGKU15, Gra15, IH18, Kos17, MM19b, SP18]. common-multiplicand [SP18]. Communication [AJLM11, SCL^{+}11, JF15, Juk12, Kor12, ML10a, Shr18, Shr19, YYZH13]. Communication-efficient [AJLM11, SCL^{+}11, ML10a]. communications [BB11, RV10]. communities [WDH16]. commutativity [Nan15]. Comorphisms [Tut13]. Compact [ZZH16, FG14, KR10, YiN10]. compacted [BV11]. comparability [Och17]. compare [BVF12]. compare-by-hash [BVF12]. Comparison [KS19, YYDL11]. comparisons [ACK11, HI11, SM17]. compartmented [EZ15]. compatible [KPSZ11]. Competitive [NZX19, AFPT10, SM13, THS12]. compiler [LWS10]. Compiling [CR10]. complement [JLMO17]. complementary [CCCN19]. complemented [JLMO17]. complements [PP10]. Complete [AdFEGRI11, AGI15, AHK^{+}17, BH17, Che11, Che17, CG10b, DFdFT16, DEL10, GN10, GI18, GZM15, HWA12, KM12, ML10b, NN17, Och17, Par16, Pér17, SeF14, WWZ15, YeCM14, ZST13]. completely [HL16, PC19]. Completeness [Luc15, GNV14, LZSX17, Mil19, Rom11, SW12, YWWW14]. completion [CK12b, FL16, HKW14, Kun17, LZ14, LH11, MY13, YW12, ZLZX19]. complex [AABB17, CNKS15, LLT14]. complexities [TWZ17]. Complexity [ALT18, Csi12, DW10b, Ivá16, LHW^{+}16, Pul16, Xu12, ACK11, AD13, ALT15, AT15, BBKS17, BBDS12, BvK15, Bol10, Bol14, Bor16, CCRS14, CHK13, CLY11, DS15, DS16, DKKY10, DW12, DCH12, DKC12, Duc18, EP16, FS14, Fin15, FG18b, GGI11, HK16, HKT17, IM17, JKS10, JT15, JS18, Juk12, JSR11, KZ12, KSD18, KLMP18, KP12, KPSZ11, Kra12, LZ10b, LT13c, LFXH17, Lof14, LXJ^{+}14, MMP15, MM15a, Miy14, MOS16, Mon10, Mon12, NVB15, PS19a, Rac10, RDX13, Kar17, Sal12, SS12, Sch13a, SDM14, Shi12, Shi18, Shr18, Shr19, Sto16, Vay13, VSP18, Wij10, WXCK19, YA13, YYK17, YKD^{+}12, Zha10a, vBCH^{+}15]. component [CGJY12, HYYZ15, MK11, NT14]. components [BKMT14, CC14, CNKS15, Pea16]. Composable [CD10]. Composing [LABKS17]. composition [YeCM14]. Compositional [FOSC18, Mos13b]. compositions [VV11]. compound [CX18]. compress [LC13]. Compressed [CD10, BGI^{+}12, SSS12, Fay16]. compressible [BFKL13]. compression [Aku10, KN13c, LK14, LKR17, LRS19, RK15, YA13]. compressive [YLGC16]. compromise [jDX11]. compromise-tolerant [jDX11]. computable [FWS13, Tys13]. computably [BHLM13]. Computation [ADKM12, BJ18, CIK^{+}13, DM16, FB10, GP17, KKSŠ11, LXX14, Nis92, QK15, SS19a, Sim16, TXQ11, WJS10]. computational [Car19, HRS13, JT15, SDM14]. computations [Bee16, CZ13, Gue12]. compute [Bon13, KM16a]. computer [MN15]. Computing [AADB^{+}18, Ber12, BHL17, CCF^{+}12, CT11, EH18, FHvtH^{+}15, FX11, FIV15, Iba17, KP13, Kob15, Kos16, LT13b, LT15a, MMM11, MTT12, TLWZ16, XWLJ16, Ale13, Ama10, BKZ15, Bir11, Bor16, CCRS14, DLRS14, Duc18, EN17, GS10a, JT16, LQL^{+}17, LT15b, MOW17, NG10, TLL16, TJ18, vIL13]. concept [Dar15, PYYC16]. concept-based [PYYC16]. concerning [KK11, Nin15]. concise [FL13a, CD10]. concisely [RJS^{+}10]. concurrency [DMS11]. Concurrent [ZC12, Att17, CVZ11, XLWZ16]. condition [BIR16, Cai15, CLZ18, HL16, Hua18, MY13, Nin13, TD14, WZL12, YH19, ZSY13]. Conditional [YLM10, YYDL11, ZX10, ZX11, CLQS12, CHF15, GXZZ13, GNV14, LWL11, LTK^{+}19, Ria19, WH16, YLG10, YZZ17, ZXY10]. conditions [Dam16, Shi18, YZRC18, YL16b, ZXS19]. configurations [BGP12]. conflicting [FT15]. Conflicts [JK18, Sun19]. confluent [SO10]. congestion [AKMR19, Che18b, FT15, MM15a]. congruence [KKS15]. Conjecturally [Tys13]. conjecture [Bol14, FL12, GV14, GR13, GR14, Lü19, Nin15, PPH18, zSfNbL11, Wu13, YKD^{+}12]. conjunctive [Ksh11, SM17, SV19, Wij10]. connected [BKMT14, BvdZ19, CLZ18, Cha11, CO13, DFMS10, GHK^{+}18, GLY11, JT15, Jor12, KP13, Koc18, LZ10a, LLW15, Mou19, NT14, Nut18, PP10, PI13, PI19, Pea16, SS12, Tak16, WZL12, Wid17, YL16b, ZF10]. connectedness [DJ15, GS12]. Connecting [SS10a]. Connection [LM15, DD14, Jan12a, Kle13, RPL19]. connections [GHK^{+}18]. Connectivity [CCYW16, CV19, CBSV11, DEG^{+}12, EKN11, GH14, GQG10, HL12, Ili13, IN12, KVK18, KLS10, LLL18a, LJL18, MZ11, Man19, NFW15, Nin16, Sch13b, WSW12, XPC^{+}10, YLM10, YZZ17]. coNP [Ama10, Che17, Par16]. coNP-complete [Che17, Par16]. conquer [AR13a]. consecutive [CCL10, KKR^{+}13, HCCG15]. consensus [SCL^{+}11, TLL18, VLV15, APR13]. consistency [CLR13, EGK13, LABKS17]. consistent [KP12, Wij10]. Constant [BK10, FG18b, Yam14, CVZ11, CO17, EFKR10, HLT10, LGT17, MMKK18]. constant-round [CVZ11]. Constant-space [Yam14]. Constrained [FR12, AGHY12, AMP13, BF17, BVDP10, Deo12, HKT17, Kis19, FR12]. constraint [Fuj16, GLMS18, GHK11, HHK17, Las17, Mar11, ZZ11]. constraints [GS17, KM13, LL13, MU15, Vou19, YYR16]. Construct [Han17, WT13]. constructed [Har18]. Constructing [GY19, NII^{+}15, LLL18b]. construction [BHM^{+}10, BJMC17, Che11, EGKL11, Joh14a, Lai16a, LNP11, LGT17, YGK12]. constructions [SLF19]. constructive [WB12]. consumption [AMN^{+}10, Oro11]. contacts [STAR15]. contagion [SF17]. contained [Tel19]. containing [BC15, MM19b, Zal11]. Containment [SM17, ADG10, ACK11, Par16, SV19]. contains [FC14]. context [CÉ13, CVV10, Ési11, EGK13, Fil11, JOS10, Luc15, RS10, SS18, VB19]. context-aware [VB19]. context-free [CÉ13, Ési11, EGK13, Fil11, JOS10, RS10, SS18]. context-sensitive [Luc15]. contextual [SSS10]. contiguous [LL13]. Continuous [CYH15, Che18b, BF19, CW15, GXZZ13, yJxW16]. Continuous-time [CYH15, GXZZ13]. contour [GHC15]. contractible [DFMS10]. Contracting [DP17]. contraction [Cou14, SS19b, GM13b]. contrast [BL10, LWL10]. control [BD16, BLM10, BNRC10, BT16b, LZ10b, Ria17, ZCT^{+}12, MGNAB11]. Controller [PP14]. Convergence [FT15, LXT19, ZDQ^{+}17]. convergences [SSS10]. convertible [LHH11]. convex [BV19, Che18b, IM12, MMM11, Oro11, STU12, Tiw14]. convexifications [ÁCLS12]. Convolutional [LCC17]. convolutions [FB10]. cooperative [RX17]. Coordination [GL13, LYF11]. Coppersmith [Dra16]. core [ZLS^{+}15]. Cores [BWZ12]. coritivity [ZLS^{+}15]. Cornaz [BGR13]. Correct [EAE21]. Correcting [Shp13, RV17]. correction [AW12, SSZW16, Sun16]. correlated [Jia16]. Correlation [CP16, DM10, yJxW16, KYC13, WWBC14]. correlations [Sar14]. Correspondence [Fin15, BDNVP15, SC12]. corridor [Xu12]. Corrigendum [GR14, KMN18, LPWZ14, Pas15a, SSW16, Zha13]. corruption [Shr19]. coset [Las17]. cost [ALT16, CDPB^{+}10, GLO12, HKT17, KT11, KT12, LLG14, MY13, YL11b]. costly [DF11]. costs [Bae10, SLC15, XZBX16]. coteries [Jia11]. COTS [AMN^{+}10].

COTS-based [AMN^{+}10]. count [ARdSP15, Bab17, JLdSO^{+}14, LR10]. Counter [GK10, Fay16]. counterexample [Bol14, GR13, GR14]. Counterexamples [Ket11, SSSM11, YKD^{+}12]. countermeasure [GJ19]. Counting [BK13, DM19, Elm15, GPR10, LS14, PT11, AGZP15, DTS15, DFRS13, Kut12, LH10, LC17, Mil19, PT12, PRM16, YK11, ZST13]. counts [DEG^{+}12]. coupon [FC14]. coupons [FC14]. Cover [BI14, AK14, BT16b, FGG^{+}10, FLRS11, Kat16, Kob15, Lam11, MM19b, Mor11, PI19, TZ11, TY13, Tu15, ZLS^{+}17, CP13, GMY13, Man19]. coverage [CdA13, GM15, Shi13, YTYZ15, ZC10]. covered [YH19, ZXS19]. Covering [Ber17, BHKK11, GWL18, KK16, KPR18, LWŻ07, LPWZ14, MÖ15, Val10, WL11b, WL12, WCJ15, ZL11]. covers [BFRV15, Che10c, Che12, Che16c, LL10c, PI13]. CP [EN17, TY16a]. CP-rank [EN17]. CPL [YWWW14]. Crash [MPRC19, AJLM11, FCLCR17, ML10a]. crash-quiescent [AJLM11]. crash-recovery [FCLCR17, ML10a]. Crash-tolerant [MPRC19]. CRC32 [Gue12]. CRC32C [Gue12]. Credibility [SWLX15]. credit [Pér17]. criterion [Mat12]. critical [FM19, Liu10, ZJ11, ZXS19]. criticality [CSX16, EGK13, RZ14]. cross [SS17]. cross-matching [SS17]. crossed [BJ18, CV19, CWYP14, CZWP17, DY10, DZFY12, NFW15, Nin16, PCY16, YDT10, ZHX13]. crossing [Bra18, DKL^{+}12, KT16]. Cryptanalysis [MZ15, SM10a, SM10b, TY16a, WWYY11, WYW14, AP11, BMB16, DMSD18, Kar13, KDH13, LFW^{+}16, SDM14, Sun11, WWBC14]. cryptographic [CR10, MMZ12, MM13, Mes15, PLPW13, SLF19, WT13]. cryptography [HH15, LWL10, VN17, LZJX10]. cryptosystem [ACD18, Gal13, GV14, MM13]. cryptosystems [FWS13]. CSAT [Mat12]. CSL [GXZZ13]. CSP [FOSC18]. CU [CWY15]. cube [BJ18, CCH14b, DY10, DZFY12, GH14, GZJ^{+}12, LWL11, NFW15, Nin16, PI19, YYYZ12, YYZH13, ZFJZ11]. cubes [AAB^{+}16, CV19, CWYP14, CCYW16, CZWP17, CN18, Fu10, LYHC10, Mol18, PCY16, PC19, PI13, YDT10, YWY12, ZHX13]. cubic [LS11, MOW17, Tak16, TY13, XXZ14]. cubic-time [MOW17]. Cuckoo [RPL19]. CuCoTrack [RPL19]. Curry [SeF14]. curve [FWS13]. curves [HZX16, Moo11, RS12b, TXQ11, ZHXS10]. cut [BES17, CCF14, GHK^{+}18, Ind15, Rus18, Sax10, Zel11, Zha18]. cuts [Ber12, KLS13, KR14, Xia10]. cutting [Brâ15b, JSO11, Juk12, LT18, M.15]. cutting-sticks [M.15]. cycle [CM17, CZZ^{+}12, DY10, DX11, GBH12, KN13c, LL10d, SS19b, Tsa11, WLW11, XL15, Yus11, Zeh16]. cycle-embedding [Tsa11]. cycle-radius [GBH12]. Cycles [Che10a, AMMY10, Cai15, CHF15, DP17, FL12, Hou10, Hua14, Hua18, JSR10, Kot12, LZCM10, LW19, MCS12, Nin15, WX13, XS15, YZRC18, ZLCM10, Zha10b, ZW10, Zha13]. cyclic [Bar13, LH10, SQLS17]. Cyclically [HN10, vZBSY16]. cyclotomic [KZ12, WXCK19].

D [EAA^{+}16, JSZC15]. DAA [CL10]. daemon [CCT14]. dag [LRS19]. DAGs [PWC^{+}15]. Dantzig [CFMT16]. dart [KS12]. Data [CZD14, SL19, ZJ16, BL10, BVF12, BOV15, CdA13, DG16, DDPBT11, ED17, FKC13, GL15, GS10b, GHL18, HLY11, IMCP15, KR10, KZP10, KWH16, SLL13, Tsu18, Vrg15, WZS^{+}18, ZMZ19]. data-flow [CdA13]. data-link [DDPBT11]. databases [GHK11, KN13a, KRV16, SS17]. dates [OZL16]. DCell [Lü19]. DDoS [AT18]. de-parallelization [YEMR19]. de-randomization [DTS15]. Deadlock [GS10c]. Deadlock-freeness [GS10c]. decay [SWLX15]. December [Ano19a]. Decidability [GS12, CS16a]. decidable [BE16, CYH15, FG18b, Oan11]. Deciding [BJKZ14, Für19]. Decision [SLdAMP17, BCS15, CWY15, GJR10, JKS10, SLC15]. decoding [Vid13, WS10]. decomposing [STMD18]. Decomposition [MRZ10, Bar10, CT16a, NHK15]. decompositions [EN17, SS19a]. decrease [RR16]. decrementation [CGG14]. decryption [SM10a, SM10b]. Decycling [LRSS17, LKF15]. DED [BB11, RV10]. deduplication [BVF12]. Deep [LCC17, YLZZ18a, YLZZ18b]. defect [MZQL14]. defense [AT18]. defined [nXlCL14]. defining [MMKK18]. definite [AMT12]. definition [LWL10]. definitive [DMS12]. defogging [GPT16]. degenerate [BS17, DGG^{+}19]. Degree [CFJ12, HL16, YZRC18, ZXS19, BM19, Cai15, CLZ18, CQ12, Dam14, FGvL11, Hoc11, HKW18, Hua18, JT15, LS15, Man19, MRZ10, Nin13, Rac10, RZ10, SW18, WW11, WLW11, YH19, ZW10]. degrees [MS18c, Zha11]. Delaunay [AP14]. delay [BEFP11, WDH16]. delayed [LMCG16]. Delayer [BGL10]. Delegation [XWLJ16, XWS17]. deleted [ZSY13]. deletion [BD11, BDNVP15, CGJY12, Tsu19b, CP13, LFW^{+}18]. delivery [LYF11, LZCX12, TFY11]. demand [AHLS18, NLXX15]. demand-aware [AHLS18]. demosaicing [YLGC16]. Denial [Bee16]. dense [Szy12, VSP18]. densest [BK10, CCH^{+}14a]. Density [Did13, AT18, BI14, SH17, SS10b]. dependence [AAC^{+}10]. dependencies [PH11]. Dependency [CZZ^{+}10]. dependent [FM18, GM12a, JC10, KMS19, KY12, LZCX12, MKI11, Mos11b, SLC15]. deposits [Bee16]. depreciable [ZZXL11]. depth [CWY15, GJ11, JSZC15, MS18c, PPN^{+}17, Pod12, Suz18]. depth-first [PPN^{+}17]. derivation [YYK17]. derived [DCH12]. deriving [CLQS12, GNV14]. descents [Che18b]. description [XZ18]. Design [PGZB19, MZQL14, VLV15]. designated [SY15a]. designs [AHLS18, GD13, KDH15]. destination [Kar17]. Desynchronization [ASA13]. Detecting [FLPS15, RS10, SML^{+}10]. detection [CWW10, Ksh11, PYYC16, SCL^{+}11, ZLM17]. detector [BR10b, MG16]. deteriorating [LJ15, MZC11]. deterioration [LL10a, YZH14]. determinacy [Had18, Le 19]. determinants [Bir11]. determination [JK19]. determine [KSBT13]. Determining [MTA10]. determinism [KZP10]. Deterministic [Cha13, DP12, ZL18, BN10a, BDPP18, FS14, FV13, JKS10, KK15b, KP14a, Kos18, NLZX14, SSSM11, TK15, YYK17]. Developing [Far10]. developments [SeF14]. DFA [Val12]. DFS [JLMO17]. DHA [AKY13]. DHA-256 [AKY13]. diagnosability [Che18a, CLQS12, CLS13, GHL18, LTK^{+}19, Tsa15, WH16, YYDL11, ZX10]. Diagnosable [TC17]. diagonal [Pie15]. diagrams [CV12, LT13c, MK14]. Diameter [CFJ12, AR18, BDF^{+}18, BK18, Che10b, DGR15, FX11, KMK18, ML10b, WL12, ZFJZ11, Zun18]. diameters [PC19]. dichotomy [BR12, BM19, DFdFT16, KP12, Mar11]. Did [Kar17]. difference [Li15, Mat19, SWF18]. different [DKNQ18, Ger19]. Differential [DMSD18, LSLY11, LYHH14, Sun11, ZW14, Blo15, Bog10, GLMS18, MNP12, SB19, SDM14, SPdR13, TS16]. Differential-linear [DMSD18]. differentially [AS18]. differentiation [LH11]. Diffie [RH10]. Diffusing [WWBB15]. diffusion [AFPT10, GHC15, MG16, SM13, THS12, WB12]. diffusion-driven [MG16]. digit [KWH16]. digraph [Bor16, GZJ^{+}12]. digraphs [AALL16, AKMR19, BJH15, BGP12, Cro15, JLMO17, LX13c, LKF15, LJL18, Nin15, PT11, RT13, SZ18, XZW15, CLM12]. dihedral [YL11a]. dilation [SI18]. Dilemma [Pel10]. Dillon [BPRMS14]. dimension [FHvtH^{+}15, Gil14, GSR^{+}14, HM13, Lai16b, MSZ11, Mos11a, ZLNB19]. dimensional [IK15b, Jan12b, KS10, LL10b, LL10c, PYHA10, PYTH11, QD16, RARM12, RV13, TNN11, WL11a, WH16, YBMK15, ZHXS10, Zho16]. dimensionality [GL15]. dimensions [Jor12]. Diophantine [Tys13]. direct [CBSV11, EGKL11, JKS10, LZ15, Mei18]. Directed [BBK17, CGLS16, DJ15, FH10, FLP^{+}18, GKPP16, KVK18, Kob15, LS14, LT15a, LZCM10, MMS15, SY15b, Zeh16, ZLCM10, WWBB15]. Direction [Win13]. Direction-reversing [Win13]. directional [JSZC15]. directions [Ang17]. discarding [DJS13]. disconnected [MK16]. discount [ZPX11]. discounted [ZLZX19]. discrepancy [Aka14]. Discrete [Rud17, CR18, CHK13, Che18b, Fil18, FIV15, yJxW16, KZP10, LFXH17, MM13, Mes15, TPL16, Nan15]. discretely [PYHA10]. discretely-scaled [PYHA10]. discriminant [ZpH15]. discriminative [HYYZ15]. diseases [VB19]. disequalities [SR11]. disequality [HHK17]. Disjoint [LS10, Che10c, Che12, Che16c, KKKS14, Kot12, KN13c, Lai16a, LYHC10, LX19, LL10c, LW19, PI13, PI19, Pul16, Sch15, SW19, YZRC18, Yan14, ZGL11, Zha11, ZZ14]. disjunction [MMP15]. disjunctive [MU15]. disk [DDK^{+}15, DGR15, NPR17, TX11]. disks [Shi13]. disparity [FBD18]. displacement [KK16]. disproving [Li15]. dissecting [YK13]. distance [AIR17, AD11, CR18, Fil18, GWL18, Hru12, JF15, Joh14a, Kun18, Luc10, LEP10, LP13b, MN19, Mil15, MOW17, QK15, Suk19, TLL16, TLL18, WL11b, WP11, XZ17a, Shi15]. distance- [Joh14a]. distances [AG19, Ars18, Hil14]. distinct [FLPS15, GPR10, ZYL15]. distinguishers [AY14, AP11]. Distinguishing [KM10, CL15, HYC12, YeCM14, YYR16]. distortion [HM10, KM12]. Distributed [Sch14, BVF12, CCT14, Far10, FKC13, FCLCR17, GS12, Kak15, Ksh11, LQL^{+}17, TNN11]. Distribution [EAA^{+}16, CKP16, LLP^{+}18, MP13, YL11b]. distributions [MHIU18, PPN^{+}17, Suz18, XTTH12]. divide [AR13a]. divisible [HM18]. Division [HZSL05, Bir11, MN14]. division-free [Bir11]. DNF [Lau18b]. DNR [MS18c]. do [SW18, Tys13]. document [Tsu13]. documentation [MK11]. documents [YLLL16]. does [AK13, CS16a]. DOM [Ama10]. DOM-trees [Ama10]. domain [BBR^{+}18, Str16]. domains [Gil14]. domatic [TLWZ16]. Dominating [BdSL15, NS11, AKH16, BU17, BYK14, BM19, CCT14, Den14, GG18, HHL15, Joh14a, KP13, Krz13, KM16b, Kun18, LZ10a, LMS14, LLW15, Nut18, RS19, RR16, SS12, Slí12, SW19, Suk19, Waw14]. Domination [MC15, ZLCM10, AR18, ALT15, BL14, BFLM15, BH17, CCKP13, gCF15, CPW11, DW10a, Den14, DLN11, FM19, Jha15, JPB19, KM14, LX17, LZCM10, PP13, Pra12, Sie19, VK18, XS15, ZLJS18, ZS18]. dominations [ALT18]. don’t [CP10, NR17, ZZH10a]. Double [ZLJS18, BS11, TX11, LST11]. Double-Base [LST11]. doubling [MSZ11]. Doubly [FR12, Laz10, PP10]. Doubly-Constrained [FR12]. doubly-exponential [Laz10]. down [LTK^{+}19]. Downhill [gCF15]. Drawing [BBD^{+}12, Fra10]. drawings [Ang17, CY17, DKL^{+}12, Did13, LM16]. Drazin [ZQJ^{+}15]. driven [MG16]. DSA [Dra16]. DSS [JK18]. dual [Har18, SQLS17]. dually [BFLM15, Lei17, PG12]. due [GM13a]. due-window [GM13a]. duplicate [CJ11]. duplicating [CGR14]. duplication [BV10]. duplication-random [BV10]. durations [VML18]. Dynamic [DJRB15, GS19, HPY^{+}10, HH15, ZJ16, AC10, JS19, KR10, OG11, RK15, RJS^{+}10, WWWZ13, WZS^{+}18, XL15, YBM^{+}10, ZLM17, ZZJ11]. Dynamics [LLWH13, Che16a, FT15, ZKXY10].

e-cash [BB15]. E-passport [LZJX10]. E0L [DK14]. EAC [LZJX10]. each [FC14, WCW11, Yus11]. ear [SS19a]. earliness [MM15b]. early [DJS13, HLY11, LZLY12]. easier [KL18, ZDQ^{+}17]. Eavesdropping [MP13]. EC [Dra16, Bae13]. Eccentric [DEG^{+}12]. eccentricity [MWZ12]. EDCA [YWcW^{+}14]. EDF [DJS13]. Edge [Che10b, FS13a, Fu10, KSK13, LX13a, LW19, WD11, BJH15, BT16b, CBSV11, CL15, Cou14, DS16, DX10, EKN11, Fie11, FM19, GLPR19, GG18, HOV13, LYHC10, LX19, MG16, MK16, Maß15, MKI11, Mou19, NT14, OM17, SS19a, Sch13b, Sun16, Tak16, VK18, WZL12, XN12, Xu11, Yan14, YeCM14, YL16b, ZL11, ZW11, ZL12, ZF10, BBKS17, LM11]. Edge-colored [WD11]. Edge-coloring [FS13a, DS16, MKI11, LM11]. edge-colouring [HOV13]. edge-connected [YL16b, ZF10]. edge-connectivity [EKN11]. Edge-disjoint [LW19, LYHC10, LX19, Yan14]. Edge-fault [LX13a]. Edge-fault-tolerant [Che10b, Fu10]. edge-vertex [VK18]. edge-weighted [BJH15]. edges [Ber17, Che10a, Che12, Che16c, CLS13, CHF15, DFMS10, Fie11, MBW19, Zha18, ZLLZ18].

Edit [Shi15, Dam16]. edit-optimal [Dam16]. editing [BD11, LSS15, Man10]. Editorial [Ano18g, Ano18h, Ano18i, Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano10m, Ano10n, Ano10o, Ano10p, Ano10q, Ano10r, Ano10s, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano11n, Ano11o, Ano11p, Ano11q, Ano11r, Ano11s, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano12r, Ano12s, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano13n, Ano14a, Ano14b, Ano14c, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano16a]. Editorial [Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18j, Ano18k, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m]. Effect [NB12, CC14, WB12]. effective [CS16b, Tsu19a]. effects [DK10, FT15, JC10]. efficiency [Bog10, SLL13, SLS16, Vou19]. Efficient [BL10, BYK14, CC12, CPHS18, CPTZ13, DGG^{+}19, DE14, Den14, DG14, GHK11, GK10, HZSL05, Ili13, IL12, yJxW16, LLP^{+}18, LZ12, MV13, MN14, OG10, PYHA10, Sim16, WJS10, XLWZ16, ZHXS10, AR18, AJLM11, BGI^{+}12, BL14, BFLM15, BM19, CJ11, CdA13, CIK^{+}13, jDX11, HM18, HHJ^{+}12, eSKAI10, Lee10, LKC^{+}12, LH10, ML10a, Mes15, Pea16, QX10, SCL^{+}11, TLL16, Tat19, Tur13, VN17, XTTH12, YA13, YiN10, YMSA14]. Efficiently [FWS13, MT10]. eigenvalues [XZ17a]. eight [Cal13, RZ10, Sun11]. eight-regular [Cal13]. eight-round [Sun11]. election [BEFP11, FCLCR17, LZ10b, MS19, Vay13]. elegant [ZYC13]. element [DZ11, Zal11]. element-wise [DZ11]. elementary [YGK12]. elements [DT10]. ELI [KD13]. Eliminability [Ind15]. elliptic [Moo11, ZHXS10]. ellipticity [ZZ13]. embeddable [WWLC14]. embedded [AHR10, ABPS15, AMN^{+}10]. Embedding [CHF15, DY10, DZFY12, RRR12, YDT10, BR10a, CZWP17, DX11, GL15, KM12, Tsa11]. embeddings [BR10a, HM10]. emerging [DWQ10]. empirical [Sar14]. employed [MGPI12]. emptiness [HHK17, Lan11]. empty [NMB10]. emulators [HP19]. enciphering [Sar11]. encoding [FG14, JS18, SPdR13, Sun16, YiN10]. encodings [GSS16, vG18]. encrypted [ED17]. Encryption [XWLJ16, CW12, Fay16, HLR11, LM91, LPdS10, LHH11, Mes15, Sar11, SY15a, TPL16, WWYY11, XWS17, YL11b, ZY17]. end [WBC^{+}12]. endomorphism [FWS13]. Energy [VN17, AMN^{+}10, AZ14, HM18, NB12, Pér17, SO19, SLL13, YA13]. energy-aware [AZ14]. energy-efficient [HM18]. Enhance [NTD16, ABPS15]. Enhanced [CD19, PP14, AC12, JZ18, TC17, YW14]. enhancement [CGLS10]. Enhancing [YBM^{+}10]. Ensemble [MS15, HHL15]. ensuing [SS17]. ensures [CZZ^{+}12]. ensuring [Nin13]. entanglement [Ito14]. entanglement-resistant [Ito14]. Entropy [BI14, CA17, DKZ18, JCC11, Tsu19a]. entropy-based [JCC11]. enumerable [BHLM13, CVV10]. Enumerating [AKH16, Dam14, VML18]. Enumeration [CGLS16, FLMQ10, Lau18a, Mat15, YL17]. envelope [Lu15]. environments [SCL^{+}11]. envy [Brâ15b]. envy-free [Brâ15b]. epidemic [YC19]. EPTAS [KS10]. equal [EO13, FPY12, LZLY12, LX17]. equal-length [FPY12, LZLY12]. equality [Für19, Kam19, Li15, SSS12]. equation [CYQ13, YCL11]. equations [Che16a, LXT19, VSP18, XLT^{+}19, ZZZ15]. equidistribution [Vaj18]. equilibria [DHW13, MM15a, PPN^{+}17, THS12, WDT13, SM13]. equilibrium [CJ11, WCJ15]. equimatchable [DE14]. equipment [ZZXL11]. equitable [FKL^{+}11, GZM15, NN17]. Equivalence [Dem18, SSS10]. Equivariant [Las17]. Erdos [FL13b]. EREW [Sor10]. Erratum [HAM11, KT12]. Error [CHP19, CZD14, GHC15, HKK12, KSD18, WXCK19]. errors [RV17, SLdAMP17]. establishment [TC11]. Estimating [BKMT14, AMN^{+}10]. estimation [LLT14, LQL^{+}17, MZQL14]. estimators [CKY15]. eternal [FGG^{+}10, RS19, BdSL15]. Euclidean [BLC10, GWL18]. Euler [DCH12]. Eulerian [CGJY12, Val10]. evaluating [YWWW14]. Evaluation [BHKK10, Car19, GHK11, SC12, TC17]. evasive [Agg15]. Even [BD11, AdFEGRI11, BKPP18, CHF15, MBW19, SSS15]. event [AL18, CS16b, YWWW14]. event-recording [AL18]. Every [ZYC13]. evidence [SWLX15]. evolution [BSM14]. evolutionary [LFZJ14]. Evolving [WMLN10]. Exact [BRFGL10, CCH^{+}14a, iP13, Rza14, vdBCDH17, AY12a, BLC10, CP13, ĎHPT10, Fin15, JSR11, Kut12, LFXH17, NS11, WP11, ZZ11]. exactly [MTT12]. Excessively [CGR14]. exchange [LWS10, XW12, YL16a]. Exchanged [BJ18, CCYW16, Jha15, KM14, LX13b, Liu15, MZ11, NFW15, Nin16]. exchanging [MOS16]. exclusion [TNN11]. executions [Ksh11]. exist [SQLS17]. Existence [BE16, AAB^{+}16, Dia15, GMY13, Mol18, YYK17]. Expansion [TS16, Vid13]. expansions [DD14]. expected [BEFP11, MMM^{+}17, VSP18]. experimentation [Car19]. experiments [YYK17]. explanation [ZKXY10]. Explicit [Aka14, Nic19]. Exploiting [MS19, SW13]. exploration [AMMY10, DP12, MPU17]. exploratory [RM18]. explore [FIPS11]. Exploring [CS16b, BvdZ19]. exponent [KKO10, SM10a]. Exponential [Bol10, BT16a, GW16, Pod12, Bae10, BRFGL10, BG11, CP13, Kut12, Laz10]. exponential-time [BRFGL10, CP13]. exponentially [BN10a]. exponentiation [SP18, Shp13, VN17]. exponents [BPRMS14, SM10b]. expressed [Att17, SV19]. expression [KSBT13, Yam19, ZZH16]. expressive [Str16]. EXPTIME [Kie13]. EXPTIME-hard [Kie13]. Extended [TC11, BKN^{+}12, Ket11, LLG10, ZW14]. extendible [Koc12]. Extending [SR11, BH11]. Extension [MK16, TWZ17, AT15, Kos17, LlCyChL10, Mor16]. extensional [PT11, RT13]. extensionality [SSSM11]. extensions [HI11]. extent [BV11]. external [Dha14]. externalities [EGK^{+}12]. extra [GH14]. Extracting [CNKS15, SML^{+}10]. extraction [JSZC15, YLLL16]. Extremal [ZS18].

F2 [GWL18]. F4 [Har18]. F4-codes [Har18]. face [EFMA10, SS10a, ZpH15]. facilitates [BTW15]. Facility [YC11, LXDX12, RX17, TL12]. fact [YL16a]. factor [BK10, MHFMSa11, SCB13, TZ11]. factoring [AH17, CD19, MM13]. factorization [Mes15, TPL16]. factorizations [SZ18]. factors [GGG^{+}14, KN13b, vdBCDH17, CT11]. Failed [CHP19]. failure [BR10b, SCL^{+}11, WWS12]. failures [NLZX14, TX11]. fair [AZ14, Ara10, Bee16, KC11]. fairness [ZZJ11]. false [CRJ10, GHC15, HKK12, JF15]. false-positive [HKK12]. false-reject [JF15]. families [LZLY12, ZLLZ18]. family [FYS10, LYF11, Mol18]. Fan [CZ16, Nin13, Bra18, Li10]. fan-crossing [Bra18]. Fan-type [CZ16, Nin13, Li10]. fans [AMRR11]. Fast [BEFP11, CY18, Joh14a, KKKS14, LMS14, NCP18, SSS12, SP18, Val12, ARdSP15, Ars18, CC13, CWY15, CNPS15, CK11b, GK10, LLT14, LTWS11, PC18, WCW11]. Faster [ASM17, Dha14, Fuj16, KP14a, LKR17, LXX14, NPR17, OZL16, TXQ11, Tsu19b, Vyg11, BD11, CPW11, FB10, GM13b, HZX16, Kat16, KT16, LSS15, SW13, Yam19, ZZH10a, ZHXS10]. Fault [BF19, IK10, Pel10, WWS12, YZ19, AMRR11, Che10b, CH12, DDPBT11, DY10, Far10, Fu10, Kim10, LSLY11, LWL11, LXDX12, LX13a, LX13b, SSK12, Tsa11, YLG10, ZXY10, ZX11, LSLY11, YC11]. fault-free [DY10]. fault-resilience [DDPBT11]. Fault-tolerant [IK10, Pel10, Far10, LXDX12, Tsa11, YC11]. Fault-tolerant-prescribed [YZ19]. faults [IS10, SM19, TVB15, WD11, YZZ17]. faulty [Che10a, Che12, Che16b, Che16c, CZWP17, CHF15, DY10, DX11, YDT10]. favorable [Dam14]. FC [CÉ13]. FC-rank [CÉ13]. feasibility [LL17, LXJ^{+}14]. feasible [HM13]. Feature [PdAL18, yJxW16, SaBG17, YLLL16, ZD18]. features [yJxW16]. Feedback [KP14a, BDF^{+}18, DLRS14, LX13c, WXZ^{+}12, XN12]. Feistel [ZW14, Bog10, BS11, Kar13, KDH15]. Fermat [DKC12]. few [Ang17, NR17, Pod12]. fewer [FA17, GKM14]. FFT [Bab17]. FFT-based [Bab17]. Fibonacci [AAB^{+}16, LLP^{+}18, Mol18, Wal10]. Fibonacci-number [LLP^{+}18]. field [CK11a]. fields [BMW13]. FIFO [DDPBT11, SD13]. FIFO-queues [SD13]. file [DHW11]. filter [CRJ10, PRM16, RPL19]. filtering [CZD14, CP10, GWJ11]. filters [Gra18]. filtration [CT16b]. final [BR12]. Find [GI18, KL18, Lee10]. Finding [BN10b, HLT10, KO16, LLW15, MT10, SH17, Xia10, ZGL11, BRFGL10, BPP^{+}18, CM17, EO13, Gil18, HWJ96, HAM11, MIM16, PI19, Pea16, Kar17, Sax10, SY15b, Yam19, ZKXY10]. fingerprints [PRM16]. Finite [Att17, BB12, CGLS16, GWL18, Had18, Li15, LXT19, RS10, SW12, Tys13, ZDQ^{+}17, HYC12]. finite-fold [Tys13]. Finite-state [Att17, HYC12]. finite-time [LXT19, ZDQ^{+}17]. finitely [AdFEGRI11]. firing [YNH^{+}14]. first [PPN^{+}17, RK15, Suz18, ZZ13, CG10b]. Fisher [ZpH15]. Fishing [BSM14]. FISM [FKC13]. Fitness [Wit14]. fitting [FV13]. fixed [CSX16, CP13, Has13, KT16, LH11, LMS13, Pér17, SH17, TL12, Tu15, vdBCDH17, CGG10]. fixed-density [SH17]. Fixed-Job [CGG10]. fixed-parameter [CP13, Tu15]. fixed-priority [CSX16, vdBCDH17]. fixpoint [BB12]. flexibility [Men12]. Flexible [LRS11]. Flip [DK14]. Flip-pushdown [DK14]. Flipping [Kal12]. flocking [Lee10]. flow [AC10, CdA13, CLP10, CC14, DM16, DJZ^{+}15, LY11, SL19]. flow-time [LY11]. Flower [WZQH16]. flows [AMO19, DS15, HKT17, Kam19, KT11, KT12, YL11a]. flowshop [FM18, GMM15, KMS19, LH11, MM15b]. flowshops [KMS19, KY12, Mos11b]. Floyd [Hou10]. fold [Tys13]. Folded [MK14, BJ18, CV19, CHF15, PCY16, YZZ17]. Fooling [KT18]. forecasting [TWW17]. Forest [LFW^{+}18, PR17]. forests [MRZ10]. form [Att17, GSA^{+}19, KL10]. former [ZKXY10]. forms [Har16, TY16a]. Formula [AB15, BV11]. Formula-trees [AB15]. Formulas [CCH14b, KS18, Mos13b, Pue11, XZ17b]. formulation [MSM14, MP15]. formulations [CG15]. Forward [XW12]. forwarding [VN17]. four [TZF16]. four-particle [TZF16]. Fourier [Nan15]. FPGA [eSKAI10]. FPT [FLP^{+}18, GM13b, Kat16, LFW^{+}18, RR16, WWBB15, XN12]. FPT-algorithm [RR16]. FPTAS [HK17, Nan13]. fractal [EAE21, MY18]. fraction [HPY^{+}10, IK15a]. Fractional [MSM14, Liu10, YHLC12, YH19, ZSY13, ZXS19]. fragment [HKK12]. frames [Ind15]. frameworks [DW10b, NAD19]. Frank [Ber17]. Fréchet [CR18, Fil18]. Free [PYW^{+}13, Ser10, Bir11, BBDS12, Bra18, BG12, BL14, BG15, BFLM15, Brâ15b, CÉ13, DP18, DM13, DY10, Ési11, EGK13, Fil11, GLPR19, GHvtHP13, HN10, JOS10, LX13c, LS15, LSK17, MNV11, Mos13a, RS10, Ria17, SS18, SS16, Sie19, SWZ12, WWW18, YL16b]. Free-gram [PYW^{+}13]. freeness [BGI^{+}12, GS10c]. freeze [YBMK15]. freeze-tag [YBMK15]. frequent [FKC13, HLT10, Lut14]. fresh [GJ19]. Frobenius [ABH^{+}14]. FSMs [YYK17]. FTP [MB14]. Fugue [AP11]. full [AABCC19, BR12, LC13, TY16b]. Fully [GP17, XPC^{+}10, FH12, Ket11]. Function [EFKR10, AKY13, AP11, CS10, FV13, Fuj16, GLW12, Han17, LL10a, LK14, Li12, LL16, MT18, MGPI12, Pie15, Sar11, VNP10, WS13, WYW14, ZDQ^{+}17]. functional [Dem18, FS12, PYYC16]. functions [Ama11, BPRMS14, Ber17, BS11, BC13, CKK13, Che18b, CGLS16, CW12, DKKY10, DZQF13, Fil13, FLMQ10, GPSD17, GNG11, IK15b, Kis19, MNV11, MS13, Oro11, PZ12, Pas15a, Pas15b, PGZB19, SWF18, Sim16, STMD18, Tys13, Vir11, WT13, ZLB11, ZQJ^{+}15]. fundamental [Sax10]. fusion [YLGC16]. Fuzzy [MGNAB11, NTD16, RAT13, TWW17].

G [Tho18]. GA [Gav17]. GA-H [Gav17]. Gabriel [KSV15]. Gallai [FL13b]. gallery [IM12]. Galois [GK10]. Game [JSR10, AB15, AH11, BH17, Kal12, Mat19, XZBX16]. games [BS12, BGL10, Che18b, FT15, Gen14, Had18, Le 19, MM15a, MN17, OBT12, Pér17, RX17, THS12, KMN18]. Gamma [XTTH12]. gap [FS11, LWXZ14, Man19, NB12]. Garden [HWA12]. gas [PC18]. gates [Pod12]. gathering [CRTU15]. Gaussian [MHIU18, XTTH12]. GCD [Sor10, vzGS15]. GCM [SKK10]. GCM/GMAC [SKK10]. GDLP [MMZ12]. General [DKNQ18, DW10a, Kos16, KMS19, KLS10, KY12, LWS10, Mos11b, NHK15, Til19, YW12, YZH14, YC19, ZGL11, ZXZL11]. generalisations [Lei17]. generalised [BS17, FCWZ18]. generalization [AT15, ACL18, SZ18]. generalizations [BT16b, DKC12]. Generalized [Che18b, LX13b, Maß15, Par11, TY16b, WLLS08, ZW14, BPRMS14, Ber12, BS11, CT16a, CCXW12, DZA15, GLW12, KSK13, KS12, Kun18, LLL18a, LKF15, Mol18, MHIU18, SQLS17, WWWZ13, WXCK19, nXlCL14, YZZX12, ZJN10, JP11]. Generalizing [CSX16, CKY15]. generals [YNH^{+}14]. generate [BP11, CVV10, GSA^{+}19]. generated [CLS13, DKC12, LlChL11, YLM10]. generating [Glü10]. generation [Ghi14, Kok17, MV13, iP13, RT13, SS16, ZMZ19]. generator [DPS11, LMCG16, MGPI12, NTD16]. generators [GR19, Nis92]. Generically [Jor10]. Genetic [GPT16, LREIMBMV16, Ghi14, JAH16, WMLN10]. genomic [WJS10]. genus [FWS13, TXQ11]. geodesic [LT13c, LFXH17]. geodesics [BP14]. geodetic [BPP^{+}18, CCCN19]. Geometric [HMSS18, AHK^{+}17, GGG^{+}14, MM19a, MM19b, MB14, YBMK15]. girth [LDW14]. given [AALL16, JLdSO^{+}14, SW18]. global [DM10, GL11, HHK17, KHKS16, YBHK13, YLLL16, ZST13]. globally [Jor10]. GLS [HZX16, ZHXS10]. GLV [ZHXS10]. GMAC [SKK10]. Good [SQLS17, GM12b, LN11, LTK^{+}19, SLF19, WT13, WH16]. Google [BvK15]. Gossip [GH15]. gossipers [GH15]. gossiping [Vay13]. GOST [LC13, WYW14]. grade [GL13]. graded [AIS10]. Gradient [JZ18, CYQ13, XLT^{+}19, YCL11, ZKXY10]. gradient-based [YCL11]. Gradient/Hessian [JZ18]. Gradient/Hessian-enhanced [JZ18]. grained [LH10]. gram [PYW^{+}13, Sal12]. grammar [Aku10]. grammar-based [Aku10]. grammars [BRMP13, CVV10, EGK13, Fil11]. graph [AX18, AMMY10, BF19, BGJ^{+}16, Ber17, BP11, BGR13, BJMC17, BPP^{+}18, CGS18, DDK^{+}15, DW10a, DGR15, Did13, EKN11, Fra10, FL13b, FS13b, GM15, GKS13, HI12, HWJ96, HHL15, HAM11, Iba17, IN12, JLMO17, KO16, KRV16, LZ15, LLL18a, LV15, MWZ12, MCS12, MHFMSa11, MPU17, NPR17, NG10, PI19, RR16, Rza14, Shi18, Suk19, XZ17a, YeCM14, YYR16, Yus11, ZLS^{+}17, Zho15]. graph-theoretic [CGS18]. graphics [MHIU18]. graphlet [ABH19]. Graphs [Hoc11, JLdSO^{+}14, LM15, MC15, ZYL15, AR18, AKH16, ADKM12, AHK^{+}17, AGHY12, Ang17, ALT15, ALT18, AMRR11, ADH14, BK10, BKPP18, BU17, Bar10, BPW12, BYK14, BS12, BvdZ19, BDF^{+}18, BDNVP15, BGP12, BES17, BdSL15, Bra18, BG12, BL14, BG15, BFLM15, BKN^{+}12, BS17, CZ16, CLZ18, CBSV11, CS16a, CKL10, CHR10, Cha11, CCKP13, CY17, CFMR18, Che10c, Che11, CQ12, CL15, gCF15, CT16a, CCXW12, CLS13, CGLS16, CG10a, CGLS10, CO13, CG10b, Cza13, DP17, DP18, Dam14, DFMS10, DDER18, DGŚN15, DE14, DZ12, Den14, DEL10, DJ15, DLN11, DX10, DRdSS12, DS14, DM19, FK10, FKL^{+}11, FS11, Feg19, FL13a, FHvtH^{+}15, Fie11, FH12, FH10, FGvL11, FM19, GLPR19, Gav17, GKPP16, GHvtHP13, Gol18, GSR^{+}14, Gro15, GLY11, GQG10, GLLX12, GLXG13, GZM15]. graphs [HL12, HHJ^{+}12, HRW11, Hou10, Hua14, Ili13, IK10, JN12, JT15, JT16, Jor12, JSR10, JSR11, KSV15, KP13, KSK13, Kob15, KHKS16, KS12, KM16a, LLL^{+}11, LN11, Lei17, Li10, LSZX15, LZSX17, LS11, LS15, LSK17, LT13b, LS14, LT15a, LSS15, LLW15, LT15b, LC17, LM16, LL10d, Liu10, LDW14, MBW19, MMS15, MS12, Mat19, ML10b, MRZ10, Mor16, Mos13a, Mou19, NN17, NHK15, Och17, PP10, PG12, PP13, PI13, PS19b, Pra12, Pul16, RRR12, RS19, RZ10, San11, SS16, Sch15, SS10a, zSfNbL11, SSW15, SSW16, STU12, Sie19, SK12, Slí12, SY15b, SWZ12, SW18, SSK12, Tak16, TLWZ16, Tsa11, Tsa15, TY13, Tur13, Vrg15, WD11, WL11b, WW11, WXZ^{+}12, WL12, WZL12, WSW12, WX13, WWLC14, WWZ15, WYLP17, WWW18, Waw14, Wid17, WLW11, XXZ14, Xu11, XLZ16, YBHK13]. graphs [YZRC18, YLM10, YLG10, YL11a, YL17, YC19, YL16b, YH19, Zel11, ZXY10, Zha10b, ZW10, ZL11, ZW11, ZGL11, ZL12, Zha13, ZLY13, ZZ18, Zha18, ZLLZ18, ZF10, ZX11, ZJ11, ZSY13, ZXS19, ZLS^{+}15, vE17, vzGS15]. Gray [GMS19, MNV11, VV11]. greater [DN12]. greatest [ZZZ15]. Greedily [BKZ15]. Greedy [JS19, Sie19, ZJN10, Ada11, FKR^{+}16, SW18, VLV15]. Green [SLL13]. grid [BV19, JKL18, RARM12, ZCC^{+}11]. grids [Cal13, DN12, Jia11]. Gromov [FIV15]. ground [AGI15]. ground-complete [AGI15]. Group [SZC^{+}17, WL11b, WL12, CGLS16, Den19, LWS10, Tsa11, Tsa15, YLZZ18b, ZXY10, ZX10]. Group-based [SZC^{+}17]. Groups [ZJ16, DZ12, HHTL10, WZS^{+}18]. growth [MNV11, MV13]. GRS [TD14]. guarantee [SW18]. guard [FS13b]. guarded [BB12]. guarding [Xu12]. guards [TJ14]. guesses [JP11]. GUI [RM18]. guillotine [ZSLW16].

Guruswami [WS10].

H [Gav17]. Hakimi [Bar12]. Hales [Shr18]. half [FC14]. halfplanes [BN10b]. halfspaces [DST12, Min11]. Halin [DS14]. Hall [Mor16]. ham [Ber12]. ham-sandwich [Ber12]. Hamilton [Hua14, Nin15]. Hamiltonian [AMRR11, CLZ18, CT16a, HCWH15, KSV15, LW19, YZ19]. hamiltonicity [CZ16, DZA15, SSK12, Che16b, CO13, Nin13, Val10]. Hammerstein [CW15]. Hamming [AG19, AIR17, Ars18, AD11, Kle13, MN19]. Hanano [LY19]. hand [JT10, SSZW16]. Hard [BNRC10, ADKM12, AD18, BvK15, FX11, FK17, HHK17, Kie13, LY19, MS18b, Pol18, SSS15, Tel19, GQ17]. Hardness [HM10, MN16, AD16, Ama10, APR13, BT16b, BK18, BPP^{+}18, Çiv13, EO13, Für19, IH18, JT16, Jan12a, Lan11, Lee17, MM19a, MM19b, MI17]. hardness-randomness [Jan12a]. Hash [BVF12, EAA^{+}16, AKY13, AP11, Han17, MS13, VNP10, WS13, WYW14]. hashing [ASM17, LWXS18]. Hasse [MK14]. HAVAL [GSY15]. HAVAL-3 [GSY15]. Havel [Bar12]. having [Krz13, MRZ10]. HC AB [ZXJ^{+}11]. head [RS12a]. heaps [NII^{+}15]. Heapsort [Kle13]. Heath [KL10]. heavy [CZ16, CLZ18, Hua14, Wid17]. hedonic [OBT12]. Helios [MS19]. Hellman [RH10]. help [AK13]. herding [Men12]. Herman [FZ13]. Hessian-enhanced [JZ18]. heterogeneous [CGG10]. heuristic [JK18, ZSLW16, ZLM17, ZXJ^{+}11]. Heuristical [LLT14]. heuristics [SW18, Wit14]. HEVC [CWY15, CY18]. hexagonal [GS10c, SWZ12]. hiding [KWH16, LXLY12, ZC12]. hierarchical [ADFM13, CPTZ13, CXD^{+}13, HL17, NZX19, QK15, WCJ15]. hierarchically [LMU15]. hierarchies [SLL13]. hierarchy [HS15]. high [KZ12]. higher [IK15b]. higher-dimensional [IK15b]. Highly [Ara10, Jor12, AABB17, BFKL13, KN13a]. highly-complex [AABB17]. Highly-fair [Ara10]. HIGHT [WWBC14]. Hill [KA17, YBC11]. hindsight [ST18]. hinge [HWJ96, HAM11, KHKS16]. hitting [CCL10, Dam17, GR19, MM19a, MM19b, SS10a]. HMMs [NVB15]. hoc [HKW18, Kor12]. hole [BG12, BG15]. hole- [BG12, BG15]. Holographic [FY14, YK11]. home [BE16]. homing [WL11a]. homogeneous [WD11]. homomorphic [CW12]. homomorphism [Rza14]. homomorphisms [FWH14, OM17]. homophonic [SPdR13]. honeycomb [DZA15]. hop [Cho12, KM16b, LZ10a]. hopsets [HP19]. Horn [BBK17, Pue11]. HORSIC [LKC^{+}12]. houses [CJ11]. housing [CJ11]. HsMM [XTTH12]. Hu [ZZ13]. hub [LDW14]. Huffman [Maß15, Sun16]. hulls [KP14b, MMM11]. Hungarian [DHW13]. Hybrid [XWLJ16, YWcW^{+}14, CPHS18, DJZ^{+}15, TNN11, XWS17, FR12, LREIMBMV16]. Hybrid-Constrained [FR12]. hybridization [vIL13]. hyperbolicity [FIV15]. hypercube [Che10a, CG10a, LX13a, MÖ15, YTN10, ZFJZ11]. hypercube-like [LX13a]. hypercubes [Che10b, Che12, Che16b, CHF15, DX11, HCWH15, Jha15, KM14, LX13b, LTK^{+}19, LL10c, Liu15, LW19, MZ11, RRR12, RARM12, WH16, YZZ17, YZ19]. hyperelliptic [FWS13, TXQ11]. hypergeometric [YL11b]. hypergraph [MTT12, SS10b]. hypergraphs [BBK17, BNRC10, BWZ12, DFRS13, Dur12, FM11, FA17, Hal10, MMM11, Szy12, Xia10]. hypermeshes [YYDL11]. hyperpath [RAT13]. hypersequent [Ind15]. Hypothesis [BT16a, GW16].

IC [LM16]. IC-planar [LM16]. ID

[EZ15, MMZ12, Mes15, PLPW13, TPL16]. ID-based [EZ15, MMZ12, Mes15, PLPW13, TPL16]. ideal [NAD19]. idealness [TD14]. identical [CK12b, GM12a, GM13a, HL17, MLW11, MM12, XZBX16, XCYZ11]. Identification [CW15, CZD14, CTHP13, GS19, PYW^{+}13]. identifying [Jia18]. Identity [Tia15, CTHP13, GR19, HLR11, Jan12a, MM13, QYWX16, WWYY11, XW12, LZJX10]. Identity-based [Tia15, CTHP13, HLR11, MM13, QYWX16, XW12, LZJX10]. IFP [MMZ12]. image [GKCK11, GPT16, KC17, LFZJ14, Sun16, YA13, YLGC16]. images [MHIU18]. Imbalance [LMS13]. immunity [GNG11]. Impact [SF12, SCB13]. implement [BR10b]. implementation [CJ11, CLY11, GK10, eSKAI10, PT12, Pie15, SF12, ZDQ^{+}17]. implicants [CKK13]. implication [GR13, GR14, LS10]. implicit [Cai15, CZ16, CLZ18, Hua14, Hua18]. implicit-heavy [CZ16]. importance [MK11]. Impossibility [CRTU15]. Impossible [Blo15, MNP12, SB19, SDM14]. imprecisions [TLS^{+}18]. impression [YLZZ18a, YLZZ18b]. improbable [TS16]. Improved [Ada11, AD16, BI14, CYQ13, Che16a, DM16, Doe13, GL11, IK15a, JKY15, KS11, LSW17, LXDX12, LHH11, LM11, LXT19, Mor11, Nut18, SF13, TS16, VSP18, WWBC14, YCL11, ZZ19, AK10, BL12, CR18, CCL10, FS12, GL15, Ger12, GM13a, GSY15, HHTL10, LFW^{+}18, Luc10, Man19, Möm15, PPH18, SS19b, Shi12, TJ18, VK18, WNFC10, YWWW14, Zha10a, ZGLY12, ZL18]. Improvement [PLPW13, BMB16, HCCG15, Wu13]. Improving [AHR10, BT16a, ĎHPT10, PC19, PRM14, PRM16, ZZJ11, ŽGH10, BDH^{+}11]. In-place [YEMR13]. inapproximability [LSW17, MI17, Man19, NLZX14, STY12]. Incidence [SSW15, SSW16]. inclusion [IL12, Kak15, SS10b]. incomparable [DK14, MS16]. incompatibility [MCS12]. incompatible [LZLY12]. incomplete [Ama10]. incompressibility [KK11]. Increasing [Men12, Pol18, Cha12, Dur13, Elm10, NB12]. Incremental [VAC13, BL10, GL15, Lof14, PR17]. IND-CCA2 [Gal13]. indeg [AX18]. indegrees [XZW15]. independence [GNV14]. Independent [BDF^{+}18, BG12, BG15, CWYP14, LYHC10, ZHX13, Che11, CK12a, DDK^{+}15, DRdSS12, DM19, GR13, GR14, Han17, HL16, Joh14a, KM16a, LX17, LS11, LS14, LMV17, LC17, Mos13a, NPR17, NG10, PC19, PPN^{+}17, Sch15, SW13, Suz18]. index [AX18, DGŚN15, HRW11, KKS15, MBW19, WWW18, ŽGH10]. Indexing [MR10, PYHA10, PYTH11]. indices [FH12]. indivisible [Lee17]. induced [AK14, BGJ^{+}16, Für19, Gav11, HOV13, LMS14, Nin13, SS12, Vág18, ZGL11]. inductive [Dem18]. industrial [MK11]. Inefficiencies [CGS18]. Inefficiency [WDT13, WCJ15, XZBX16]. Inequalities [Fil13, PPH18, SWW16]. inequality [DZ11, Möm15]. inequivalence [OKM13]. Inexact [LL17]. Inf [KMN18, KMS19, LPWZ14, SSW16]. infeasibility [LG18]. infeasible [Car19]. infection [MS18b]. inference [TWW17]. infinitary [Ket11]. Infinite [BC15, Fin15, JKL18, Le 19, Mis15, YC19]. Infinitely [ZQJ^{+}15]. infinity [Mis15]. influence [EDS11]. Inform [BB11]. Information [GR14, HAM11, KT12, Pas15a, YHLC12, Zha13, Ars19, BBR^{+}18, CH12, GS19, HYYZ15, JSZC15, MZQL14, Pel18, RZ14, ST18, SM13, YTYZ15, YLLL16]. inhabitation [AB15]. initial [Pér17]. injective [Rza14]. inputs [HYC12, SKN11]. inserting [ZCC^{+}11]. insertion [Mon15]. inside-out [AP11]. insights [Gav17]. inspired [LFZJ14]. instance [DST12]. instances [BI14, SD19]. instantiation [EP11]. institutions [Tut13]. instruction [Gue12, SF12]. Integer [BB11, CD10, RV10, RV17, AS18, AABB17, CPTZ13, Ger12, JSO11, JT10, LKR17, LL17, MM13, Mes15, MN14]. integers [CGG14, Han17, KQ11, Mat15]. Integral [AY14]. integrator [MK11]. integrity [ED17, IMCP15]. Intel [Gue12, SF12]. interaction [PYYC16]. interactive [BD16, Ito14, QX10, Yam14]. interchangeable [ŽGH10]. interchanges [HHTL10]. interconnection [BJ18, CLQS12, GCH^{+}10, TC17]. Interconnections [Mos13b]. interest [CX18, YZZX12]. interference [Kor12, MSZ11]. interleaving [SPdR13]. Internal [MGPI12, BV11]. interpretation [DLMV10]. interpretations [GCNR10]. Intersection [Vág18, BN10b, Cai15, CP11, Dar15, OG11, STY12, Xav12]. intersection-closed [Dar15]. interval [BK10, BES17, BdSL15, CHK13, Iba17, JLdSO^{+}14, KK16, LSS15, LLW15, PP13, RS19, YK15]. interval-valued [CHK13]. intervals [Dam17, FPY12, vBCH^{+}15]. intra [CdA13, CY18]. intra-procedural [CdA13]. Intractability [Zel11]. Introducing [DMS11, Fay16]. intrusion [CWW10]. Intuitionistic [RAT13]. intuitive [Fre14]. invariance [HM13]. invariant [LR18, ZZ13]. invariants [Li15, ZCC18]. inventory [YXW13]. Inverse [RSwW11, AX18, CP15, ZQJ^{+}15]. inversion [LlCyChL10, TLL16, TLL18, ZQJ^{+}15]. inversions [Elm15, GFG11]. Invertible [PRM14]. involutions [GMS19, YEMR13]. involutory [GSA^{+}19]. involving [Cai15, TWZ17]. irreducibility [ABH^{+}14]. islands [AADB^{+}18]. iso [Ale13, TS16, WWBC14]. iso-oriented [Ale13]. isogenies [Moo11]. isometric [GL15]. isomorphism [JKSS19]. isomorphisms [Sim16]. isosceles [Fra10]. Issuer [Ria17]. Issuer-free [Ria17]. issues [BB15]. item [FKC13, GWJ11]. item-based [GWJ11]. items [BTW15, HPY^{+}10, HLT10, KKZ13, Kok17, Lee17, Doe13]. itemsets [DWQ10]. Iterated [Pel10]. iteration [GY19, ZKXY10]. iterative [CZD14]. IV [FK17]. IV-matching [FK17].

Jacobi [Kim10]. Java [DMS11]. Jewett [Shr18]. job [FM18, GM12a, Hav10, JC10, KMS19, LH11, ZvdV10, CGG10]. job-dependent [JC10]. job-rejection [FM18, GM12a, KMS19]. jobs [EZH16, GM13a, GMM15, LYF11, LZY12, LZLY12, LJ15, MZC11, MM12, Sun19]. joining [AY12b]. Joltik [LJP19]. Joltik-BC [LJP19]. Jost [BGR13]. JPEG [GHC15]. jumbled [BFKL13, GG13]. jump [Tur12, YK15]. juntas [AW12]. just [GMM15]. just-in-time [GMM15].

Karatsuba [LlChL11, nXlCL14]. Kautz [LKF15]. KCI [TC11]. KCube [GCH^{+}10]. Keccak [MS13]. Kenyon [Svi12]. kernel [BGJ^{+}16, EFMA10, Lam11, LX19, Luc10, SS19b, Yan14, vIL13]. Kernelization [BCKM15, AK10, GG18, WNFC10]. kernels [Cro15, GMY13]. Key [EAA^{+}16, Kis19, LREIMBMV16, LJP19, ADFM13, DSPSHJNA18, jDX11, DMSD18, GLMS18, KDH15, LLP^{+}18, LWS10, LPdS10, MP13, MNP12, Nos11, Nos14, QX10, RS15, SB19, Sar14, SY15a, TC11, VN17, Wij10, XW12, YL16a, ZY17]. key-correlations [Sar14]. keying [GJ19]. keys [HBL14, IK15a]. keystream [MGPI12]. keyword [SY15a, SZC^{+}17]. kind [YLG10, ZXY10]. kinship [XZ18]. Kleene [Str16]. Kleenean [Pal10]. Knapsack [CFMT16, ACD18, GHRT17, HK17, IZ10, KS10]. König [Mis15]. Kraus [GQ17]. Krawtchouk [ZCC18]. Kronecker [GQG10, WSW12].

Label [Man19, ZMZ19]. label-unknown [ZMZ19]. labeling [Bar10, Cal13, IN12, JSR11, KSK13, PG12, Slí12, WL11b, WP11]. labelings [DL13, FS11, JN12, WL12]. laceability [YZ19]. laden [BKN^{+}12]. lags [ZvdV10]. Lai [LLG10]. lambda [SSSM11]. lambda-calculi [SSSM11]. language [AL18, CÉ13, CVV10]. languages [CGR14, FS14, FYS10, HS19, JOS10, Lan11, RS10, SS18, YK13]. large [LNP11, NII^{+}15, SLL13, TFY11, YWC11, YL17]. large-scale [YWC11]. larger [Tsu19a]. larger/smaller [Tsu19a]. largest [HK16, NMB10, SH17, Tiw14, XZ17a, ZLY13, Zha18, MS15]. last [DK10]. latin [Koc12, GH15]. lattice [DL13, Dra16, Tho12]. lattices [BP14, CGG14, Tia15, XLWZ16]. law [Mos11a, YBC11]. laws [DMS11]. layer [KT16]. layered [GJ11]. LBlock [KDH13, MNP12]. LCM [vzGS15]. LCP [Fis10, LGT17]. LCS [Deo12, FR12, WWWZ13]. LDA [EFMA10]. LDAG [WCZZ12]. leader [BEFP11, FCLCR17, Vay13]. leaf [GHK^{+}18]. Leakage [ZY17, GV14, YL16a]. Leakage-resilient [ZY17, GV14]. leaker [GS19]. learn [YBC11]. Learning [BGSM10, Pue11, SF17, DST12, GS10b, JC10, Min11, Mon12, PR17, PH11, SB13, YLZZ18a, YLZZ18b, Zha10a]. leasing [ZZXL11]. least [BF17, BMI14, CQ12, CZD14, HHK17, JZ18, Min11, ZW10]. least-squares [BMI14]. leaves [DM13]. left [Bae13, Kim10]. lemma [Mis15, KK11]. length [BFKL13, Cho12, DG14, EP16, FPY12, Ghi14, GKM14, GSS16, HYC12, KZ12, LZY12, LZLY12, LLL18b, MCS12, Mil15, Vaj18, WX13, ZZH10a, ZZH10b]. length-bounded [LLL18b]. length-three [Vaj18]. lengths [Maß15]. less [BF19, GK10]. letrec [SSSM11]. Lett [BB11, KMN18, KMS19, LPWZ14, SSW16]. letter [CK11a, Dur13]. Letters [GR14, HAM11, KT12, Pas15a, Zha13]. Level [MS16, DEH^{+}19, EP16, KYC13, Mar11, MOW17, SD13]. level-1 [MOW17]. Level-ordered [MS16]. leveled [DKL^{+}12]. levels [Wit14]. leverage [Kun17]. liar [PP13]. Lie [YLZZ18b]. LIFO [SD13]. LIFO-stacks [SD13]. Light [ZLLZ18]. lightweight [KDH15, MNP12, WWBC14]. like [BGL10, BGL13, Lau18a, LX13a, LLWH13, MS16]. likelihood [GP17]. LILI [MGPI12]. LILI-128 [MGPI12]. Limit [MS18c]. Limit-depth [MS18c]. Limited [WWBB15, DLN11, FCNY10, XZBX16]. Limits [GV14, BVF12]. Lin [CGJ10]. Lin-2 [CGJ10]. Linda [AGI15]. Line [GKS13, KPR18, AD18, BR10a, BGI^{+}12, BS11, CW13, Did13, FIPS11, FPY12, Jan12b, LLL^{+}11, LNP11, Man10, SSS12, TFY11, ZvdV10, ZZXL11, Zho15]. Linear [BGJ^{+}16, DG16, DCH12, DKC12, FGMP12, LC17, LFW^{+}16, XZ17b, AABB17, BHKK11, Bog10, CQ12, Dur13, DMSD18, EP16, Ési11, GWL18, Gio18, GE12, GGI11, GFG11, GPR10, HI11, HWJ96, HAM11, Iba17, Ind15, KZ12, KM16a, KKR^{+}13, KM16b, Kun18, Liu15, LJ15, LXT19, NII^{+}15, PP13, PI19, PR17, SS19b, SWZ12, SWW16, VSP18, WWLC14, WXCK19, XLT^{+}19, YYZH13, YL17, ZDQ^{+}17, Zho16, vdBCDH17]. Linear-time [LC17, PI19, PR17, vdBCDH17]. Linear-vertex [BGJ^{+}16]. linearizability [FG18a]. linearly [Fie11]. lines [KM12, TJ18]. link [DDPBT11, KM11, Rus18, VML18, ZKXY10, ZQJ^{+}15]. link-cut [Rus18]. Linkage [JKL18]. linked [SD13]. links [LTK^{+}19, MM15a]. LIS [AR13a]. List [DLO12, GLO12, Qi17, SD13]. lists [SL19]. livelock [FOSC18]. liveness [BJKZ14]. load [AY12b, DHW11, FM18, GJ14, KMS19, YZH14]. loading [LLG14]. lobster [ZYC13]. Local [AW12, KK11, AALL16, CC12, CLS13, FOSC18, GJR10, HCCG15, Jia11, Kak15, LSZX15, LZSX17, Shi18, Waw14, YWWW14, YLLL16]. local-improvement [HCCG15]. locality [DLO12, LWXS18]. localized [SLS16]. Locally [CCYW16, PP10, Ama10, IS10, JSR10, Koc18, LYHC10, PC19, Rza14, YWY12, KR16]. Locating [IMCP15, vISS10, HHL15, XS15]. locating-dominating [HHL15]. location [jDX11, JK19, LXDX12, RX17, TL12, XL15]. location-based [jDX11]. locked [DSPSHJNA18]. log [BH17, HPP17, Shr19, RJS^{+}10]. log-complete [BH17]. log-rank [Shr19]. Logarithm [Rud17, MM13, Mes15, TPL16]. logging [CZZ^{+}10, RJS^{+}10]. logic [BB12, MMP15, ZST13, BGM16, DRD11]. logical [LG18, TNN11]. logics [Ind15]. logs [YWWW14]. Long [FLP^{+}18, BGJ^{+}16, Bol14, Cai15, CKK13, DY10, YKD^{+}12, Zeh16]. long-standing [Bol14, YKD^{+}12]. Longest [FGKU15, Pol18, BBDS12, BVDP10, Cha12, DG14, Dur13, Elm10, Gra15, IH18, Kos17, RK15, SY15b, CT11]. look [LP13b, ZWWL12]. lookahead [LZY12]. Lookup [PRM14]. Loop [MNV11, Li15, YEMR19]. Loop-free [MNV11]. loops [Far10]. loose [HS15]. Loss [DJS13, BV10]. Lossy [CW12]. lost [CBHW10]. Lot [HYK14, ZLZX19]. Lovász [KK11]. Low [AT18, MSZ11, Mar11, YA13, BB11, BI14, CLY11, EN17, Kun17, RV10]. low-complexity [CLY11]. low-density [BI14]. Low-interference [MSZ11]. Low-level [Mar11]. low-rank [Kun17]. Lower [Fil11, LEP10, MO12, AS18, AGW13, Ber11, BGL10, BCKM15, BT16a, CA17, CP16, DE19, GY10, GW16, Kos17, Lu15, Pod12, PS19b, Pud12, Zha10a, Zho15]. lower-variance [AGW13]. lowest [NB12]. LP [DGK^{+}17, GKW15, Vid13]. LP-rounding [GKW15]. lp231 [LK14]. LPmkr [LP10]. LTL [MDB14]. lucky [ADKM12]. Lunch [Ser10]. Lyapunov [CYQ13, YCL11]. Lyndon [SH17]. LZ77 [KS19]. LZ77-type [KS19]. LZ78 [FB10, NII^{+}15].

M [Tho18]. M/G/1 [Tho18]. Ma [EAE21]. MAC [EMS15, OPS14, ZWWL12]. Machine [LL13, DKNQ18, FL16, FTYL14, GLW12, HKW14, HYK14, LL10a, LZCX12, LJ15, MY13, Oro11, OZL16, TCXT10, TFY11, WCJ15, YW12, YZH14, ZvdV10, ZLZX19]. machines [ACL18, CK12b, EZH16, FCNY10, FPY12, GM12a, GM13a, HM18, HL17, LY11, LZY12, LZLY12, LZ14, MZC11, MLW11, MM12, WCJ15, XZBX16, XCYZ11, HYC12]. Mahonian [KV16]. main [EAE21]. maintenance [LJ15]. Maiorana [GNG11]. Majority [KdER^{+}14, STMD18]. majority-decomposing [STMD18]. makes [BvK15]. makespan [CLP10, FCNY10, FTYL14, KM13, SF13]. malicious [ZLM17]. Malik [MMKK18]. malleable [CVZ11, Hav10, ZC12]. management [ABS13, jDX11, NZX19]. Manhattan [CDM^{+}11, SSKA17]. manifold [LT13c]. Manoussakis [Nin15]. Many [LL10c, Che10c, Che12, Dam17, FIPS11, HRS13, Krz13, PI19, YYR16, ZQJ^{+}15]. Many-to-many [LL10c, Che10c, Che12]. map [LMCG16, STAR15]. Mapping [Mis15, KSBT13]. MapReduce [PT12]. maps [GQ17]. Marek [Tar14]. margin [EMS15]. markdown [NLXX15]. market [DHW13, TWW17]. markets [CJ11]. Marking [CS16a, HKK12]. Markov [AAOW15, BCS15, CHK13, GXZZ13, GP17]. markup [NLXX15]. marriage [Miy14]. Marty [YBC11]. masking [CH12]. Massey [LLG10]. massive [HLY11]. Mastermind [JP11]. match [Bab17]. match-count [Bab17]. matches [Yam19]. matching [Ada11, ASM17, AD11, AGW13, BFKL13, CFMR18, CT16b, CNPS15, CS10, DGG^{+}19, ĎHPT10, FK17, Fre10, Fuj16, Für19, GG18, GGF13, GG13, GFG11, KKR^{+}13, LS10, LRSS17, LMS14, MKI11, Miy14, NR17, Pet12, Pul16, SS17, Sal12, Sch13a, YGK12, ZZH10a, ZZH10b, ZLY13, ZYL15, ZZH16, ZA17, Zha18, BCKM15]. matchings [HOV13, Zho15]. matrices [GSA^{+}19, Hru12]. Matrix [KKSŠ11, CPHS18, CYQ13, DZ11, GSA^{+}19, Kun17, MCS12, OKM13, SWW16, YCL11, ZQJ^{+}15]. matrix-valued [DZ11]. matrix-vector [CPHS18]. Max [JN12, BNRC10, HKW18, KKZ13, Zel11, XGX11, CGJ10, Doe13, LM11]. max-algebra [BNRC10]. max-cut [Zel11]. Max-optimal [JN12]. max-power [HKW18]. maximal [BP11, Dam14, GY10, KKO10, LS14, MKI11, NHK15, Sch15, Shi12, VML18, YL17]. maximally [WZL12]. Maximin [LZ10b]. maximization [DM10]. maximize [LZLY12]. maximizing [EDS11, Fuj16, Lee17, GMM15]. Maximum [AK14, BCNPL14, CP11, FS11, GG18, Mos13a, AMO19, BES17, BN10b, CQ12, DDK^{+}15, DMS12, GHK^{+}18, GKW15, GP17, Hoc11, Joh14b, KTUY17, KM16a, LL14, LY11, MBW19, MMM^{+}17, MU15, MRZ10, MM15b, NPR17, NG10, Pet12, Rac10, RZ10, STY12, SW18, WW11, WLW11, XZW15, ZW10, BG12, BG15, Shi15, Xav12]. Maximum-weight [BCNPL14]. may [SSS10]. may- [SSS10]. McFarland [GNG11]. MDS [GSA^{+}19]. mean [CM17, Gen14, LLWH13]. mean-payoff [Gen14]. mean-shift-like [LLWH13]. means [JKY15, LSW17]. measure [CH12, yJxW16, JCC11]. measures [LX13b, MTT12, TLS^{+}18, YHLC12]. mechanism [CWY15, GL13, PLPW13, WB12, YWcW^{+}14]. mechanisms [MMZ12, Vou19]. median [Cha13, Cha19, Wu14]. medians [HMSS18]. meet [GH15]. meet-all [GH15]. meeting [JK19]. meets [CP13]. membership [AJLM11, Pue11]. membranes [AC12].

Memory [DW12, SD13, Sto16].

Memory-restricted [DW12]. Memoryless [Had18, LT13a, Le 19]. Menger [YZZ17]. Menu [LCC17]. Merging [Ehl17]. Merlin [Ned17]. mesh [DZFY12, SHFMSA11, WWS12]. meshes [YDT10]. meshes/tori [YDT10]. Meshram [PLPW13]. message [BR10b, Chr19, CZZ^{+}10, Jia16, KVK18, MSV14]. message-passing [BR10b]. messages [MPRC19]. meta [BBR^{+}18, JK18]. meta-heuristic [JK18]. meta-ontology [BBR^{+}18]. Method [BPPS18, AT18, Ber11, CT16b, CY18, Dra16, KSBT13, Lof14, Mei12, WT13, ZHXS10, DHW13]. methods [Kle13, KK11, PGZB19]. metric [BC13, Cha13, Cha19, DGKS14, FHvtH^{+}15, FIV15, GS10a, KLS10, Lai16b, MSZ11, Wu14]. metrics [KLS13, KSD18, ŽGH10]. Metropolis [NB12]. middle [Gio18]. millionaire [MN15]. mimicking [KR14]. min [DKZ18, KKZ13, Zel11, LMSN16, Doe13, Man19]. min- [Zel11]. min-entropy [DKZ18]. min-max [KKZ13, Doe13]. Min-Rep [Man19]. miner [ABP11]. minimal [AKH16, BPW12, BYK14, BP11, CCT14, DMS12, Fre10, KVK18, Krz13, SW19, YBHK13]. minimally [GLY11]. Minimax [XL15]. minimising [MMM^{+}17]. minimization [BBK17, CLP10, KT16, KM13, LH11, Val12, HCCG15]. minimize [FL16, FCNY10, FTYL14, LY11, LZ14, MY13, YW12, YZH14, ZLZX19]. Minimizing [AGÁRCS^{+}19, FM18, HKW14, HKW18, Kor12, KMS19, CK12b, NLZX14, MM15b]. Minimum [AAB^{+}19, CFJ12, KT11, KT12, LX13c, LL10d, RARM12, Rud17, BSM14, BDNVP15, CM17, FX11, HM10, HKW18, HKT17, KS11, KP13, KR14, LZ10a, LLL18b, MKI11, Mil15, Mor11, RR16, SI18, Shi13, Tak16, Waw14, WCW11, Xia10, Xu12, ZY12, ZLS^{+}17, BI14]. minimum-sum [Shi13]. Minimum-width [AAB^{+}19]. Mining [FKC13, BL10, CZZ^{+}10, TLS^{+}18]. mining-based [CZZ^{+}10]. Minkowski [Tiw14]. minor [WWW18]. Minterm [Ama11]. Minterm-transitive [Ama11]. mirror [BV10, Che18b]. mismatch [CP10]. mismatches [FGKU15, GGF13, Gra15, NR17]. missing [CLS13, LTK^{+}19]. mistake [BGSM10]. mistake-bound [BGSM10]. MitM [TY16b]. mix [MM15a]. mix-weighted [MM15a]. MixColumns [DK10]. mixed [CSX16, JF15, LL17]. mixed-criticality [CSX16]. Mixing [ABH19, NB12]. MJRTY [AR13b]. MM [WH16]. MMH [BKS16]. mobile [AC12, Ria19, YMSA14]. mobility [CDM^{+}11]. Möbius [GZJ^{+}12]. Mod [LLL^{+}11]. modal [Ind15, ZST13]. mode [Fay16, GK10]. Model [DBB12, GXZZ13, BV10, BBD^{+}12, BGSM10, CA17, CHK13, Che18a, CX18, DLMV10, GLO12, LWL11, LQL^{+}17, MMKK18, ML10a, Mas17, MM13, MIM16, NLZX14, QX10, SWLX15, SSZW16, Tho18, Tsa15, VB19, WH16, XMLL11, YYDL11, YBC11, YW14, ZZXL11, ZLM17]. Modeling [RZ14, PYW^{+}13]. models [Ars19, CGS18, CA17, EAE21, JCC11, LL10b, LG18, MY18, RM18, TK15, ZQJ^{+}15]. modes [CY18, SKK10]. modification [RR16]. modifications [IMCP15]. modified [CTHP13, JSO11, WS10]. modify [FG18a]. modifying [JC10, RR16]. Modular [EP11, NCP18, SP18, VN17, ZpH15]. modularity [DD14]. modules [PYYC16]. modulo [Agg15, BHL17, Shp13]. modulus [BKS16, DCH12]. molecular [Jor12]. moment [ZZ13]. moments [ZCC18]. monitoring [LLK10, BBKS17]. Monochromatic [GGG^{+}14]. monomials [FLPS15]. Monotone [Ang17, Fuj16, HP18, KKKS14, LWŻ07, LPWZ14, Lu15, SV19]. monotonic [Par11, PP14, WLLS08, XZ10]. Monte [Lof14]. Moore [AR13b]. Morpion [KTUY17]. Morrissey [CL10]. most [BG11, GFG11, LS15]. Motzkin [BP14]. move [CCT14]. movement [CDM^{+}11]. moves [ÁCLS12, FS13b]. MRM [GLO12]. MSC [AGHY12]. MSO [CS16a, MS18a]. MTS [JK18]. Multi [YZH14, Cho12, CZZ^{+}12, DEH^{+}19, Ito14, LMCG16, Mas17, MZC11, PPN^{+}17, RS12a, Ser10, XLWZ16, ZMZ19]. multi-branching [PPN^{+}17]. multi-cycle [CZZ^{+}12]. multi-delayed [LMCG16]. multi-head [RS12a]. multi-hop [Cho12]. multi-level [DEH^{+}19]. Multi-machine [YZH14]. multi-objective [Ser10]. multi-party [XLWZ16]. multi-prover [Ito14]. multi-server [DEH^{+}19]. multi-stage [Mas17]. multi-view [ZMZ19]. Multicasting [CG10a]. multiclass [FK10]. multicoloring [BKPP18]. multicover [vBCH^{+}15]. Multicut [KV10]. Multidimensional [WWBC14, ADF13, Che16c]. multigraphs [FS13a]. Multilevel [HF14]. multilinear [MST16, Mon12]. multiobjective [LFZJ14]. multipartite [WWZ15]. multiparty [Shr18]. Multiple [HZSL05, DST12, DLN11, JC10, LZLY12, MN14, PZ12, STAR15, WDH16, Yam14, ZGL11, ZPX11, vIL13]. multiple-instance [DST12]. Multiple-Precision [HZSL05, MN14]. Multiplexer [CLY11]. multiplicand [SP18]. multiplication [Bae13, BG11, CC13, HZX16, OKM13, SP18, SF12, ZHXS10]. Multiplicative [Gil18, SK12]. multiplicity [BKPP18]. multiplier [CPHS18, CLY11, GK10, LlChL11, PCH13, nXlCL14]. multiprocessor [CGG10]. multiprocessors [KC11]. multisecret [FGMP12]. multiset [FL12, KN13b, KR10]. multisignatures [QX10]. multiterminal [CCF14]. multivalued [PH11]. must [SSS10]. must-convergences [SSS10]. Mutual [ADH14, Kak15, TNN11]. Mycielski [LDW14]. n [CD10]. Naccache [ACD18]. Naor [GGI11]. Nash [Lee17, OBT12, SM13, THS12, WDT13, WCJ15]. natural [CGG14, MHIU18, SML^{+}10]. navigational [KRV16]. NC [DDER18]. near [LT13a]. near-collisions [LT13a]. nearest [AMMY10, BJMC17, Lee10, Tsu18]. nearly [ZZ18]. necessarily [GM12b]. necessary [BIR16, TD14]. necessity [BR10b]. necklace [RRR12, SH17]. need [BG11, SSSM11]. negated [SM17]. Negative [PH11, Hou10]. neighbor [AMMY10, BJMC17, LTK^{+}19, WH16, BSM14]. Neighbor-Nets [BSM14]. neighborhood [Cai15, LMV17, WZL12]. neighbors [Lee10]. neither [WX13, WLW11, Zha10b, Zha13]. nested [Hen16]. nets [AC12, BE16, Ezr10, FS14, Oan11, BSM14]. Network [LCC17, AC10, AHLS18, BJ18, CGS18, CA17, CS16b, DP12, EGK^{+}12, FKC13, JAH16, KK15b, LXT19, MTA10, PYYC16, RAT13, SF17, TWW17, WDH16, XLT^{+}19, YA13, YTYZ15, YYYZ12, vISS10]. network-based [TWW17]. networking [CFJ12]. networks [AMT12, AFPT10, AC10, Bog10, BS11, CCRS14, Che16c, CLQS12, Cho12, CDPB^{+}10, Den19, DZA15, jDX11, EDS11, GH14, GHL18, GCH^{+}10, HKW18, KK15a, Kam19, Kar13, KR14, KK15b, KVK18, Kor12, Lai16a, LLT14, LKC^{+}12, LWL11, LX13a, Liu15, MSZ11, MHFMSa11, MOW17, RARM12, SM19, SZC^{+}17, SWW16, SKK10, TC17, Vay13, VLV15, WWS12, XPC^{+}10, XL15, YWC11, YWL15, YCL11, YYZH13, ZCT^{+}12, ZX10, ZGLY12, Che18a, PP14, ZW14]. Neural [LCC17, CYQ13, Che16a, JAH16, LXT19, SWW16, XLT^{+}19, YCL11]. Newton [ZKXY10]. NExpTime [ZST13]. NExpTime-complete [ZST13]. next [Tsu19a]. next/previous [Tsu19a]. nine [LYHH14]. Nisan [DPS11]. NLM [OPS14]. NN [KZP10]. no [KS10, Pel18, Ser10]. node [BDPP18, DGR15, HKW18, HBL14, Lai16a, Tsu18, WWS12, YBM^{+}10]. node-disjoint [Lai16a]. nodes [DY10, LTK^{+}19, Mis15, WWBB15]. noise [SB13]. noisy [JF15, KN13a, Shp13, TK15]. Non [DFMS10, GMY13, Koc12, QX10, RX17, Sun19, Suz18, TK15, ADF13, Ale13, AAB^{+}16, CVZ11, CFJ12, Che18b, CX18, CGR14, CT11, DDPBT11, FS14, GS10a, GPSD17, GWL18, HYC12, Ket11, KZP10, MCS12, MP15, OBT12, SO10, SSSM11, TLL16, TLL18, ZC12, ZY12, vdBCDH17]. Non- [GMY13, AAB^{+}16]. non-additive [CX18]. non-C1P [MCS12]. Non-clairvoyant [Sun19]. non-collapsing [SO10]. Non-contractible [DFMS10]. non-convex [Che18b]. Non-cooperative [RX17]. Non-depth-first [Suz18]. non-determinism [KZP10]. Non-deterministic [TK15, FS14, SSSM11]. non-edges [DFMS10]. Non-extendible [Koc12]. non-FIFO [DDPBT11]. non-fully-extended [Ket11]. Non-interactive [QX10]. non-malleable [CVZ11, ZC12]. non-metric [GS10a]. non-observable [HYC12]. non-ordered [KZP10]. non-overlapping [CT11, TLL16, TLL18]. non-preemptive [vdBCDH17]. non-regular [CGR14]. non-splitting [GPSD17]. non-terminal [ZY12]. non-termination [MP15]. non-trivial [Ale13, OBT12]. non-uniform [CFJ12]. non-wandering [ADF13]. Nonadaptive [Mon10]. noncommunicable [VB19]. nondeterminism [RS12a]. nondeterministic [HYC12]. Nondominated [Jia11]. nonexistence [Pas15a, Pas15b]. nonisomorphism [HI12]. Nonlinear [XLT^{+}19, BD16, DLRS14, GL15, LL10a, LL17]. nonlinearity [CW15]. nonnegative [Hru12]. nonprime [BMW13]. nonrandom [NTD16]. nonrepresentable [Mat15]. nonsquare [ABH^{+}14]. nonstochastic [FYS10]. nonterminals [CVV10]. nonuniform [WWS12]. nor [WX13, WLW11, Zha10b, Zha13]. Normal [KL10, Val10, Att17, Har16, TY16a]. Normalised [CHP19]. NORX [NA19]. Note [CGJ10, GY10, LWŻ07, LPWZ14, WWZ15, AFPT10, AB15, ADFM13, BLYL17, BDH^{+}11, BGR13, Bor16, BMI14, BL14, Brâ15b, BK18, Cha12, CM17, CL10, CW13, CZWP17, CLP10, Çiv13, CK12b, CIK^{+}13, Cro15, Cza13, DZQF13, DT10, DZ11, DX11, Duc18, EDS11, Ezr10, FKL^{+}11, Fra10, FM19, GLW12, Gal13, GPSD17, GY15, Gen14, GMM15, GR19, Gra15, Gro15, GJ15, Har16, HP18, IN12, Kam19, KYC13, KMN18, Kot12, Kra12, KY12, Lau18b, LL14, LSZX15, Lib10, LWXZ14, Man19, MN17, Mon15, MM15b, Ned17, Pas15a, Pas15b, Pie15, RS15, Sch13a, SeF14, Shi13, Shr18, Svi12, Tsa11, Vou19, WP11, Xav12, YYR16, YW12, YC19]. Notes [CP15, GLXG13, HL17, KLMP18, Nin15, vE17]. notions [IM17]. Novel [nXlCL14, FBD18, GCH^{+}10, MGPI12, SaBG17, Sun16, YGK12, YWcW^{+}14]. November [Ano19n]. Nowhere [YL11a]. Nowhere-zero [YL11a]. NP} [Tel19, ADKM12, BT16b, Che17, DFdFT16, FK17, GQ17, GI18, HHK17, HWA12, LZSX17, LY19, MM19b, MI17, Och17, Rom11, Zha11]. NP-complete [Che17, DFdFT16, GI18, HWA12, Och17]. NP-completeness [LZSX17, Rom11]. NP-hard [GQ17, ADKM12, FK17, HHK17, LY19]. NP-hardness [BT16b, MM19b, MI17]. NP-pairs [Zha11]. Number [LST11, ADKM12, AD13, AH11, BKMT14, Bor16, BS17, BPP^{+}18, CCCN19, CY18, Den14, DN12, DS14, EAE21, Fie11, Für19, GMM15, HKW18, HYC12, IM12, JP11, Jha15, JSR10, KM14, LLP^{+}18, LZLY12, LLL18b, LKF15, LZCM10, LDW14, MY18, MBW19, MKI11, Mat19, iP13, SLdAMP17, TLWZ16, VK18, WWBB15, XS15, XZW15, YK15, ZLCM10, Zho15, ZS18, vIL13, vZBSY16]. numbering [MMS15]. numberings [SS19a]. Numbers [LM15, CKL10, CZ13, LX17, WXZ^{+}12]. numeral [GKCK11].

OBDD [Bol10]. OBDD-based [Bol10]. OBDDs [BG11]. object [Zun18]. objective [CG15, GLW12, Ser10]. objects [Lai16b]. oblivious [CDPB^{+}10, FIPS11, Ria17]. obnoxious [TL12]. observable [HYC12]. observation [NA19, Pér17]. observations [LR18]. occlusion [ZpH15]. October [Ano19o]. odd [BC15, KN13c, MCS12, ZYC13]. odd-elegant [ZYC13]. Offline [DHW11]. offs [LST11]. old [VV11]. Omega [AJLM11, BHLM13, ML10a]. omission [DK10, SCL^{+}11]. omissive [FCLCR17]. omni [JSZC15]. omni-directional [JSZC15]. On-line [FPY12, Jan12b, LNP11, ZvdV10, TFY11, ZZXL11]. on-stack [YEMR19]. One [zSfNbL11, SS10b, XW12, BN10a, BDNVP15, CLY11, Chr19, GJ15, HHK17, JK19, LKC^{+}12, MRZ10, Mos11a, PI19, PYHA10, PYTH11, SM10a, WBC^{+}12, nXlCL14, BvK15]. one-dimensional [PYHA10, PYTH11]. One-inclusion [SS10b]. One-round [XW12, Chr19, JK19]. one-space [GJ15]. one-time [LKC^{+}12]. one-to-many [PI19]. one-to-one [BDNVP15]. OneMax [DW12]. ones [CCL10]. Online [AMP13, EZH16, FL16, FCNY10, FTYL14, GSZ13, Hal10, Hav10, IZ10, LY11, LZY12, LZLY12, MY13, WBC^{+}12, YXW13, ZXZL11, ZCT^{+}12, CXD^{+}13, Che16a, DKNQ18, DHW11, EGK^{+}12, EL10, FK10, HMS16, MO12, MLW11, Miy14, NLXX15, NZX19, Qi17, SF13, TCXT10, YCL11, YW12]. only [KA17, Sar11]. ontology [BBR^{+}18]. opacity [BCS15]. open [CZCD13, MSV14, SB13]. operation [Fay16, GKS13, SKK10, TX11]. operations [FG18a, YKD^{+}12]. operator [BRMP13]. operators [ZST13, ZV14]. opinion [YLLL16]. OPT [JSO11]. optical [Liu15, TC17, YYYZ12, YYZH13]. Optimal [Cal13, DM10, EGK^{+}12, FC14, GLO12, GMS19, HM18, JKS10, KLS13, KM12, LGT17, MLW11, NLXX15, PPN^{+}17, YWY12, YZZX12, AR13a, Ale13, Che11, CN18, Dam16, Dar15, DDPBT11, EP16, Gou15, HI12, HHL15, HP19, JN12, KMMN15, KM16b, Kun18, LJL18, LRS19, Suk19, TCXT10, Tsa11, Tsu13, Yan14, YW12, Zho16]. optimality [LMSN16, MPU17, XMLL11, CLM12]. optimally [HHJ^{+}12]. optimiser [LTWS11]. optimization [AMP13, FS12, GL11, Gav11, Lib10, Ser10]. optimizer [JK18]. Optimizing [GN10]. Optimum [TJ14, Vyg11, PR17]. Optimum-Path [PR17]. option [CX18]. options [ZPX11]. oracles [SD19, WWYY11]. orbit [CYH15]. Order [KP14b, AKY13, AAC^{+}10, CT16b, CNPS15, Elm12, Lam11, MP15, YL11b]. Order- [KP14b]. order-preserving [CT16b, CNPS15, YL11b]. orderable [PG12]. ordered [Aku10, CG15, CCF^{+}12, KZP10, LABKS17, Li10, MS16]. ordering [WCZZ12]. orderings [CD19, GKPP16]. orders [Ési11, FMHL11, Tur12, YXW13, YK15]. ordinals [JKSS19, vB12]. Ore [SSK12]. organizing [BvK15]. orientation [AD13]. orientations [GBH12, LLL^{+}11, ML10b]. Oriented [DN12, OM17, Ale13, AH11, DS14, GLLX12, LKR17]. orienteering [GKM^{+}15]. orthoconvex [NMB10]. orthogonal [AGÁRCS^{+}19, LWŻ07, LPWZ14, SZ18]. other [BNRC10, JKY15, Kle13]. OTIS [MHFMSa11, SHFMSA11]. OTIS-mesh [SHFMSA11]. out-arcs [GLLX12]. outdegrees [XZW15]. outer [CCXW12, KP13, LLW15, ZLLZ18]. outer- [ZLLZ18]. outer-connected [KP13, LLW15]. outer-planar [CCXW12]. outerplanar [LL10d]. outerplane [CY17]. outliers [AAB^{+}19]. output [CZD14, LXX14, NG10, iP13, PZ12]. output-sensitive [LXX14]. outputs [HYC12]. overflow [ABS13]. overlapping [CT11, SO10, TLL16, TLL18].

P [AT15]. P2P [CFJ12, FKC13]. P2P-FISM [FKC13]. PAC [Dar15, Zha10a]. package [FC14]. packed [GGF13]. packet [ZGLY12]. Packing [AHK^{+}17, BTW15, BGJ^{+}16, BHKK11, Bra17, CP15, DLN11, DE19, GJ15, HPY^{+}10, HMS16, Jan12b, KK15a, KLMP18, LMU15, LX19, Shi12, Svi12, WNFC10, Yan14, ZSLW16, AK10]. packings [MÖ15]. PageRank [Gro15, Lof14]. Pages [Ano19o, Ars19]. paging [DEH^{+}19, KMMN15]. Paillier [Gal13]. Paillier-based [Gal13]. pair [DGKS14]. Paired [Che12, Che16c, LTWS11]. Paired-bacteria [LTWS11]. pairing [TXQ11, ZY17]. pairings [QYWX16]. Pairs [STMD18, Ars18, KYC13, MN19, OG10, Zha11]. palindromes [GPR10]. Palindromic [GSS16, IH18]. palstars [RSwW11]. pancake [IK10, DZ12]. Pancyclic [GLLX12]. Pancyclicity [LWL11, MHFMSa11, Fu10, GLXG13, SHFMSA11, Wid17]. paper [KK11]. Parallel [BRMP13, Che11, PCH13, SM19, AR13a, ACL18, BJ18, CZCD13, CK12b, DLRS14, Ezr10, FL16, FCNY10, FTYL14, FPY12, GY15, Ger12, GM12a, GM13a, GBH12, HKW14, Hil14, LY11, LlChL11, LYF11, LZY12, LZ14, MM19a, MM15a, MZC11, MM12, Sor10, TFY11, WDT13, nXlCL14, YSGY10]. parallel-batch [FL16, FTYL14, GY15, LY11, LYF11, LZY12, MZC11]. parallel-batching [FCNY10, TFY11]. parallelepipeds [Koc12]. parallelism [ASM17]. Parallelization [Ito14, YEMR19]. parameter [CZD14, CP13, GPT16, IS10, KT16, LMS13, MZ15, Tu15, Vid13]. Parameterized [CGJY12, FGG^{+}10, GJ14, LZ10b, AK14, BBKS17, BBDS12, BD11, CCF14, CLR13, GR19, LNP11, MS18b, PS19a, Tsu19b, vBCH^{+}15]. parameterizing [XTTH12]. parameters [GD13]. Parametric [Che17, AL18, DM16, DKNQ18, GHRT17, HK17, DRD11]. parametrization [RS12b]. Parikh [EGKL11]. parities [BGSM10]. Parity [BS12, Had18, Le 19, SB13, TX11]. parsing [BRMP13, FB10]. parsings [KS19]. Partial [DLMV10, Mac19, BR10a, Elm12, FLRS11, IMCP15, MS12, Par16, Pér17, Qi17, ST18]. partial-observation [Pér17]. partially [JLMO17]. particle [TZF16]. partite [BWZ12, ML10b]. partition [GSR^{+}14, JLdSO^{+}14, ZLB11]. partitionable [BPW12]. partitioned [Mor16]. Partitioning [BKN^{+}12, MO12, AP11]. partitions [AS18, KLS13, OBT12]. party [TC11, XLWZ16]. Pascal [KK11]. passing [BR10b, Che10a]. passive [ABPS15]. passport [LZJX10]. password [LWS10]. password-authenticated [LWS10]. past [LZCX12]. past-sequence-dependent [LZCX12]. Path [BKPP18, Che10a, Che10c, Che12, Che16c, CZWP17, CDM^{+}11, FLP^{+}18, Gol18, Kat16, LS14, LT15a, LL10c, MMM11, PI13, PI19, SY15b, WL11b, WL12, ZLS^{+}17, ZLNB19, PR17]. pathbreadth [Duc18]. Paths [Feg19, AHK^{+}17, BGJ^{+}16, Bon13, CS16a, DP17, Ghi14, Lai16a, XS15, ZJN10, ZGL11, ZLLZ18]. pathwidth [BvdZ19, Kob15]. Pattern [AD11, BCKM15, ZZH10b, ACK11, BV10, BPPS18, CNPS15, DGG^{+}19, GGF13, GG13, KKR^{+}13, LS10, NR17, ZA17]. patterns [BL10, CGR14, JOS10, RS10, SKN11, Sch13a, Vaj18, WYLP17, YYZH13, ZZH10a]. paucity [Lut14]. pay [Ria19]. pay-TV [Ria19]. payoff [Gen14]. PCA [EFMA10, HYYZ15, MI17]. PCA/LDA [EFMA10]. Pearson [DM10]. peeling [MN16]. peg [JP11]. penalty [LLG14]. perceivable [BBB^{+}17]. Percolation [KdER^{+}14]. Peres [iP13]. Perfect [LZ15, TZF16, AAB^{+}16, CFMR18, Han17, Mol18, YEMR13, Zho15, XW12]. Perfectly [San11]. Performance [YWC11, AMN^{+}10, BLM10, BDH^{+}11, PRM14, PRM16, RS12b, SCB13]. perimeter [Gil14]. period [BC15, Par11, WLLS08, WXCK19]. periodic [AMT12, ABP11, KC11]. Periodicity [GRSS17, Jia18]. periods [CIK^{+}13]. permanent [BHL17, Jan12a]. permanents [BHKK10]. permutation [BK10, HWJ96, HAM11, HHTL10, KO16, KKR^{+}13, LK14, LC17, vZBSY16, BCKM15]. permutation-based [LK14]. permutations [BV10, Bar13, Kar13, Kok17, MR10, Rus18, VV11, Zal11]. Permuted [CS10]. permuting [YEMR13]. Perona [MMKK18]. Perron [ABH^{+}14]. person [BBB^{+}17]. Personalized [HPP17]. perspective [CGS18, LWXS18, MK11, SO19]. perturbation [ABS12]. pessimistic [Che18a, GHL18, JP11, Tsa15]. Petersen [KSK13]. Petri [AC12, BE16, FS14]. phase [LQL^{+}17, ZZ11]. phrase [PYW^{+}13]. phylogenetic [CDD^{+}11, HMSS18, vISS10]. Physarum [Bon13]. PI [MBW19]. pieces [LLL18b]. piecewise [HS19, KKS15]. piecewise-testable [HS19]. pigeonhole [BGL10]. Pisot [CZ13]. pitfalls [YMSA14]. place [YEMR13, vGGS11]. place/transition [vGGS11]. placed [KK16]. placement [HM18, YC11]. plain [QX10]. Planar [BR10a, Lu15, WX13, AR18, ADKM12, BCNPL14, BV19, CHR10, CQ12, CL15, CCXW12, Cza13, DG16, Did13, Fra10, FM19, HRW11, LL14, LSK17, LM16, Mou19, Och17, RZ10, SS10a, WW11, Waw14, XLZ16, Zha10b, ZW11, ZL12, Zha13, ZLLZ18]. planarity [AHR10]. plane [AHK^{+}17, AADB^{+}18, DX10, FMHL11, Jor10, Juk12, Pel18, WLW11, YiN10, ZW10]. planes [LT18]. plus [MY13]. PMC [Che18a, Tsa15]. point [AADB^{+}18, BR10a, DBFMPLV17, FMHL11, FV13, GGG^{+}14, LL14, MM19b, NMB10, NTD16, ZHXS10, vGLM12]. point-set [BR10a, FV13]. pointing [SSZW16]. points [GGG^{+}14, Man10, Moo11, ZCC^{+}11]. policies [NLXX15]. Policy [XWLJ16, DJS13, LMSN16, TY16a, WDT13, XWS17]. Pollination [WZQH16]. Pollinator [WZQH16]. poly [Tel19]. polygon [BBB^{+}17, KM11, NMB10]. polygonal [Lu15]. polygons [AAJ15, EH18, IM12, LWŻ07, LPWZ14, TJ14]. polyhedra [AGÁRCS^{+}19, Has13]. polylogarithmic [GS10a]. Polynomial [AR18, BFLM15, HCCG15, BES17, Brâ15b, CYH15, CLY11, CCH14b, DFdFT16, EN17, EGK13, GR19, Gil18, Gio18, GP17, GMY13, HYC12, KLS13, LlCyChL10, LT15b, MS18a, Man10, Mil19, Tat19, WWBB15, nXlCL14]. Polynomial-time [AR18, BFLM15, HCCG15, BES17, CYH15, EN17, GP17, LT15b]. polynomials [AH17, GJR10, LPdS10, Mon12]. polyomino [Bra17]. polytope [KT18]. pooling [KD13]. popularity [YBM^{+}10]. poset [Kal12]. Posets [HI11]. position [FM18, GM12a, KMS19, KY12, Mac19, Mos11b, NII^{+}15, Vou19, YZH14]. position-based [YZH14]. position-dependent [FM18, GM12a, KMS19, KY12, Mos11b]. Positional [MN19]. positive [CRJ10, HKK12, JT10, Mat15]. positively [EH18]. positivity [GQ17]. possible [BR12, SKN11]. Post [Fin15, Tho12]. potential [BDNVP15, Che18b]. Power [BLM10, AT18, BB11, BHL17, DCH12, HKW18, IM17, KSBT13, RV10, Str16, YK13]. powerful [YBHK13]. powers [BFP18, FL12, XZ17a]. PPP [BJP^{+}19]. Practical [OPS14, ĎHPT10, LWXS18, Val12, YTYZ15]. PRAM [Sor10]. prBPP [Tel19]. pre [AH17, BBR^{+}18]. pre-processing [BBR^{+}18]. pre-test [AH17]. precedence [BRMP13]. Precision [HZSL05, MN14, TLS^{+}18]. preclusion [YWL15]. precomputation [LZ12]. predicates [Ksh11]. Predictability [CGG10]. Prediction [ST18, CY18, FK10, STAR15, YMJ17]. predictor [ARdSP15]. preemptive [vdBCDH17]. preferences [GS17]. Prefix [DHR13, OG10]. Preimage [LP10, WS13, GSY15, MS13]. prenex [Har16]. prescribed [Che10a, YZ19]. presence [FG18a, TVB15]. preservation [AL18, SV19]. preserve [CS16a]. preserving [ÁCLS12, Bar13, CT16b, CNPS15, FWH14, GBH12, YL11b]. prevention [VB19]. previous [Tsu19a, CT11]. Price [YSGY10, DE19]. pricing [DJRB15, EGK^{+}12, NLXX15]. primal [MMP15]. primary [Wij10]. prime [Agg15, BHL17, CKK13, Den14, DCH12, Shp13]. prime-power [DCH12]. primitivity [ZV14]. principle [BGL10]. Priority [BC13, CSX16, ZSLW16, vdBCDH17, CGG10]. prisms [CCCN19]. Prisoner [Pel10]. Privacy [SS17]. private [AS18, DSPSHJNA18, IK15a]. private-keys [IK15a]. Probabilistic [BCS15, TL12, CDPB^{+}10, DLO12, DBB12, Ger12, Gio18, HS15]. probability [JF15, Til19, WWS12]. probe [CGLS10, DFdFT16, Nic19]. Problem [JK18, AR18, AD16, AR13a, APR13, AL18, AMMY10, Bab17, BK10, BLC10, BL12, BLYL17, BGJ^{+}16, BHM^{+}10, BBKS17, BR12, BTW15, BBDS12, BGR13, BDNVP15, BES17, BdSL15, BH17, CV12, CCF14, Cha12, CCL10, CCH^{+}14a, CFJ12, gCF15, CCT14, CGJY12, DFdFT16, Deo12, Doe13, DKNQ18, FKR^{+}16, GLW12, GKM^{+}15, GH15, GHRT17, GLO12, Gra15, GMY13, GJ14, HK16, HKW14, HL17, HHK17, Hen16, Hil14, HK17, IH18, JKSS19, JT15, JPB19, Juk12, KKZ13, KS11, KV10, Kos17, LMU15, LHW^{+}16, Lan11, Laz10, LXDX12, LLG14, LS15, LSK17, LSS15, LFW^{+}18, LXJ^{+}14, LM11, M.15, Man10, MU15, MÖ15, MSM14, MKI11, Mes15, Min11, MS15, Möm15, Mon15, Mor11, Nan13, OG10, PP13, PC18, Pér17, Pul16, RS19, RH10, SS19b, TLL18, Tak16, TCXT10, TZ11, TY13, Tu15, WWWZ13]. problem [WS16, WWBB15, Waw14, Xav12, XL15, XZ18, YNH^{+}14, YC11, YBMK15, YK15, ZST13, ZC10, ZXZL11, ZPX11, ZY12, ZSLW16, ZZ11, BI14, CFMT16, Fin15, GW16, LREIMBMV16, Rom11]. problems [AAOW15, AY12a, BNRC10, BFLM15, CCH^{+}14a, CXD^{+}13, CYH15, CP15, DDER18, DJ15, DLN11, DTS15, FR12, FH10, FLRS11, GHK^{+}18, Gav11, GBH10, HK16, JKY15, KK15b, Las17, Lib10, LL10b, LMS14, LZ10b, LWXZ14, LP13a, Man19, Mar11, MT18, MLW11, NVB15, Nut18, RDX13, STY12, TPL16, Xu12, YK11, ZS18]. procedural [CdA13]. Process [BB11, KMN18, KMS19, LPWZ14, SSW16, JCC11, MG16, TLS^{+}18]. processes [BCS15, EGK13, FCWZ18, vGGS11]. Processing [GR14, HAM11, KT12, Pas15a, Zha13, BBR^{+}18, FM18, GM12a, HLY11, KMS19, KY12, LL13, Mos11b, TCXT10, Tho18, WDT13, YSGY10, ZZJ11]. product [CPHS18, Che11, CCXW12, Gio18, GM12b, GZJ^{+}12, LJL18, LZCM10, SSW15, SSW16, XS15, YLLL16]. products [CBSV11, GQG10, HS19, LLL18a, Mei12, TWZ17, WSW12, ZLCM10]. Program [MIM16, CPTZ13, Glü10]. programming [GLMS18, GE12, JS19, LL17, MSM14, WMLN10, WWWZ13]. programs [Att17, BGI^{+}12, CC14, Dem18, JT10]. projection [Has13]. promise [AY12a]. Proof [Zha11, ADF13, Ber17, Bon13, BR10b, CVZ11, EAE21, FL12, FL13a, GR13, GR14, Had18, Kra12, Le 19, LMSN16, Pel15, ST18, SW12]. proofs [FK10, Ito14, Juk12, KS18, Lau18a, Mei12, Pud12, Yam14]. Propagation [CHP19, TVB15]. proper [AD13, BK10, BES17, BdSL15, CL15, LSS15, PP13, YeCM14, YYR16]. Properties [AC12, BD16, Den19, HCWH15, LP13a, MHFMSa11, SHFMSA11, SLF19, WT13]. property [CCL10, Ési11, FMHL11, JSO11, WWZ15]. proportionally [PYTH11]. proportionally-scaled [PYTH11]. Proportionate [KMS19, KY12, Mos11b, CLP10, DJZ^{+}15, FM18, GMM15, MM15b]. proposal [LM91]. protein [STAR15]. protocol [ASA13, BB15, CZZ^{+}12, JB16, JK19, LWS10, MP13, YL16a, YWcW^{+}14]. protocols [CR10, JF15, Ned17, Nos11, Nos14, TC11, TK15]. Prototype [BD16]. provable [HRS13, LHH11, WB12]. prover [Ito14, BGL10]. provers [Dem18, Yam14]. Proving [Sar14, Tel19, BD16, Li15, LG18]. provision [GL13]. Proximity [MWZ12, ADH14]. proxy [Tia15, YMSA14]. prP} [Tel19]. pruned [RM18]. pruning [HLY11]. pseudo [DPS11, KM10, WS13]. pseudo-collision [WS13]. pseudo-random [DPS11, KM10]. pseudometrics [vB12]. Pseudorandom [Nis92, DKC12, Kis19, LMCG16]. Pseudorandomness [LLG10]. PSPACE [BH17, GN10]. PSPACE-complete [GN10]. PTAS [JKY15, MB14, ZLS^{+}17]. Public [LPdS10, QX10, SY15a, VN17, ZY17]. Public-key [LPdS10, QX10, VN17, ZY17]. publications [BvK15]. pump [LL17]. pumpkin [Tsu19b]. pure [BR12, JS19, MM15a, THS12]. pure-strategy [THS12]. purely [XLZ16]. Pushdown [CO17, CS16a, DBB12, DK14, Lan11].

Puzzle [HWA12, LY19]. pyramid [TWZ17].

QoS [ZZJ11]. QPTAS [LL14]. Quadratic [Deo12, CK12b, vIL13, LREIMBMV16]. Quadratic-time [Deo12]. quality [CV12, ŽGH10]. Quantified [KS18, JS18, Mar11]. quantities [KT11, KT12]. quantum [AY12a, Ber11, DT10, Har18, LLP^{+}18, LFZJ14, LQL^{+}17, Mon10, Mon12, QD16, RS15, TZF16, Yam14, Zha10a]. quantum-inspired [LFZJ14]. quartet [CLR13]. quartets [DMS12]. Quasi [DF11, DKZ18, Cro15, GD13, Pue11, SQLS17, Win13]. quasi-cyclic [SQLS17]. quasi-kernels [Cro15]. Quasi-random [DF11, Win13]. quasi-symmetric [GD13]. quasiperiodic [AIR17]. quaternary [KZ12]. queries [Fre10, GHK11, KP12, Pue11, PH11, SGM13, SM17, SSKA17, SV19, Tsu19a]. Query [ADG10, DGKS14, ED17, HLY11, KR16, KP12, Mon10, Mon12, Wij10, YiN10, Zha10a]. querying [SZC^{+}17, Vrg15]. question [SB13]. queue [Cho12]. queueing [AZ14]. queues [MSV14, SD13, Tho18]. Quicksort [Fre14]. quiescent [AJLM11]. quintuple [Moo11]. quotients [DCH12, DKC12].

R [WYW14]. RAC [DEL10]. radiation [BDH^{+}11]. radio [Vay13]. radius [GWL18, GBH12, KM11, Kor12]. radix [GKCK11]. Radon [DRdSS12]. Radon-independent [DRdSS12]. RAID5 [Tho18]. Rainbow [LM15, HL12, KS11]. rainbow- [HL12]. RAMs [Bra15a]. Ramsey [Pud12]. Random [EAA^{+}16, LREIMBMV16, ABH19, BV10, BWZ12, DPS11, DF11, FP18, HL12, KM10, MT18, iP13, RT13, WWYY11, Win13, XGX11, ZZ11, CDM^{+}11]. randomization [DTS15]. Randomized [AY12b, BG11, EL10, KK15b, AGW13, Cha12, HI12, JKS10, MO12, Sor10, Wit14, YZZX12, Zeh16]. Randomly [FM11, FA17, KK16]. randomness [DF11, Jan12a]. range [DG16, SGM13, Zho16]. range-aggregate [SGM13]. rank [CÉ13, DDER18, EN17, FY14, Hru12, Kun17, Shr19]. rank-width [DDER18]. Ranking [CKL10, RT13, CC12, PCK10]. rankings [ZJN10]. Raphson [ZKXY10]. RAPP [ASA13]. Rate [AZ14, XZ10, AT18, CX18, CRJ10, HKK12, JC10, iP13, Par11, PP14, WLLS08, YA13, YZZX12]. Rate-adaptive [AZ14]. rate-modifying [JC10]. rate-monotonic [PP14]. rates [DJRB15]. Rating [YMJ17]. Ratio [AMO19, DJS13, Nan13]. Ratio-balanced [AMO19]. ratios [Par11, WLLS08]. RBAC [LXJ^{+}14, SL19]. RC4 [Sar14]. rDAN [AHLS18]. RDO [CY18]. re [GJ19, MS19, Tia15, XZ10]. re-analysed [XZ10]. re-keying [GJ19]. re-signatures [Tia15]. re-voting [MS19]. reach [Kar17]. Reachability [AAOW15, Bol10, CO17, Laz10]. reachable [Bor16]. Reaching [Pel18]. reactive [FCWZ18]. read [FG18a]. readers [Tho18]. ready [AdFEGRI11]. real [AMN^{+}10, BLM10, CGG10, HP18, LLK10, Man10, ZZJ11, ZGLY12]. real-time [AMN^{+}10, BLM10, CGG10, LLK10, ZZJ11, ZGLY12]. rearrangement [WBC^{+}12]. rearrangements [CXD^{+}13]. reasoning [CV12]. reassignment [MLW11]. rebuild [Tho18]. recently [FKC13]. recoding [Bae13]. Recognition [DGK^{+}17, LCC17, NMB10, DE14, DRdSS12, EFMA10, ZpH15]. Recognizability [KQ11]. recognizable [CK11a, KQ11]. Recognizing [GI18]. recommendation [GWJ11, HPP17]. recommendations [CS16b]. reconciliation [HMSS18]. reconstruction [HH15, IK15a]. recording [AL18]. recourse [AMP13]. Recovering [AABB17]. recovery [FCLCR17, LJP19, ML10a, MGPI12]. rectangle [DGKS14, Mei12]. rectangles [Ale13, MM19a, MM19b]. rectangular [AAB^{+}19, GRSS17, ZSLW16]. rectilinear [KM11]. recurrences [AABB17]. recursion [GY19]. Recursive [CFMT16, CCXW12, LXLY12, CT16a]. recursively [CVV10]. Red [HBL14, CCL10]. Red-black [HBL14]. Reduced [JUY15, AKY13, AY14, DMSD18, Gav17, LJP19, LFW^{+}16, MS13, WS13]. reduced-round [AY14, LJP19, LFW^{+}16]. Reducing [DGR15, KZP10, DF11]. reduction [GL15, GK10, GHC15, IM17, KKSŠ11, NCP18, Pet12, SeF14, SC12, YK11, ZZ19]. Reductions [BJP^{+}19]. redundant [KHKS16]. Reed [WS10]. Reference [LR10, DLO12, LH10]. Refined [Dam17, YK15]. Refining [GCNR10]. reflection [Kar13]. reflex [IM12]. reflexive [Vág18]. Reformulation [CFMT16]. region [YC19, GN10]. register [BR10b]. registers [DLRS14]. regression [BMI14, JZ18, Kis19]. regret [KKZ13, XL15]. Regular [AGHY12, JOS10, Bae13, BJMC17, Cal13, CGG14, CO13, CGR14, DFRS13, Kim10, LFXH17, RS10, SM19, Slí12, YK13, Yam19, YC19, ZZH16]. regularized [Min11]. Reingold [GGI11]. reject [JF15]. rejection [EZH16, FM18, GM12a, KMS19, MY13, OZL16]. related [CCH^{+}14a, DJ15, DMSD18, GLMS18, HK16, MNP12, PPH18, RS15, SB19, STU12]. related-key [GLMS18, RS15, SB19]. Relating [CB18]. Relations [HLR11, Mei18, Vág18]. Relationship [Suk19]. relative [HBL14, IM17, Par11, WLLS08]. relaxations [GE12]. Relaxed [Kun17, DL13, FKL^{+}11, Möm15]. release [AS18, LZCX12, OZL16]. Reliability [CWW10, CCRS14, Ger12, LT13b, LT15a, LT15b]. reliable [Chr19]. Remainder [HF14]. remark [Wij10]. Remila [Svi12]. rendezvous [BDPP18]. rental [CX18, YZZX12, ZPX11]. reoptimization [Mon15]. reordering [SF13]. Rep [Man19]. Repeated [Ksh11]. repeater [BHM^{+}10]. repeats [CCF^{+}12]. repetition [BBDS12, HN10]. repetition-free [HN10]. repetitions [BC15, KKO10, Pel15]. replacement [BOV15, LMSN16, YEMR19]. replenishment [YXW13]. replication [RJS^{+}10]. reporting [DGKS14, DG16]. represent [CGR14]. representation [JAH16, Pas15a, Pas15b, SD13, Tys13, Wal10, ZpH15]. representations [Zha11, ZZH16]. representatives [AGHY12]. representing [BBB^{+}17, KR10, KR14]. reproducible [CW12]. requests [Hil14]. required [WCW11]. requirement [MZQL14]. requirements [CV12, HKW18, KVK18, ZZJ11]. requires [Laz10]. Research [ZLM17]. reset [JUY15]. residual [LT15b, ZpH15]. Residuation [GS17]. residues [Bar12]. resilience [ABPS15, DDPBT11]. resilient [GV14, YL16a, ZY17]. resistance [LP10]. resistant [Ito14]. resolution [BGL13, BT16a, CB18, DKL^{+}12, Lau18a, Lau18b, MS16, Pud12, BGL10]. resolving [Lai16b]. Resource [AIS10, eSKAI10, IZ10, Jia11, Oro11, XCYZ11]. resources [NZX19]. respect [SeF14]. response [FT15]. Responsive [Lei18]. restarts [FCNY10, Win13]. Restricted [VV11, DW12, Für19, GKW15, GBH10, LJL18, MNV11, MV13, Nin13, Pal10, RS12a, WZL12, WS16, YL16b]. result [EAE21, EZH16, ED17, IH18, Li10, SV19, Til19, Tsa11, WYLP17]. resulting [ZQJ^{+}15]. results [AD16, AX18, BFRV15, DBFMPLV17, DRD11, DX10, JKS10, PH11]. RETRACTED [Zha10b]. retransmission [TK15]. retrieval [Ars19, Tsu13]. reusable [CNKS15]. revenue [NZX19]. reversal [BN10a]. reversals [BN10a, DK14, FKR^{+}16, Rus18]. Reverse [Ger19]. reversible [ASTD14, GY19]. reversing [Win13]. review [YMJ17]. revised [WYLP17]. Revisit [FL13b]. revisited [FR12, HPY^{+}10, LT13a, MZ15, NT14, Ria17, Ria19, RV13, SS10b, TS16, YZH14, ZCT^{+}12, WBC^{+}12]. Revisiting [GLMS18, Nic19]. revocation [Ria19]. reward [CX18, ZZXL11]. rewiring [SW18]. rewrite [Vág18]. rewriting [Fil13, FG18b, Ket11, Luc15, LG18, SO10, Vág18]. RFID [JF15, JB16]. rich [GSS16]. right [Bae13, JT10, Kim10]. right-hand [JT10]. right-to-left [Bae13, Kim10]. rigid [Jor10, Jor12]. rigidity [KHKS16]. ring [GWL18, LLG14]. rings [BEFP11, BHKK10]. risk [CX18, ZZXL11]. risk-reward [CX18, ZZXL11]. risks [SS17]. RNS [Pie15]. road [FBD18, SZC^{+}17, Rom11]. robot [CDPB^{+}10]. robots [FIPS11]. Robust [MG16, AHLS18, EAA^{+}16, LXDX12]. robustly [GM12b]. robustness [Lib10, Mei12]. role [LXJ^{+}14, MGNAB11]. Role-Based [MGNAB11]. rollback [CZZ^{+}12]. Roman [TLWZ16, ZLJS18, ZS18]. Root [Pal10, Gil18, ZLY13]. Root-restricted [Pal10]. rooted [Gol18]. roots [CCH14b, Mil19, ZKXY10, ZYL15, Zha18]. rotation [ALT16, Luc10, LEP10]. rotations [LP13b, Pal10]. rotor [MPU17]. rotor-router [MPU17]. Round [Vay13, AY14, Blo15, CVZ11, Chr19, DK10, DMSD18, JK19, JSO11, Kar13, LJP19, LFW^{+}16, Sun11, XW12, ZL17, ZL18, ZLNB19]. round-reduced [DMSD18]. round-trip [ZL17, ZL18, ZLNB19]. round-up [JSO11]. rounding [GKW15]. rounds [LYHH14, MNP12, YYZH13]. route [TJ18]. router [MPU17]. Routing [AKMR19, Liu15, YTN10, YYYZ12, YYZH13, CN18, IK10, SM19]. RSA [IK15a, MZ15, SM10a, SM10b]. RSBFs [SLF19]. Rule [MS15, DKZ18]. rules [BR12, Ket11, TWW17]. rumor [DF11, Win13]. run [BFKL13, GSS16]. run-length [BFKL13, GSS16]. running [AHR10]. runs [Kos16]. runtime [Che17, FG18b, YEMR19].

S [KSBT13]. S-boxes [KSBT13]. safety [BJKZ14, BD16]. sales [Gou15]. salesman [BL12, BLYL17, Möm15, Mon15]. samples [Til19]. sampling [BOV15, JKY15, Kun17]. sandwich [Ber12]. SAR [LFZJ14]. Sarkar [BB15]. SAT [Kra12, MS13, PS19a, SW13, XGX11, XZ17b]. SAT-based [MS13, XZ17b]. satisfaction [DGK^{+}17, Las17, Mar11, ZZ11]. Satisfiability [GW16, ZST13, BB12, LP13a]. satisfying [SSK12]. saturated [GNV14]. Scalable [LLK10, SS16, SLL13, VB19]. Scalar [Bae13, HZX16]. scale [SS16, YWC11, ZCC18]. scale-free [SS16]. scaled [PYHA10, PYTH11]. scanning [ZLM17]. Scattered [CVV10]. schedulability [ABPS15, CSX16, Par11, WLLS08, vdBCDH17]. schedulers [CGG10]. Scheduling [GM12a, JC10, LJ15, Oro11, XCYZ11, AZ14, ACL18, CZCD13, CXD^{+}13, CSX16, DJS13, DKNQ18, DJZ^{+}15, EZH16, FL16, FCNY10, FTYL14, FS12, FPY12, GLW12, GY15, GL13, Hav10, HL17, HYK14, KC11, LL10a, LY11, LYF11, LZY12, LZLY12, LZ14, LL13, LZCX12, MY13, MZC11, MLW11, MM12, OZL16, PP14, RDX13, SF13, Sun19, TCXT10, Tho18, TFY11, WBC^{+}12, XZBX16, XZ10, XMLL11, YW12, YXW13, YZH14, ZvdV10, ZLZX19, ZZJ11, ZGLY12, Zhu12, vdBCDH17]. schema [WYLP17]. Scheme [ZJ16, CL10, CTHP13, CKY15, DWQ10, DKNQ18, jDX11, EAA^{+}16, EZ15, GS19, HKK12, KDH15, KWH16, LXLY12, LKC^{+}12, LZJX10, LHH11, LWL10, LLG10, Nos14, OPS14, SW13, SPdR13, TY16a, Tat19, TD14, WWYY11, WZS^{+}18, WCW11, YA13, YMSA14, YW14, ZY17]. schemes [ADFM13, EP11, FGMP12, GHRT17, IN12, LZ12, Nic19, QYWX16, Sar11]. Scholar [BvK15]. Schweitzer [KK11]. score [KTUY17]. scoring [BR12]. scream [DMSD18]. SCS [GKM14]. SD [BS11]. SD-functions [BS11]. search [CBHW10, CYQ13, DT10, DHW13, JB16, LH10, SY15a, SO19, Suz18, VAC13, Wit14, ZXZL11, CG10b]. Searching [DLRS14, CW13, KZP10]. SEC [BB11, RV10]. Second [AKY13, MP15]. second-order [MP15]. Secondary [SLF19]. secret [EZ15, HF14, HH15, LXLY12, Mas17, QD16, TD14]. secrets [HRS13]. Secure [MC15, ZJ16, Gal13, JPB19, KVK18, LX17, WZS^{+}18, ZY17]. securing [CR10]. Security [BB15, JB16, LK14, Nos11, Nos14, SB19, YMSA14, CTHP13, EMS15, HLR11, HRS13, JK19, LHH11, NTD16, QYWX16, RH10, SY15a, YL16a, YW14, ZGLY12, XW12]. security-aware [ZGLY12]. sedenions [CC13]. SEED [LYHH14, Sun11]. segment [KPR18]. segmentation [FBD18, LFZJ14, Tat19]. segments [AD18]. segregation [ABS13]. Selecting [Doe13, KKZ13, Til19]. selection [AC10, Cha13, Cha19, CA17, GE12, GPT16, HYYZ15, yJxW16, PdAL18]. selector [SaBG17]. Self [Glü10, Kak15, SW19, Tur13, YBHK13, BYK14, CCT14, EAE21, FZ13, Har18, HHJ^{+}12, Joh14a, MY18, MMKK18, NHK15, Shi12, SQLS17]. self-adjusting [MMKK18]. self-dual [Har18, SQLS17]. Self-generating [Glü10]. self-similar [EAE21, MY18]. self-stabilization [FZ13]. Self-stabilizing [Kak15, SW19, Tur13, YBHK13, BYK14, CCT14, HHJ^{+}12, Joh14a, NHK15, Shi12]. selfish [DE19, GL13, WCJ15]. semantic [HLR11, LT18]. semantics [AdFEGRI11, ADG10, AB15, BGM16, DW10b, KL10].

Semi [CXD^{+}13, CNKS15, Dha14, DW10b, FG18b, LQL^{+}17, MLW11, PGZB19, SD19, SF13, TCXT10]. semi-automated [CNKS15]. semi-bent [PGZB19]. semi-decidable [FG18b]. semi-external [Dha14]. Semi-online [CXD^{+}13, MLW11, SF13, TCXT10]. semi-quantum [LQL^{+}17]. semi-stable [DW10b]. semi-supervised [SD19]. semigraphoid [GNV14]. semiquantum [MP13]. semiring [KQ11]. semirings [BHKK10]. sensing [YLGC16, Fay16]. sensitive [Ars19, LXX14, LWXS18, Luc15, NG10]. Sensitivity [Vir11, Ama11, Li12, LL16]. sensor [jDX11, LKC^{+}12, SKK10, VLV15, YA13, YTYZ15]. sensors [CWW10, KK16]. sentiments [YMJ17]. separable [BTW15]. separator [Iba17]. separators [BP11]. Sequence [NVB15, Bra15a, GGI11, Gou15, KZ12, KYC13, LZCX12, WJS10]. sequences [AG19, AABB17, BH11, BMW13, Dam14, DCH12, DKC12, EP16, Ehl17, HYC12, JUY15, KPSZ11, LH11, WXCK19]. sequential [DMS11, LABKS17]. sequentially [Mos13b]. Serial [LZ14]. Series [GBH12, CK11a, KQ11]. Series-parallel [GBH12]. server [BLM10, DEH^{+}19, EFKR10, HKW14, Hil14, SY15a, Tho18]. service [GL13, Bee16]. Set [BI14, KP14a, ADG10, AGHY12, BR10a, BU17, BV19, BIR16, CCL10, CCH^{+}14a, CCT14, DDK^{+}15, DD14, DMS12, FV13, GG18, GR19, Joh14a, KP13, KM16a, KM16b, Kun18, LZ10a, LX13c, LS15, LSK17, MM19a, MM19b, Mor11, NMB10, NPR17, NS11, NG10, Nut18, RS19, RR16, SWW16, Tsu19b, Tur12, Waw14, XN12, ZZH10a, ZC10, ZY12, ZMZ19, AK10, BdSL15, CD10]. Sets [AG19, AKH16, Agg15, AADB^{+}18, Aka14, BHLM13, BYK14, BDF^{+}18, BM19, BKN^{+}12, CGG14, CK12a, Den14, DBFMPLV17, DRdSS12, DM19, FMHL11, FKC13, FX11, GGG^{+}14, HHJ^{+}12, KKKS14, KT18, Krz13, LL14, LS11, LS14, LLW15, LMV17, LC17, MÖ15, Mos13a, Sax10, SS12, Sch15, SS10a, SW19, SF12, Suk19, Tur13, YBHK13, ZZ14, ŽGH10, vGLM12, BG12, BG15]. setting [LJP19]. Settling [KM11]. setup [Jia16]. seven [HI11]. several [BTW15, Lai16b]. shallow [LR10, SO10]. Shannon [PS19b]. Shape [ZZ13, Zun18, AP14, MMKK18]. shape-defining [MMKK18]. shaped [LWŻ07]. shapes [KP14b]. Sharing [HRS13, ZJ16, EZ15, FGMP12, HF14, LXLY12, Mas17, QD16, TD14, WZS^{+}18]. shift [DLRS14, LLWH13]. shifted [CCH14b]. shop [CZCD13, DJZ^{+}15, ZvdV10]. shops [CLP10]. Short [FK10, KS18, AB15, Ber17, Bon13, Duc18, KY12, Ned17, ZW10]. Shorter [Zal11, ZZH10b]. shortest [AD16, Bon13, CS16a, Lai16a, RAT13, TJ18, WJS10, XZ18, Yam19, Yus11, ZLNB19, FKR^{+}16]. should [SSS10]. should- [SSS10]. shrinking [Koc18]. shuffle [HS19]. shuffling [YEMR13]. side [ABPS15]. side-channel [ABPS15]. Sidelńikov [BMW13]. sides [JT10]. sign [AMT12]. sign-definite [AMT12]. signals [DM10]. signature [LKC^{+}12, Nos14, QYWX16, XLWZ16, YMSA14, YW14]. signatures [Tia15]. signcryption [EZ15]. Signed [KWH16, DW10a, IMCP15, Rus18]. significant [BG11]. Signum [ZDQ^{+}17]. Signum-function [ZDQ^{+}17]. silent [Joh14a]. SIMECK [SB19]. similar [EAE21, MY18]. similarities [KK11]. similarity [PCK10]. Simon [KKS15]. Simple [JLMO17, Nan13, SS19a, BBB^{+}17, BP11, Bir11, BR10b, EH18, EGKL11, Fre14, FM11, FA17, Had18, IL12, IM12, JSO11, KM11, Le 19, LMSN16, LTWS11, MMM11, ML10a, NT14, Pet12, Sch13b, TJ14, Wal10]. simple-path [MMM11]. simpler [Bol14]. simplicial [Has13]. simplification [Fil18]. simplified [LSW17, LL16]. simply [Sch14]. simulation [AdFEGRI11, SSSM11]. Simultaneous [YLGC16]. Single [LL10a, LZCX12, PI13, ZLZX19, ÁCLS12, CK11a, DJ15, DP12, FL16, GLW12, GPT16, HKW14, HYK14, LSLY11, LJP19, LJ15, MY13, MZC11, OZL16, RV17, SD13, TCXT10, TFY11, TL12, YW12]. single-key [LJP19]. single-level [SD13]. Single-machine [LL10a, LZCX12]. Single-source [PI13]. single-vertex [ÁCLS12]. Singleton [AMT12, AdFEGRI11, MTA10]. sink [XL15]. site [CS16b]. six [ZYL15]. Size [LRS19, Rud17, BGL13, CB18, DMS12, Elm12, GJ11, Han17, Har16, Has13, Luc10, Pud12, RR16, Sun19]. Size-optimal [LRS19]. sizes [HM18]. skeletons [EH18, MK16]. skeptical [NAD19]. ski [CX18, YZZX12, ZPX11]. ski-rental [YZZX12, ZPX11]. skyline [LXX14, SSKA17]. slide [LC13]. sliding [BKZ15, HLT10]. slopes [AD18]. Slot [YWcW^{+}14, WCW11]. SM3 [WS13]. Small [Bra17, AG19, Aka14, BvdZ19, DD14, GKS13, HN10, Kob15, LT18, MZ15]. smaller [Tsu19a, ZFJZ11]. smallest [Ama11]. smart [LLK10, CL10]. Smooth [CDM^{+}11]. SMS4 [LSLY11]. snake [RRR12]. snow [SML^{+}10]. social [AFPT10, CS16b, EDS11, Lee17, ZD18]. soft [BLM10, GS17, SS17]. software [ABPS15, MZQL14, MK11, SF12]. solid [AGÁRCS^{+}19]. Solitaire [KTUY17]. Solomon [WS10]. solution [CYQ13, JAH16, WWWZ13, YCL11, ZZZ15]. solutions [BLC10, BDNVP15, Car19]. solve [BGR13]. Solving [CFMT16, GKM14, SB13, Che16a, LXT19, XLT^{+}19, ZDQ^{+}17]. Some [AX18, BS17, BFRV15, Den19, DRD11, DX10, Dur12, HZX16, Joh14b, PPH18, vE17, ALT18, CFMR18, DJ15, Har18, Ind15, KS18, LMS14, LV15, Mat15, Vaj18, WWZ15, WS16, YK11, ZHXS10]. sort [zSfNbL11, WL11a, YWL15, vZBSY16]. sortable [HI11]. sorted [Ehl17]. sorter [Ehl17]. Sorting [Rus18, Dha14, Ger12, HHTL10, LGT17, SKN11]. sorts [PWC^{+}15]. Source [ZL17, PI13, ŽGH10]. Source-wise [ZL17]. Space [LFXH17, BHKK11, Bol10, BG11, Cha12, DG16, DTS15, Ezr10, FIV15, GJ15, HMS16, Jan12b, Laz10, LLK10, MPU17, Nis92, PCK10, Pea16, Tsu13, WWBB15, Yam14, Zho16]. space-bounded [DTS15, Nis92]. space-efficient [Pea16]. spaces [HMSS18, KZP10, KLS10, MSZ11]. spam [Ars19]. span [CPTZ13, JT16]. Spanners [FGvL11, BF17, ZL17, ZL18]. Spanning [CFJ12, AHK^{+}17, CWYP14, Che11, Dat15, EAE21, FX11, GR13, GR14, HL16, KMK18, LYHC10, MY18, PC19, PP10, Suk19, ZY12, ZHX13, KT18]. SPARQL [ZV14]. Sparse [GE12, BV19, BDNVP15, BMI14, CHR10, Çiv13, DGŚN15, GJR10, MI17, Mil19, MRZ10, YL17, ZpH15]. sparsification [DZ11]. spatial [SSKA17]. spatio [HPP17]. spatio-temporal [HPP17]. special [SI18, WS16]. specializers [Glü10]. specific [GHK11, LLL18b]. specifications [Dia15, DBB12]. specified [Ama10, Hua18, SS12]. SPECK [LFW^{+}16]. spectral [AT18, GS10b]. spectrum [AC10]. speed [XCYZ11]. speed-up [XCYZ11]. Speeding [Gue12]. speeds [DKNQ18]. Speedup [ABP11, LlChL11, vdBCDH17]. Sphere [MS15]. spider [BKPP18]. spin [ZLB11]. spined [ZFJZ11]. Split [CP13, DRdSS12, LN11, Che18a]. Split-Star [Che18a]. splitting [AY12b, GPSD17]. Sponsored [DHW13]. spread [EDS11]. spreading [DF11, Win13]. squad [YNH^{+}14]. Square [Kar13, AAB^{+}19, BGI^{+}12, DL13, GH15, GJ15, JZ18, Mac19, ZKXY10, ZCC^{+}11]. square-freeness [BGI^{+}12]. squares [BMI14, CZD14, LN11, Min11]. squaring [NCP18]. Squeeziness [CH12, CHP19]. SRPT [FS12]. stability [ABS12]. stabilization [FZ13]. Stabilizing [DDPBT11, BYK14, CCT14, HHJ^{+}12, Joh14a, Kak15, NHK15, Shi12, SW19, Tur13, YBHK13]. stable [BKN^{+}12, DW10b, Miy14, OBT12, SD19]. stack [YEMR19]. stacks [SD13]. stage [DJZ^{+}15, DW10b, LH11, Mas17, Zhu12]. standard [FTYL14, LM91, Mas17, TS16, WWBC14, YW14]. standing [Bol14, YKD^{+}12]. Star [XXZ14, CLS13, Lai16a, LWŻ07, NHK15, RSwW11, WXZ^{+}12, YLG10, YeCM14, Che18a]. star-shaped [LWŻ07]. stars [BGJ^{+}16]. starting [Mac19]. starwidth [vE17]. stashes [MN16]. STAT [KV16]. state [Att17, MGPI12, QD16, TZF16, ZLB11, HYC12]. stateless [AGI15]. statements [GNV14]. states [BE16, NB12]. Static [KRV16, OG11, RK15]. station [PC18]. stationary [CKP16]. statistical [Car19, KN13a, KD13]. statistically [ZC12]. Statistics [Bar13]. Statistics-preserving [Bar13]. Statman [SW12]. stealing [SLS16]. steganography [GKCK11, Sun16]. Steiner [WWBB15, AABCC19, BLC10, GS10a, Vyg11]. step [AKY13, BR12, FV13, WS13]. step-reduced [WS13]. steps [GKM14]. Stern [ACD18]. sticks [M.15]. stochastic [Ada11, EGK13, SS18]. stock [JSO11, LLL18b, TWW17]. straight [BR10a, BGI^{+}12, Did13, EH18, SSS12]. straight-line [BR10a, BGI^{+}12, Did13, SSS12]. strategies [GLO12, Pel10]. strategy [FC14, KD13, THS12, BGM16]. stratified [CCKP13]. stream [KM10, Til19, LKR17]. streaming [BBD^{+}12, Cha12, Zel11]. streams [BOV15, FKC13, VML18]. strengthened [Waw14]. Stretching [BVF12]. Stribog [AY14]. strict [CZZ^{+}12, GQ17, KdER^{+}14]. String [GFG11, ASM17, AGW13, BFKL13, CD19, CIK^{+}13, Deo12, ĎHPT10, FB10, Sal12, TLL18, APR13]. strings [AIR17, BFP18, BGI^{+}12, DGG^{+}19, DHR13, GKM14, LS10, Lut14, MR10, PYHA10, PYTH11, SSS12, Zal11, ZZ14]. strip [Svi12]. strip-packing [Svi12]. striping [EMS15]. strips [Jia18]. Strong [CLS13, HOV13, ML10b, WWW18, YZZ17, AG15, CGJY12, DPS11, DGŚN15, Duc18, NN17, BT16a, GW16]. Strongly [Tat19, EGK13, FK17, LN11, PG12, Pea16]. structure [CPTZ13, JAH16, KR10, KMMN15, LXLY12]. structured [Dia15, LMU15, Tut13]. Structures [ZZ14, AALL16, CNKS15, Csi12, DG16, SF17, Tsu18, WMLN10]. study [MK11, SS17]. Sturmian [Pel15]. sub [YNH^{+}14]. sub-generals [YNH^{+}14]. subclass [GNG11, Pue11]. subclasses [ALT18]. subcoloring [Och17]. subcubic [BPP^{+}18, FL13a, JT16]. subdivided [FH12]. SUBEXP [Mos11a]. SUBEXP-dimension [Mos11a]. Subexponential [FLRS11]. subformulas [SW13]. subgraph [AK14, ABP11, BK10, GKW15, KS11, MU15, Slí12, Tak16]. subgraphs [CZ16, Koc18, KL18, LLL18b, MT10, Nin13, SS12, Wid17]. subgroups [ASTD14]. subject [Fuj16, SeF14]. sublinear [BKMT14, Cha13, Sor10, Wu14]. sublinear-time [Cha13]. submodular [Fil13, Fuj16]. Subnetwork [YWL15]. subposets [BP14]. Subsequence [Pol18, BBDS12, BVDP10, Cha12, DG14, Dur13, Elm10, IH18]. Subset [Xav12, CP11, Elm12, FC14, Fre10, HK16, Nan13, Ned17, XN12]. subset-sums [Nan13]. subsets [EO13, SS18, Tiw14, vzGS15]. subspace [LR18]. substitutability [CM18]. substitution [RK15]. substring [Gra15, OG10, Yam19]. substring-prefix [OG10]. substrings [DG14, FGKU15]. subtree [CCF^{+}12, Shi15]. subwords [Wal10]. successor [Zho16]. Succinct [Tsu18, BN10a]. succinctly [Att17]. Sudan [WS10]. Sufficient [Dam16, YL16b, Cai15, TD14]. suffix [Dha14, KTV13, LNP11, LGT17, OG10, RK15]. suffix-prefix [OG10]. sum [AGÁRCS^{+}19, AX18, BDNVP15, CCXW12, CKY15, CK12b, DS15, DS16, Hua18, JKS10, JN12, Li12, LL16, Mei18, Ned17, Shi13, YZRC18]. sum-optimal [JN12]. sum-product [CCXW12]. summarization [Lei18]. summation [KKKS14]. sums [Nan13, Tiw14]. sumset [BFRV15]. Super [GQG10, ZF10, CV19, CCYW16, MZ11, Nin16, WSW12, YL16b]. Super-connected [ZF10]. superclass [BL14]. Supereulerian [AALL16]. Superiority [AY12a]. supermodular [Ber17]. Superstring [FKR^{+}16]. supervised [GS10b, HYYZ15, SD19, SaBG17]. supply [RDX13]. support [JZ18, KQ11, OG11]. Supporting [HLY11]. supports [CK11a, YiN10]. Suppression [QK15]. surfaces [LT13c, WWLC14]. Surjective [ADF13]. swap [FG14]. swapped [MHFMSa11, XPC^{+}10]. sweep [GM15]. sweeps [TJ14]. switching [WB12]. Sylvester [Che16a]. symbol [Kim10]. Symbolic [VB15]. symmetric [Ara10, Ber17, DKKY10, FLMQ10, GD13, KS18, OBT12, SKK10, XPC^{+}10]. synchronization [YNH^{+}14]. synchronizing [YYK17]. synchronous [vG18]. synthesis [GS12, MIM16, Sch14]. system [CC14, GKCK11, LLK10, LXT19, MK11, MS19, Ria19, STAR15, TWW17, TC17, XLT^{+}19, ZLB11, Zha11, BD16]. systems [AMN^{+}10, ABH^{+}14, BR10b, CVZ11, CZD14, CW15, CSX16, CGG10, DBB12, DK14, FCLCR17, Laz10, MSV14, Ria19, SO10, Vág18, VSP18, ZDQ^{+}17, ZLNB19, vGGS11, LST11]. systolic [GS10c].

T [eSKAI10]. T-Boxes [eSKAI10]. Tables [PRM14]. tactics [CV12]. tag [JB16, PCK10, YBMK15]. tagged [HKK12]. tail [Wit14]. Taking [BR12]. Tamari [BP14]. tarai [IK15b]. tardiness [GLW12]. target [Pel18]. tasks [KC11, ZZJ11]. technique [DJS13, PYYC16, TS16, ZLM17]. techniques [HZX16, Li15, VN17]. teleportation [TZF16]. Template [BPPS18]. templates [Las17, XZ17b]. Temporal [DRD11, BvdZ19, HPP17, KK15a, Mos13b, SWLX15, CFMT16]. temporary [YXW13]. tensor [GM12b, Mei12]. term [Ars19, Fil13, FG18b, LG18, SO10, Vág18]. term-weighting [Ars19]. terminal [GS10a, KR14, LT13b, LT15a, LT15b, ZY12]. termination [IK15b, MP15]. terms [IM12]. ternary [DHR13, LWL11]. tessellations [AP14]. test [AH17, BGI^{+}12, BPPS18, Ghi14, GJR10, SSS12, Sch13b, GMY13]. testability [KKS15]. Testable [KR16, GM12b, HS19]. tester [SY15a]. Testing [ABH^{+}14, CKP16, CLR13, AHR10, AGZP15, CdA13, CM18, DJ15, GR19, IL12, Jan12a, RM18, SLC15, SSZW16]. tests [KR16, vdBCDH17]. text [GGF13, Lei18, PYW^{+}13, RK15, ZD18]. texts [BFKL13, YMJ17]. th [HK16]. their [BJMC17, DKC12, IK15a, NVB15]. theorem [BD16, Ber17, Dem18, EGKL11, FL13b, GSZ13, IM12, Mor16, Pud12, Ser10, Shr18, SW12, SSK12, HF14, KS12]. theoretic [CGS18, CH12]. theories [FCWZ18]. theory [BNRC10]. therapy [BDH^{+}11]. There [KS10]. Thin [Mou19]. Thorup [HP19]. Three [Koc18, AR18, EP16, EZH16, HI11, Jor12, PI13, Vaj18]. three-disjoint [PI13]. three-level [EP16]. Threshold [ZZ11, CPTZ13, FGMP12, GS19, HF14, HH15, KLS10, NLZX14, Pod12, QD16, Tho18, TD14, YMSA14]. thresholds [AD11, ZA17]. Thue [KMN18, GSZ13, MN17]. Tight [BU17, GW16, HYC12, Kos17, EZH16, GY10, Jan12a, MCS12, vZBSY16]. tighter [FZ13, XGX11]. Time [BT16a, GW16, MPU17, ZKXY10, AHR10, AR18, ABH19, AGHY12, AMN^{+}10, ADFM13, BCNPL14, BKMT14, BLM10, BRFGL10, BES17, BFLM15, Cha13, CBHW10, CHK13, CYH15, Che18b, CG10b, CGG10, CP13, Deo12, EN17, FL16, FZ13, FCWZ18, GXZZ13, GM13a, GMM15, Gio18, GFG11, GP17, GPR10, HCCG15, Han17, HKW14, HWJ96, HAM11, HLT10, Iba17, KLS13, KM16a, KKR^{+}13, KM16b, Kun18, Kut12, LLK10, LKC^{+}12, LY11, LZ14, LH11, LT15b, LC17, LXT19, MY13, MZQL14, MMM^{+}17, MS18a, Man10, MS18b, MTA10, MOW17, NB12, NII^{+}15, PP13, PI19, PR17, RZ14, Sor10, SWZ12, SWW16, TCXT10, WCW11, Wu14, YW12, YL17, ZvdV10, ZQJ^{+}15, ZDQ^{+}17, ZLZX19, Zho16, ZZJ11, ZGLY12, vdBCDH17]. time-bound [ADFM13]. time-constrained [AGHY12]. Time-varying [ZKXY10, LXT19, SWW16, ZQJ^{+}15]. times [CK12b, FM18, GM12a, KMS19, KY12, LZCX12, Mos11b, TFY11]. Toeplitz [CPHS18]. token [TNN11]. token-based [TNN11]. tokens [Ara10, DP12]. tolerance [LX13a, LX13b, WWS12, YLG10, ZXY10, ZX11]. tolerances [Lib10]. tolerancy [BF19]. tolerant [Che10b, jDX11, Far10, Fu10, IK10, LXDX12, MPRC19, Pel10, Tsa11, WDH16, YZ19, YC11]. tool [PdAL18]. tools [JK18]. Top [SSKA17, Tsu13, HLY11, LLT14, LRS19, SGM13]. Top- [SSKA17, Tsu13, HLY11, LLT14, SGM13]. Topological [GKPP16, MMS15, PWC^{+}15, WCZZ12, FS14, MDB14]. topologies [CLR13]. topology [TNN11, VLV15]. tori [YDT10]. toroidal [Xu11]. torus [Che16c, DZA15]. Total [Gol18, LH11, RZ10, WW11, AG19, ALT18, BYK14, BH17, BS17, CHR10, Cho12, Cza13, FL16, FL13a, FM18, GLW12, HKW14, KSK13, KMS19, LZ14, MY13, Pra12, SS12, WLW11, XS15, YYR16, YW12, YZH14, ZW10, ZLZX19]. Toughness [ZJ11, Liu10, ZSY13]. tournaments [CGJY12]. TPTL [BJKZ14].

Trace [ABPS15, Pas15a, Pas15b]. Trace-based [ABPS15]. traceable [Tur12]. traces [MK14]. tracing [MK16]. tracking [RPL19]. tractable [LMS13]. trade [LST11]. trade-offs [LST11]. trails [BK13]. train [Kar17]. trajectories [GP17]. transactions [DSPSHJNA18]. transducer [TK15]. transductions [MS18a]. transfer [Ria17]. transform [CK11b, DGG^{+}19, MN19, Nan15]. transformation [BGR13, FBD18, SKN11]. transition [vGGS11]. transitions [ZZ11]. transitive [Ama11, Vág18]. transitivity [Lü19, PCY16]. Translation [ZCC18]. translations [Dia15]. translocations [GFG11]. transmission [Chr19, KVK18]. transportation [Zhu12]. transpose [TC17]. transposition [TLL16, TLL18, YLM10]. transpositions [DHR13, vZBSY16]. transversal [KN13c, LS11, LS15, LSK17]. transversals [LV15]. trapdoor [CW12]. trapezoid [CG10b, Ili13, LT13b, LT15b]. traveling [BL12, BLYL17, Möm15, Mon15]. traversals [GY19]. TrCBC [ZWWL12]. Tree [WWBB15, ZLS^{+}15, ACK11, BLC10, BHM^{+}10, BGL10, BGL13, DKL^{+}12, Feg19, FS13b, GY19, GBH10, HHK17, JKS10, JKSS19, KLS13, KM16b, Kun18, Lau18a, LP13b, MS16, SLdAMP17, Suz18, Tsu18, VB15, VK18, WYLP17, XLZ16, YBM^{+}10, ZY12, vISS10, CFJ12, KT18, Shi15]. tree-automatic [JKSS19]. tree-colorable [XLZ16]. Tree-core [ZLS^{+}15]. tree-coritivity [ZLS^{+}15]. tree-like [BGL10, BGL13, Lau18a, MS16]. tree-width [Feg19]. Trees [Krz13, AABCC19, AHK^{+}17, Aku10, AB15, ALT16, Ama10, BR10a, BSM14, BBD^{+}12, CWYP14, Che11, CLS13, CCF^{+}12, DM13, DZFY12, EAE21, FH10, GS10a, GHK^{+}18, GI18, GR13, GR14, HBL14, HL16, HMSS18, KM12, LNP11, LX17, LYHC10, Luc10, LEP10, MY18, Maß15, Mis15, Mou19, PC19, PP10, PPN^{+}17, Rus18, SLC15, SM13, Suk19, Tur13, Vyg11, YLM10, ZHX13, ZLJS18, vGLM12, vIL13]. treespace [CDD^{+}11]. treewidth [CO17, MS18b, RR16]. Triangle [ZZ14, LX19, Möm15, PT12, SWZ12, Yan14, YL16b]. triangle-free [SWZ12, YL16b]. triangles [HRW11, Koc18, Zha10b, ZL12, Zha13]. triangulated [LT13c]. triangulation [AAJ15, LL14, SI18]. triangulations [LFXH17, YiN10, ZLSX16]. tries [BV11, NII^{+}15]. Triggering [Cha11]. trinet [MOW17]. trinomials [LlChL11, PCH13]. trip [ARdSP15, ZL17, ZL18, ZLNB19]. tripartite [GZM15, NN17]. triple [Att17]. triples [Nin13]. trivial [Ale13, Lut14, OBT12, Tur12]. TSP [BC13]. tuple [ALT15, ALT18, Pra12]. Turski [Tar14]. TV [Ria19]. Tweakable [Sar11]. TWINE [KDH13, TY16b]. twisted [CCYW16, LYHC10, PC19, YWY12]. Two [AIR17, BN10a, LZ10a, LL10b, TLL18, WL11a, Zho16, AADB^{+}18, BT16b, CX18, CK12b, CVV10, Dem18, DKNQ18, DJZ^{+}15, FCNY10, GM12b, GBH10, JLdSO^{+}14, KYC13, KT16, KP12, KS10, MLW11, MRZ10, NB12, Nic19, QYWX16, RDX13, SM10b, SI18, SW19, TNN11, TJ14, TC11, TX11, Vág18, WL11b, WL12, WCJ15, XZ17a, XCYZ11, YZRC18, ZvdV10, ZLB11, Zhu12]. Two-dimensional [WL11a, Zho16, KS10, LL10b, TNN11]. two-layer [KT16]. two-level [KYC13]. two-machine [ZvdV10]. two-option [CX18]. two-party [TC11]. two-probe [Nic19]. two-stage [DJZ^{+}15]. two-state [ZLB11]. Two-string [TLL18]. Two-way [BN10a]. type [AB15, ACD18, CZ16, CC14, GJ19, KS19, Li10, LH11, Nin13, SC12, WB12]. type-inhabitation [AB15]. types [OG11]. typing [OG11]. TYT [LlCyChL10].

ubiquitous [LLK10]. ultralightweight [ASA13]. unambiguous [HS15, IL12].

Unbounded [LYF11, FL16, FTYL14, GY15, LY11, MSV14, TFY11]. uncertainty [JCC11, NLXX15]. unconditionally [Tys13]. Undecidability [MSV14]. undecidable [AL18, Ési11, KQ11, Sch14]. underlying [YMJ17]. undirected [BS12, Cha11, Gro15]. unequal [Til19]. unfair [CCT14]. unicyclic [ZLY13]. unidirectional [HCWH15]. unified [CKY15]. uniform [CKP16, CFJ12, DKNQ18, KLS10, LZ14, WWS12, WCJ15]. Uniformly [ZCC^{+}11]. Unifying [FCWZ18]. unigraphs [Bar12]. Union [GI18, OG11, Ale13]. Union-Find [GI18]. unique [AD16, CFMR18, KL18, Mac19, ZLNB19]. uniquely [Für19]. unit [DDK^{+}15, DGR15, EZH16, GM13a, HPY^{+}10, Hil14, KK16, LZY12, NPR17, Sun19]. unit-size [Sun19]. unit-time [GM13a]. Universal [Fil18, SPdR13, vGLM12, BKS16, CY17, Csi12, Mei18]. universally [HP19]. unknown [AJLM11, ZMZ19]. Unpaired [Che10c]. unpartitioned [DFdFT16]. unranking [RT13]. unrelated [ACL18]. unreliable [JUY15]. Unrestricted [CG10b]. unsigncryption [EZ15]. unsupervised [PdAL18]. update [DLO12, GLO12, HLT10]. updates [Gil18]. updating [LXJ^{+}14]. Upper [ASTD14, CA17, DW10a, DEH^{+}19, GNG11, SI18, CP16, DKKY10, GD13, KTUY17, VK18, VSP18, XGX11, vZBSY16]. upward [AHR10]. usability [ŽGH10]. usage [HPP17, Sto16]. Use [LG18]. Userrank [GWJ11]. users [CS16b, HKW18]. Using [Moo11, RK15, SLL13, Vrg15, ÁCLS12, Ara10, Bee16, BV19, CLY11, CCT14, CCH14b, DM16, Dem18, FB10, Ghi14, GR19, GK10, GHC15, HPP17, JUY15, Kim10, LXLY12, LLP^{+}18, LSLY11, NTD16, Par11, QD16, RR16, Rus18, Sar11, SPdR13, TNN11, TVB15, WLLS08, nXlCL14, YEMR13, YMJ17, YEMR19, ZZH10b]. uSVP [ZZ19]. utilities [OBT12]. utilization [PP14].

Vacationing [Tho18]. validity [vG18].

valuations [Brâ15b]. value [SLC15, Tsu19a]. valued [CHK13, CGLS16, DZ11, KM13]. variable [Ghi14, LJ15, Vrg15, ZZH10a]. variables [FLPS15, Sch13a]. variance [AGW13]. variant [ZFJZ11]. Variants [BVDP10, BJMC17, CB18, VNP10]. variation [Mis15]. various [ZQJ^{+}15]. Varying [GKCK11, LXT19, SWW16, ZKXY10, ZQJ^{+}15]. vault [NTD16]. VByte [LKR17]. VC [Gil14, Joh14b, ZLNB19]. VC-dimension [Gil14, ZLNB19]. vector [AD16, CPHS18, JZ18, JS18, Laz10, PCK10, RJS^{+}10]. vectorial [DZQF13, Pas15a, Pas15b]. vectors [RJS^{+}10, YGK12]. Vegas [Cha19]. velocities [BDPP18]. Verifiable [QD16, XWLJ16, XWS17]. Verification [BD16, Che17, KPSZ11, MMM^{+}17, QX10, XZ17b]. Verifying [BD16, Gio18]. version [GSZ13, RJS^{+}10]. versus [DFdFT16, Vir11, WWS12, ZKXY10]. Vertex [BJH15, DKL^{+}12, KP14a, PCY16, ÁCLS12, AK14, BGJ^{+}16, BDF^{+}18, Che10c, CL15, FGG^{+}10, Fu10, GLLX12, GLXG13, GZM15, Ili13, Kat16, Kob15, Lam11, LX19, Lü19, MSM14, Mon15, NN17, OM17, Sch13b, SK12, Tsu19b, TZ11, TY13, Tu15, VK18, YeCM14, Yus11, ZZ14, ZLS^{+}17, CP13]. vertex-colouring [SK12]. vertex-disjoint [Che10c, ZZ14]. vertex-distinguishing [YeCM14]. vertex-pancyclicity [Fu10]. Vertex-transitivity [PCY16, Lü19]. vertices [Bor16, BM19, Che16b, CZWP17, DX11, Fie11, HWJ96, HAM11, Hua18, IM12]. very [LT18, YA13]. vF2 [GWL18]. via [AC12, CFMT16, CK11b, DLMV10, DGR15, DT10, DZ11, DMS11, Fil13, HYYZ15, Mor11, SC12, YLGC16, ZKXY10]. video [DWQ10]. videos [SML^{+}10]. view [ZMZ19]. vincular [Vaj18]. violations [FG18a, Wij10]. Virtex [eSKAI10]. Virtex-5 [eSKAI10]. virtual [HM18]. Visibility [ÁCLS12, AGZP15, Gil14, LM16]. Visibility-preserving [ÁCLS12]. visibly [Lan11]. visual [DWQ10, LWL10, YLZZ18a]. Viterbi [SS18]. Vizing [PPH18, Wu13]. VNP [MST16]. volatility [TWW17]. Voronoi [LT13c, MK16]. voting [BT16b, GE12, MS19]. VP [MST16]. vs [AB15, LYHC10, Man19].

walk [DT10, FP18]. Walks [CDD^{+}11, ABH19]. wandering [ADF13]. wants [MN15]. Warehousing [JK18]. Warshall [Hou10]. watchman [TJ18]. wavelength [Liu15, YYYZ12, YYZH13]. way [BN10a, Xia10]. Waypoint [CDM^{+}11]. Weak [Lag14, Par16, AAC^{+}10, Ezr10, RH10, ZS18]. Weakly [Pol18, SO10, Dur13]. Weakly-non-overlapping [SO10]. weakness [MGPI12]. weaknesses [Nos11, Nos14]. Web [Ars19, BLM10, SLL13]. Wee [Fis10]. Weight [BG12, BG15, ZMZ19, BCNPL14, LL14, Mos13a, Pet12]. Weight-based [ZMZ19]. Weighted [AMMY10, BOV15, AR18, AZ14, BJH15, BFLM15, BM19, CG15, EH18, FL16, FV13, GKPP16, GMM15, GE12, HKW14, KM16a, Kos18, Kun18, Li12, LZLY12, LL16, LMV17, LL10d, MS12, MM15a, Suk19, ZLZX19, ZpH15]. weighting [Ars19, DWQ10, PdAL18]. weightings [SK12]. weights [JAH16]. welfare [Lee17]. well [GS12]. well-connectedness [GS12]. wheel [KdER^{+}14]. Wheeler [MN19, DGG^{+}19, SKN11]. wheels [AMRR11]. whether [CKP16]. which [Tur12, Tys13]. Who [MN15]. Whole [BV10]. width [AAB^{+}19, BT16a, CB18, Cou14, DDER18, Feg19, MTT12]. wildcards [ZZH10b]. windmill [RRR12]. window [GM13a]. windows [BKZ15, GS10b, HLT10]. winner [BR12, Kal12]. Wirelength [AMRR11, RARM12]. wireless [CWW10, Cho12, HKW18, KK15b, LKC^{+}12, SKK10, VLV15, YA13, YWC11, YTYZ15, ZGLY12]. wiring [SW18]. wise [DZ11, EFMA10, ZL17]. within [RV17]. without [BGJ^{+}16, BOV15, WWYY11, ZW10, ZL12, ZY17]. witness [ADH14, Kar17]. Władyslaw [Tar14]. Wolfe [CFMT16]. word [DWQ10, GPR10, Han17, Hen16, KN13b, SH17, Wal10]. words [BC15, GSS16, HN10, KM10, KV16, Mac19, MV13, Par16, Pel15, WJS10, ZZH10b]. work [AR13a, SLS16, EFKR10]. work-optimal [AR13a]. Workflow [GW16]. works [LWXS18]. world [MST16]. worlds [GKS13]. wraparound [TNN11]. write [FG18a]. writers [Tho18].

XPath [KRV16].

Yao [BF19, EAE21]. yields [Ehl17]. Young [ASTD14].

Zen [HWA12]. zeolites [Jor10]. zero [DS15, DS16, Mos11a, SWF18, WWBC14, YL11a]. zero-correlation [WWBC14]. zerodifference [SWF18]. zero-one [Mos11a]. zero-sum [DS15, DS16]. zeta [CK11b]. Zhang [LXT19, SWW16, ZKXY10, ZQJ^{+}15]. ZNN [ZQJ^{+}15, ZDQ^{+}17]. Zwick [HP19].

References

Ashrafi:2016:NEP

[AAB^{+}16] Ali Reza Ashrafi, Jernej Azarija, Azam Babai, Khadijeh Fathalikhani, and Sandi Klavzar. The (non- )existence of perfect codes in Fibonacci cubes. Information Processing Letters, 116 (5):387–390, May 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019016000120.

Ahn:2019:MWA

[AAB^{+}19] Hee-Kap Ahn, Taehoon Ahn, Sang Won Bae, Jongmin Choi, Mincheol Kim, Eunjin Oh, Chan-Su Shin, and Sang Duk Yoon. Minimum-width annulus with outliers: Circular, square, and rectangular cases. Information Processing Letters, 145(??):16– 23, May 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300067.

Aleksandrowicz:2017:RHC

[AABB17] Gadi Aleksandrowicz, Andrei Asinowski, Gill Barequet, and Ronnie Barequet. Recovering highly-complex linear recurrences of integer sequences. Information Processing Letters, 127(??):62– 66, November 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901730128X.

Abu-Affash:2019:BBF

[AABCC19] A. Karim Abu-Affash, Sujoy Bhore, Paz Carmi, and Dibyayan Chakraborty. Bottleneck bichromatic full Steiner trees. Information Processing Letters, 142(??): 14–19, February 2019. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019018301881.

Amtoft:2010:ACW

[AAC^{+}10] Torben Amtoft, Kelly Androutsopoulos, David Clark, Mark Harman, and Zheng Li. An alternative characterization of weak order dependence. Information Processing Letters, 110(21): 939–943, October 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aichholzer:2018:CBI

[AADB^{+}18] Oswin Aichholzer, Nieves Atienza, José M. Díaz-Báñez, Ruy Fabila-Monroy, David Flores-Peñaloza, Pablo Pérez-Lantero, Birgit Vogtenhuber, and Jorge Urrutia. Computing balanced islands in two colored point sets in the plane. Information Processing Letters, 135(??): 28–32, July 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300371.

Aigner:2015:TAP

[AAJ15] Wolfgang Aigner, Franz Aurenhammer, and Bert Jüttler. On triangulation axes of polygons. Information Processing Letters, 115(1):45–51, January 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001689.

Algefari:2016:SDG

[AALL16] Mansour J. Algefari, Khalid A.

Alsatami, Hong-Jian Lai, and Juan Liu. Supereulerian digraphs with given local structures. Information Processing Letters, 116(5): 321–326, May 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015002318.

Akshay:2015:RPM

[AAOW15] S. Akshay, Timos Antonopoulos, Joël Ouaknine, and James Worrell. Reachability problems for Markov chains. Information Processing Letters, 115(2):155– 158, February 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014001781.

Alves:2015:SNT

[AB15] S. Alves and S. Broda. A short note on typeinhabitation: Formula-trees vs. game semantics. Information Processing Letters, 115(11):908–911, November 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000745.

Avin:2014:TIN

[ABH^{+}14] C. Avin, M. Borokhovich, Y. Haddad, E. Kantor, Z. Lotker, M. Parter, and D. Peleg. Testing the irreducibility of nonsquare Perron–Frobenius systems. Information Processing Letters, 114(12):728–733, December 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014000994.

Agostini:2019:MTB

[ABH19] Matteo Agostini, Marco Bressan, and Shahrzad Haddadan. Mixing time bounds for graphlet random walks. Information Processing Letters, 152(??):Article 105851, December 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019301346.

Apostolico:2011:SPS

[ABP11] Alberto Apostolico, Manuel Barbares, and Cinzia Pizzi. Speedup for a periodic subgraph miner. Information Processing Letters, 111(11): 521–523, May 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Agosta:2015:TBS

[ABPS15] Giovanni Agosta, Alessandro Barenghi, Gerardo Pelosi, and Michele Scandale. Trace-based schedulability analysis to enhance passive side-channel attack resilience of embedded software. Information Processing Letters, 115(2):292–297, February 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002130.

Awasthi:2012:CBC

[ABS12] Pranjal Awasthi, Avrim Blum, and Or Sheffet. Center-based clustering under perturbation stability. Information Processing Letters, 112(1–2):49–54, January 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011002778.

Al-Bawani:2013:BOM

[ABS13] Kamal Al-Bawani and Alexander Souza. Buffer overflow management with class segregation. Information Processing Letters, 113(4):145– 150, February 28, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019013000057.

Anand:2010:NFB

[AC10] S. Anand and R. Chandramouli. A network flow based approach for network selection in dynamic spectrum access networks. Information Processing Letters, 110(3):104–107, January 1, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aman:2012:PEM

[AC12] Bogdan Aman and Gabriel Ciobanu. Properties of enhanced mobile membranes via coloured Petri nets. Information Processing Letters, 112(6):243–248, March 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011003255.

Anastasiadis:2018:BTA

[ACD18] M. Anastasiadis, N. Chatzis, and K. A. Draziotis. Birthday type attacks to the Naccache–Stern knapsack cryptosystem. Information Processing Letters, 138(??): 35–38, October 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019018301273.

Afrati:2011:CTP

[ACK11] Foto N. Afrati, Sara Cohen, and Gabriel Kuper. On the complexity of tree pattern containment with arithmetic comparisons. Information Processing Letters, 111(15):754–760, August 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001207.

Azar:2018:AAG

[ACL18] Yossi Azar, Jaya Prakash Champati, and Ben Liang. 2-approximation algorithm for a generalization of scheduling on unrelated parallel machines. Information Processing Letters, 139(??): 35–38, November 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019018301522.

Abrego:2012:VPC

[ÁCLS12] B. M. Ábrego, M. Cetina, J. Leaños, and G. Salazar. Visibility-preserving convexifications using singlevertex moves. Information Processing Letters, 112(5):161–163, February 28, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011003164.

Atallah:2011:PMH

[AD11] Mikhail J. Atallah and Timothy W. Duket. Pattern matching in the Hamming distance with thresholds. Information Processing Letters, 111(14):674–677, July 31, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011001049.

Ahadi:2013:CPO

[AD13] A. Ahadi and A. Dehghan. The complexity of the proper orientation number. Information Processing Letters, 113 (19–21):799–803, September/October 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013002068.

Aggarwal:2016:IHR

[AD16] Divesh Aggarwal and Chandan Dubey. Improved hardness results for unique shortest vector problem. Information Processing Letters, 116(10):631–637, October 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300746.

Angelini:2018:CAL

[AD18] Patrizio Angelini and Giordano Da Lozzo. 3-coloring arrangements of line segments with 4 slopes is hard. Information Processing Letters, 137(??):47–50, September 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301054.

Adamczyk:2011:IAG

[Ada11] Marek Adamczyk. Improved analysis of the greedy algorithm for stochastic matching. Information Processing Letters, 111(15):731– 737, August 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901100127X.

Acerbi:2013:SMC

[ADF13] Luigi Acerbi, Alberto Dennunzio, and Enrico Formenti. Surjective multidimensional cellular automata are non-wandering: a combinatorial proof. Information Processing Letters, 113(5–6):156–159, March 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012003195.

Aceto:2011:CRS

[AdFEGRI11] Luca Aceto, David de Frutos Escrig, Carlos Gregorio-Rodríguez, and Anna Ingolfsdottir. Complete and ready simulation semantics are not finitely based over BCCSP, even with a singleton alphabet. Information Processing Letters, 111 (9):408–413, April 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ateniese:2013:NTB

[ADFM13] Giuseppe Ateniese, Alfredo De Santis, Anna Lisa Ferrara, and Barbara Masucci. A note on timebound hierarchical key assignment schemes. Information Processing Letters, 113(5–6):151–155, March 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000124.

Afrati:2010:QCU

[ADG10] Foto N. Afrati, Matthew Damigos, and Manolis Gergatsoulis. Query containment under bag and bagset semantics. Information Processing Letters, 110(10): 360–369, April 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aronov:2014:MWP

[ADH14] Boris Aronov, Muriel Dulieu, and Ferran Hurtado. Mutual witness proximity graphs. Information Processing Letters, 114(10): 519–523, October 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901400060X.

Ahadi:2012:CLN

[ADKM12] A. Ahadi, A. Dehghan, M. Kazemi, and E. Mollaahmadi. Computation of lucky number of planar graphs is NP-hard. Information Processing Letters, 112(4):109– 112, February 15, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011003048.

Alon:2010:NCD

[AFPT10] Noga Alon, Michal Feldman, Ariel D. Procaccia, and Moshe Tennenholtz. A note on competitive diffusion through social networks. Information Processing Letters, 110(6):221–225, February 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ashok:2015:SC

[AG15] Pradeesha Ashok and Sathish Govindarajan. On strong centerpoints. Information Processing Letters, 115(3): 431–434, March 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019014002373.

Abdel-Ghaffar:2019:SBS

[AG19] Khaled A. S. Abdel-Ghaffar. Sets of binary sequences with small total Hamming distances. Information Processing Letters, 142(??):27– 29, February 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901830190X.

Aldana-Galvan:2019:MSA

[AGÁRCS^{+}19] I. Aldana-Galván, J. L. Álvarez-Rebollar, J. C. Catana-Salazar, M. Jiménez-Salinas, E. Solís-Villarreal, and J. Urrutia. Minimizing the solid angle sum of orthogonal polyhedra. Information Processing Letters, 143(??):47–50, March 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901830231X.

Aggarwal:2015:AES

[Agg15] Divesh Aggarwal. Affineevasive sets modulo a prime. Information Processing Letters, 115(2):382–385, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002300.

Akshay:2012:RSR

[AGHY12] S. Akshay, Blaise Genest, Loïc Hélouët, and Shaofa Yang. Regular set of representatives for timeconstrained MSC graphs. Information Processing Letters, 112(14–15):592–598, August 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001202.

Aceto:2015:GCA

[AGI15] Luca Aceto, Eugen-Ioan Goriac, and Anna Ingolfsdottir. A ground-complete axiomatization of stateless bisimilarity over Linda. Information Processing Letters, 115(2):212–220, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014001902.

Atallah:2013:LVR

[AGW13] Mikhail J. Atallah, Elena Grigorescu, and Yi Wu. A lower-variance randomized algorithm for approximate string matching. Information Processing Letters, 113(18):690–692, September 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001622.

Alipour:2015:VTC

[AGZP15] Sharareh Alipour, Mohammad Ghodsi, Alireza Zarei, and Maryam Pourreza. Visibility testing and counting. Information Processing Letters, 115(9):649–654, September 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015000484.

Andres:2011:GCN

[AH11] Stephan Dominique Andres and Winfried Hochstättler. The game chromatic number and the game colouring number of classes of oriented cactuses. Information Processing Letters, 111(5): 222–226, February 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Allem:2017:PTF

[AH17] Luiz Emilio Allem and Carlos Hoppen. A pre-test for factoring bivariate polynomials with coefficients in F_{2}. Information Processing Letters, 121(??):22–28, May 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300157.

Aichholzer:2017:PPS

[AHK^{+}17] Oswin Aichholzer, Thomas Hackl, Matias Korman, Marc van Kreveld, Maarten Löffler, Alexander Pilz, Bettina Speckmann, and Emo Welzl. Packing plane spanning trees and paths in complete geometric graphs. Information Processing Letters, 124(??):35–41, August 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300637.

Avin:2018:RTR

[AHLS18] Chen Avin, Alexandr Hercules, Andreas Loukas, and Stefan Schmid. rDAN: Toward robust demand-aware network designs. Information Processing Letters, 133 (??):5–9, May 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019017302223.

Abbasi:2010:IRT

[AHR10] Sarmad Abbasi, Patrick Healy, and Aimal Rextin. Improving the running time of embedded upward planarity testing. Information Processing Letters, 110 (7):274–278, March 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Amir:2017:TSH

[AIR17] Amihood Amir, Costas S. Iliopoulos, and Jakub Radoszewski. Two strings at Hamming distance 1 cannot be both quasiperiodic. Information Processing Letters, 128(??):54–57, December 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017301473.

Aceto:2010:RBG

[AIS10] Luca Aceto, Anna Ingolfsdottir, and Joshua Sack. Resource bisimilarity and graded bisimilarity coincide. Information Processing Letters, 111(2):68–76, December 31, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arevalo:2011:CEC

[AJLM11] Sergio Arévalo, Ernesto Jiménez, Mikel Larrea, and Luis Mengual. Communicationefficient and crash-quiescent Omega with unknown membership. Information Processing Letters, 111(4):194– 199, January 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Abu-Khzam:2010:IKA

[AK10] Faisal N. Abu-Khzam. An improved kernelization algorithm for r-Set Packing. Information Processing Letters, 110(16):621–624, July 31, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Avni:2013:WDA

[AK13] Guy Avni and Orna Kupferman. When does abstraction help? Information Processing Letters, 113 (22–24):901–905, November/December 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013002445.

Abu-Khzam:2014:MCI

[AK14] Faisal N. Abu-Khzam. Maximum common induced subgraph parameterized by vertex cover. Information Processing Letters, 114(3):99– 103, March 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002755.

Akavia:2014:ESS

[Aka14] Adi Akavia. Explicit small sets with ϵ-discrepancy on Bohr sets. Information Processing Letters, 114(10): 564–567, October 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014000970.

Abu-Khzam:2016:EMD

[AKH16] Faisal N. Abu-Khzam and Pinar Heggernes. Enumerating minimal dominating sets in chordal graphs. Information Processing Letters, 116(12):739–743, December 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301004.

Amiri:2019:RCA

[AKMR19] Saeed Akhoondian Amiri, Stephan Kreutzer, Dániel Marx, and Roman Rabinovich. Routing with congestion in acyclic digraphs. Information Processing Letters, 151(??):Article 105836, November 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019301073.

Akutsu:2010:BAG

[Aku10] Tatsuya Akutsu. A bisection algorithm for grammarbased compression of ordered trees. Information Processing Letters, 110(18– 19):815–820, September 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

AlTawy:2013:SOC

[AKY13] Riham AlTawy, Aleksandar Kircanski, and Amr Youssef. Second order collision for the 42-step reduced DHA-256 hash function. Information Processing Letters, 113 (19–21):764–770, September/October 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002044.

Andre:2018:LPP

[AL18] Étienne André and Shang-Wei Lin. The language preservation problem is undecidable for parametric event-recording automata. Information Processing Letters, 136(??):17–20, August 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300681.

Alevizos:2013:OAC

[Ale13] Panagiotis D. Alevizos. An optimal algorithm for computing the non-trivial circuits of a union of isooriented rectangles. Information Processing Letters, 113(3):55–59, February 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012002992.

Argiroffo:2015:CDT

[ALT15] Gabriela Argiroffo, Valeria Leoni, and Pablo Torres. On the complexity of {k}-domination and k-tuple domination in graphs. Information Processing Letters, 115(6–8):556–561, June/ August 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000162.

Amani:2016:ARC

[ALT16] Mahdi Amani, Kevin A.

Lai, and Robert E. Tarjan. Amortized rotation cost in AVL trees. Information Processing Letters, 116(5):327– 330, May 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901500232X.

Argiroffo:2018:CTT

[ALT18] G. Argiroffo, V. Leoni, and P. Torres. Complexity of k-tuple total and total kdominations for some subclasses of bipartite graphs. Information Processing Letters, 138(??):72–74, October 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301327.

Amano:2010:CHC

[Ama10] Shun’ichi Amano. On the coNP hardness of computing certain answers over locally specified incomplete DOM-trees. Information Processing Letters, 110(17): 753–756, August 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Amano:2011:MTF

[Ama11] Kazuyuki Amano. Mintermtransitive functions with asymptotically smallest block sensitivity. Information Processing Letters, 111(23– 24):1081–1084, December 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011002493.

Asahiro:2010:WNN

[AMMY10] Yuichi Asahiro, Eiji Miyano, Shuichi Miyazaki, and Takuro Yoshimuta. Weighted nearest neighbor algorithms for the graph exploration problem on cycles. Information Processing Letters, 110 (3):93–98, January 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Andrade:2010:CBA

[AMN^{+}10] Ermeson Andrade, Paulo Maciel, Bruno Nogueira, Carlos Araújo, and Gustavo Callou. A COTS-based approach for estimating performance and energy consumption of embedded realtime systems. Information Processing Letters, 110(14– 15):525–534, July 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Akrami:2019:RBM

[AMO19] Hannaneh Akrami, Kurt Mehlhorn, and Tommy Odland. Ratio-balanced maximum flows. Information Processing Letters, 150(??): 13–17, October 2019. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0020019019301024.

Avitabile:2013:OCO

[AMP13] T. Avitabile, C. Mathieu, and L. Parkinson. Online constrained optimization with recourse. Information Processing Letters, 113(3):81–86, February 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002980.

Arockiaraj:2011:WFH

[AMRR11] M. Arockiaraj, Paul Manuel, Indra Rajasingh, and Bharati Rajan. Wirelength of 1fault Hamiltonian graphs into wheels and fans. Information Processing Letters, 111(18):921–925, September 30, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001827.

Akutsu:2012:SPA

[AMT12] Tatsuya Akutsu, Avraham A. Melkman, and Takeyuki Tamura. Singleton and 2-periodic attractors of sign-definite Boolean networks. Information Processing Letters, 112(1–2):35–38, January 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011002675.

Angelini:2017:MDG

[Ang17] Patrizio Angelini. Monotone drawings of graphs with few directions. Information Processing Letters, 120(??):16– 22, April 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016301855.

Anonymous:2010:EBa

[Ano10a]  Anonymous.  Editorial

Board. Information Processing Letters, 110(3):??, January 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2010:EBb

[Ano10b]  Anonymous.  Editorial

Board. Information Processing Letters, 110(4):??, January 16, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2010:EBc

[Ano10c]  Anonymous.  Editorial

Board. Information Processing Letters, 110(5):??, February 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2010:EBd

[Ano10d]  Anonymous.  Editorial

Board. Information Processing Letters, 110(6):??, February 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2010:EBe

[Ano10e]  Anonymous.  Editorial

Board. Information Processing Letters, 110(7):??, March 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2010:EBf

[Ano10f]  Anonymous.  Editorial

Board. Information Processing Letters, 110(8–9):??, April 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2010:EBg

[Ano10g]  Anonymous.  Editorial

Board. Information Processing Letters, 110(10):??, April 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2010:EBh

[Ano10h]  Anonymous.  Editorial

Board. Information Processing Letters, 110(11):??, May 16, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2010:EBi

[Ano10i]  Anonymous.  Editorial

Board. Information Processing Letters, 110(12–13): ??, June 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2010:EBj

[Ano10j]  Anonymous.  Editorial

Board. Information Processing Letters, 110(14–15): ??, July 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2010:EBk

[Ano10k]  Anonymous.  Editorial

Board. Information Processing Letters, 110(16):??, July 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2010:EBl

[Ano10l]  Anonymous.  Editorial

Board. Information Processing Letters, 110(17):??, August 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2010:EBm

[Ano10m]  Anonymous.  Editorial

Board. Information Processing Letters, 110(18–19): ??, September 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2010:EBn

[Ano10n]  Anonymous.  Editorial

Board. Information Processing Letters, 110(20): ??, September 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2010:EBo

[Ano10o]  Anonymous.  Editorial

Board. Information Processing Letters, 110(21):??, October 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2010:EBp

[Ano10p]  Anonymous.  Editorial

Board. Information Processing Letters, 110(22):??, October 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2010:EBq

[Ano10q]  Anonymous.  Editorial

Board. Information Processing Letters, 110(23): ??, November 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2010:EBr

[Ano10r]  Anonymous.  Editorial

Board. Information Processing Letters, 111(1):??, December 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2010:EBs

[Ano10s]  Anonymous.  Editorial

Board. Information Processing Letters, 111(2):??, December 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2011:EBa

[Ano11a]  Anonymous.  Editorial

Board. Information Processing Letters, 111(3):??, January 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2011:EBb

[Ano11b]  Anonymous.  Editorial

Board. Information Processing Letters, 111(4):??, January 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2011:EBc

[Ano11c]  Anonymous.  Editorial

Board. Information Processing Letters, 111(5):??, February 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2011:EBd

[Ano11d]  Anonymous.  Editorial

Board. Information Processing Letters, 111(6):??, February 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2011:EBe

[Ano11e]  Anonymous.  Editorial

Board. Information Processing Letters, 111(7):??, March 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2011:EBf

[Ano11f]  Anonymous.  Editorial

Board. Information Processing Letters, 111(8):??, March 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2011:EBg

[Ano11g]  Anonymous.  Editorial

Board. Information Processing Letters, 111(9):??, April 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2011:EBh

[Ano11h]  Anonymous.  Editorial

Board. Information Processing Letters, 111(10):??, April 30, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2011:EBi

[Ano11i]  Anonymous.  Editorial

Board. Information Processing Letters, 111(12):ifc, June 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011000950.

Anonymous:2011:EBj

[Ano11j]  Anonymous.  Editorial

Board. Information Processing Letters, 111(13):??, July 1, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001098.

Anonymous:2011:EBk

[Ano11k]  Anonymous.  Editorial

Board. Information Processing Letters, 111(14):??, July 31, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001311.

Anonymous:2011:EBl

[Ano11l]  Anonymous.  Editorial

Board. Information Processing Letters, 111(15):??, August 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001463.

Anonymous:2011:EBm

[Ano11m]  Anonymous.  Editorial

Board. Information Processing Letters, 111(16):??, August 30, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001608.

Anonymous:2011:EBn

[Ano11n]  Anonymous.  Editorial

Board. Information Processing Letters, 111(17):??, September 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901100175X.

Anonymous:2011:EBo

[Ano11o]  Anonymous.  Editorial

Board. Information Processing Letters, 111(18):??, September 30, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011001931.

Anonymous:2011:EBp

[Ano11p]  Anonymous.  Editorial

Board. Information Processing Letters, 111(19):??, October 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011002122.

Anonymous:2011:EBq

[Ano11q]  Anonymous.  Editorial

Board. Information Processing Letters, 111(20):??, October 31, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011002274.

Anonymous:2011:EBr

[Ano11r]  Anonymous.  Editorial

Board. Information Processing Letters, 111(21–22):??, November 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011002523.

Anonymous:2011:EBs

[Ano11s]  Anonymous.  Editorial

Board. Information Processing Letters, 111(23–24):??, December 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011002699.

Anonymous:2012:EBa

[Ano12a]  Anonymous.  Editorial

Board. Information Processing Letters, 112(1–2):??, January 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011002894.

Anonymous:2012:EBb

[Ano12b]  Anonymous.  Editorial

Board. Information Processing Letters, 112(3):??, January 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011003292.

Anonymous:2012:EBc

[Ano12c]  Anonymous.  Editorial

Board. Information Processing Letters, 112(4):??, February 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011003401.

Anonymous:2012:EBd

[Ano12d]  Anonymous.  Editorial

Board. Information Processing Letters, 112(5):??, February 28, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012000038.

Anonymous:2012:EBe

[Ano12e]  Anonymous.  Editorial

Board. Information Processing Letters, 112(6):??, March 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012000129.

Anonymous:2012:EBf

[Ano12f]  Anonymous.  Editorial

Board. Information Processing Letters, 112(7):??, March 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012000294.

Anonymous:2012:EBg

[Ano12g]  Anonymous.  Editorial

Board. Information Processing Letters, 112(8–9):??, April 30, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012000476.

Anonymous:2012:EBh

[Ano12h]  Anonymous.  Editorial

Board. Information Processing Letters, 112(10):??, May 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000646.

Anonymous:2012:EBi

[Ano12i]  Anonymous.  Editorial

Board. Information Processing Letters, 112(11):??, June 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000877.

Anonymous:2012:EBj

[Ano12j]  Anonymous.  Editorial

Board. Information Processing Letters, 112(12):??, June 30, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000981.

Anonymous:2012:EBk

[Ano12k]  Anonymous.  Editorial

Board. Information Processing Letters, 112(13):??, July 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001135.

Anonymous:2012:EBl

[Ano12l]  Anonymous.  Editorial

Board. Information Processing Letters, 112(14–15):??, August 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012001329.

Anonymous:2012:EBm

[Ano12m]  Anonymous.  Editorial

Board. Information Processing Letters, 112(16):??, August 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012001469.

Anonymous:2012:EBn

[Ano12n]  Anonymous.  Editorial

Board. Information Processing Letters, 112(17–18):??, September 30, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019012001743.

Anonymous:2012:EBo

[Ano12o]  Anonymous.  Editorial

Board. Information Processing Letters, 112(19):??, October 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012001834.

Anonymous:2012:EBp

[Ano12p]  Anonymous.  Editorial

Board. Information Processing Letters, 112(20):??, October 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002049.

Anonymous:2012:EBq

[Ano12q]  Anonymous.  Editorial

Board. Information Processing Letters, 112(21):??, November 15, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019012002219.

Anonymous:2012:EBr

[Ano12r]  Anonymous.  Editorial

Board. Information Processing Letters, 112(23):??, December 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002530.

Anonymous:2012:EBs

[Ano12s]  Anonymous.  Editorial

Board. Information Processing Letters, 112(24):??, December 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002724.

Anonymous:2013:EBa

[Ano13a]  Anonymous.  Editorial

Board. Information Processing Letters, 113(1–2): ??, January 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012002852.

Anonymous:2013:EBb

[Ano13b]  Anonymous.  Editorial

Board. Information Processing Letters, 113(3):??, February 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012003122.

Anonymous:2013:EBc

[Ano13c]  Anonymous.  Editorial

Board. Information Processing Letters, 113(4):??, February 28, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013000197.

Anonymous:2013:EBd

[Ano13d]  Anonymous.  Editorial

Board. Information Processing Letters, 113(5–6):??, March 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013000318.

Anonymous:2013:EBe

[Ano13e]  Anonymous.  Editorial

Board. Information Processing Letters, 113(7):??, April 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000410.

Anonymous:2013:EBf

[Ano13f]  Anonymous.  Editorial

Board. Information Processing Letters, 113(8):??, April 30, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000586.

Anonymous:2013:EBg

[Ano13g]  Anonymous.  Editorial

Board. Information Processing Letters, 113(9):??, May 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000756.

Anonymous:2013:EBh

[Ano13h]  Anonymous.  Editorial

Board. Information Processing Letters, 113(10–11):??, May/June 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901300094X.

Anonymous:2013:EBi

[Ano13i]  Anonymous.  Editorial

Board. Information Processing Letters, 113(12):??, June 30, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001063.

Anonymous:2013:EBj

[Ano13j]  Anonymous.  Editorial

Board. Information Processing Letters, 113(13):??, July 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001245.

Anonymous:2013:EBk

[Ano13k]  Anonymous.  Editorial

Board. Information Processing Letters, 113(14–16):??, July/August 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001440.

Anonymous:2013:EBl

[Ano13l]  Anonymous.  Editorial

Board. Information Processing Letters, 113(17):??, August 30, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001646.

Anonymous:2013:EBm

[Ano13m]  Anonymous.  Editorial

Board. Information Processing Letters, 113(18):??, September 15, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013001841.

Anonymous:2013:EBn

[Ano13n]  Anonymous.  Editorial

Board. Information Processing Letters, 113(19–21):??, September/October 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019013002275.

Anonymous:2014:EBa

[Ano14a]  Anonymous.  Editorial

Board. Information Processing Letters, 114(10):ifc, October 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901400115X.

Anonymous:2014:EBb

[Ano14b]  Anonymous.  Editorial

Board. Information Processing Letters, 114(11):IFC, November 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014001343.

Anonymous:2014:EBc

[Ano14c]  Anonymous.  Editorial

Board. Information Processing Letters, 114(12):IFC, December 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014001574.

Anonymous:2015:EBa

[Ano15a]  Anonymous.  Editorial

Board. Information Processing Letters, 115(1):IFC, January 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014002063.

Anonymous:2015:EBb

[Ano15b]  Anonymous.  Editorial

Board. Information Processing Letters, 115(2):IFC, February 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014002397.

Anonymous:2015:EBc

[Ano15c]  Anonymous.  Editorial

Board. Information Processing Letters, 115(3):IFC, March 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002695.

Anonymous:2015:EBd

[Ano15d]  Anonymous.  Editorial

Board. Information Processing Letters, 115(4):IFC, April 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000034.

Anonymous:2015:EBe

[Ano15e]  Anonymous.  Editorial

Board. Information Processing Letters, 115(5):IFC, May 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000265.

Anonymous:2015:EBf

[Ano15f]  Anonymous.  Editorial

Board. Information Processing Letters, 115(6–8):ifc, June/August 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000514.

Anonymous:2015:EBg

[Ano15g]  Anonymous.  Editorial

Board. Information Processing Letters, 115(9):ifc, September 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015000782.

Anonymous:2015:EBh

[Ano15h]  Anonymous.  Editorial

Board. Information Processing Letters, 115(10):ifc, October 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015000939.

Anonymous:2015:EBi

[Ano15i]  Anonymous.  Editorial

Board. Information Processing Letters, 115(11):ifc, November 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001192.

Anonymous:2015:EBj

[Ano15j]  Anonymous.  Editorial

Board. Information Processing Letters, 115(12):ifc, December 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001441.

Anonymous:2016:EBa

[Ano16a]  Anonymous.  Editorial

Board. Information Processing Letters, 116(1):ifc, January 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001799.

Anonymous:2016:EBb

[Ano16b]  Anonymous.  Editorial

Board. Information Processing Letters, 116(2):ifc, February 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001969.

Anonymous:2016:EBc

[Ano16c]  Anonymous.  Editorial

Board. Information Processing Letters, 116(3):ifc, March 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015002173.

Anonymous:2016:EBd

[Ano16d]  Anonymous.  Editorial

Board. Information Processing Letters, 116(4):ifc, April 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015002264.

Anonymous:2016:EBe

[Ano16e]  Anonymous.  Editorial

Board. Information Processing Letters, 116(5):ifc, May 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016300023.

Anonymous:2016:EBf

[Ano16f]  Anonymous.  Editorial

Board. Information Processing Letters, 116(6):ifc, June 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016300175.

Anonymous:2016:EBg

[Ano16g]  Anonymous.  Editorial

Board. Information Processing Letters, 116(7):ifc, July 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016300370.

Anonymous:2016:EBh

[Ano16h]  Anonymous.  Editorial

Board. Information Processing Letters, 116(8):ifc, August 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016300503.

Anonymous:2016:EBi

[Ano16i]  Anonymous.  Editorial

Board. Information Processing Letters, 116(9):ifc, September 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300667.

Anonymous:2016:EBj

[Ano16j]  Anonymous.  Editorial

Board. Information Processing Letters, 116(10):ifc, October 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300849.

Anonymous:2016:EBk

[Ano16k]  Anonymous.  Editorial

Board. Information Processing Letters, 116(11):ifc, November 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901630103X.

Anonymous:2016:EBl

[Ano16l]  Anonymous.  Editorial

Board. Information Processing Letters, 116(12):ifc, December 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301144.

Anonymous:2017:EBa

[Ano17a]  Anonymous.  Editorial

Board. Information Processing Letters, 117(??):ifc, January 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016301223.

Anonymous:2017:EBb

[Ano17b]  Anonymous.  Editorial

Board. Information Processing Letters, 118(??):ifc, February 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016301557.

Anonymous:2017:EBc

[Ano17c]  Anonymous.  Editorial

Board. Information Processing Letters, 119(??):ifc, March 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016301764.

Anonymous:2017:EBd

[Ano17d]  Anonymous.  Editorial

Board. Information Processing Letters, 120(??):ifc, April 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017300066.

Anonymous:2017:EBe

[Ano17e]  Anonymous.  Editorial

Board. Information Processing Letters, 121(??):ifc, May 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300248.

Anonymous:2017:EBf

[Ano17f]  Anonymous.  Editorial

Board. Information Processing Letters, 122(??):ifc, June 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300418.

Anonymous:2017:EBg

[Ano17g]  Anonymous.  Editorial

Board. Information Processing Letters, 123(??):ifc, July 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300558.

Anonymous:2017:EBh

[Ano17h]  Anonymous.  Editorial

Board. Information Processing Letters, 124(??):ifc, August 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300820.

Anonymous:2017:EBi

[Ano17i]  Anonymous.  Editorial

Board. Information Processing Letters, 125(??):ifc, September 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017300947.

Anonymous:2017:EBj

[Ano17j]  Anonymous.  Editorial

Board. Information Processing Letters, 126(??):ifc, October 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017301126.

Anonymous:2017:EBk

[Ano17k]  Anonymous.  Editorial

Board. Information Processing Letters, 127(??):ifc, November 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017301369.

Anonymous:2017:EBl

[Ano17l]  Anonymous.  Editorial

Board. Information Processing Letters, 128(??):ifc, December 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017301540.

Anonymous:2018:EBa

[Ano18a]  Anonymous.  Editorial

Board. Information Processing Letters, 129(??):ifc, January 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301734.

Anonymous:2018:EBb

[Ano18b]  Anonymous.  Editorial

Board. Information Processing Letters, 130(??):ifc, February 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301904.

Anonymous:2018:EBc

[Ano18c]  Anonymous.  Editorial

Board. Information Processing Letters, 131(??):ii, March 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017302107.

Anonymous:2018:EBd

[Ano18d]  Anonymous.  Editorial

Board. Information Processing Letters, 132(??):ii, April 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300073.

Anonymous:2018:EBe

[Ano18e]  Anonymous.  Editorial

Board. Information Processing Letters, 133(??):ii, May 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901830022X.

Anonymous:2018:EBf

[Ano18f]  Anonymous.  Editorial

Board. Information Processing Letters, 134(??):ii, June 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901830053X.

Anonymous:2018:EBg

[Ano18g]  Anonymous.  Editorial

board. Information Processing Letters, 135(??):ii, July 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300723.

Anonymous:2018:EBh

[Ano18h]  Anonymous.  Editorial

board. Information Processing Letters, 136(??):ii, August 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018301029.

Anonymous:2018:EBi

[Ano18i]  Anonymous.  Editorial

board. Information Processing Letters, 137(??):ii, September 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301194.

Anonymous:2018:EBj

[Ano18j]  Anonymous.  Editorial

Board. Information Processing Letters, 138(??):1– 88, October 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301431.

Anonymous:2018:EBk

[Ano18k]  Anonymous.  Editorial

Board. Information Processing Letters, 140(??):1–48, December 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301741.

Anonymous:2019:D

[Ano19a]  Anonymous.  December

2019. Information Processing Letters, 152(??):??, December 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2019:EBa

[Ano19b]  Anonymous.  Editorial

Board. Information Processing Letters, 141(??):ii, January 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301959.

Anonymous:2019:EBb

[Ano19c]  Anonymous.  Editorial

Board. Information Processing Letters, 142(??):ii, February 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018302199.

Anonymous:2019:EBc

[Ano19d]  Anonymous.  Editorial

Board. Information Processing Letters, 143(??):ii, March 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018302448.

Anonymous:2019:EBd

[Ano19e]  Anonymous.  Editorial

Board. Information Processing Letters, 144(??):ii, April 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300122.

Anonymous:2019:EBe

[Ano19f]  Anonymous.  Editorial

Board. Information Processing Letters, 145(??):ii, May 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300377.

Anonymous:2019:EBf

[Ano19g]  Anonymous.  Editorial

Board. Information Processing Letters, 146(??):ii, June 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300614.

Anonymous:2019:EBg

[Ano19h]  Anonymous.  Editorial

Board. Information Processing Letters, 147(??):ii, July 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300766.

Anonymous:2019:EBh

[Ano19i]  Anonymous.  Editorial

Board. Information Processing Letters, 148(??):ii, August 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300869.

Anonymous:2019:EBi

[Ano19j]  Anonymous.  Editorial

Board. Information Processing Letters, 149(??):ii, September 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019300973.

Anonymous:2019:EBj

[Ano19k]  Anonymous.  Editorial

Board. Information Processing Letters, 150(??):ii, October 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019301176.

Anonymous:2019:EBk

[Ano19l]  Anonymous.  Editorial

Board. Information Processing Letters, 151(??): Article 105845, November 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019301280.

Anonymous:2019:EBl

[Ano19m]  Anonymous.  Editorial

Board. Information Processing Letters, 152(??): Article 105855, December 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019301383.

Anonymous:2019:N

[Ano19n]  Anonymous.  November

2019. Information Processing Letters, 151(??):??, November 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2019:PO

[Ano19o] Anonymous. Pages 1–22 (October 2019). Information Processing Letters, 150 (??):1–22, October 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aumasson:2011:CHF

[AP11] Jean-Philippe Aumasson and Raphael C.-W. Phan. On the cryptanalysis of the hash function Fugue: Partitioning and inside-out distinguishers. Information Processing Letters, 111(11): 512–515, May 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aurenhammer:2014:SDT

[AP14] Franz Aurenhammer and Günter Paulini. On shape Delaunay tessellations. Information Processing Letters, 114(10):535–541, October 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014000660.

Amir:2013:HCS

[APR13] Amihood Amir, Haim Paryenty, and Liam Roditty. On the hardness of the Consensus String problem. Information Processing Letters, 113(10–11):371–374, May/ June 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000707.

Alam:2013:DCA

[AR13a] Muhammad Rashed Alam and M. Sohel Rahman. A divide and conquer approach and a work-optimal parallel algorithm for the LIS problem. Information Processing Letters, 113(13):470– 476, July 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000902.

Alonso:2013:ABM

[AR13b] Laurent Alonso and Edward M. Reingold. Analysis of Boyer and Moore’s MJRTY algorithm. Information Processing Letters, 113(13):495–497, July 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001166.

Abrishami:2018:PTA

[AR18] G. Abrishami and F. Rahbarnia. Polynomial-time algorithm for weighted efficient domination problem on diameter three planar graphs. Information Processing Letters, 140(??):18– 24, December 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019018301595.

Aravind:2010:HFB

[Ara10] Alex A. Aravind. Highlyfair bakery algorithm using symmetric tokens. Information Processing Letters, 110(23):1055–1060, November 15, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alves:2015:CFT

[ARdSP15] Péricles R. O. Alves, Raphael E.

Rodrigues, Rafael Martins de Sousa, and Fernando Magno Quintão Pereira. A case for a fast trip count predictor. Information Processing Letters, 115(2):146–150, February 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014001707.

Arslan:2018:FAA

[Ars18] Abdullah N. Arslan. A fast algorithm for all-pairs Hamming distances. Information Processing Letters, 139(??): 44–48, November 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019018301534.

Arslan:2019:HST

[Ars19]  Ahmet Arslan.  How

sensitive are the termweighting models of information retrieval to spam Web pages? Information Processing Letters, 144(??): 16–24, April 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018302357.

Alda:2018:LBR

[AS18] Francesco Aldà and Hans Ulrich Simon. A lower bound on the release of differentially private integer partitions. Information Processing Letters, 129(??):1– 4, January 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017301503.

Ahmadian:2013:DAR

[ASA13] Zahra Ahmadian, Mahmoud Salmasizadeh, and Mohammad Reza Aref. Desynchronization attack on RAPP ultralightweight authentication protocol. Information Processing Letters, 113(7):205–209, April 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000094.

Al-Ssulami:2017:FSM

[ASM17] Abdulrakeeb M. Al-Ssulami and Hassan Mathkour. Faster string matching based on hashing and bitparallelism. Information Processing Letters, 123(??): 51–55, July 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300492.

Abdessaied:2014:UBR

[ASTD14] Nabila Abdessaied, Mathias Soeken, Michael Kirkedal Thomsen, and Rolf Drechsler. Upper bounds for reversible circuits based on Young subgroups. Information Processing Letters, 114 (6):282–286, June 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019014000040.

Avis:2015:GEC

[AT15] David Avis and Hans Raj Tiwary. A generalization of extension complexity that captures P. Information Processing Letters, 115(6–8):588–593, June/ August 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015000228.

Agrawal:2018:LRC

[AT18] Neha Agrawal and Shashikala Tapaswi. Low rate cloud DDoS attack defense method based on power spectral density analysis. Information Processing Letters, 138(??):39–43, October 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018301261.

Attie:2017:FSC

[Att17]  Paul C. Attie.  Finite-

state concurrent programs can be expressed succinctly in triple normal form. Information Processing Letters, 123(??):8– 13, July 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300315.

Alon:2012:LCJ

[AW12] Noga Alon and Amit Weinstein. Local correction of juntas. Information Processing Letters, 112(6):223–226, March 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011003279.

An:2018:SRI

[AX18] Mingqiang An and Liming Xiong. Some results on the inverse sum indeg index of a graph. Information Processing Letters, 134(??):42– 46, June 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300358.

Ambainis:2012:SEQ

[AY12a] Andris Ambainis and Abuzer Yakaryilmaz. Superiority of exact quantum automata for promise problems. Information Processing Letters, 112(7):289–291, March 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000087.

Aspnes:2012:RLB

[AY12b] James Aspnes and Yitong Yin. Randomized load balancing by joining and splitting bins. Information Processing Letters, 112(8–9):309–313, April 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012000208.

AlTawy:2014:IDR

[AY14] Riham AlTawy and Amr M. Youssef. Integral distinguishers for reduced-round Stribog. Information Processing Letters, 114(8):426– 431, August 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014000428.

Andrews:2014:RAW

[AZ14] Matthew Andrews and Lisa Zhang. Rate-adaptive weighted fair queueing for energy-aware scheduling. Information Processing Letters, 114(5):247–251, May 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013003128.

Baba:2017:AFB

[Bab17] Kensuke Baba. An acceleration of FFT-based algorithms for the match-count problem. Information Processing Letters, 125(??):1–4, September 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300728.

Baer:2010:ACE

[Bae10] Michael B. Baer. Alphabetic coding with exponential costs. Information Processing Letters, 110(4): 139–142, January 16, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baek:2013:SRR

[Bae13] Yoo-Jin Baek. Scalar recoding and regular 2^{w}-ary right-to-left EC scalar multiplication algorithm. Information Processing Letters, 113(10–11):357–360, May/ June 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000471.

Barrus:2010:ALC

[Bar10] Michael D. Barrus. Antimagic labeling and canonical decomposition of graphs. Information Processing Letters, 110(7):261–263, March 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barrus:2012:HHR

[Bar12] Michael D. Barrus. Havel– Hakimi residues of unigraphs. Information Processing Letters, 112(1–2): 44–48, January 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019011002821.

Baril:2013:SPB

[Bar13] Jean-Luc Baril. Statisticspreserving bijections between classical and cyclic permutations. Information Processing Letters, 113 (1–2):17–22, January 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019012002839.

Bajic:2011:CIS

[BB11] Dragana Bajic and Alister Burr. Comments on “Integer SEC–DED codes for low power communications” [Inform. Process. Lett. 110 (2010) 518–520]. Information Processing Letters, 111(9):414–415, April 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [RV10].

Barany:2012:FSG

[BB12] Vince Bárány and Mikolaj Bojańczyk. Finite satisfiability for guarded fixpoint logic. Information Processing Letters, 112(10):371– 375, May 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000543.

Barguil:2015:SIS

[BB15] Joao M. M. Barguil and Paulo S. L. M. Barreto. Security issues in Sarkar’s e-cash protocol. Information Processing Letters, 115(11):801–803, November 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001039.

Banerjee:2017:RSP

[BBB^{+}17] Sandip Banerjee, Bhargab B. Bhattacharya, Binay Bhattacharya, Arindam Biswas, Sandip Das, Ritankar Mandal, and Sasanka Roy. On representing a simple polygon perceivable to a blind person. Information Processing Letters, 120(??): 1–5, April 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301703.

Binucci:2012:DTS

[BBD^{+}12] Carla Binucci, Ulrik Brandes, Giuseppe Di Battista, Walter Didimo, Marco Gaertler, Pietro Palladino, Maurizio Patrignani, Antonios Symvonis, and Katharina Zweig. Drawing trees in a streaming model. Information Processing Letters, 112(11):418–422, June 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012000609.

Blin:2012:PCR

[BBDS12] Guillaume Blin, Paola Bonizzoni, Riccardo Dondi, and Florian Sikora. On the parameterized complexity of the repetition free longest common subsequence problem. Information Processing Letters, 112(7):272–276, March 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011003371.

Berczi:2017:DHH

[BBK17] Kristóf Bérczi and Erika R. Bérczi-Kovács. Directed hypergraphs and Horn minimization. Information Processing Letters, 128(??):32– 37, December 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019017301424.

Baste:2017:PCE

[BBKS17] Julien Baste, Fairouz Beggas, Hamamache Kheddouci, and Ignasi Sau. On the parameterized complexity of the Edge Monitoring problem. Information Processing Letters, 121(??): 39–44, May 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300145.

Biletskiy:2018:BBD

[BBR^{+}18] Yevgen Biletskiy, J. Anthony Brown, Girish R. Ranganathan, Ebrahim Bagheri, and Ismail Akbari. Building a business domain metaontology for information pre-processing. Information Processing Letters, 138(??): 75–80, October 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019018301340.

Bollig:2013:PFA

[BC13] Beate Bollig and Michael Capelle. Priority functions for the approximation of the metric TSP. Information Processing Letters, 113(14–16):584–591, July/ August 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001361.

Badkobeh:2015:IBW

[BC15] Golnaz Badkobeh and Maxime Crochemore. Infinite binary words containing repetitions of odd period. Information Processing Letters, 115(5): 543–547, May 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019015000137.

Bliznets:2015:KLB

[BCKM15] Ivan Bliznets, Marek Cygan, Pawel Komosa, and Lukás Mach. Kernelization lower bound for Permutation Pattern Matching. Information Processing Letters, 115(5): 527–531, May 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019015000101.

Barbay:2014:MWP

[BCNPL14] Jérémy Barbay, Timothy M. Chan, Gonzalo Navarro, and Pablo Pérez-Lantero. Maximum-weight planar boxes in O(n^{2}) time (and better). Information Processing Letters, 114(8):437– 445, August 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014000507.

Berard:2015:POM

[BCS15] Béatrice Bérard, Krishnendu Chatterjee, and Nathalie Sznajder. Probabilistic opacity for Markov decision processes. Information Processing Letters, 115(1):52–59, January 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001756.

Bocker:2011:EFP

[BD11] Sebastian Böcker and Peter Damaschke. Even faster parameterized cluster deletion and cluster editing. Information Processing Letters, 111(14):717–721, July 31, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001232.

Bernardeschi:2016:VSP

[BD16] Cinzia Bernardeschi and Andrea Domenici. Verifying safety properties of a nonlinear control by interactive theorem proving with the Prototype Verification System. Information Processing Letters, 116(6):409– 415, June 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016300072.

Bonamy:2018:IFV

[BDF^{+}18] Marthe Bonamy, Konrad K. Dabrowski, Carl Feghali, Matthew Johnson, and Daniël Paulusma. Independent feedback vertex sets for graphs of bounded diameter. Information Processing Letters, 131(??):26– 32, March 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017301989.

Biedl:2011:NIP

[BDH^{+}11] Therese Biedl, Stephane Durocher, Holger H. Hoos, Shuang Luan, Jared Saia, and Maxwell Young. A note on improving the performance of approximation algorithms for radiation therapy. Information Processing Letters, 111(7):326–333, March 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bonomo:2015:OOC

[BDNVP15] Flavia Bonomo, Guillermo Duran, Amedeo Napoli, and Mario Valencia-Pabon. A one-to-one correspondence between potential solutions of the cluster deletion problem and the minimum sum coloring problem, and its application to P_{4}-sparse graphs. Information Processing Letters, 115(6–8):600–603, June/ August 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000241.

Bouchard:2018:DRN

[BDPP18] Sébastien Bouchard, Yoann Dieudonné, Andrzej Pelc, and Franck Petit. On deterministic rendezvous at a node of agents with arbitrary velocities. Information Processing Letters, 133(??): 39–43, May 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300127.

Braga:2015:EDS

[BdSL15] Andrei Braga, Cid C.

de Souza, and Orlando Lee. The Eternal Dominating Set problem for proper interval graphs. Information Processing Letters, 115(6–8):582–587, June/ August 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015000216.

Best:2016:EHS

[BE16] Eike Best and Javier Esparza. Existence of home states in Petri nets is decidable. Information Processing Letters, 116(6):423– 427, June 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016000132.

Beekman:2016:DSA

[Bee16] Jethro G. Beekman. A Denial of Service attack against fair computations using Bitcoin deposits. Information Processing Letters, 116(2):144–146, February 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001647.

Bakhshi:2011:FLE

[BEFP11] Rena Bakhshi, Jörg Endrullis, Wan Fokkink, and Jun Pang. Fast leader election in anonymous rings with bounded expected delay. Information Processing Letters, 111(17):864–870, September 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011001682.

Bera:2011:LBM

[Ber11] Debajyoti Bera. A lower bound method for quantum circuits. Information Processing Letters, 111(15):723–726, August 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001220.

Bereg:2012:CGH

[Ber12] Sergey Bereg. Computing generalized ham-sandwich cuts. Information Processing Letters, 112(13):532– 534, July 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000816.

Bernath:2017:CSS

[Ber17]  Attila Bernáth.  Cover-

ing symmetric supermodular functions with graph edges: a short proof of a theorem of Benczúr and Frank. Information Processing Letters, 128(??):49–53, December 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901730145X.

Boyaci:2017:PTA

[BES17] Arman Boyaci, Tinaz Ekim, and Mordechai Shalom. A polynomial-time algorithm for the maximum cardinality cut problem in proper interval graphs. Information Processing Letters, 121(??): 29–33, May 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017300133.

Bakhshesh:2017:ACS

[BF17] Davood Bakhshesh and Mohammad Farshi. Angleconstrained spanners with angle at least π/3. Information Processing Letters, 120 (??):44–46, April 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019017300029.

Bakhshesh:2019:FTC

[BF19] Davood Bakhshesh and Mohammad Farshi. Fault tolerancy of continuous Yao graph of angle less than 2π/5. Information Processing Letters, 148(??):13– 18, August 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019300717.

Badkobeh:2013:BJS

[BFKL13] Golnaz Badkobeh, Gabriele Fici, Steve Kroon, and Zsuzsanna Lipták. Binary jumbled string matching for highly run-length compressible texts. Information Processing Letters, 113(17):604–608, August 30, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001415.

Brandstadt:2015:PTA

[BFLM15] Andreas Brandstädt, Pavel Ficur, Arne Leitert, and Martin Milanic. Polynomialtime algorithms for weighted efficient domination problems in AT-free graphs and dually chordal graphs. Information Processing Letters, 115(2):256–262, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002014.

Badkobeh:2018:AAP

[BFP18] Golnaz Badkobeh, Gabriele Fici, and Simon J. Puglisi. Algorithms for anti-powers in strings. Information Processing Letters, 137(??):57– 60, September 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001901830108X.

Bulteau:2015:SAR

[BFRV15] Laurent Bulteau, Guillaume Fertin, Romeo Rizzi, and Stéphane Vialette. Some algorithmic results for [2]- sumset covers. Information Processing Letters, 115(1): 1–5, January 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901400146X.

Bollig:2011:ROM

[BG11] Beate Bollig and Marc Gillé. Randomized OBDDs for the most significant bit of multiplication need exponential space. Information Processing Letters, 111(4):151–155, January 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brandstadt:2012:MWI

[BG12] Andreas Brandstädt and Vassilis Giakoumakis. Maximum Weight Independent Sets in hole- and co-chair-free graphs. Information Processing Letters, 112(3):67–71, January 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901100264X.

Brandstadt:2015:AMW

[BG15] Andreas Brandstädt and Vassilis Giakoumakis. Addendum to: Maximum Weight Independent Sets in hole- and co-chair-free graphs. Information Processing Letters, 115(2):345– 350, February 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014001963.

Bannai:2012:EAT

[BGI^{+}12] Hideo Bannai, Travis Gagie, Tomohiro I., Shunsuke Inenaga, Gad M. Landau, and Moshe Lewenstein. An efficient algorithm to test square-freeness of strings compressed by straightline programs. Information Processing Letters, 112(19):711–714, October 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001688.

Barbero:2016:LVK

[BGJ^{+}16] Florian Barbero, Gregory Gutin, Mark Jones, Bin Sheng, and Anders Yeo. Linear-vertex kernel for the problem of packing r-stars into a graph without long induced paths. Information Processing Letters, 116(6): 433–436, June 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019016000090.

Beyersdorff:2010:LBP

[BGL10] Olaf Beyersdorff, Nicola Galesi, and Massimo Lauria. A lower bound for the pigeonhole principle in tree-like Resolution by asymmetric Prover–Delayer games. Information Processing Letters, 110(23):1074– 1077, November 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Beyersdorff:2013:CTL

[BGL13] Olaf Beyersdorff, Nicola Galesi, and Massimo Lauria. A characterization of treelike resolution size. Information Processing Letters, 113(18):666–671, September 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001592.

Bouyer:2016:SSL

[BGM16] Patricia Bouyer, Patrick Gardy, and Nicolas Markey. On the semantics of Strategy Logic. Information Processing Letters, 116(2):75– 79, February 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001763.

Borowiecki:2012:CCC

[BGP12] Mieczyslaw Borowiecki, Jaroslaw Grytczuk, and Monika Pilśniak. Coloring chip configurations on graphs and digraphs. Information Processing Letters, 112(1–2):1– 4, January 15, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011002602.

Bonomo:2013:NCJ

[BGR13] Flavia Bonomo, Monia Giandomenico, and Fabrizio Rossi. A note on the Cornaz–Jost transformation to solve the graph coloring problem. Information Processing Letters, 113 (18):649–652, September 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001567.

Buhrman:2010:LPM

[BGSM10] Harry Buhrman, David García-Soriano, and Arie Matsliah. Learning parities in the mistake-bound model. Information Processing Letters, 111(1):16–21, December 15, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Becher:2011:EBS

[BH11] Verónica Becher and Pablo Ariel Heiber. On extending de Bruijn sequences. Information Processing Letters, 111(18):930–932, September 30, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011001840.

Bresar:2017:GTD

[BH17] Bostjan Bresar and Michael A. Henning. The game total domination problem is logcomplete in PSPACE. Information Processing Letters, 126(??):12–17, October 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017300996.

Bjorklund:2010:EPR

[BHKK10] Andreas Björklund, Thore Husfeldt, Petteri Kaski, and Mikko Koivisto. Evaluation of permanents in rings and semirings. Information Processing Letters, 110 (20):867–870, September 30, 2010. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Bjorklund:2011:CPL

[BHKK11] Andreas Björklund, Thore Husfeldt, Petteri Kaski, and Mikko Koivisto. Covering and packing in linear space. Information Processing Letters, 111(21–22):1033–1036, November 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011002237.

Bjorklund:2017:CPM

[BHL17] Andreas Björklund, Thore Husfeldt, and Isak Lyckberg. Computing the permanent modulo a prime power. Information Processing Letters, 125(??):20–25, September 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300741.

Barmpalias:2013:ACO

[BHLM13] G. Barmpalias, R. Hölzl, A. E. M. Lewis, and W. Merkle. Analogues of Chaitin’s Omega in the computably enumerable sets. Information Processing Letters, 113(5–6):171–178, March 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013000136.

Bartoschek:2010:RTC

[BHM^{+}10] C. Bartoschek, S. Held, J. Maßberg, D. Rautenbach, and J. Vygen. The repeater tree construction problem. Information Processing Letters, 110(24): 1079–1083, November 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bonchis:2014:IAA

[BI14] Cosmin Bonchis and Gabriel Istrate. Improved approximation algorithms for lowdensity instances of the Minimum Entropy Set Cover Problem. Information Processing Letters, 114(7):360– 364, July 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014000246.

Bird:2011:SDF

[Bir11] Richard S. Bird. A simple division-free algorithm for computing determinants. Information Processing Letters, 111(21–22):1072–1074, November 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019011002353.

Bouzid:2016:NCB

[BIR16] Zohir Bouzid, Damien Imbs, and Michel Raynal. A necessary condition for Byzantine k-set agreement. Information Processing Letters, 116(12):757–759, December 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901630093X.

Bhavani:2018:EFC

[BJ18] K. Bhavani and Sudarson Jena. Exchanged folded crossed cube: a new interconnection network for parallel computation. Information Processing Letters, 137(??):40–46, September 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300978.

Bang-Jensen:2015:VCE

[BJH15] Jørgen Bang-Jensen and Magnús M. Halldórsson. Vertex coloring edge-weighted digraphs. Information Processing Letters, 115(10): 791–796, October 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015000848.

Basin:2014:DSL

[BJKZ14] David Basin, Carlos Cotrini Jiménez, Felix Klaedtke, and Eugen Zalinescu. Deciding safety and liveness in TPTL. Information Processing Letters, 114(12):680–688, December 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001136.

Broelemann:2017:VRN

[BJMC17] Klaus Broelemann, Xiaoyi Jiang, Sudipto Mukherjee, and Ananda S. Chowdhury. Variants of k-regular nearest neighbor graph and their construction. Information Processing Letters, 126(??): 18–23, October 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019017301023.

Ban:2019:RP

[BJP^{+}19] Frank Ban, Kamal Jain, Christos H. Papadimitriou, Christos-Alexandros Psomas, and Aviad Rubinstein. Reductions in PPP. Information Processing Letters, 145(??):48–52, May 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300018.

Backer:2010:CFA

[BK10] Jonathan Backer and J. Mark Keil. Constant factor approximation algorithms for the densest k-subgraph problem on proper interval graphs and bipartite permutation graphs. Information Processing Letters, 110 (16):635–638, July 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bjorklund:2013:CCT

[BK13] Andreas Björklund and Petteri Kaski. Counting closed trails. Information Processing Letters, 113(1–2):1– 3, January 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002633.

Bringmann:2018:NHD

[BK18] Karl Bringmann and Sebastian Krinninger. A note on hardness of diameter approximation. Information Processing Letters, 133(??): 10–15, May 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300024.

Berenbrink:2014:ENC

[BKMT14] Petra Berenbrink, Bruce Krayenhoff, and Frederik Mallmann-Trenn. Estimating the number of connected components in sublinear time. Information Processing Letters, 114(11):639– 642, November 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014001033.

Bravo:2012:PEL

[BKN^{+}12] Raquel S. F. Bravo, Sulamita Klein, Loana T. Nogueira, Fábio Protti, and Rudini M. Sampaio. Partitioning extended P_{4}- laden graphs into cliques and stable sets. Information Processing Letters, 112(21):829–834, November 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001998.

Bampas:2018:PMS

[BKPP18] Evangelos Bampas, Christina Karousatou, Aris Pagourtzis, and Katerina Potika. Path multicoloring in spider graphs with even color multiplicity. Information Processing Letters, 133(??):1– 4, May 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300012.

Bibak:2016:MAM

[BKS16] Khodakhast Bibak, Bruce M.

Kapron, and Venkatesh Srinivasan. MMH^{∗} with arbitrary modulus is always almost-universal. Information Processing Letters, 116 (7):481–483, July 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019016300333.

Basin:2015:GCA

[BKZ15] David Basin, Felix Klaedtke, and Eugen Zalinescu. Greedily computing associative aggregations on sliding windows. Information Processing Letters, 115(2):186–192, February 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001859.

Bailey:2010:EIM

[BL10] James Bailey and Elsa Loekito. Efficient incremental mining of contrast patterns in changing data. Information Processing Letters, 110 (3):88–92, January 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bao:2012:IAA

[BL12] Xiaoguang Bao and Zhaohui Liu. An improved approximation algorithm for the clustered traveling salesman problem. Information Processing Letters, 112(23):908–910, December 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012002475.

Brandstadt:2014:NED

[BL14] Andreas Brandstädt and Van Bang Le. A note on efficient domination in a superclass of P_{5}-free graphs. Information Processing Letters, 114(7):357– 359, July 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014000258.

Bae:2010:ESE

[BLC10] Sang Won Bae, Chunseok Lee, and Sunghee Choi. On exact solutions to the Euclidean bottleneck Steiner tree problem. Information Processing Letters, 110 (16):672–678, July 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertini:2010:PPC

[BLM10] Luciano Bertini, Julius C. B. Leite, and Daniel Mossé. Power and performance control of soft realtime Web server clusters. Information Processing Letters, 110(17):767–773, August 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blondeau:2015:IDA

[Blo15] Céline Blondeau. Impossible differential attack on 13round Camellia-192. Information Processing Letters, 115(9):660–666, September 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000472.

Bao:2017:NAA

[BLYL17] Xiaoguang Bao, Zhaohui Liu, Wei Yu, and Ganggang Li. A note on approximation algorithms of the clustered traveling salesman problem. Information Processing Letters, 127(??):54– 57, November 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019017301254.

Brandstadt:2019:DWE

[BM19] Andreas Brandstädt and Martin Milanic. A dichotomy for weighted efficient dominating sets with bounded degree vertices. Information Processing Letters, 142(??):30–34, February 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901830200X.

Benamara:2016:ICA

[BMB16] Oualid Benamara, Fatiha Merazka, and Kamel Betina. An improvement of a cryptanalysis algorithm. Information Processing Letters, 116(2):192–196, February 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001398.

Boutsidis:2014:NSL

[BMI14] Christos Boutsidis and Malik Magdon-Ismail. A note on sparse least-squares regression. Information Processing Letters, 114(5):273– 276, May 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002913.

Brandstatter:2013:ASS

[BMW13] Nina Brandstätter, Wilfried Meidl, and Arne Winterhof. Addendum to Sidelńikov sequences over nonprime fields. Information Processing Letters, 113(9):332–336, May 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000549.

Balcerzak:2010:TWD

[BN10a] Marcin Balcerzak and Damian Niwiński. Two-way deterministic automata with two reversals are exponentially more succinct than with one reversal. Information Processing Letters, 110(10): 396–398, April 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brass:2010:FMB

[BN10b] Peter Brass and Hyeon-Suk Na. Finding the maximum bounded intersection of k out of n halfplanes. Information Processing Letters, 110(3):113–115, January 1, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bezem:2010:HPM

[BNRC10] Marc Bezem, Robert Nieuwenhuis, and Enric Rodríguez-Carbonell. Hard problems in max-algebra, control theory, hypergraphs and other areas. Information Processing Letters, 110(4):133–138, January 16, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bogdanov:2010:DLE

[Bog10] Andrey Bogdanov. On the differential and linear efficiency of balanced Feistel networks. Information Processing Letters, 110 (20):861–866, September 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bollig:2010:ESC

[Bol10] Beate Bollig. Exponential space complexity for OBDD-based reachability analysis. Information Processing Letters, 110(21): 924–927, October 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bollig:2014:SCL

[Bol14] Beate Bollig. A simpler counterexample to a longstanding conjecture on the complexity of Bryant’s apply algorithm. Information Processing Letters, 114(3): 124–129, March 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013002718.

Bonifaci:2013:PCC

[Bon13] Vincenzo Bonifaci. Physarum can compute shortest paths: a short proof. Information Processing Letters, 113(1– 2):4–7, January 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019012002621.

Borassi:2016:NCC

[Bor16] Michele Borassi. A note on the complexity of computing the number of reachable vertices in a digraph. Information Processing Letters, 116(10):628–630, October 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300734.

Braverman:2015:WSR

[BOV15] Vladimir Braverman, Rafail Ostrovsky, and Gregory Vorsanger. Weighted sampling without replacement from data streams. Information Processing Letters, 115(12):923–926, December 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001295.

Berry:2011:SAG

[BP11] Anne Berry and Romain Pogorelcnik. A simple algorithm to generate the minimal separators and the maximal cliques of a chordal graph. Information Processing Letters, 111(11):508– 511, May 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baril:2014:MSM

[BP14] J.-L. Baril and J.-M. Pallo. Motzkin subposets and Motzkin geodesics in Tamari lattices. Information Processing Letters, 114(1–2): 31–37, January/February 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019013002512.

Bueno:2018:HFG

[BPP^{+}18] Letícia R. Bueno, Lucia D. Penso, Fábio Protti, Victor R. Ramos, Dieter Rautenbach, and Uéverton S. Souza. On the hardness of finding the geodetic number of a subcubic graph. Information Processing Letters, 135(??):22– 27, July 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300413.

Bijlsma:2018:TMT

[BPPS18] A. Bijlsma, H. J. M. Passier, H. J. Pootjes, and S. Stuurman. Template method test pattern. Information Processing Letters, 139(??):1–7, November 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301339.

Bajric:2014:GBF

[BPRMS14] Samed Bajrić, Enes Pasalic, Amela Ribić-Muratović, and Gangopadhyay Sugata. On generalized bent functions with Dillon’s exponents. Information Processing Letters, 114(4):222– 227, April 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002688.

Baudon:2012:MAP

[BPW12] Olivier Baudon, Jakub Przybylo, and Mariusz Woźniak. On minimal arbitrarily partitionable graphs. Information Processing Letters, 112(17–18):697–700, September 30, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019012001615.

Bagheri:2010:PSL

[BR10a] Alireza Bagheri and Mohammadreza Razzazi. Planar straight-line point-set embedding of trees with partial embeddings. Information Processing Letters, 110 (12–13):521–523, June 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bonnet:2010:SPN

[BR10b] François Bonnet and Michel Raynal. A simple proof of the necessity of the failure detector Σ to implement an atomic register in asynchronous message-passing systems. Information Processing Letters, 110(4):153– 157, January 16, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baumeister:2012:TFS

[BR12] Dorothea Baumeister and Jörg Rothe. Taking the final step to a full dichotomy of the possible winner problem in pure scoring rules. Information Processing Letters, 112(5):186–190, February 28, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011003218.

Brand:2015:ASR

[Bra15a]  Michael Brand.  Arbi-

trary sequence RAMs. Information Processing Letters, 115(2):193–198, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001860.

Branzei:2015:NEF

[Brâ15b] Simina Brânzei. A note on envy-free cake cutting with polynomial valuations. Information Processing Letters, 115(2):93–95, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001422.

Brand:2017:SPP

[Bra17] Michael Brand. Small polyomino packing. Information Processing Letters, 126(??): 30–34, October 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019017301059.

Brandenburg:2018:FCF

[Bra18] Franz J. Brandenburg.

On fan-crossing and fancrossing free graphs. Information Processing Letters, 138(??):61–66, October 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301315.

Binkele-Raible:2010:EET

[BRFGL10] Daniel Binkele-Raible, Henning Fernau, Serge Gaspers, and Mathieu Liedloff. Exact exponential-time algorithms for finding bicliques. Information Processing Letters, 111(2):64–67, December 31, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barenghi:2013:PPO

[BRMP13] Alessandro Barenghi, Stefano Crespi Reghizzi, Dino Mandrioli, and Matteo Pradella. Parallel parsing of operator precedence grammars. Information Processing Letters, 113(7):245–249, April 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000148.

Bogdanov:2011:ALG

[BS11] Andrey Bogdanov and Kyoji Shibutani. Analysis of 3-line generalized Feistel networks with double SD-functions. Information Processing Letters, 111(13): 656–660, July 1, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011000937.

Berwanger:2012:PGU

[BS12] Dietmar Berwanger and Olivier Serre. Parity games on undirected graphs. Information Processing Letters, 112(23):928–932, December 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002487.

Broere:2017:SBG

[BS17] Izak Broere and Gabriel Semanisin. Some bounds on the generalised total chromatic number of degenerate graphs. Information Processing Letters, 122(??):30– 33, June 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300352.

Bastkowski:2014:FME

[BSM14] Sarah Bastkowski, Andreas Spillner, and Vincent Moulton. Fishing for minimum evolution trees with Neighbor-Nets. Information Processing Letters, 114(1–2):13–18, January/ February 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013002603.

Bonacina:2016:IRW

[BT16a] Ilario Bonacina and Navid Talebanfard. Improving resolution width lower bounds for k-CNFs with applications to the Strong Exponential Time Hypothesis. Information Processing Letters, 116(2):120–124, February 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001684.

Bredereck:2016:NHT

[BT16b] Robert Bredereck and Nimrod Talmon. NP-hardness of two edge cover generalizations with applications to control and bribery for approval voting. Information Processing Letters, 116(2):147–152, February 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001635.

Bender:2015:PIS

[BTW15] Marco Bender, Clemens Thielen, and Stephan Westphal. Packing items into several bins facilitates approximating the separable assignment problem. Information Processing Letters, 115(6–8):570–575, June/ August 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000186.

Bansal:2017:TAB

[BU17] Nikhil Bansal and Seeun William Umboh. Tight approximation bounds for dominating set on graphs of bounded arboricity. Information Processing Letters, 122(??):21– 24, June 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300200.

Baril:2010:WMD

[BV10] Jean-Luc Baril and Rémi Vernay. Whole mirror duplication-random loss model and pattern avoiding permutations. Information Processing Letters, 110 (11):474–480, May 16, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boldi:2011:IEF

[BV11] Paolo Boldi and Sebastiano Vigna. E = I + T: the internal extent formula for compacted tries. Information Processing Letters, 111 (7):310–313, March 1, 2011.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bhaskara:2019:APC

[BV19] Aditya Bhaskara and Antoine Vigneron. Approximating a planar convex set using a sparse grid. Information Processing Letters, 149(??):10–13, September 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019300900.

Bonizzoni:2010:VCL

[BVDP10] Paola Bonizzoni, Gianluca Della Vedova, Riccardo Dondi, and Yuri Pirola. Variants of constrained longest common subsequence. Information Processing Letters, 110(20): 877–881, September 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bodlaender:2019:EAC

[BvdZ19] Hans L. Bodlaender and Tom C. van der Zanden. On exploring always-connected temporal graphs of small pathwidth. Information Processing Letters, 142(??): 68–71, February 2019. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019018302102.

Barreto:2012:HCS

[BVF12] João Barreto, Luís Veiga, and Paulo Ferreira. Hash challenges: Stretching the limits of compare-by-hash in distributed data deduplication. Information Processing Letters, 112(10):380– 385, May 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000385.

Bodlaender:2015:GSM

[BvK15] Hans L. Bodlaender and Marc van Kreveld. Google Scholar makes it hard — the complexity of organizing one’s publications. Information Processing Letters, 115(12):965–968, December 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001167.

Botelho:2012:CRP

[BWZ12] Fabiano C. Botelho, Nicholas Wormald, and Nivio Ziviani. Cores of random rpartite hypergraphs. Information Processing Letters, 112(8–9):314–319, April 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011002985.

Belhoul:2014:ESS

[BYK14] Yacine Belhoul, Saïd Yahiaoui, and Hamamache Kheddouci. Efficient selfstabilizing algorithms for minimal total k-dominating sets in graphs. Information Processing Letters, 114(7): 339–343, July 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019014000209.

Chehreghani:2017:ULB

[CA17] Mostafa Haghir Chehreghani and Talel Abdessalem. Upper and lower bounds for the q-entropy of network models with application to network model selection. Information Processing Letters, 119 (??):1–8, March 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019016301533.

Cai:2015:SCI

[Cai15] Junqing Cai. A sufficient condition involving implicit degree and neighborhood intersection for long cycles. Information Processing Letters, 115(2):225–227, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001926.

Calamoneri:2013:OLE

[Cal13] Tiziana Calamoneri. Optimal L(δ_{1}, δ_{2}, 1)-labeling of eight-regular grids. Information Processing Letters, 113(10–11):361–364, May/ June 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901300080X.

Carvalho:2019:SEA

[Car19] Iago A. Carvalho. On the statistical evaluation of algorithmic’s computational experimentation with infeasible solutions. Information Processing Letters, 143(??): 24–27, March 2019. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019018302266.

Clymo:2018:RSW

[CB18] Judith Clymo and Olaf Beyersdorff. Relating size and width in variants of Qresolution. Information Processing Letters, 138(??):ii, October 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018301108.

Charron-Bost:2010:SLT

[CBHW10] Bernadette Charron-Bost, Martin Hutle, and Josef Widder. In search of lost time. Information Processing Letters, 110(21): 928–933, October 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cao:2011:ECD

[CBSV11] Xiang-Lan Cao, Spela Brglez, Simon Spacapan, and Elkin Vumar. On edge connectivity of direct products of graphs. Information Processing Letters, 111(18):899–902, September 30, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011001724.

Chang:2012:EAL

[CC12] Chia-Jung Chang and Kun-Mao Chao. Efficient algorithms for local ranking. Information Processing Letters, 112(13):517–522, July 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012000798.

Cariow:2013:AFM

[CC13] Aleksandr Cariow and Galina Cariowa. An algorithm for fast multiplication of sedenions. Information Processing Letters, 113(9):324–331, May 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000653.

Choi:2014:TES

[CC14] Kwanghoon Choi and Byeong-Mo Chang. A type and effect system for activation flow of components in Android programs. Information Processing Letters, 114(11):620–627, November 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001069.

Castonguay:2019:GNC

[CCCN19] Diane Castonguay, Erika M. M. Coelho, Hebert Coelho, and Julliano R. Nascimento. On the geodetic number of complementary prisms. Information Processing Letters, 144(??): 39–42, April 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018302394.

Christou:2012:CAS

[CCF^{+}12] Michalis Christou, Maxime Crochemore, Tomás Flouri, Costas S. Iliopoulos, Jan Janousek, Borivoj Melichar, and Solon P. Pissis. Computing all subtree repeats in ordered trees. Information Processing Letters, 112(24):958–962, December 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012002505.

Cao:2014:IPA

[CCF14] Yixin Cao, Jianer Chen, and J.-H. Fan. An O^{∗}(1.84^{k}) parameterized algorithm for the multiterminal cut problem. Information Processing Letters, 114(4):167–173, April 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013003049.

Chang:2014:EAP

[CCH^{+}14a] Maw-Shang Chang, Li-Hsuan Chen, Ling-Ju Hung, Peter Rossmanith, and Guan-Han Wu. Exact algorithms for problems related to the densest k-set problem. Information Processing Letters, 114(9):510–513, September 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014000684.

Cho:2014:FCR

[CCH14b] Young In Cho, Nam Su Chang, and Seokhie Hong. Formulas for cube roots in F_{3^{m}} using shifted polynomial basis. Information Processing Letters, 114(6):331– 337, June 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014000027.

Chang:2013:AAS

[CCKP13] Gerard Jennhwa Chang, Chan-Wei Chang, David Kuo, and Sheung-Hung Poon. Algorithmic aspect of stratified domination in graphs. Information Processing Letters, 113 (22–24):861–865, November/December 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901300224X.

Chang:2010:IAR

[CCL10] Maw-Shang Chang, Hsiao-Han Chung, and Chuang-Chieh Lin. An improved algorithm for the red–blue hitting set problem with the consecutive ones property. Information Processing Letters, 110(20):845– 848, September 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Canale:2014:CCK

[CCRS14] Eduardo Canale, Héctor Cancela, Franco Robledo, and Pablo Sartor. The complexity of computing the 2-K-reliability in networks. Information Processing Letters, 114(9):457–461, September 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014000532.

Chiu:2014:MSS

[CCT14] Well Y. Chiu, Chiuyuan Chen, and Shih-Yu Tsai. A 4n-move self-stabilizing algorithm for the minimal dominating set problem using an unfair distributed daemon. Information Processing Letters, 114(10): 515–518, October 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019014000702.

Cheng:2012:RSP

[CCXW12] Qiang Cheng, Feng Chen, Wenli Xu, and Song Wang. Recursive sum-product algorithm for generalized outer-planar graphs. Information Processing Letters, 112(11):449–456, June 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000695.

Chang:2016:LET

[CCYW16] Jou-Ming Chang, Xiang-Rui Chen, Jinn-Shyong Yang, and Ro-Yu Wu. Locally exchanged twisted cubes: Connectivity and super connectivity. Information Processing Letters, 116(7):460– 466, July 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300266.

Colantonio:2010:CCC

[CD10] Alessandro Colantonio and Roberto Di Pietro. Concise: Compressed ‘n’ Composable Integer Set. Information Processing Letters, 110 (16):644–650, July 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Clare:2019:ESF

[CD19] Amanda Clare and Jacqueline W. Daykin. Enhanced string factoring from alphabet orderings. Information Processing Letters, 143(??): 4–7, March 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018302059.

Chaim:2013:EBA

[CdA13] Marcos Lordello Chaim and Roberto Paulo Andrioli de Araujo. An efficient bitwise algorithm for intraprocedural data-flow testing coverage. Information Processing Letters, 113(8):293– 300, April 30, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019013000537.

Caceres:2011:WPT

[CDD^{+}11] Alan Joseph J. Caceres, Samantha Daley, John De-Jesus, Michael Hintze, Diquan Moore, and Katherine St. John. Walks in phylogenetic treespace. Information Processing Letters, 111(12): 600–604, June 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019011000780.

Crescenzi:2011:SMM

[CDM^{+}11] Pilu Crescenzi, Miriam Di Ianni, Andrea Marino, Donatella Merlini, Gianluca Rossi, and Paola Vocca. Smooth movement and Manhattan path based Random Waypoint mobility. Information Processing Letters, 111(5):239–246, February 1, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Clement:2010:CPA

[CDPB^{+}10] Julien Clement, Xavier Défago, Maria Gradinariu Potop-Butucaru, Taisuke Izumi, and Stephane Messika. The cost of probabilistic agreement in oblivious robot networks. Information Processing Letters, 110 (11):431–438, May 16, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carayol:2013:FRC

[CÉ13] Arnaud Carayol and Zoltán Ésik. The FC-rank of a context-free language. Information Processing Letters, 113(8):285–287, April 30, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000112.

Chawachat:2012:NUB

[CFJ12] Jakarin Chawachat, Jittat Fakcharoenphol, and Wattana Jindaluang. The nonuniform Bounded Degree Minimum Diameter Spanning Tree problem with an application in P2P networking. Information Processing Letters, 112(24):937–941, December 31, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019012002347.

Chaplick:2018:SGU

[CFMR18] Steven Chaplick, Maximilian Fürst, Frédéric Maffray, and Dieter Rautenbach. On some graphs with a unique perfect matching. Information Processing Letters, 139(??):53–59, November 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018301558.

Caprara:2016:STK

[CFMT16] Alberto Caprara, Fabio Furini, Enrico Malaguti, and Emiliano Traversi. Solving the Temporal Knapsack Problem via recursive Dantzig–Wolfe reformulation. Information Processing Letters, 116(5):379– 386, May 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016000107.

Cipriano:2010:MHC

[CG10a] Christopher C. Cipriano and Teofilo F. Gonzalez. Multicasting in the hypercube, chord and binomial graphs. Information Processing Letters, 110(17):774–777, August 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crespelle:2010:UCB

[CG10b] Christophe Crespelle and Philippe Gambette. Unrestricted and complete Breadth-First Search of trapezoid graphs in O(n) time. Information Processing Letters, 110(12–13):497– 502, June 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chassein:2015:AFO

[CG15] André Chassein and Marc Goerigk. Alternative formulations for the ordered weighted averaging objective. Information Processing Letters, 115(6–8):604–608, June/August 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000319.

Cucu-Grosjean:2010:PFJ

[CGG10] Liliana Cucu-Grosjean and Joël Goossens. Predictability of Fixed-Job Priority schedulers on heterogeneous multiprocessor realtime systems. Information Processing Letters, 110(10):

399–402, April 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cegielski:2014:LRS

[CGG14] Patrick Cégielski, Serge Grigorieff, and Irène Guessarian. On lattices of regular sets of natural integers closed under decrementation. Information Processing Letters, 114(4):197– 202, April 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013002937.

Crowston:2010:NML

[CGJ10] Robert Crowston, Gregory Gutin, and Mark Jones. Note on Max Lin-2 above Average. Information Processing Letters, 110(11): 451–454, May 16, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crowston:2012:PES

[CGJY12] R. Crowston, G. Gutin, M. Jones, and A. Yeo. Parameterized Eulerian strong component arc deletion problem on tournaments. Information Processing Letters, 112(6):249–251, March 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011003188.

Cohen:2010:BEC

[CGLS10] Elad Cohen, Martin Charles Golumbic, Marina Lipshteyn, and Michal Stern. On the bi-enhancement of chordal-bipartite probe graphs. Information Processing Letters, 110(5):193– 197, February 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cherniavsky:2016:EBF

[CGLS16] Yonah Cherniavsky, Avraham Goldstein, Vadim E. Levit, and Robert Shwartz. Enumeration of balanced finite group valued functions on directed graphs. Information Processing Letters, 116(7):484–488, July 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019016300084.

Creus:2014:EDP

[CGR14] Carles Creus, Guillem Godoy, and Lander Ramos. Excessively duplicating patterns represent non-regular languages. Information Processing Letters, 114(3):85– 93, March 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901300286X.

Cenciarelli:2018:INM

[CGS18] Pietro Cenciarelli, Daniele Gorla, and Ivano Salvo. Inefficiencies in network models: A graph-theoretic perspective. Information Processing Letters, 131(??):44– 50, March 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017301862.

Clark:2012:SIT

[CH12] David Clark and Robert M. Hierons. Squeeziness: an information theoretic measure for avoiding fault masking. Information Processing Letters, 112(8–9):335–340, April 30, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901200021X.

Chang:2011:TCU

[Cha11] Ching-Lueh Chang. Triggering cascades on undirected connected graphs. Information Processing Letters, 111(19):973–978, October 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011002006.

Chakrabarti:2012:NRS

[Cha12]  Amit Chakrabarti.  A

note on randomized streaming space bounds for the longest increasing subsequence problem. Information Processing Letters, 112(7):261–263, March 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901100336X.

Chang:2013:DST

[Cha13] Ching-Lueh Chang. Deterministic sublinear-time approximations for metric 1-median selection. Information Processing Letters, 113(8):288–292, April 30, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000483.

Chang:2019:VAM

[Cha19] Ching-Lueh Chang. On Las Vegas approximations for metric 1-median selection. Information Processing Letters, 146(??):44–48, June 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300444.

Chen:2010:CPT

[Che10a] Xie-Bin Chen. Cycles passing through a prescribed path in a hypercube with faulty edges. Information Processing Letters, 110(16): 625–629, July 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2010:EFT

[Che10b] Xie-Bin Chen. Edge-fault-tolerant diameter and bipanconnectivity of hypercubes. Information Processing Letters, 110(24):1088– 1092, November 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2010:UMM

[Che10c] Xie-Bin Chen. Unpaired many-to-many vertex-disjoint path covers of a class of bipartite graphs. Information Processing Letters, 110(6): 203–205, February 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2011:PCO

[Che11] Xie-Bin Chen. Parallel construction of optimal independent spanning trees on Cartesian product of complete graphs. Information Processing Letters, 111(5): 235–238, February 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2012:PMM

[Che12] Xie-Bin Chen. Paired manyto-many disjoint path covers of hypercubes with faulty edges. Information Processing Letters, 112(3):61–66, January 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901100281X.

Chen:2016:IND

[Che16a] Ke Chen. Improved neural dynamics for online Sylvester equations solving. Information Processing Letters, 116(7):455–459, July 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300278.

Chen:2016:HHF

[Che16b] Xie-Bin Chen. Hamiltonicity of hypercubes with faulty vertices. Information Processing Letters, 116(5):343– 346, May 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016000041.

Chen:2016:PDP

[Che16c] Xie-Bin Chen. Paired 2disjoint path covers of multidimensional torus networks with faulty edges. Information Processing Letters, 116(2):107–110, February 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001714.

Chen:2017:PRV

[Che17]  Zhe Chen.  Paramet-

ric runtime verification is NP-complete and coNPcomplete. Information Processing Letters, 123(??):14– 20, July 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300339.

Chen:2018:PDS

[Che18a] Jing Chen. The pessimistic diagnosability of Split-Star Networks under the PMC model. Information Processing Letters, 136(??):80– 82, August 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300632.

Chen:2018:GMD

[Che18b] Po-An Chen. Generalized mirror descents with nonconvex potential functions in atomic congestion games: Continuous time and discrete time. Information Processing Letters, 130(??):36– 39, February 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301813.

Cheng:2015:EEC

[CHF15] Dongqin Cheng, Rong-Xia Hao, and Yan-Quan Feng. Embedding even cycles on folded hypercubes with conditional faulty edges. Information Processing Letters, 115(12):945–949, December 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001374.

Chen:2013:CMC

[CHK13] Taolue Chen, Tingting Han, and Marta Kwiatkowska. On the complexity of model checking interval-valued discrete time Markov chains. Information Processing Letters, 113(7):210–216, April 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000100.

Choi:2012:ATA

[Cho12] Jin-Ghoo Choi. Analysis of total average queue length in multi-hop wireless networks. Information Processing Letters, 112(3):90–94, January 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011002808.

Clark:2019:NSF

[CHP19] David Clark, Robert M. Hierons, and Krishna Patel. Normalised squeeziness and failed error propagation. Information Processing Letters, 149(??):6–9, September 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019300729.

Chang:2010:TCP

[CHR10] Gerard Jennhwa Chang, Jianfeng Hou, and Nicolas Roussel. On the total choosability of planar graphs and of sparse graphs. Information Processing Letters, 110(20):849–853, September 30, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Christensen:2019:ORR

[Chr19] René Bødker Christensen. On one-round reliable message transmission. Information Processing Letters, 147 (??):22–26, July 2019. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019300481.

Crochemore:2013:NEC

[CIK^{+}13] M. Crochemore, C. S. Iliopoulos, T. Kociumaka, M. Kubica, J. Pachocki, J. Radoszewski, W. Rytter, W. Tyczyński, and T. Waleń. A note on efficient computation of all Abelian periods in a string. Information Processing Letters, 113(3):74–77, February 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002943.

Civril:2013:NHS

[Çiv13]  A. Çivril.  A note on

the hardness of sparse approximation. Information Processing Letters, 113(14–16):543–545, July/ August 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001312.

Cechlarova:2011:EIE

[CJ11] Katarína Cechlárová and Eva Jelínková. An efficient implementation of the equilibrium algorithm for housing markets with duplicate houses. Information Processing Letters, 111(13): 667–670, July 1, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019011001050.

Chapuy:2011:SRS

[CK11a] Guillaume Chapuy and Ines Klimann. On the supports of recognizable series over a field and a single letter alphabet. Information Processing Letters, 111(23–24):1096–1098, December 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011002596.

Cygan:2011:CAF

[CK11b] Marek Cygan and Lukasz Kowalik. Channel assignment via fast zeta transform. Information Processing Letters, 111(15):727– 730, August 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019011001281.

Couturier:2012:BIS

[CK12a]  Jean-François  Couturier

and Dieter Kratsch. Bicolored independent sets and bicliques. Information Processing Letters, 112(8–9):329–334, April 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000270.

Croce:2012:NMS

[CK12b] F. Della Croce and C. Koulamas. A note on minimizing the sum of quadratic completion times on two identical parallel machines. Information Processing Letters, 112(19):738–742, October 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001627.

Cepek:2013:BFL

[CKK13] Ondrej Cepek, Petr Kucera, and Stanislav Kurík. Boolean functions with long prime implicants. Information Processing Letters, 113(19– 21):698–703, September/ October 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001804.

Chang:2010:RNG

[CKL10] Chan-Wei Chang, David Kuo, and Hsing-Ching Lin.

Ranking numbers of graphs. Information Processing Letters, 110(16):711–716, July 31, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chakraborty:2016:TWU

[CKP16] Sourav Chakraborty, Akshay Kamath, and Rameshwar Pratap. Testing whether the uniform distribution is a stationary distribution. Information Processing Letters, 116(7):475– 480, July 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016300308.

Cohen:2015:USG

[CKY15] Reuven Cohen, Liran Katzir, and Aviv Yehezkel. A unified scheme for generalizing cardinality estimators to sum aggregation. Information Processing Letters, 115(2):336–342, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014002245.

Chen:2010:NCM

[CL10] Liqun Chen and Jiangtao Li. A note on the Chen–Morrissey–Smart DAA scheme. Information Processing Letters, 110(12– 13):485–488, June 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2015:AVD

[CL15] Xiang’en Chen and Zepeng Li. Adjacent-vertex-distinguishing proper edge colorings of planar bipartite graphs with Δ = 9, 10, or 11. Information Processing Letters, 115(2):263–268, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002026.

Chen:2012:OBD

[CLM12] Xing Chen, Juan Liu, and Jixiang Meng. λ′- Optimality of Bipartite Digraphs. Information Processing Letters, 112(17– 18):701–705, September 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001214.

Choi:2010:NMM

[CLP10] Byung-Cheon Choi, Joseph Y.-T. Leung, and Michael L. Pinedo. A note on makespan minimization in proportionate flow shops. Information Processing Letters, 111(2): 77–81, December 31, 2010.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheng:2012:DCD

[CLQS12] E. Cheng, L. Lipták, K. Qiu, and Z. Shen. On deriving conditional diagnosability of interconnection networks. Information Processing Letters, 112(17–18):674–677, September 30, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019012001597.

Chang:2013:TCQ

[CLR13] Maw-Shang Chang, Chuang-Chieh Lin, and Peter Rossmanith. Testing consistency of quartet topologies: a parameterized approach. Information Processing Letters, 113(22– 24):852–857, November/ December 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013002226.

Cheng:2013:SLD

[CLS13] Eddie Cheng, László Lipták, and Daniel E. Steffy. Strong local diagnosability of (n, k)- star graphs and Cayley graphs generated by 2-trees with missing edges. Information Processing Letters, 113(12):452–456, June 30, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000732.

Chiou:2011:MIL

[CLY11] Che Wun Chiou, Chiou-Yng Lee, and Yun-Chi Yeh. Multiplexer implementation of low-complexity polynomial basis multiplier in GF(m^{2}) using all one polynomial. Information Processing Letters, 111(21–22):1044–1047, November 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011002250.

Cai:2018:IDC

[CLZ18] Junqing Cai, Hao Li, and Yuzhong Zhang. An implicit degree condition for kconnected 2-heavy graphs to be Hamiltonian. Information Processing Letters, 134 (??):9–13, June 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019018300255.

Chaturvedi:2017:NFM

[CM17] Mmanu Chaturvedi and Ross M. McConnell. A note on finding minimum mean cycle. Information Processing Letters, 127(??):21–22, November 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017301084.

Croitoru:2018:TS

[CM18] Cosmina Croitoru and Kurt Mehlhorn. On testing substitutability. Information Processing Letters, 138(??): 12–18, October 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019018301133.

Chen:2018:ORA

[CN18] Meirun Chen and Reza Naserasr. The optimal routing of augmented cubes. Information Processing Letters, 136(??):59–63, August 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300838.

Constantinou:2015:ERC

[CNKS15] Eleni Constantinou, Athanasios Naskos, George Kakarontzas, and Ioannis Stamelos. Extracting reusable components: a semi-automated approach for complex structures. Information Processing Letters, 115(3):414– 417, March 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002579.

Cho:2015:FAO

[CNPS15] Sukhyeun Cho, Joong Chae Na, Kunsoo Park, and Jeong Seop Sim. A fast algorithm for order-preserving pattern matching. Information Processing Letters, 115(2):397–402, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002336.

Cranston:2013:HCR

[CO13] Daniel W. Cranston and Suil O. Hamiltonicity in connected regular graphs. Information Processing Letters, 113(22– 24):858–860, November/ December 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002214.

Chatterjee:2017:PRC

[CO17] Krishnendu Chatterjee and Georg Osang. Pushdown reachability with constant treewidth. Information Processing Letters, 122(??):25– 29, June 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017300224.

Courcelle:2014:CWE

[Cou14] Bruno Courcelle. Cliquewidth and edge contraction. Information Processing Letters, 114(1–2):42–44, January/February 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019013002494.

Clifford:2010:FAM

[CP10] Raphaël Clifford and Ely Porat. A filtering algorithm for k-mismatch with don’t cares. Information Processing Letters, 110(22):1021– 1025, October 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Clifford:2011:MSI

[CP11] Raphaël Clifford and Alexandru Popa. Maximum subset intersection. Information Processing Letters, 111 (7):323–325, March 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cygan:2013:SVD

[CP13] Marek Cygan and Marcin Pilipczuk. Split Vertex Deletion meets Vertex Cover: New fixed-parameter and exact exponentialtime algorithms. Information Processing Letters, 113(5–6):179–182, March 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000069.

Chung:2015:NIP

[CP15] Yerim Chung and Myoung-Ju Park. Notes on inverse bin-packing problems. Information Processing Letters, 115(1):60–68, January 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901400180X.

Chen:2016:CLB

[CP16] Shiteng Chen and Periklis A. Papakonstantinou. Correlation lower bounds from correlation upper bounds. Information Processing Letters, 116(8):537– 540, August 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300424.

Chang:2018:EMB

[CPHS18] Ku-Young Chang, Sun-Mi Park, Dowon Hong, and Changho Seo. Efficient multiplier based on hybrid approach for Toeplitz matrixvector product. Information Processing Letters, 131(??): 33–38, March 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019017302004.

Chen:2013:EIS

[CPTZ13] Qi Chen, Dingyi Pei, Chunming Tang, and Gansen Zhao. Efficient integer span program for hierarchical threshold access structure. Information Processing Letters, 113(17):621–627, August 30, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001518.

Cygan:2011:CDF

[CPW11] Marek Cygan, Marcin Pilipczuk, and Jakub Onufry Wojtaszczyk. Capacitated domination faster than O(n^{2}). Information Processing Letters, 111(23– 24):1099–1103, December 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011002390.

Chen:2012:LAP

[CQ12] Hong-Yu Chen and Jian-Ming Qi. The linear arboricity of planar graphs with maximum degree at least 5. Information Processing Letters, 112(20):767–771, October 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001585.

Chevalier:2010:CSC

[CR10] Yannick Chevalier and Michaël Rusinowitch. Compiling and securing cryptographic protocols. Information Processing Letters, 110(3):116–122, January 1, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chan:2018:IAA

[CR18] Timothy M. Chan and Zahed Rahmati. An improved approximation algorithm for the discrete Fréchet distance. Information Processing Letters, 138(??):67– 71, October 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301364.

Christensen:2010:NAF

[CRJ10] Ken Christensen, Allen Roginsky, and Miguel Jimeno. A new analysis of the false positive rate of a Bloom filter. Information Processing Letters, 110(21):944–949, October 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Croitoru:2015:NQK

[Cro15]  Cosmina Croitoru.  A

note on quasi-kernels in digraphs. Information Processing Letters, 115(11): 863–865, November 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0020019015001131.

Courtieu:2015:IGC

[CRTU15] Pierre Courtieu, Lionel Rieg, Sébastien Tixeuil, and Xavier Urbain. Impossibility of gathering, a certification. Information Processing Letters, 115(3):447– 452, March 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014002348.

Clifford:2010:PFM

[CS10] Raphaël Clifford and Benjamin Sach. Permuted function matching. Information Processing Letters, 110 (22):1012–1015, October 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carayol:2016:MSP

[CS16a] Arnaud Carayol and Olivier Serre. Marking shortest paths on pushdown graphs does not preserve MSO decidability. Information Processing Letters, 116(10): 638–643, October 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019016300758.

Chen:2016:EAS

[CS16b] Chien Chin Chen and Yu-Chun Sun. Exploring acquaintances of social network site users for effective social event recommendations. Information Processing Letters, 116(3):227– 236, March 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015002070.

Csirmaz:2012:CUA

[Csi12] Laszlo Csirmaz. Complexity of universal access structures. Information Processing Letters, 112(4):149–152, February 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011003085.

Chen:2016:GFP

[CSX16] Yao Chen, Kang G. Shin, and Huagang Xiong. Generalizing fixed-priority scheduling for better schedulability in mixed-criticality systems. Information Processing Letters, 116(8):508–512, August 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016300151.

Crochemore:2011:CLP

[CT11] Maxime Crochemore and German Tischler. Computing Longest Previous nonoverlapping Factors. Information Processing Letters, 111(6):291–295, February 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2016:HDG

[CT16a] Y-Chuang Chen and Tsung-Han Tsai. Hamiltonian decomposition of generalized recursive circulant graphs. Information Processing Letters, 116(9):585– 589, September 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001901630045X.

Chhabra:2016:FMO

[CT16b] Tamanna Chhabra and Jorma Tarhio. A filtration method for order-preserving matching. Information Processing Letters, 116(2):71– 74, February 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001775.

Chin:2013:SMB

[CTHP13] Ji-Jian Chin, Syh-Yuan Tan, Swee-Huay Heng, and Raphael C.-W. Phan. On the security of a modified Beth identity-based identification scheme. Information Processing Letters, 113(14–16):580–583, July/ August 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001324.

Canete-Valdeon:2012:APD

[CV12] José M. Cañete-Valdeón.

Annotating problem diagrams with architectural tactics for reasoning on quality requirements. Information Processing Letters, 112(16):656–661, August 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001421.

Cai:2019:SCF

[CV19] Xuepeng Cai and Elkin Vumar. The super connectivity of folded crossed cubes. Information Processing Letters, 142(??):52–56, February 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018302072.

Csuhaj-Varju:2010:SCG

[CVV10] Erzsébet Csuhaj-Varjú and György Vaszil. Scattered context grammars generate any recursively enumerable language with two nonterminals. Information Processing Letters, 110 (20):902–907, September 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cao:2011:CRC

[CVZ11] Zhenfu Cao, Ivan Visconti, and Zongyang Zhang. On constant-round concurrent non-malleable proof systems. Information Processing Letters, 111 (18):883–890, September 30, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011001694.

Choi:2012:LTF

[CW12] Seung Geol Choi and Hoeteck Wee. Lossy trapdoor functions from homomorphic reproducible encryption. Information Processing Letters, 112(20):794–798, October 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002001.

Chen:2013:NSL

[CW13] Danny Z. Chen and Haitao Wang. A note on searching line arrangements and applications. Information Processing Letters, 113(14–16):518–521, July/ August 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001221.

Chen:2015:IHS

[CW15] Jing Chen and Xiuping Wang. Identification of Hammerstein systems with continuous nonlinearity. Information Processing Letters, 115(11):822– 827, November 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015001003.

Chen:2010:RWS

[CWW10] Ing-Ray Chen, Yating Wang, and Ding-Chau Wang. Reliability of wireless sensors with code attestation for intrusion detection. Information Processing Letters, 110(17):778–786, August 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cen:2015:FCD

[CWY15] Yue-Feng Cen, Wan-Liang Wang, and Xin-Wei Yao. A fast CU depth decision mechanism for HEVC. Information Processing Letters, 115(9):719–724, September 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015000617.

Chang:2014:CIS

[CWYP14] Jou-Ming Chang, Jhen-Ding Wang, Jinn-Shyong Yang, and Kung-Jui Pai. A comment on “Independent spanning trees in crossed cubes”. Information Processing Letters, 114(12):734–739, December 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901400129X.

See [ZHX13].

Chen:2018:RRM

[CX18] Xiaoli Chen and Weijun Xu.

A risk-reward model with compound interest rate for non-additive two-option ski rental. Information Processing Letters, 135(??):9– 13, July 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300425.

Chen:2013:SOH

[CXD^{+}13] Xin Chen, Zhenzhen Xu, György Dósa, Xin Han, and He Jiang. Semionline hierarchical scheduling problems with buffer or rearrangements. Information Processing Letters, 113(4):127–131, February 28, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012003171.

Chang:2017:AUD

[CY17] Yi-Jun Chang and Hsu-Chun Yen. Area-universal drawings of biconnected outerplane graphs. Information Processing Letters, 118(??):1–5, February 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301302.

Chung:2018:FIP

[CY18] Byungjin Chung and Changhoon Yim. Fast intra prediction method by adaptive number of candidate modes for RDO in HEVC. Information Processing Letters, 131(??):20– 25, March 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017301990.

Chen:2015:CTO

[CYH15] Taolue Chen, Nengkun Yu, and Tingting Han. Continuous-time orbit problems are decidable in polynomial-time. Information Processing Letters, 115(1):11–14, January 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014001628.

Chen:2013:INS

[CYQ13] Yuhuan Chen, Chenfu Yi, and Dengyu Qiao. Improved neural solution for the Lyapunov matrix equation based on gradient search. Information Processing Letters, 113(22– 24):876–881, November/ December 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901300238X.

Cheng:2013:CCP

[CZ13] Qi Cheng and Jincheng Zhuang. On certain computations of Pisot numbers. Information Processing Letters, 113(8):271–275, April 30, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000367.

Cai:2016:FTI

[CZ16] Junqing Cai and Yuzhong Zhang. Fan-type implicitheavy subgraphs for hamiltonicity of implicit clawheavy graphs. Information Processing Letters, 116(11):668–673, November 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300965.

Czap:2013:NTC

[Cza13] Július Czap. A note on total colorings of 1-planar graphs. Information Processing Letters, 113(14–16):516–517, July/August 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001208.

Chen:2013:AAP

[CZCD13] Yong Chen, An Zhang, Guangting Chen, and Jianming Dong. Approximation algorithms for parallel open shop scheduling. Information Processing Letters, 113(7):220–224, April 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901300015X.

Chen:2014:DFB

[CZD14] Huibo Chen, Wenge Zhang, and Feng Ding. Data filtering based least squares iterative algorithm for parameter identification of output error autoregressive systems. Information Processing Letters, 114(10):573–578, October 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001021.

Chen:2017:NPE

[CZWP17] Hon-Chan Chen, Yun-Hao Zou, Yue-Li Wang, and Kung-Jui Pai. A note on path embedding in crossed cubes with faulty vertices. Information Processing Letters, 121(??):34– 38, May 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300121.

Ci:2010:DMB

[CZZ^{+}10] Yi-Wei Ci, Zhan Zhang, De-Cheng Zuo, Zhi-Bo Wu, and Xiao-Zong Yang. Dependency mining-based causal message logging. Information Processing Letters, 110(5):182–187, February 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ci:2012:MCC

[CZZ^{+}12] Yi-Wei Ci, Zhan Zhang, De-Cheng Zuo, Zhi-Bo Wu, and Xiao-Zong Yang. A multi-cycle checkpointing protocol that ensures strict 1-rollback. Information Processing Letters, 112(20):788–793, October 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001962.

Damaschke:2014:EMB

[Dam14] Peter Damaschke. Enumerating maximal bicliques in bipartite graphs with favorable degree sequences. Information Processing Letters, 114(6):317–321, June 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0020019014000192.

Damaschke:2016:SCE

[Dam16] Peter Damaschke. Sufficient conditions for edit-optimal clusters. Information Processing Letters, 116(4):267– 272, April 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901500215X.

Damaschke:2017:RAH

[Dam17]  Peter Damaschke.  Re-

fined algorithms for hitting many intervals. Information Processing Letters, 118(??):117–122, February 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016301521.

Darnstadt:2015:OPB

[Dar15]  Malte Darnstädt.  The

optimal PAC bound for intersection-closed concept classes. Information Processing Letters, 115(4):458– 461, April 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014002658.

Datta:2015:ASC

[Dat15] Alak Kumar Datta. Approximate spanning cactus. Information Processing Letters, 115(11):828– 832, November 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015001052.

Dubslaff:2012:MCP

[DBB12] Clemens Dubslaff, Christel Baier, and Manuela Berg. Model checking probabilistic systems against pushdown specifications. Information Processing Letters, 112(8–9):320–328, April 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000233.

Diaz-Banez:2017:NRC

[DBFMPLV17] J. M. Díaz-Báñez, R. Fabila-Monroy, P. Pérez-Lantero, and I. Ventura. New results on the coarseness of bicolored point sets. Information Processing Letters, 123(??): 1–7, July 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300340.

Du:2012:LCB

[DCH12] Xiaoni Du, Zhixiong Chen, and Lei Hu. Linear complexity of binary sequences derived from Euler quotients with prime-power modulus. Information Processing Letters, 112(14–15):604–609, August 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012001226.

DasGupta:2014:CBS

[DD14] Bhaskar DasGupta and Devendra Desai. On a connection between small set expansions and modularity clustering. Information Processing Letters, 114(7):349– 352, July 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014000222.

Das:2018:NAP

[DDER18] Bireswar Das, Anirban Dasgupta, Murali Krishna Enduri, and I. Vinod Reddy. On NC algorithms for problems on bounded rank-width graphs. Information Processing Letters, 139(??):60– 63, November 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019018301546.

Das:2015:AAM

[DDK^{+}15] Gautam K. Das, Minati De, Sudeshna Kolay, Subhas C. Nandy, and Susmita Sur-Kolay. Approximation algorithms for maximum independent set of a unit disk graph. Information Processing Letters, 115(3):439– 446, March 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901400235X.

Dolev:2011:SDL

[DDPBT11] Shlomi Dolev, Swan Dubois, Maria Potop-Butucaru, and Sébastien Tixeuil. Stabilizing data-link over non-FIFO channels with optimal fault-resilience. Information Processing Letters, 111(18):912–920, September 30, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001815.

Demange:2014:ERE

[DE14] Marc Demange and Tinaz Ekim. Efficient recognition of equimatchable graphs. Information Processing Letters, 114(1–2):66–71, January/February 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013002184.

Dosa:2019:NLB

[DE19] György Dósa and Leah Epstein. A new lower bound on the price of anarchy of selfish bin packing. Information Processing Letters, 150(??): 6–12, October 2019. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0020019019301012.

Dankelmann:2012:ECC

[DEG^{+}12] Peter Dankelmann, David Erwin, Wayne Goddard, Simon Mukwembi, and Henda C. Swart. Eccentric counts, connectivity and chordality. Information Processing Letters, 112(24):944–947, December 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002499.

Dolev:2019:UBM

[DEH^{+}19] Shlomi Dolev, Anat Eyal, Danny Hendler, Philip Derbeko, and Marina Kogan-Sadetsky. Upper bounds for multi-level multi-server paging. Information Processing Letters, 142(??):72– 76, February 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018302060.

Didimo:2010:CCB

[DEL10] Walter Didimo, Peter Eades, and Giuseppe Liotta. A characterization of complete bipartite RAC graphs. Information Processing Letters, 110(16):687–691, July 31, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Demba:2018:ECT

[Dem18] Moussa Demba. Equivalence checking of two functional programs using inductive theorem provers. Information Processing Letters, 129(??):16–24, January 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301448.

Deng:2014:EDS

[Den14] Yun-Ping Deng. Efficient dominating sets in circulant graphs with domination number prime. Information Processing Letters, 114(12):700–702, December 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001227.

Deng:2019:SPA

[Den19] Yun-Ping Deng. Some properties of alternating group networks. Information Processing Letters, 150(??):18– 21, October 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0020019019301036.

Deorowicz:2012:QTA

[Deo12] Sebastian Deorowicz. Quadratictime algorithm for a string constrained LCS problem. Information Processing Letters, 112(11):423–426, June 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000567.

Doerr:2011:QRR

[DF11] Benjamin Doerr and Mahmoud Fouz. Quasi-random rumor spreading: Reducing randomness can be costly. Information Processing Letters, 111(5):227–230, February 1, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dantas:2016:UPP

[DFdFT16] Simone Dantas, Luerbio Faria, Celina M. H. de Figueiredo, and Rafael B. Teixeira. The (k, l) unpartitioned probe problem NP-complete versus polynomial dichotomy. Information Processing Letters, 116(4):294–298, April 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002001901500188X.

Das:2010:NCN

[DFMS10] Anita Das, Mathew C.

Francis, Rogers Mathew, and N. Sadagopan. Noncontractible non-edges in 2connected graphs. Information Processing Letters, 110(23):1044–1048, November 15, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dudek:2013:ACR

[DFRS13] Andrzej Dudek, Alan Frieze, Andrzej Ruciński, and Matas Sileikis. Approximate counting of regular hypergraphs. Information Processing Letters, 113 (19–21):785–788, September/October 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901300207X.

Deorowicz:2014:EAL

[DG14] Sebastian Deorowicz and Szymon Grabowski. Efficient algorithms for the longest common subsequence in k-length substrings. Information Processing Letters, 114(11):

634–638, November 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0020019014001045.

Das:2016:LSA

[DG16] Ananda Swarup Das and Prosenjit Gupta. Linear space adaptive data structures for planar range reporting. Information Processing Letters, 116(5):361– 366, May 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016000028.

Daykin:2019:EPM

[DGG^{+}19] J. W. Daykin, R. Groult, Y. Guesnet, T. Lecroq, A. Lefebvre, M. Léonard, L. Mouchard, É. Prieur-Gaston, and B. Watson. Efficient pattern matching in degenerate strings with the Burrows–Wheeler transform. Information Processing Letters, 147(??):82–87, July 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019300535.

Demaine:2017:ASR

[DGK^{+}17] Erik D. Demaine, Varun Ganesan, Vladislav Kontsevoi, Qipeng Liu, Quanquan Liu, Fermi Ma, Ofir Nachum, Aaron Sidford, Erik Waingarten, and Daniel Ziegler. Arboral satisfaction: Recognition and LP approximation. Information Processing Letters, 127(??):1–5, November 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300716.

Das:2014:RIM

[DGKS14] Ananda Swarup Das, Prosenjit Gupta, Kishore Kothapalli, and Kannan Srinathan. On reporting the L_{1} metric closest pair in a query rectangle. Information Processing Letters, 114(5):256– 263, May 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013003098.

DiIanni:2015:RDU

[DGR15] Miriam Di Ianni, Luciano Gualà, and Gianluca Rossi. Reducing the diameter of a unit disk graph via node addition. Information Processing Letters, 115(11):845– 850, November 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019015001118.

Debski:2015:SCI

[DGŚN15] Michał Debski, Jaroslaw Grytczuk, and Małgorzata Śleszyńska-Nowak. The strong chromatic index of sparse graphs. Information Processing Letters, 115(2):326–330, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901400221X.

Dhaliwal:2014:FSE

[Dha14] Jasbir Dhaliwal. Faster semi-external suffix sorting. Information Processing Letters, 114(4):174–178, April 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013002974.

Durian:2010:IPE

[ĎHPT10] Branislav Ďurian, Jan Holub, Hannu Peltola, and Jorma Tarhio. Improving practical exact string matching. Information Processing Letters, 110(4):148– 152, January 16, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dutta:2013:PTB

[DHR13] Amit Kumar Dutta, Masud Hasan, and M. Sohel Rahman. Prefix transpositions on binary and ternary strings. Information Processing Letters, 113(8):265– 270, April 30, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901300029X.

Dutting:2011:OFA

[DHW11] Paul Dütting, Monika Henzinger, and Ingmar Weber. Offline file assignments for online load balancing. Information Processing Letters, 111(4):178–183, January 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dutting:2013:SSM

[DHW13] Paul Dütting, Monika Henzinger, and Ingmar Weber. Sponsored search, market equilibria, and the Hungarian Method. Information Processing Letters, 113(3):67–73, February 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012003031.

Diaconescu:2015:ETS

[Dia15] Razvan Diaconescu. On the existence of translations of structured specifications. Information Processing Letters, 115(1):15– 22, January 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014001690.

Didimo:2013:DSL

[Did13] Walter Didimo. Density of straight-line 1-planar graph drawings. Information Processing Letters, 113(7):236– 240, April 15, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019013000252.

Dietzfelbinger:2015:TSC

[DJ15] Martin Dietzfelbinger and Raed Jaberi. On testing single connectedness in directed graphs and some related problems. Information Processing Letters, 115(9):684–688, September 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901500068X.

Dugar:2015:DPC

[DJRB15] Chetan Dugar, Anuj Jain, Amit Rajawat, and Swapan Bhattacharya. Dynamic pricing of call rates: Bayesian approach. Information Processing Letters, 115(2):237–242, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001975.

Das:2013:LRE

[DJS13] Sudipta Das, Lawrence Jenkins, and Debasis Sengupta. Loss ratio of the EDF scheduling policy with early discarding technique. Information Processing Letters, 113(5–6):165–170, March 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000082.

Dong:2015:AAP

[DJZ^{+}15] Jianming Dong, Yiwei Jiang, An Zhang, Jueliang Hu, and Hui Luo. An approximation algorithm for proportionate scheduling in the two-stage hybrid flow shop. Information Processing Letters, 115(4):475–480, April 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002646.

Dunkelman:2010:EOL

[DK10] Orr Dunkelman and Nathan Keller. The effects of the omission of last round’s MixColumns on AES. Information Processing Letters, 110(8–9):304–308, April 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Duris:2014:FPA

[DK14] Pavol Duris and Marek Kosta. Flip-pushdown automata with k pushdown reversals and E0L systems are incomparable. Information Processing Letters, 114(8): 417–420, August 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019014000404.

Du:2012:LCP

[DKC12] Xiaoni Du, Andrew Klapper, and Zhixiong Chen. Linear complexity of pseudorandom sequences generated by Fermat quotients and their generalizations. Information Processing Letters, 112(6):233–237, March 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901100322X.

Demenkov:2010:NUB

[DKKY10] E. Demenkov, A. Kojevnikov, A. Kulikov, and G. Yaroslavtsev. New upper bounds on the Boolean circuit complexity of symmetric functions. Information Processing Letters, 110 (7):264–267, March 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Didimo:2012:VAC

[DKL^{+}12] Walter Didimo, Michael Kaufmann, Giuseppe Liotta, Yoshio Okamoto, and Andreas Spillner. Vertex angle and crossing angle resolution of leveled tree drawings. Information Processing Letters, 112(16):630– 635, August 31, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019012001299.

Dolgui:2018:GPS

[DKNQ18] Alexandre Dolgui, Vladimir Kotov, Aliaksandr Nekrashevich, and Alain Quilliot. General parametric scheme for the online uniform machine scheduling problem with two different speeds. Information Processing Letters, 134(??):18– 23, June 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300188.

Dziembowski:2018:QCR

[DKZ18] Stefan Dziembowski, Tomasz Kazana, and Maciej Zdanowicz. Quasi chain rule for min-entropy. Information Processing Letters, 134(??): 62–66, June 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901830036X.

Dai:2013:RLS

[DL13] Benqiu Dai and Wensong Lin. On (s, t)- relaxed L(2, 1)-labelings of the square lattice. Information Processing Letters, 113 (19–21):704–709, September/October 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001798.

DeFrancesco:2010:PMC

[DLMV10] N. De Francesco, G. Lettieri, L. Martini, and G. Vaglini. Partial model checking via abstract interpretation. Information Processing Letters, 110(3):99–103, January 1, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dobson:2011:MDL

[DLN11] M. P. Dobson, V. Leoni, and G. Nasini. The multiple domination and limited packing problems in graphs. Information Processing Letters, 111(23–24):1108–1113, December 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011002377.

Dorrigiv:2012:LUP

[DLO12] Reza Dorrigiv and Alejandro López-Ortiz. List update with probabilistic locality of reference. Information Processing Letters, 112(13):540–543, July 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001044.

Dabrowski:2014:SNF

[DLRS14] Przemyslaw Dabrowski, Grzegorz Labuzek, Tomasz Rachwalik, and Janusz Szmidt. Searching for nonlinear feedback shift registers with parallel computing. Information Processing Letters, 114(5):268–272, May 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013003050.

DiLena:2010:OGA

[DM10] Pietro Di Lena and Luciano Margara. Optimal global alignment of signals by maximization of Pearson correlation. Information Processing Letters, 110(16):679– 686, July 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dean:2013:BCT

[DM13] Brian C. Dean and Raghuveer Mohan. Building Cartesian trees from free trees with k leaves. Information Processing Letters, 113(9):345–349, May 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013000689.

Darwish:2016:IBF

[DM16] Omar Darwish and Kurt Mehlhorn. Improved balanced flow computation using parametric flow. Information Processing Letters, 116(9):560–563, September 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016300564.

Dyer:2019:CIS

[DM19] Martin Dyer and Haiko Müller. Counting independent sets in cocomparability graphs. Information Processing Letters, 144(??):31– 36, April 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018302370.

Duarte:2011:ICS

[DMS11] Rafael Duarte, Alexandre Mota, and Augusto Sampaio. Introducing concurrency in sequential Java via laws. Information Processing Letters, 111(3):129–134, January 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dietrich:2012:BMS

[DMS12] Max Dietrich, Catherine McCartin, and Charles Semple. Bounding the maximum size of a minimal definitive set of quartets. Information Processing Letters, 112(16):651–655, August 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001408.

Dwivedi:2018:DLR

[DMSD18] Ashutosh Dhar Dwivedi, Pawel Morawiecki, Rajani Singh, and Shalini Dhar. Differential-linear and related key cryptanalysis of round-reduced scream. Information Processing Letters, 136(??):5– 8, August 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300656.

Dybizbanski:2012:OCN

[DN12] Janusz Dybizbański and Anna Nenca. Oriented chromatic number of grids is greater than 7. Information Processing Letters, 112(4):113–117, February 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011003000.

Doerr:2013:IAA

[Doe13] Benjamin Doerr. Improved approximation algorithms for the Min-Max Selecting Items problem. Information Processing Letters, 113 (19–21):747–749, September/October 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013002007.

Dieudonne:2012:DNE

[DP12] Yoann Dieudonné and Andrzej Pelc. Deterministic network exploration by a single agent with Byzantine tokens. Information Processing Letters, 112(12): 467–470, June 30, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019012000853.

Dabrowski:2017:CBG

[DP17] Konrad K. Dabrowski and Daniël Paulusma. Contracting bipartite graphs to paths and cycles. Information Processing Letters, 127(??): 37–42, November 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019017301229.

Dabrowski:2018:CFF

[DP18] Konrad K. Dabrowski and Daniël Paulusma. On colouring (2P_{2}, H)-free and (P_{5}, H)-free graphs. Information Processing Letters, 134(??):35–41, June 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019018300322.

David:2011:HSN

[DPS11] Matei David, Periklis A. Papakonstantinou, and Anastasios Sidiropoulos. How strong is Nisan’s pseudorandom generator? Information Processing Letters, 111(16):804–808, August 30, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011001190.

See [Nis92].

Draziotis:2016:EDL

[Dra16] Konstantinos A. Draziotis. (EC)DSA lattice attacks based on Coppersmith’s method. Information Processing Letters, 116(8):541– 545, August 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016300436.

Dixit:2011:SRP

[DRD11] Manoj G. Dixit, S. Ramesh, and Pallab Dasgupta. Some results on Parametric Temporal Logic. Information Processing Letters, 111(20):994–998, October 31, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011001888.

Dourado:2012:CRR

[DRdSS12] Mitre C. Dourado, Dieter Rautenbach, Vinícius Fernandes dos Santos, and Jayme L. Szwarcfiter. Characterization and recognition of Radon-independent sets in split graphs. Information Processing Letters, 112(24):948–952, December 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002323.

Dybizbanski:2014:OCN

[DS14] Janusz Dybizbański and Andrzej Szepietowski. The oriented chromatic number of Halin graphs. Information Processing Letters, 114(1–2):45–49, January/ February 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002482.

Dehghan:2015:CZS

[DS15] Ali Dehghan and Mohammad-Reza Sadeghi. The complexity of the zero-sum 3flows. Information Processing Letters, 115(2):316–320, February 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002191.

Dehghan:2016:ACZ

[DS16] Ali Dehghan and Mohammad-Reza Sadeghi. On the algorithmic complexity of zerosum edge-coloring. Information Processing Letters, 116(11):660–667, November 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016300941.

Delgado-Segura:2018:BPK

[DSPSHJNA18] Sergi Delgado-Segura, Cristina Pérez-Solà, Jordi Herrera-Joancomartí, and Guillermo Navarro-Arribas. Bitcoin private key locked transactions. Information Processing Letters, 140(??):34–36, December 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018301698.

Diochnos:2012:MIL

[DST12] D. I. Diochnos, R. H. Sloan, and Gy. Turán. On multiple-instance learning of halfspaces. Information Processing Letters, 112(23):933–936, December 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901200244X.

Dorn:2010:NSE

[DT10] Sebastian Dörn and Thomas Thierauf. A note on the search for k elements via quantum walk. Information Processing Letters, 110(22):

975–978, October 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Doron:2015:RSB

[DTS15] Dean Doron and Amnon Ta-Shma. On the derandomization of spacebounded approximate counting problems. Information Processing Letters, 115(10): 750–753, October 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015000435.

Ducoffe:2018:SNC

[Duc18] Guillaume Ducoffe. A short note on the complexity of computing strong pathbreadth. Information Processing Letters, 133(??):56– 58, May 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300140.

Duris:2012:SCA

[Dur12] David Duris. Some characterizations of γ and βacyclicity of hypergraphs. Information Processing Letters, 112(16):617–620, August 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012001287.

Duraj:2013:LAL

[Dur13] Lech Duraj. A linear algorithm for 3-letter longest common weakly increasing subsequence. Information Processing Letters, 113(3):94–99, February 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012003080.

Delic:2010:USD

[DW10a] Dejan Delić and Changping Wang. Upper signed k-domination in a general graph. Information Processing Letters, 110(16):662– 665, July 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dvorak:2010:CSS

[DW10b] Wolfgang Dvořák and Stefan Woltran. Complexity of semi-stable and stage semantics in argumentation frameworks. Information Processing Letters, 110(11): 425–430, May 16, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Doerr:2012:MRB

[DW12] Benjamin Doerr and Carola Winzen. Memoryrestricted black-box complexity of OneMax. Information Processing Letters, 112(1–2):32–34, January 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011002754.

Ding:2010:VWW

[DWQ10] Guiguang Ding, Jianmin Wang, and Kai Qin. A visual word weighting scheme based on emerging itemsets for video annotation. Information Processing Letters, 110(16):692–696, July 31, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dong:2010:SRA

[DX10] Wei Dong and Baogang Xu. Some results on acyclic edge coloring of plane graphs. Information Processing Letters, 110 (20):887–892, September 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Du:2011:NCE

[DX11] Zheng-Zhong Du and Jun-Ming Xu. A note on cycle embedding in hypercubes with faulty vertices. Information Processing Letters, 111(12):557–560, June 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dong:2010:ELF

[DY10] Qiang Dong and Xiaofan Yang. Embedding a long fault-free cycle in a crossed cube with more faulty nodes. Information Processing Letters, 110(11): 464–468, May 16, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Drineas:2011:NEW

[DZ11] Petros Drineas and Anastasios Zouzias. A note on element-wise matrix sparsification via a matrix-valued Bernstein inequality. Information Processing Letters, 111(8):385–389, March 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Deng:2012:AGP

[DZ12] Yun-Ping Deng and Xiao-Dong Zhang. Automorphism groups of the Pancake graphs. Information Processing Letters, 112(7):264– 266, March 31, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019011003383.

Dong:2015:HGH

[DZA15] Qiang Dong, Qian Zhao, and Yahui An. The hamiltonicity of generalized honeycomb torus networks. Information Processing Letters, 115(2):104–111, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001495.

Dong:2012:EMT

[DZFY12] Qiang Dong, Junlin Zhou, Yan Fu, and Xiaofan Yang. Embedding a mesh of trees in the crossed cube. Information Processing Letters, 112(14–15):599–603, August 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001251.

Dong:2013:NVB

[DZQF13] Deshuai Dong, Xue Zhang, Longjiang Qu, and Shaojing Fu. A note on vectorial bent functions. Information Processing Letters, 113 (22–24):866–870, November/December 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013002135.

Ehdaie:2016:HCR

[EAA^{+}16] Mohammad Ehdaie, Nikos Alexiou, Mahmoud Ahmadian, Mohammad Reza Aref, and Panos Papadimitratos. 2D hash chain robust random key distribution scheme. Information Processing Letters, 116(5): 367–372, May 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015002239.

ElAtik:2021:CPM

[EAE21] Abd El Fattah A. El Atik, A. W. Aboutahoun, and A. Elsaid. Correct proof of the main result in “The number of spanning trees of a class of self-similar fractal models” by Ma and Yao. Information Processing Letters, 170(??):Article 106117, September 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019021000314.

Esiner:2017:QRI

[ED17] Ertem Esiner and Anwitaman Datta. On query result integrity over encrypted data. Information Processing Letters, 122(??):34–39, June 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300327.

Even-Dar:2011:NMS

[EDS11] Eyal Even-Dar and Asaf Shapira. A note on maximizing the spread of influence in social networks. Information Processing Letters, 111(4):184–187, January 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Emek:2010:ACS

[EFKR10] Yuval Emek, Pierre Fraigniaud, Amos Korman, and Adi Rosén. On the additive constant of the kserver Work Function Algorithm. Information Processing Letters, 110(24):1120– 1123, November 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Eftekhari:2010:BWK

[EFMA10] Armin Eftekhari, Mohamad Forouzanfar, Hamid Abrishami Moghaddam, and Javad Alirezaie. Blockwise 2D kernel PCA/LDA for face recognition. Information Processing Letters, 110(17):761–766, August 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ehsani:2012:OOP

[EGK^{+}12] Shayan Ehsani, Mohammad Ghodsi, Ahmad Khajenezhad, Hamid Mahini, and Afshin Nikzad. Optimal online pricing with network externalities. Information Processing Letters, 112(4):118–123, February 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011003061.

Esparza:2013:SPA

[EGK13] Javier Esparza, Andreas Gaiser, and Stefan Kiefer. A strongly polynomial algorithm for criticality of branching processes and consistency of stochastic context-free grammars. Information Processing Letters, 113(10–11):381–385, May/June 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000690.

Esparza:2011:PTS

[EGKL11] Javier Esparza, Pierre Ganty, Stefan Kiefer, and Michael Luttenberger. Parikh’s theorem: a simple and direct automaton construction. Information Processing Letters, 111(12):614– 619, June 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011000822.

Eder:2018:CPW

[EH18] Günther Eder and Martin Held. Computing positively weighted straight skeletons of simple polygons based on a bisector arrangement. Information Processing Letters, 132(??):28– 32, April 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017302156.

Ehlers:2017:MAS

[Ehl17] Thorsten Ehlers. Merging almost sorted sequences yields a 24-sorter. Information Processing Letters, 118(??):17–20, February 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301193.

Even:2011:AAA

[EKN11] Guy Even, Guy Kortsarz, and Zeev Nutov. A 1.5-approximation algorithm for augmenting edgeconnectivity of a graph from 1 to 2. Information Processing Letters, 111(6):296–300, February 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Epstein:2010:RAO

[EL10] Leah Epstein and Asaf Levin. Randomized algorithms for online bounded bidding. Information Processing Letters, 110(12–13): 503–506, June 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Elmasry:2010:LAI

[Elm10] Amr Elmasry. The longest almost-increasing subsequence. Information Processing Letters, 110(16): 655–658, July 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Elmasry:2012:SSP

[Elm12] Amr Elmasry. On the size of the subset partial order. Information Processing Letters, 112(12):487– 489, June 30, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012000737.

Elmasry:2015:CIA

[Elm15] Amr Elmasry. Counting inversions adaptively. Information Processing Letters, 115(10):769–772, October 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000733.

Eisenbarth:2015:SMM

[EMS15] T. Eisenbarth, A. Meyerowitz, and R. Steinwandt. On the security margin of MAC striping. Information Processing Letters, 115(11):899–902, November 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000861.

Elbassioni:2017:PTA

[EN17] Khaled Elbassioni and Trung Thanh Nguyen. A polynomial-time algorithm for computing low CPrank decompositions. Information Processing Letters, 118(??):10–14, February 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301314.

Elkind:2013:HFS

[EO13] Edith Elkind and James B.

Orlin. On the hardness of finding subsets with equal average. Information Processing Letters, 113(13): 477–480, July 15, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001901300104X.

Echenim:2011:MIS

[EP11] Mnacho Echenim and Nicolas Peltier. Modular instantiation schemes. Information Processing Letters, 111(20):989–993, October 31, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901100189X.

Edemskiy:2016:LCB

[EP16] V. Edemskiy and A. Palvinskiy. The linear complexity of binary sequences of length 2p with optimal three-level autocorrelation. Information Processing Letters, 116(2):153–156, February 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001623.

Esik:2011:UPC

[Ési11] Z. Ésik. An undecidable property of context-free linear orders. Information Processing Letters, 111(3): 107–109, January 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kundi:2010:REI

[eSKAI10] Dur e Shahwar Kundi, Arshad Aziz, and Nasar Ikram. Resource efficient implementation of T-Boxes in AES on Virtex-5 FPGA. Information Processing Letters, 110(10):373–377, April 30, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Enos:2015:IBS

[EZ15] Graham Enos and Yuliang Zheng. An ID-based signcryption scheme with compartmented secret sharing for unsigncryption. Information Processing Letters, 115(2):128–133, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001537.

Epstein:2016:OSU

[EZH16] Leah Epstein and Hanan Zebedat-Haider. Online scheduling of unit jobs on three machines with rejection: a tight result. Information Processing Letters, 116(3):252–255, March 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015002069.

Ezra:2010:NAW

[Ezr10] Esther Ezra. A note about weak ϵ-nets for axis-parallel boxes in d-space. Information Processing Letters, 110 (18–19):835–840, September 15, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Frieze:2017:RCS

[FA17] Alan Frieze and Michael Anastos. Randomly coloring simple hypergraphs with fewer colors. Information Processing Letters, 126(??): 39–42, October 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019017301060.

Farrag:2010:DFT

[Far10] A. A. Farrag. Developing fault-tolerant distributed loops. Information Processing Letters, 111(2):97– 101, December 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fay:2016:ICM

[Fay16]  Robin Fay.  Introduc-

ing the counter mode of operation to Compressed Sensing based encryption. Information Processing Letters, 116(4):279– 283, April 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001945.

Freschi:2010:FAC

[FB10] Valerio Freschi and Alessandro Bogliolo. A faster algorithm for the computation of string convolutions using LZ78 parsing. Information Processing Letters, 110(14– 15):609–613, July 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fan:2018:NDT

[FBD18] Rui Fan, Mohammud Junaid Bocus, and Naim Dahnoun. A novel disparity transformation algorithm for road segmentation. Information Processing Letters, 140(??):13–17, December 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301583.

Fang:2014:OSC

[FC14] Chengfang Fang and Ee-Chien Chang. Optimal strategy of coupon subset collection when each package contains half of the coupons. Information Processing Letters, 114(12):703–705, December 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014001264.

Fernandez-Campusano:2017:DLE

[FCLCR17] Christian Fernández-Campusano, Mikel Larrea, Roberto Cortiñas, and Michel Raynal. A distributed leader election algorithm in crashrecovery and omissive systems. Information Processing Letters, 118(??):100– 104, February 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001901630148X.

Fu:2010:OST

[FCNY10] Ruyan Fu, T. C. E. Cheng, C. T. Ng, and Jinjiang Yuan. Online scheduling on two parallel-batching machines with limited restarts to minimize the makespan. Information Processing Letters, 110(11):444–450, May 16, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Foster:2018:UTT

[FCWZ18] Simon Foster, Ana Cavalcanti, Jim Woodcock, and Frank Zeyda. Unifying theories of time with generalised reactive processes. Information Processing Letters, 135 (??):47–52, July 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019018300462.

Feghali:2019:PBC

[Feg19] Carl Feghali. Paths between colourings of graphs with bounded tree-width. Information Processing Letters, 144(??):37–38, April 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019018302382.

Fredriksson:2014:CES

[FG14] Kimmo Fredriksson and Emanuele Giaquinta. On a compact encoding of the swap automaton. Information Processing Letters, 114 (7):392–396, July 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014000052.

Fan:2018:ALV

[FG18a] Hua Fan and Wojciech Golab. Analyzing linearizability violations in the presence of read–modify–write operations. Information Processing Letters, 138(??):51– 56, October 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018301297.

Frohn:2018:CRC

[FG18b] Florian Frohn and Jürgen Giesl. Constant runtime complexity of term rewriting is semi-decidable. Information Processing Letters, 139(??):13–17, November 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018301376.

Fomin:2010:PAE

[FGG^{+}10] Fedor V. Fomin, Serge Gaspers, Petr A. Golovach, Dieter Kratsch, and Saket Saurabh. Parameterized algorithm for eternal vertex cover. Information Processing Letters, 110(16):702– 706, July 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Flouri:2015:LCS

[FGKU15] Tomas Flouri, Emanuele Giaquinta, Kassian Kobert, and Esko Ukkonen. Longest common substrings with k mismatches. Information Processing Letters, 115(6–8):643–647, June/ August 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000459.

Farras:2012:LTM

[FGMP12] Oriol Farràs, Ignacio Gracia, Sebastià Martín, and Carles Padró. Linear threshold multisecret sharing schemes. Information Processing Letters, 112 (17–18):667–673, September 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001378.

Fomin:2011:SBD

[FGvL11] Fedor V. Fomin, Petr A.

Golovach, and Erik Jan van Leeuwen. Spanners of bounded degree graphs. Information Processing Letters, 111(3):142–144, January 1, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fischer:2010:NCA

[FH10] Johannes Fischer and Daniel H.

Huson. New common ancestor problems in trees and directed acyclic graphs. Information Processing Letters, 110(8–9):331–335, April 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fiedorowicz:2012:ACI

[FH12] Anna Fiedorowicz and Mariusz Haluszczak. Acyclic chromatic indices of fully subdivided graphs. Information Processing Letters, 112(13):557–561, July 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001068.

Fernau:2015:CMD

[FHvtH^{+}15] Henning Fernau, Pinar Heggernes, Pim van ’t Hof, Daniel Meister, and Reza Saei. Computing the metric dimension for chain graphs. Information Processing Letters, 115(9):671– 676, September 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015000666.

Fiedorowicz:2011:AEC

[Fie11] Anna Fiedorowicz. Acyclic edge colourings of graphs with the number of edges linearly bounded by the number of vertices. Information Processing Letters, 111(6):287–290, February 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Filmus:2011:LBC

[Fil11]  Yuval Filmus.  Lower

bounds for context-free grammars. Information Processing Letters, 111 (18):895–898, September 30, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001712.

Filmus:2013:ISF

[Fil13] Yuval Filmus. Inequalities on submodular functions via term rewriting. Information Processing Letters, 113(13): 457–464, July 15, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013001038.

Filtser:2018:UAS

[Fil18] Omrit Filtser. Universal approximate simplification under the discrete Fréchet distance. Information Processing Letters, 132(??):22– 27, April 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301801.

Finkel:2015:ECI

[Fin15] Olivier Finkel. The exact complexity of the infinite Post Correspondence Problem. Information Processing Letters, 115(6–8):609–611, June/August 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000320.

Flocchini:2011:HMO

[FIPS11] Paola Flocchini, David Ilcinkas, Andrzej Pelc, and Nicola Santoro. How many oblivious robots can explore a line. Information Processing Letters, 111(20):1027– 1031, October 31, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011002109.

Fischer:2010:WL

[Fis10]  Johannes Fischer.  Wee

LCP. Information Processing Letters, 110(8–9):317– 320, April 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fournier:2015:CGH

[FIV15] Hervé Fournier, Anas Ismail, and Antoine Vigneron. Computing the Gromov hyperbolicity of a discrete metric space. Information Processing Letters, 115(6–8):576–579, June/ August 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000198.

Fakcharoenphol:2010:SPO

[FK10] Jittat Fakcharoenphol and Boonserm Kijsirikul. Short proofs for online multiclass prediction on graphs. Information Processing Letters, 110(8–9):309–311, April 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Folwarczny:2017:IMS

[FK17] Lukás Folwarczný and Dusan Knop. IV-matching is strongly NP-hard. Information Processing Letters, 125(??):5–8, September 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901730073X.

Farzanyar:2013:PFM

[FKC13] Zahra Farzanyar, Mohammadreza Kangavari, and Nick Cercone. P2P-FISM: Mining (recently) frequent item sets from distributed data streams over P2P network. Information Processing Letters, 113(19– 21):793–798, September/ October 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002056.

Fan:2011:NRE

[FKL^{+}11] Hao Fan, H. A. Kierstead, Guizhen Liu, Theodore Molla, Jian-Liang Wu, and Xin Zhang. A note on relaxed equitable coloring of graphs. Information Processing Letters, 111 (21–22):1062–1066, November 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011002225.

Fici:2016:GAS

[FKR^{+}16] Gabriele Fici, Tomasz Kociumaka, Jakub Radoszewski, Wojciech Rytter, and Tomasz Waleń. On the greedy algorithm for the Shortest Common Superstring problem with reversals. Information Processing Letters, 116(3): 245–251, March 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015002094.

Feng:2012:PCM

[FL12] Yun Feng and Wensong Lin. A proof of a conjecture on multiset coloring the powers of cycles. Information Processing Letters, 112 (17–18):678–682, September 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001445.

Feng:2013:CPT

[FL13a] Yun Feng and Wensong Lin. A concise proof for total coloring subcubic graphs. Information Processing Letters, 113 (18):664–665, September 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001609.

Fujita:2013:REG

[FL13b] Shinya Fujita and Linda Lesniak. Revisit of Erdős– Gallai’s theorem on the circumference of a graph. Information Processing Letters, 113(17):646–648, August 30, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001579.

Fang:2016:OPB

[FL16] Yang Fang and Xiwen Lu. Online parallel-batch scheduling to minimize total weighted completion time on single unbounded machine. Information Processing Letters, 116(8):526– 531, August 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016300345.

Fu:2010:EBS

[FLMQ10] Shaojing Fu, Chao Li, Kanta Matsuura, and Longjiang Qu. Enumeration of balanced symmetric functions over GF(p). Information Processing Letters, 110(14– 15):544–548, July 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fomin:2018:LDP

[FLP^{+}18] Fedor V. Fomin, Daniel Lokshtanov, Fahad Panolan, Saket Saurabh, and Meirav Zehavi. Long directed (s, t)- path: FPT algorithm. Information Processing Letters, 140(??):4–7, December 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901830111X.

Floderus:2015:DMD

[FLPS15] Peter Floderus, Andrzej Lingas, Mia Persson, and Dzmitry Sledneu. Detecting monomials with k distinct variables. Information Processing Letters, 115(2):82– 86, February 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014001392.

Fomin:2011:SAP

[FLRS11] Fedor V. Fomin, Daniel Lokshtanov, Venkatesh Raman, and Saket Saurabh. Subexponential algorithms for partial cover problems. Information Processing Letters, 111(16):814–818, August 30, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901100144X.

Frieze:2011:RCS

[FM11] Alan Frieze and Páll Melsted. Randomly coloring simple hypergraphs. Information Processing Letters, 111(17):848–853, September 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901100158X.

Fiszman:2018:MTL

[FM18] Shir Fiszman and Gur Mosheiov. Minimizing total load on a proportionate flowshop with positiondependent processing times and job-rejection. Information Processing Letters, 132 (??):39–43, April 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019017302181.

See comments [KMS19].

Furuya:2019:NDE

[FM19] Michitaka Furuya and Naoki Matsumoto. A note on domination 3-edge-critical planar graphs. Information Processing Letters, 142(??): 64–67, February 2019. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019018302084.

Fabila-Monroy:2011:CPA

[FMHL11] Ruy Fabila-Monroy, Clemens Huemer, and Dolores Lara. A combinatorial property on angular orders of plane point sets. Information Processing Letters, 111(12): 591–594, June 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001901100072X.

Filho:2018:CLL

[FOSC18] M. S. Conserva Filho, M. V. M. Oliveira, A. Sampaio, and Ana Cavalcanti. Compositional and local livelock analysis for CSP. Information Processing Letters, 133(??):21–25, May 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300036.

Frieze:2018:BAT

[FP18] Alan Frieze and Samantha Petti. Balanced allocation through random walk. Information Processing Letters, 131(??):39–43, March 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017302041.

Fung:2012:LSE

[FPY12] Stanley P. Y. Fung, Chung Keung Poon, and Duncan K. W. Yung. On-line scheduling of equal-length intervals on parallel machines. Information Processing Letters, 112(10):376– 379, May 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000415.

Farhana:2012:DCL

[FR12] Effat Farhana and M. Sohel Rahman. Doublyconstrained LCS and Hybrid-Constrained LCS problems revisited. Information Processing Letters, 112(13): 562–565, July 15, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019012001093.

Frati:2010:NIP

[Fra10] Fabrizio Frati. A note on isosceles planar graph drawing. Information Processing Letters, 110(12–13):507– 509, June 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fredriksson:2010:BMA

[Fre10] Kimmo Fredriksson. On building minimal automaton for subset matching queries. Information Processing Letters, 110(24): 1093–1098, November 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fredman:2014:ISB

[Fre14] Michael L. Fredman. An intuitive and simple bounding argument for Quicksort. Information Processing Letters, 114(3):137–139, March 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013002676.

Feder:2011:MGL

[FS11] Tomás Feder and Carlos Subi. Maximum gap labelings of graphs. Information Processing Letters, 111(4): 169–173, January 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fujiwara:2012:IAS

[FS12] Hiroshi Fujiwara and Yoshiyuki Sekiguchi. An improved analysis of SRPT scheduling algorithm on the basis of functional optimization. Information Processing Letters, 112(23):911–915, December 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002281.

Feder:2013:ECA

[FS13a] Tomás Feder and Carlos Subi. Edge-coloring almost bipartite multigraphs. Information Processing Letters, 113(18): 685–689, September 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019013001737.

Fujito:2013:HGG

[FS13b] Toshihiro Fujito and Takayoshi Sakamaki. How to guard a graph against tree moves. Information Processing Letters, 113(19–21):844–847, September/October 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019013002147.

Finkel:2014:TCL

[FS14] Olivier Finkel and Michal Skrzypczak. On the topological complexity of ωlanguages of non-deterministic Petri nets. Information Processing Letters, 114(5):229– 233, May 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013003104.

Feldman:2015:CBR

[FT15] Michal Feldman and Tami Tamir. Convergence of bestresponse dynamics in games with conflicting congestion effects. Information Processing Letters, 115(2):112– 118, February 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014001501.

Fu:2014:OSU

[FTYL14] Ruyan Fu, Ji Tian, Jinjiang Yuan, and Ya Li. Online scheduling on an unbounded parallel-batch machine and a standard machine to minimize makespan. Information Processing Letters, 114(4):179–184, April 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019013002962.

Fu:2010:EFT

[Fu10] Jung-Sheng Fu. Edge-fault-tolerant vertex-pancyclicity of augmented cubes. Information Processing Letters, 110(11):439–443, May 16, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fujii:2016:FAA

[Fuj16] Kaito Fujii. Faster approximation algorithms for maximizing a monotone submodular function subject to a b-matching constraint. Information Processing Letters, 116(9):578– 584, September 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019016300461.

Furst:2019:HDE

[Für19] M. Fürst. On the hardness of deciding the equality of the induced and the uniquely restricted matching number. Information Processing Letters, 147(??): 77–81, July 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300687.

Fournier:2013:DAF

[FV13] Hervé Fournier and Antoine Vigneron. A deterministic algorithm for fitting a step function to a weighted point-set. Information Processing Letters, 113(3):51– 54, February 15, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019012003006.

Fan:2014:BPH

[FWH14] Chen-Ming Fan, Jen-Tse Wang, and Cheng-Chih Huang. Borderednesspreserving homomorphisms. Information Processing Letters, 114(10):529–534, October 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014000672.

Feng:2013:ECE

[FWS13] Jun Feng, Xueming Wang, and Hong Sun. Efficiently computable endomorphism for genus 3 hyperelliptic curve cryptosystems. Information Processing Letters, 113(12):405–408, June 30, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000914.

Fleischer:2011:CMD

[FX11] Rudolf Fleischer and Xiaoming Xu. Computing minimum diameter colorspanning sets is hard. Information Processing Letters, 111(21–22):1054–1056, November 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011002079.

Fu:2014:HAB

[FY14] Zhiguo Fu and Fengqin Yang. Holographic algorithms on bases of rank

2. Information Processing Letters, 114(11):585– 590, November 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014001203.

Freivalds:2010:NFN

[FYS10] Rūsiņš Freivalds, Abuzer Yakaryılmaz, and A. C. Cem Say. A new family of nonstochastic languages. Information Processing Letters, 110(10):410–413, April 30, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Feng:2013:TBS

[FZ13] Yuan Feng and Lijun Zhang.

A tighter bound for the self-stabilization time in Herman’s algorithm. Information Processing Letters, 113(13):486–488, July 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001178.

Galindo:2013:NIC

[Gal13] David Galindo. A note on an IND-CCA2 secure Paillier-based cryptosystem. Information Processing Letters, 113(22– 24):913–914, November/ December 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013002457.

Gavril:2011:AIB

[Gav11] Fanica Gavril. Algorithms for induced biclique optimization problems. Information Processing Letters, 111(10):469–473, April 30, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gavril:2017:NIG

[Gav17] Fanica Gavril. New insights on GA-H reduced graphs. Information Processing Letters, 128(??):58–61, December 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017301497.

Guttmann-Beck:2010:TRA

[GBH10] Nili Guttmann-Beck and Refael Hassin. On two restricted ancestors tree problems. Information Processing Letters, 110(14–15):570– 575, July 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guttmann-Beck:2012:SPO

[GBH12] Nili Guttmann-Beck and Refael Hassin. Seriesparallel orientations preserving the cycle-radius. Information Processing Letters, 112(4):153–160, February 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011003012.

Chen:2015:DDP

[gCF15] Xue gang Chen and Shinya Fujita. Downhill domination problem in graphs. Information Processing Letters, 115(6–8):580–581, June/ August 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000204.

Guo:2010:KNA

[GCH^{+}10] Deke Guo, Hanhua Chen, Yuan He, Hai Jin, Chao Chen, Honghui Chen, Zhen Shu, and Guangqi Huang. KCube: a novel architecture for interconnection networks. Information Processing Letters, 110(18–19):821– 825, September 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gulavani:2010:RAI

[GCNR10] Bhargav S. Gulavani, Supratik Chakraborty, Aditya V. Nori, and Sriram K. Rajamani. Refining abstract interpretations. Information Processing Letters, 110 (16):666–671, July 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ghosh:2013:NUB

[GD13] Debashis Ghosh and Lakshmi Kanta Dey. A new upper bound on the parameters of quasi-symmetric designs. Information Processing Letters, 113(12):444– 446, June 30, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000872.

Goldberg:2012:SWV

[GE12] Noam Goldberg and Jonathan Eckstein. Sparse weighted voting classifier selection and its linear programming relaxations. Information Processing Letters, 112(12):

481–486, June 30, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019012000725.

Gentilini:2014:NAM

[Gen14] Raffaella Gentilini. A note on the approximation of mean-payoff games. Information Processing Letters, 114(7):382–386, July 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019014000349.

Gerbessiotis:2012:IRB

[Ger12] Alexandros V. Gerbessiotis. An improved reliability bound of a probabilistic parallel integer sorting algorithm. Information Processing Letters, 112(24):976–979, December 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002608.

Gerstgrasser:2019:RAD

[Ger19] Matthias Gerstgrasser. Reverse auctions are different from auctions. Information Processing Letters, 147(??): 49–54, July 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019300560.

Grabowski:2011:SMI

[GFG11] Szymon Grabowski, Simone Faro, and Emanuele Giaquinta. String matching with inversions and translocations in linear average time (most of the time). Information Processing Letters, 111(11):516–520, May 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Giaquinta:2013:NAB

[GG13] Emanuele Giaquinta and Szymon Grabowski. New algorithms for binary jumbled pattern matching. Information Processing Letters, 113(14–16):538–542, July/ August 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001300.

Gao:2018:MMK

[GG18] Hang Gao and Wenyu Gao. Maximum matching and kernelization of edge dominating set. Information Processing Letters, 136(??):21– 24, August 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300802.

Giaquinta:2013:APM

[GGF13] Emanuele Giaquinta, Szymon Grabowski, and Kimmo Fredriksson. Approximate pattern matching with k-mismatches in packed text. Information Processing Letters, 113(19– 21):693–697, September/ October 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001919.

Garijo:2014:MGF

[GGG^{+}14] D. Garijo, M. A. Garrido, C. I. Grima, A. Márquez, A. Moreno-González, J. R. Portillo, P. Reyes, R. Robles, and J. Valenzuela. Monochromatic geometric k-factors for bicolored point sets with auxiliary points. Information Processing Letters, 114(1–2):19–24, January/February 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013002597.

Gomez:2011:LCN

[GGI11] Domingo Gómez, Jaime Gutierrez, and Álvar Ibeas. On the linear complexity of the Naor–Reingold sequence. Information Processing Letters, 111 (17):854–856, September 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011001517.

Gu:2014:ECA

[GH14] Mei-Mei Gu and Rong-Xia Hao. 3-extra connectivity of 3-ary n-cube networks. Information Processing Letters, 114(9):486–491, September 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014000623.

Gelernter:2015:GLS

[GH15] Nethanel Gelernter and Amir Herzberg. Gossip Latin square and the meetall gossipers problem. Information Processing Letters, 115(10):738–743, October 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015000393.

Guo:2015:JFC

[GHC15] Jing-Ming Guo, Chih-Hsien Hsia, and Chia-Hao Chang. JPEG false contour reduction using error diffusion. Information Processing Letters, 115(3):403–407, March 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002610.

Ghiduk:2014:AGB

[Ghi14] Ahmed S. Ghiduk. Automatic generation of basis test paths using variable length genetic algorithm. Information Processing Letters, 114(6):304– 316, June 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014000180.

Grimson:2011:EES

[GHK11] Rafael Grimson, Joos Heintz, and Bart Kuijpers. Efficient evaluation of specific queries in constraint databases. Information Processing Letters, 111(19):941–944, October 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001864.

Gandhi:2018:MLT

[GHK^{+}18] Rajiv Gandhi, Mohammad Taghi Hajiaghayi, Guy Kortsarz, Manish Purohit, and Kanthi Sarpatwar. On maximum leaf trees and connections to connected maximum cut problems. Information Processing Letters, 129(??):31–34, January 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017301035.

Gu:2018:PDD

[GHL18] Mei-Mei Gu, Rong-Xia Hao, and Jian-Bing Liu. The pessimistic diagnosability of data center networks. Information Processing Letters, 134(??):52– 56, June 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300383.

Giudici:2017:ASP

[GHRT17] Alberto Giudici, Pascal Halffmann, Stefan Ruzika, and Clemens Thielen. Approximation schemes for the parametric knapsack problem. Information Processing Letters, 120(??):11–15, April 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016301843.

Golovach:2013:CFG

[GHvtHP13] Petr A. Golovach, Pinar Heggernes, Pim van ’t Hof, and Daniël Paulusma. Choosability on H-free graphs. Information Processing Letters, 113(4):107– 110, February 28, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019012003067.

Gelle:2018:RUF

[GI18] Kitti Gelle and Szabolcs Iván. Recognizing unionfind trees is NP-complete. Information Processing Letters, 131(??):7–14, March 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019017301977.

Gilbers:2014:VDP

[Gil14] Alexander Gilbers. VCdimension of perimeter visibility domains. Information Processing Letters, 114(12):696–699, December 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001252.

Gillis:2018:MUP

[Gil18] Nicolas Gillis. Multiplicative updates for polynomial root finding. Information Processing Letters, 132(??): 14–18, April 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017302028.

Giorgi:2018:PAV

[Gio18] Pascal Giorgi. A probabilistic algorithm for verifying polynomial middle product in linear time. Information Processing Letters, 139(??): 24–29, November 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019018301479.

Gal:2011:SDL

[GJ11] Anna Gál and Jing-Tang Jang. The size and depth of layered Boolean circuits. Information Processing Letters, 111(5):213–217, February 1, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gutin:2014:PAL

[GJ14] Gregory Gutin and Mark Jones. Parameterized algorithms for load coloring problem. Information Processing Letters, 114(8):446– 449, August 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014000519.

Grzegorek:2015:NOS

[GJ15] Paulina Grzegorek and Janusz Januszewski. A note on one-space bounded square packing. Information Processing Letters, 115(11):872–876, November 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000915.

Guo:2019:NBT

[GJ19] Qian Guo and Thomas Johansson. A new birthdaytype algorithm for attacking the fresh re-keying countermeasure. Information Processing Letters, 146(??):30– 34, June 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300420.

Grigorescu:2010:LDT

[GJR10] Elena Grigorescu, Kyomin Jung, and Ronitt Rubinfeld. A local decision test for sparse polynomials. Information Processing Letters, 110(20):898–901, September 30, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gueron:2010:EIG

[GK10] Shay Gueron and Michael Kounavis. Efficient implementation of the Galois Counter Mode using a carryless multiplier and a fast reduction algorithm. Information Processing Letters, 110(14–15):549–553, July 1, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Geetha:2011:VRN

[GKCK11] S. Geetha, V. Kabilan, S. P. Chockalingam, and N. Kamaraj. Varying radix numeral system based adaptive image steganography. Information Processing Letters, 111(16):792–797, August 30, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001414.

Golovnev:2014:SSB

[GKM14] Alexander Golovnev, Alexander S. Kulikov, and Ivan Mihajlin. Solving SCS for bounded length strings in fewer than 2^{n} steps. Information Processing Letters, 114(8):421–425, August 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014000416.

Gavalas:2015:AAA

[GKM^{+}15] Damianos Gavalas, Charalampos Konstantopoulos, Konstantinos Mastakas, Grammati Pantziou, and Nikolaos Vathis. Approximation algorithms for the arc orienteering problem. Information Processing Letters, 115(2):313–315, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901400218X.

Gerbner:2016:TOW

[GKPP16] Dániel Gerbner, Balázs Keszegh, Cory Palmer, and Dömötör Pálvölgyi. Topological orderings of weighted directed acyclic graphs. Information Processing Letters, 116(9):564– 568, September 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019016300552.

Govorcin:2013:LGO

[GKS13] Jelena Govorcin, Martin Knor, and Riste Skrekovski. Line graph operation and small worlds. Information Processing Letters, 113(5–6):196–200, March 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000045.

Grandoni:2015:LRI

[GKW15] Fabrizio Grandoni, Tomasz Kociumaka, and Michal Wlodarczyk. An LProunding 2\sqrt{2}-approximation for restricted maximum acyclic subgraph. Information Processing Letters, 115(2):182–185, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014001847.

Gao:2011:IAB

[GL11] Weifeng Gao and Sanyang Liu. Improved artificial bee colony algorithm for global optimization. Information Processing Letters, 111(17):871–882, September 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011001670.

Guan:2013:CMS

[GL13] Li Guan and Jianping Li. Coordination mechanism for selfish scheduling under a grade of service provision. Information Processing Letters, 113(8):251–254, April 30, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013000264.

Gao:2015:IIN

[GL15] Xiaofang Gao and Jiye Liang. An improved incremental nonlinear dimensionality reduction for isometric data embedding. Information Processing Letters, 115(4):492–501, April 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019014002749.

Guo:2012:PAV

[GLLX12] Qiaoping Guo, Shengjia Li, Ruijuan Li, and Gaokui Xu. Pancyclic out-arcs of a vertex in oriented graphs. Information Processing Letters, 112(20):759–761, October 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001810.

Gerault:2018:RAR

[GLMS18] David Gérault, Pascal Lafourcade, Marine Minier, and Christine Solnon. Revisiting AES related-key differential attacks with constraint programming. Information Processing Letters, 139(??):18–23, November 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901830139X.

Golynski:2012:OSL

[GLO12] Alexander Golynski and Alejandro López-Ortiz. Optimal strategies for the list update problem under the MRM alternative cost model. Information Processing Letters, 112(6):218–222, March 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011003231.

Galby:2019:CEC

[GLPR19] Esther Galby, Paloma T. Lima, Daniël Paulusma, and Bernard Ries. Classifying k-edge colouring for H-free graphs. Information Processing Letters, 146(??):39– 43, June 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019300432.

Gluck:2010:SGP

[Glü10]  Robert Glück.  Self-

generating program specializers. Information Processing Letters, 110(17): 787–793, August 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gafarov:2012:NSM

[GLW12] Evgeny R. Gafarov, Alexander A. Lazarev, and Frank Werner. A note on a single machine scheduling problem with generalized total tardiness objective function. Information Processing Letters, 112(3):72–76, January 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011002845.

Guo:2013:NVP

[GLXG13] Qiaoping Guo, Shengjia Li, Gaokui Xu, and Yubao Guo. Notes on vertex pancyclicity of graphs. Information Processing Letters, 113 (19–21):710–713, September/October 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001828.

Gu:2011:CMC

[GLY11] Xiaofeng Gu, Hong-Jian Lai, and Senmei Yao. Characterization of minimally (2, ℓ)-connected graphs. Information Processing Letters, 111(23–24):1124–1129, December 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011002614.

Gerstl:2012:SPI

[GM12a] Enrique Gerstl and Gur Mosheiov. Scheduling on parallel identical machines with job-rejection and position-dependent processing times. Information Processing Letters, 112(19):743–747, October 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012001603.

Goldreich:2012:TPT

[GM12b] Oded Goldreich and Or Meir. The tensor product of two good codes is not necessarily robustly testable. Information Processing Letters, 112(8–9):351–355, April 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012000245.

Gerstl:2013:IAD

[GM13a] Enrique Gerstl and Gur Mosheiov. An improved algorithm for due-window assignment on parallel identical machines with unittime jobs. Information Processing Letters, 113(19– 21):754–759, September/ October 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001816.

Guillemot:2013:FFA

[GM13b] Sylvain Guillemot and Dániel Marx. A faster FPT algorithm for Bipartite Contraction. Information Processing Letters, 113 (22–24):906–912, November/December 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901300241X.

Gorain:2015:AAS

[GM15] Barun Gorain and Partha Sarathi Mandal. Approximation algorithm for sweep coverage on graph. Information Processing Letters, 115(9):712– 718, September 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015000605.

Gerstl:2015:NMW

[GMM15] Enrique Gerstl, Baruch Mor, and Gur Mosheiov. A note: Maximizing the weighted number of just-in-time jobs on a proportionate flowshop. Information Processing Letters, 115(2):159– 162, February 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019014001793.

Gutierres:2019:OGC

[GMS19] Gonçalo Gutierres, Ricardo Mamede, and José Luis Santos. Optimal Gray code for involutions. Information Processing Letters, 148(??): 19–22, August 2019. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019019300808.

Gutin:2013:NEP

[GMY13] G. Gutin, G. Muciaccia, and A. Yeo. (non-)existence of polynomial kernels for the Test Cover problem. Information Processing Letters, 113(4):123–126, February 28, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012003183.

Gelade:2010:ORA

[GN10] Wouter Gelade and Frank Neven. Optimizing the Region Algebra is PSPACEcomplete. Information Processing Letters, 110(16): 639–643, July 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gupta:2011:UBA

[GNG11] Kishan Chand Gupta, Yassir Nawaz, and Guang Gong. Upper bound for algebraic immunity on a subclass of Maiorana McFarland class of bent functions. Information Processing Letters, 111(5):247–249, February 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gyssens:2014:CSA

[GNV14] Marc Gyssens, Mathias Niepert, and Dirk Van Gucht. On the completeness of the semigraphoid axioms for deriving arbitrary from saturated conditional independence statements. Information Processing Letters, 114(11):628– 633, November 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014001057.

Golumbic:2018:TCR

[Gol18] Martin Charles Golumbic.

Total coloring of rooted path graphs. Information Processing Letters, 135(??): 73–76, July 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300498.

Gourves:2015:AOS

[Gou15] Laurent Gourvès. Approximating the optimal sequence of acquisitions and sales with a capped budget. Information Processing Letters, 115(10):760–764, October 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901500071X.

Grinberg:2017:FPT

[GP17] Yuri Grinberg and Theodore J. Perkins. Fully polynomial-time computation of maximum likelihood trajectories in Markov chains. Information Processing Letters, 118(??):53– 57, February 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016301399.

Groult:2010:CDP

[GPR10] Richard Groult, Élise Prieur, and Gwénaël Richomme. Counting distinct palindromes in a word in linear time. Information Processing Letters, 110(20):908– 912, September 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gangopadhyay:2017:NNS

[GPSD17] Sugata Gangopadhyay, Enes Pasalic, Pantelimon Stanica, and Saral Datta. A note on non-splitting Z-bent functions. Information Processing Letters, 121(??):1– 5, May 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300017.

Guo:2016:GAB

[GPT16] Fan Guo, Hui Peng, and Jin Tang. Genetic algorithmbased parameter selection approach to single image defogging. Information Processing Letters, 116(10): 595–602, October 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019016300618.

Gaubert:2017:CSP

[GQ17] Stéphane Gaubert and Zheng Qu. Checking strict positivity of Kraus maps is NP-hard. Information Processing Letters, 118(??):35– 43, February 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301351.

Guo:2010:SCK

[GQG10] Litao Guo, Chengfu Qin, and Xiaofeng Guo. Super connectivity of Kronecker products of graphs. Information Processing Letters, 110(16):659–661, July 31, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gopalan:2013:CPI

[GR13] Abishek Gopalan and Srinivasan Ramasubramanian. A counterexample for the proof of implication conjecture on independent spanning trees. Information Processing Letters, 113(14–16):522–526, July/ August 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001191.

See corrigendum [GR14].

Gopalan:2014:CCP

[GR14] Abishek Gopalan and Srinivasan Ramasubramanian. Corrigendum to: “A counterexample for the proof of implication conjecture on independent spanning trees” [Information Processing Letters 113 (14–

16) (2013) 522–526]. Information Processing Letters, 114(1–2):84, January/ February 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002421.

See [GR13].

Ghosal:2019:NPP

[GR19] Purnata Ghosal and B. V. Raghavendra Rao. A note on parameterized polynomial identity testing using hitting set generators. Information Processing Letters, 151 (??):Article 105839, November 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901930122X.

Grabowski:2015:NLC

[Gra15]  Szymon Grabowski.  A

note on the longest common substring with kmismatches problem. Information Processing Letters, 115(6–8):640–642, June/ August 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000411.

Grandi:2018:ABF

[Gra18] Fabio Grandi. On the analysis of Bloom filters. Information Processing Letters, 129(??):35–39, January 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017301606.

Grolmusz:2015:NPU

[Gro15] Vince Grolmusz. A note on the PageRank of undirected graphs. Information Processing Letters, 115(6–8):633–634, June/ August 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015000381.

Gamard:2017:PRA

[GRSS17] Guilhem Gamard, Gwenaël Richomme, Jeffrey Shallit, and Taylor J. Smith. Periodicity in rectangular arrays. Information Processing Letters, 118(??):58–63, February 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016301387.

Gamzu:2010:PAC

[GS10a] Iftah Gamzu and Danny Segev. A polylogarithmic approximation for computing non-metric terminal Steiner trees. Information Processing Letters, 110 (18–19):826–829, September 15, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gnecco:2010:SWS

[GS10b] Giorgio Gnecco and Marcello Sanguineti. On spectral windows in supervised learning from data. Information Processing Letters, 110(23):1031–1036, November 15, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gruner:2010:DFH

[GS10c] S. Gruner and T. J. Steyn.

Deadlock-freeness of hexagonal systolic arrays. Information Processing Letters, 110(14–15):539–543, July 1, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gastin:2012:DWC

[GS12] Paul Gastin and Nathalie Sznajder. Decidability of well-connectedness for distributed synthesis. Information Processing Letters, 112(24):963–968, December 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002451.

Gadducci:2017:RBP

[GS17] Fabio Gadducci and Francesco Santini. Residuation for bipolar preferences in soft constraints. Information Processing Letters, 118(??): 69–74, February 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019016301430.

Gupta:2019:DTB

[GS19] Ishu Gupta and Ashutosh Kumar Singh. Dynamic threshold based information leaker identification scheme. Information Processing Letters, 147(??):69– 73, July 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300559.

Guzel:2019:NMF

[GSA^{+}19] Gülsüm Gözde Güzel, Muharrem Tolga Sakalli, Sedat Akleylek, Vincent Rijmen, and Yasemin Çengellenmis. A new matrix form to generate all 3×3 involutory MDS matrices over F_{2m}. Information Processing Letters, 147(??): 61–68, July 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300511.

Grigorious:2014:PDC

[GSR^{+}14] Cyriac Grigorious, Sudeep Stephen, Bharati Rajan, Mirka Miller, and Albert William. On the partition dimension of a class of circulant graphs. Information Processing Letters, 114(7): 353–356, July 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014000234.

Guo:2016:PRW

[GSS16] Chuan Guo, Jeffrey Shallit, and Arseny M. Shur. Palindromic rich words and run-length encodings. Information Processing Letters, 116(12):735–738, December 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300990.

Guo:2015:IPA

[GSY15] Jian Guo, Chunhua Su, and Wun-She Yap. An improved preimage attack against HAVAL-3. Information Processing Letters, 115(2):386– 393, February 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014002312.

Grytczuk:2013:OVT

[GSZ13] Jaroslaw Grytczuk, Piotr Szafruga, and Michal Zmarz. Online version of the theorem of Thue. Information Processing Letters, 113(5–6):193–195, March 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012003043.

Gueron:2012:SCC

[Gue12] Shay Gueron. Speeding up CRC32C computations with Intel CRC32 instruction. Information Processing Letters, 112(5):179–185, February 28, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901100319X.

Galindo:2014:LCL

[GV14] David Galindo and Srinivas Vivek. Limits of a conjecture on a leakage-resilient cryptosystem. Information Processing Letters, 114(4): 192–196, April 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019013002949.

Gutin:2016:TLB

[GW16] Gregory Gutin and Magnus Wahlström. Tight lower bounds for the Workflow Satisfiability Problem based on the Strong Exponential Time Hypothesis. Information Processing Letters, 116(3):223–226, March 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001921.

Gao:2011:UIB

[GWJ11] Min Gao, Zhongfu Wu, and Feng Jiang. Userrank for item-based collaborative filtering recommendation. Information Processing Letters, 111(9):440–446, April 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gao:2018:BCR

[GWL18] Jian Gao, Yongkang Wang, and Juan Li. Bounds on covering radius of linear codes with Chinese Euclidean distance over the finite non chain ring F2+vF2. Information Processing Letters, 138(??):19–21, October 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301145.

Gao:2013:MCC

[GXZZ13] Yang Gao, Ming Xu, Naijun Zhan, and Lijun Zhang. Model checking conditional CSL for continuoustime Markov chains. Information Processing Letters, 113(1–2):44–50, January 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012002803.

Gutin:2010:NMB

[GY10] Gregory Gutin and Anders Yeo. Note on maximal bisection above tight lower bound. Information Processing Letters, 110(21): 966–969, October 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Geng:2015:NUP

[GY15] Zhichao Geng and Jinjiang Yuan. A note on unbounded parallel-batch scheduling. Information Processing Letters, 115(12):969–974, December 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001155.

Gluck:2019:CBT

[GY19] Robert Glück and Tetsuo Yokoyama. Constructing a binary tree from its traversals by reversible recursion and iteration. Information Processing Letters, 147(??): 32–37, July 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019300523.

Guo:2012:MBP

[GZJ^{+}12] Deke Guo, Guiming Zhu, Hai Jin, Panlong Yang, Yingwen Chen, Xianqing Yi, and Junxian Liu. Möbius–de Bruijn: The product of Möbius cube and de Bruijn digraph. Information Processing Letters, 112(5):205–211, February 28, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011003103.

Guo:2015:EVA

[GZM15] Zhiwei Guo, Haixing Zhao, and Yaping Mao. The equitable vertex arboricity of complete tripartite graphs. Information Processing Letters, 115(12):977–982, December 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901500112X.

Haddad:2018:MDF

[Had18] Serge Haddad. Memoryless determinacy of finite parity games: Another simple proof. Information Processing Letters, 132(??):19– 21, April 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017302065.

Halldorsson:2010:OCH

[Hal10] Magnús M. Halldórsson. Online coloring of hypergraphs. Information Processing Letters, 110(10): 370–372, April 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Honma:2011:EAL

[HAM11] Hirotoshi Honma, Kodai Abe, and Shigeru Masuyama. Erratum and addendum to “A linear time algorithm for finding all hinge vertices of a permutation graph” [Information Processing Letters 59 (2) (1996) 103–107]. Information Processing Letters, 111(18):891–894, September 30, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011000238.

See [HWJ96].

Han:2017:CPW

[Han17] Yijie Han. Construct a perfect word hash function in time independent of the size of integers. Information Processing Letters, 128(??): 5–10, December 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019017301266.

Harwath:2016:NSP

[Har16] Frederik Harwath. A note on the size of prenex normal forms. Information Processing Letters, 116(7):443– 446, July 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901630028X.

Harada:2018:NQC

[Har18] Masaaki Harada. New quantum codes constructed from some self-dual additive F4codes. Information Processing Letters, 138(??):31– 34, October 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301157.

Hasan:2013:FSP

[Has13] Masud Hasan. On fixed size projection of simplicial polyhedra. Information Processing Letters, 113 (22–24):882–887, November/December 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013002378.

Havill:2010:OMJ

[Hav10] Jessen T. Havill. Online malleable job scheduling for m ≤ 3. Information Processing Letters, 111(1):31– 35, December 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Holenderski:2014:RBT

[HBL14] Mike Holenderski, Reinder J. Bril, and Johan J. Lukkien. Red-black trees with relative node keys. Information Processing Letters, 114(11):591– 596, November 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014001124.

Haddadi:2015:PTL

[HCCG15] S. Haddadi, S. Chenche, M. Cheraitia, and F. Guessoum. Polynomial-time local-improvement algorithm for Consecutive Block Minimization. Information Processing Letters, 115(6–8):612–617, June/ August 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000332.

Hung:2015:HPU

[HCWH15] Chun-Nan Hung, Eddie Cheng, Tao-Ming Wang, and Lih-Hsing Hsu. On Hamiltonian properties of unidirectional hypercubes. Information Processing Letters, 115(6–8):551–555, June/ August 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000150.

Henry:2016:NWP

[Hen16] Christopher S. Henry. The (nested) word problem. Information Processing Letters, 116(11):729–734, November 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300977.

Harn:2014:MTS

[HF14] Lein Harn and Miao Fuyou.

Multilevel threshold secret sharing based on the Chinese Remainder Theorem. Information Processing Letters, 114(9):504–509, September 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014000659.

Harn:2015:DTS

[HH15] Lein Harn and Ching-Fang Hsu. Dynamic threshold secret reconstruction and its application to the threshold cryptography. Information Processing Letters, 115(11):851–857, November 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001106.

Hedetniemi:2012:SSA

[HHJ^{+}12] Sandra M. Hedetniemi, Stephen T. Hedetniemi, Hao Jiang, K. E. Kennedy, and Alice A. McRae. A self-stabilizing algorithm for optimally efficient sets in graphs. Information Processing Letters, 112(16):621–623, August 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901200124X.

Heam:2017:EPT

[HHK17] P.-C. Héam, V. Hugot, and O. Kouchnarenko. The emptiness problem for tree automata with at least one global disequality constraint is NP-hard. Information Processing Letters, 118(??): 6–9, February 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001901630134X.

Honkala:2015:EOD

[HHL15] Iiro Honkala, Olivier Hudry, and Antoine Lobstein. On the ensemble of optimal dominating and locatingdominating codes in a graph. Information Processing Letters, 115(9):699–702, September 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000654.

Huang:2010:IAS

[HHTL10] Yen-Lin Huang, Cheng-Chen Huang, Chuan Yi Tang, and Chin Lung Lu. An improved algorithm for sorting by blockinterchanges based on permutation groups. Information Processing Letters, 110(8–9):345–350, April 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hanamura:2011:PSL

[HI11] Satoshi Hanamura and Shigeki Iwata. Posets with seven linear extensions sortable by three comparisons. Information Processing Letters, 111(8):365–369, March 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hirsch:2012:ORA

[HI12] Edward A. Hirsch and Dmitry Itsykson. On an optimal randomized acceptor for graph nonisomorphism. Information Processing Letters, 112(5):166–171, February 28, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011003176.

Hildenbrandt:2014:SPP

[Hil14] R. Hildenbrandt. A kserver problem with parallel requests and unit distances. Information Processing Letters, 114(5):239– 246, May 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013003207.

Haase:2016:CLS

[HK16] Christoph Haase and Stefan Kiefer. The complexity of the K th largest subset problem and related problems. Information Processing Letters, 116(2):111–115, February 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001702.

Holzhauser:2017:FPK

[HK17] Michael Holzhauser and Sven O. Krumke. An FPTAS for the parametric knapsack problem. Information Processing Letters, 126(??):43–47, October 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301072.

Hwang:2012:AFP

[HKK12] Jinsoo Hwang, Jeankyung Kim, and Kichang Kim. Analysis of the false-positive error rate of tagged fragment marking scheme. Information Processing Letters, 112(6):227–232, March 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011003346.

Holzhauser:2017:CAB

[HKT17] Michael Holzhauser, Sven O.

Krumke, and Clemens Thielen. On the complexity and approximability of budgetconstrained minimum cost flows. Information Processing Letters, 126(??):24– 29, October 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301047.

Hasani:2014:MTW

[HKW14] Keramat Hasani, Svetlana A. Kravchenko, and Frank Werner. Minimizing total weighted completion time approximately for the parallel machine problem with a single server. Information Processing Letters, 114(9):500–503, September 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014000647.

Hoffmann:2018:MNM

[HKW18] Stefan Hoffmann, Thomas Kampermann, and Egon Wanke. Minimizing the number of max-power users in ad-hoc wireless networks with minimum node degree requirements. Information Processing Letters, 136(??): 25–29, August 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019018300760.

He:2012:RCR

[HL12] Jing He and Hongyu Liang. On rainbow-k-connectivity of random graphs. Information Processing Letters, 112(10):406–410, May 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012000403.

Hong:2016:DCC

[HL16] Xia Hong and Qinghai Liu. Degree condition for completely independent spanning trees. Information Processing Letters, 116(10): 644–648, October 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901630076X.

He:2017:NHS

[HL17] Cheng He and Hao Lin.

Notes on a hierarchical scheduling problem on identical machines. Information Processing Letters, 120(??): 6–10, April 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301740.

Herranz:2011:RBS

[HLR11] Javier Herranz, Fabien Laguillaumie, and Carla Ràfols. Relations between semantic security and anonymity in identity-based encryption. Information Processing Letters, 111(10): 453–460, April 30, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hung:2010:FFI

[HLT10] Regant Y. S. Hung, Lap-Kei Lee, and H. F. Ting. Finding frequent items over sliding windows with constant update time. Information Processing Letters, 110 (7):257–260, March 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Han:2011:SEP

[HLY11] Xixian Han, Jianzhong Li, and Donghua Yang. Supporting early pruning in top-k query processing on massive data. Information Processing Letters, 111 (11):524–532, May 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Heggernes:2010:HAM

[HM10] Pinar Heggernes and Daniel Meister. Hardness and approximation of minimum distortion embeddings. Information Processing Letters, 110(8–9):312– 316, April 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hitchcock:2013:BIF

[HM13] John M. Hitchcock and Elvira Mayordomo. Base invariance of feasible dimension. Information Processing Letters, 113(14–16):546– 551, July/August 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019013001154.

Halacsy:2018:OEE

[HM18] Gergely Halácsy and Zoltán Ádám Mann. Optimal energyefficient placement of virtual machines with divisible sizes. Information Processing Letters, 138(??):44– 50, October 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301285.

Hokama:2016:BSA

[HMS16] Pedro Hokama, Flávio K.

Miyazawa, and Rafael C. S. Schouery. A bounded space algorithm for online circle packing. Information Processing Letters, 116(5):337– 342, May 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015002240.

Huber:2018:GMR

[HMSS18] Katharina T. Huber, Vincent Moulton, Marie-France Sagot, and Blerina Sinaimeri. Geometric medians in reconciliation spaces of phylogenetic trees. Information Processing Letters, 136(??): 96–101, August 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019018300814.

Harju:2010:CRF

[HN10] Tero Harju and Dirk Nowotka.

Cyclically repetition-free words on small alphabets. Information Processing Letters, 110(14–15):591–595, July 1, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hocquard:2011:GMD

[Hoc11] Hervé Hocquard. Graphs with maximum degree 6 are acyclically 11-colorable. Information Processing Letters, 111(15):748–753, August 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011001256.

Hougardy:2010:FWA

[Hou10]  Stefan Hougardy.  The

Floyd–Warshall algorithm on graphs with negative cycles. Information Processing Letters, 110(8–9):279– 281, April 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hocquard:2013:SEC

[HOV13] Hervé Hocquard, Pascal Ochem, and Petru Valicov. Strong edge-colouring and induced matchings. Information Processing Letters, 113(19–21):836–843, September/October 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019013002159.

Hrubes:2018:NMR

[HP18] Pavel Hrubes and Pavel Pudlák. A note on monotone real circuits. Information Processing Letters, 131(??):15–19, March 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019017301965.

Huang:2019:TZE

[HP19] Shang-En Huang and Seth Pettie. Thorup–Zwick emulators are universally optimal hopsets. Information Processing Letters, 142(??): 9–13, February 2019. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019018301868.

Han:2017:PAR

[HPP17] Yong-Jin Han, Seong-Bae Park, and Se-Young Park. Personalized app recommendation using spatiotemporal app usage log. Information Processing Letters, 124(??):15–20, August 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300534.

Han:2010:DPU

[HPY^{+}10] Xin Han, Chao Peng, Deshi Ye, Dahai Zhang, and Yan Lan. Dynamic bin packing with unit fraction items revisited. Information Processing Letters, 110(23): 1049–1054, November 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Herranz:2013:SMS

[HRS13] Javier Herranz, Alexandre Ruiz, and Germán Sáez. Sharing many secrets with computational provable security. Information Processing Letters, 113(14–16):572– 579, July/August 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019013001373.

Hrubes:2012:NRD

[Hru12]  Pavel Hrubes.  On the

nonnegative rank of distance matrices. Information Processing Letters, 112(11): 457–461, June 15, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019012000580.

Hou:2011:ACI

[HRW11] Jianfeng Hou, Nicolas Roussel, and Jianliang Wu. Acyclic chromatic index of planar graphs with triangles. Information Processing Letters, 111(17):836–840, September 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011001578.

Hirsch:2015:PCL

[HS15] Edward A. Hirsch and Dmitry Sokolov. On the probabilistic closure of the loose unambiguous hierarchy. Information Processing Letters, 115(9):725–730, September 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000708.

Halfon:2019:SPA

[HS19] S. Halfon and Ph. Schnoebelen. On shuffle products, acyclic automata and piecewise-testable languages. Information Processing Letters, 145(??):68– 73, May 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300201.

Huang:2014:HCI

[Hua14]  Xing Huang.  Hamilton

cycles in implicit clawheavy graphs. Information Processing Letters, 114(12):676–679, December 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014001215.

Huang:2018:IDS

[Hua18] Xing Huang. An implicit degree sum condition for cycles through specified vertices. Information Processing Letters, 129(??):40–43, January 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901730162X.

Houston:2012:ZPG

[HWA12] Robin Houston, Joseph White, and Martyn Amos. Zen Puzzle Garden is NPcomplete. Information Processing Letters, 112(3):106– 108, January 31, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019011002870.

Ho:1996:LTA

[HWJ96] Ting-Yem Ho, Yue-Li Wang, and Ming-Tsan Juan. A linear time algorithm for finding all hinge vertices of a permutation graph. Information Processing Letters, 59(2):103–107, July 22, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See errata [HAM11].

Hwang:2012:TBL

[HYC12] Iksoon Hwang, Nina Yevtushenko, and Ana Cavalli. Tight bound on the length of distinguishing sequences for non-observable nondeterministic Finite-State Machines with a polynomial number of inputs and outputs. Information Processing Letters, 112(7):298–301, March 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011003516.

Hou:2014:LSS

[HYK14] Yung-Tsung Hou, Dar-Li Yang, and Wen-Hung Kuo. Lot scheduling on a single machine. Information Processing Letters, 114(12):718–722, December 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001306.

Huang:2015:CSI

[HYYZ15] Sheng Huang, Dan Yang, Ge Yongxin, and Xiaohong Zhang. Combined supervised information with PCA via discriminative component selection. Information Processing Letters, 115(11):812–816, November 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001064.

Huang:2005:EMP

[HZSL05] Liusheng Huang, Hong Zhong, Hong Shen, and Yonglong Luo. An efficient multiple-precision division algorithm. In Hong Shen and Koji Nakano, editors, Sixth International Conference on Parallel and Distributed Computing, Applications and Technologies, 2005. PDCAT 2005: 5– 8 December 2005, Dalian, China, pages 971–974. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 0-7695-2405-2. LCCN QA76.58 .I5752 2005. The authors present an integerdivision algorithm that runs three to five times faster than Knuth’s 1981 original. However, there is an error in the renormalization algorithm that is corrected in [MN14], while retaining the speedup.

Hu:2016:STF

[HZX16] Zhi Hu, Guoliang Zhang, and Maozhi Xu. Some techniques for faster scalar multiplication on GLS curves. Information Processing Letters, 116(1):41–46, January 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001556.

Ibarra:2017:CCS

[Iba17] Louis Ibarra. Computing the clique-separator graph for an interval graph in linear time. Information Processing Letters, 124(??):21– 25, August 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300649.

Inenaga:2018:HRN

[IH18] Shunsuke Inenaga and Heikki Hyyrö. A hardness result and new algorithm for the longest common palindromic subsequence problem. Information Processing Letters, 129(??):11–15, January 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301485.

Iwasaki:2010:FTR

[IK10] Tatsuya Iwasaki and Keiichi Kaneko. Fault-tolerant routing in burnt pancake graphs. Information Processing Letters, 110(14–15): 535–538, July 1, 2010. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Imai:2015:IRR

[IK15a] Shigeyoshi Imai and Kaoru Kurosawa. Improved reconstruction of RSA privatekeys from their fraction. Information Processing Letters, 115(6–8):630–632, June/ August 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015000368.

Ishiu:2015:THD

[IK15b] Tetsuya Ishiu and Masashi Kikuchi. The termination of the higher-dimensional tarai functions. Information Processing Letters, 115(2):125– 127, February 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019014001525.

Isaak:2012:EIT

[IL12] Dimitri Isaak and Christof Löding. Efficient inclusion testing for simple classes of unambiguous ω-automata. Information Processing Letters, 112(14–15):578–582, August 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001184.

Ilic:2013:EAV

[Ili13]  Aleksandar Ilić.  Effi-

cient algorithm for the vertex connectivity of trapezoid graphs. Information Processing Letters, 113(10–11):398–404, May/ June 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000665.

Iwerks:2012:AGT

[IM12] Justin Iwerks and Joseph S. B. Mitchell. The art gallery theorem for simple polygons in terms of the number of reflex and convex vertices. Information Processing Letters, 112(20):778–782, October 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001937.

Ikenmeyer:2017:RPR

[IM17] Christian Ikenmeyer and Stefan Mengel. On the relative power of reduction notions in arithmetic circuit complexity. Information Processing Letters, 130(??): 7–10, February 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019017301655.

Idalino:2015:LMS

[IMCP15] Thaís Bardini Idalino, Lucia Moura, Ricardo Felipe Custódio, and Daniel Panario. Locating modifications in signed data for partial data integrity. Information Processing Letters, 115(10):731–737, October 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901500037X.

Izsak:2012:NLS

[IN12] Rani Izsak and Zeev Nutov. A note on labeling schemes for graph connectivity. Information Processing Letters, 112(1–2):39–43, January 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011002663.

Indrzejczak:2015:ECH

[Ind15] Andrzej Indrzejczak. Eliminability of cut in hypersequent calculi for some modal logics of linear frames. Information Processing Letters, 115(2):75–81, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901400132X.

Pae:2013:EOR

[iP13] Sung il Pae. Exact output rate of Peres’s algorithm for random number generation. Information Processing Letters, 113(5–6):160– 164, March 15, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013000033.

Ichimura:2010:NPB

[IS10] Akira Ichimura and Maiko Shigeno. A new parameter for a broadcast algorithm with locally bounded Byzantine faults. Information Processing Letters, 110(12–13):514–517, June 15, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ito:2014:PER

[Ito14]  Tsuyoshi Ito.  Paral-

lelization of entanglementresistant multi-prover interactive proofs. Information Processing Letters, 114(10): 579–583, October 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014001008.

Ivan:2016:CAC

[Ivá16] Szabolcs Iván. Complexity of atoms, combinatorially. Information Processing Letters, 116(5):356–360, May 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016000053.

Iwama:2010:OKR

[IZ10] Kazuo Iwama and Guochuan Zhang. Online knapsack with resource augmentation. Information Processing Letters, 110(22):1016– 1020, October 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jaddi:2016:SRG

[JAH16] Najmeh Sadat Jaddi, Salwani Abdullah, and Abdul Razak Hamdan. A solution representation of genetic algorithm for neural network weights and structure. Information Processing Letters, 116(1):22–25, January 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001386.

Jansen:2012:TTH

[Jan12a] Maurice Jansen. Towards a tight hardness-randomness connection between permanent and arithmetic circuit identity testing. Information Processing Letters, 112(24):969–975, December 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002128.

Januszewski:2012:LAS

[Jan12b] Janusz Januszewski. Online algorithms for 2-space bounded 2-dimensional bin packing. Information Processing Letters, 112(19): 719–722, October 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019012001676.

Jannati:2016:SAR

[JB16] Hoda Jannati and Behnam Bahrak. Security analysis of an RFID tag search protocol. Information Processing Letters, 116(10):618– 622, October 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300643.

Ji:2010:SJD

[JC10] Min Ji and T. C. E.

Cheng. Scheduling with jobdependent learning effects and multiple rate-modifying activities. Information Processing Letters, 110(11): 460–463, May 16, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jung:2011:EBU

[JCC11] Jae-Yoon Jung, Chang-Ho Chin, and Jorge Cardoso. An entropy-based uncertainty measure of process models. Information Processing Letters, 111(3): 135–141, January 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Duan:2011:ELB

[jDX11] Mei jiao Duan and Jing Xu. An efficient locationbased compromise-tolerant key management scheme for sensor networks. Information Processing Letters, 111 (11):503–507, May 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jannati:2015:AFR

[JF15] Hoda Jannati and Abolfazl Falahati. Analysis of false-reject probability in distance bounding protocols with mixed challenges over RFID noisy communication channel. Information Processing Letters, 115(6–8):623–629, June/ August 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000356.

Jha:2015:CDN

[Jha15] Pranava K. Jha. A comment on “The domination number of exchanged hypercubes”. Information Processing Letters, 115(2):343– 344, February 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014001938.

Jiang:2011:NLC

[Jia11] Jehn-Ruey Jiang. Nondominated local coteries for resource allocation in grids and clouds. Information Processing Letters, 111(8): 379–384, March 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jiang:2016:MAC

[Jia16] Shaoquan Jiang. On message authentication with a correlated setup. Information Processing Letters, 116(4):289–293, April 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019015001908.

Jiang:2018:PIC

[Jia18] Minghui Jiang. Periodicity of identifying codes in strips. Information Processing Letters, 135(??):77– 84, July 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300620.

Jouida:2018:DBO

[JK18] Sihem Ben Jouida and Saoussen Krichen. A DSS based on optimizer tools and MTS meta-heuristic for the warehousing problem with conflicts. Information Processing Letters, 135(??):14– 21, July 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300309.

Jannati:2019:SOR

[JK19] Hoda Jannati and Ramtin Khosravi. On the security of one-round meeting location determination protocol. Information Processing Letters, 146(??):35–38, June 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019300274.

Jobson:2018:LIG

[JKL18] Adam S. Jobson, André E.

Kézdy, and Jenő Lehel. Linkage on the infinite grid. Information Processing Letters, 137(??):51–56, September 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901830098X.

Jain:2010:ODS

[JKS10] Rahul Jain, Hartmut Klauck, and Miklos Santha. Optimal direct sum results for deterministic and randomized decision tree complexity. Information Processing Letters, 110(20):893– 897, September 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jain:2019:IPT

[JKSS19] Sanjay Jain, Bakhadyr Khoussainov, Philipp Schlicht, and Frank Stephan. The isomorphism problem for treeautomatic ordinals with addition. Information Processing Letters, 149(??):19–24, September 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300924.

Jaiswal:2015:IAI

[JKY15] Ragesh Jaiswal, Mehul Kumar, and Pulkit Yadav. Improved analysis of D^{2}- sampling based PTAS for kmeans and other clustering problems. Information Processing Letters, 115(2):100– 103, February 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014001471.

Joos:2014:GIC

[JLdSO^{+}14] Felix Joos, Christian Löwenstein, Fabiano de S. Oliveira, Dieter Rautenbach, and Jayme L. Szwarcfiter. Graphs of interval count two with a given partition. Information Processing Letters, 114(10): 542–546, October 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014000611.

Joeris:2017:SDC

[JLMO17] Benson Joeris, Nathan Lindzey, Ross M. Mc-Connell, and Nissa Osheim. Simple DFS on the complement of a graph and on partially complemented digraphs. Information Processing Letters, 117(??):35– 39, January 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901630120X.

Jamison:2012:MOS

[JN12] Robert E. Jamison and Darren A. Narayan. Maxoptimal and sum-optimal labelings of graphs. Information Processing Letters, 112(1–2):26–31, January 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901100247X.

Johnen:2014:FSS

[Joh14a] Colette Johnen. Fast, silent self-stabilizing distance-k independent dominating set construction. Information Processing Letters, 114(10): 551–555, October 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014000945.

Johnson:2014:SNM

[Joh14b] Hunter R. Johnson. Some new maximum VC classes. Information Processing Letters, 114(6):294–298, June 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019014000076.

Jordan:2010:GGR

[Jor10] Tibor Jordán. Generically globally rigid zeolites in the plane. Information Processing Letters, 110(18–19):841– 844, September 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jordan:2012:HCM

[Jor12] Tibor Jordán. Highly connected molecular graphs are rigid in three dimensions. Information Processing Letters, 112(8–9):356– 359, April 30, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019012000397.

Jain:2010:RPR

[JOS10] Sanjay Jain, Yuh Shin Ong, and Frank Stephan. Regular patterns, regular languages and context-free languages. Information Processing Letters, 110(24): 1114–1119, November 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jager:2011:NPG

[JP11] Gerold Jäger and Marcin Peczarski. The number of pessimistic guesses in Generalized Black-peg Mastermind. Information Processing Letters, 111(19):933– 940, October 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011001803.

Jha:2019:SDP

[JPB19] Anupriya Jha, D. Pradhan, and S. Banerjee. The secure domination problem in cographs. Information Processing Letters, 145(??):30– 38, May 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300079.

Jonas:2018:CQB

[JS18] M. Jonás and J. Strejcek.

On the complexity of the quantified bit-vector arithmetic with binary encoding. Information Processing Letters, 135(??):57–61, July 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300474.

Jukna:2019:GCB

[JS19] Stasys Jukna and Hannes Seiwert. Greedy can beat pure dynamic programming. Information Processing Letters, 142(??):90–95, February 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018302126.

Jansen:2011:SOA

[JSO11] Klaus Jansen and Roberto Solis-Oba. A simple OPT +1 algorithm for cutting stock under the modified integer round-up property assumption. Information Processing Letters, 111(10): 479–482, April 30, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Junosza-Szaniawski:2010:GCN

[JSR10] Konstanty Junosza-Szaniawski and Łukasz Rożej. Game chromatic number of graphs with locally bounded number of cycles. Information Processing Letters, 110(17): 757–760, August 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Junosza-Szaniawski:2011:CEA

[JSR11] Konstanty Junosza-Szaniawski and Pawel Rzazewski. On the complexity of exact algorithm for L(2, 1)-labeling of graphs. Information Processing Letters, 111(14): 697–701, July 31, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019011001165.

Jia:2015:DIE

[JSZC15] Tong Jia, Yan Shi, ZhongXuan Zhou, and DongYue Chen. 3D depth information extraction with omnidirectional camera. Information Processing Letters, 115(2):285–291, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002129.

Jonsson:2010:AIP

[JT10] Peter Jonsson and Johan Thapper. Approximating integer programs with positive right-hand sides. Information Processing Letters, 110(10):351–355, April 30, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Janczewski:2015:CCB

[JT15] Robert Janczewski and Krzysztof Turowski. The computational complexity of the backbone coloring problem for bounded-degree graphs with connected backbones. Information Processing Letters, 115(2):232–236, February 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001951.

Janczewski:2016:HCS

[JT16] Robert Janczewski and Krzysztof Turowski. On the hardness of computing span of subcubic graphs. Information Processing Letters, 116(1):26–32, January 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001520.

Jukna:2012:CPC

[Juk12]  Stasys Jukna.  Clique

problem, cutting plane proofs and communication complexity. Information Processing Letters, 112(20):772–777, October 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012001913.

Jourdan:2015:RCS

[JUY15] Guy-Vincent Jourdan, Hasan Ural, and Hüsnü Yenigün. Reduced checking sequences using unreliable reset. Information Processing Letters, 115(5):532–535, May 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0020019015000113.

Jiang:2018:GHE

[JZ18] Ting Jiang and XiaoJian Zhou. Gradient/Hessianenhanced least square support vector regression. Information Processing Letters, 134(??):1–8, June 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019018300292.

Khazaei:2017:COA

[KA17] Shahram Khazaei and Siavash Ahmadi. Ciphertext-only attack on d × d Hill in O(d13^{d}). Information Processing Letters, 118(??):25– 29, February 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301338.

Kakugawa:2015:SSD

[Kak15] Hirotsugu Kakugawa. Selfstabilizing distributed algorithm for local mutual inclusion. Information Processing Letters, 115(6–8):562–569, June/ August 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000174.

Kalinich:2012:FWP

[Kal12] Adam O. Kalinich. Flipping the winner of a poset game. Information Processing Letters, 112(3):86–89, January 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011002651.

Kamiyama:2019:NBF

[Kam19] Naoyuki Kamiyama. A note on balanced flows in equality networks. Information Processing Letters, 145(??): 74–76, May 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300225.

Kara:2013:SRC

[Kar13] Orhun Kara. Square reflection cryptanalysis of 5round Feistel networks with permutations. Information Processing Letters, 113 (19–21):827–831, September/October 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002172.

S:2017:DTR

[Kar17] Karthik C. S. Did the train reach its destination: the complexity of finding a witness. Information Processing Letters, 121(??):17– 21, May 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300042.

Katrenic:2016:FFA

[Kat16] Ján Katrenic. A faster FPT algorithm for 3-path vertex cover. Information Processing Letters, 116(4):273– 278, April 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015002136.

Kim:2011:NFS

[KC11] Heeheon Kim and Yookun Cho. A new fair scheduling algorithm for periodic tasks on multiprocessors. Information Processing Letters, 111(7):301–309, March 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kwon:2017:CBI

[KC17] Shinil Kwon and Sungdeok Cha. CAPTCHA-based image annotation. Information Processing Letters, 128(??): 27–31, December 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901730131X.

Kocabas:2013:NSS

[KD13] Ilker Kocabas and Bekir Taner Dinçer. A new statistical strategy for pooling: ELI. Information Processing Letters, 113(19– 21):739–746, September/ October 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001968.

Kiwi:2014:SMB

[KdER^{+}14] M. Kiwi, P. Moisset de Espanés, I. Rapaport, S. Rica, and G. Theyssier. Strict Majority Bootstrap Percolation in the r-wheel. Information Processing Letters, 114(6): 277–281, June 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014000064.

Karakoc:2013:BCL

[KDH13] F. Karakoç, H. Demirci, and A. E. Harmanci. Biclique cryptanalysis of LBlock and TWINE. Information Processing Letters, 113(12): 423–429, June 30, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013000884.

Karakoc:2015:AKA

[KDH15] F. Karakoç, H. Demirci, and A. E. Harmanci. AKF: a key alternating Feistel scheme for lightweight cipher designs. Information Processing Letters, 115(2):359–367, February 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002257.

Ketema:2011:CIR

[Ket11] Jeroen Ketema. Counterexamples in infinitary rewriting with non-fully-extended rules. Information Processing Letters, 111(13):642– 646, July 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011000925.

Kobayashi:2016:CRR

[KHKS16] Yuki Kobayashi, Yuya Higashikawa, Naoki Katoh, and Adnan Sljoka. Characterizing redundant rigidity and redundant global rigidity of body-hinge graphs. Information Processing Letters, 116(2):175–178, February 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001544.

Kiefer:2013:BBE

[Kie13] Stefan Kiefer. BPA bisimilarity is EXPTIME-hard. Information Processing Letters, 113(4):101–106, February 28, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012003079.

Kim:2010:NFA

[Kim10] Chong Hee Kim. New fault attacks using Jacobi symbol and application to regular right-to-left algorithms. Information Processing Letters, 110(20): 882–886, September 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kissel:2019:KRC

[Kis19] Zachary A. Kissel. Key regression from constrained pseudorandom functions. Information Processing Letters, 147(??):10–13, July 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0020019019300493.

Kulich:2011:PPS

[KK11] Tomáš Kulich and Miroslava Kemeňová. On the paper of Pascal Schweitzer concerning similarities between incompressibility methods and the Lovász Local Lemma. Information Processing Letters, 111(9):436– 439, April 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kamiyama:2015:PAT

[KK15a] Naoyuki Kamiyama and Yasushi Kawase. On packing arborescences in temporal networks. Information Processing Letters, 115(2):321– 325, February 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014002208.

Kiraly:2015:RDA

[KK15b] Zoltán Király and Erika R.

Kovács. Randomized and deterministic algorithms for network coding problems in wireless networks. Information Processing Letters, 115(4):507– 511, April 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002622.

Kapelko:2016:DCU

[KK16] Rafal Kapelko and Evangelos Kranakis. On the displacement for covering a unit interval with randomly placed sensors. Information Processing Letters, 116(11):710–717, November 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016300825.

Kaski:2014:FMS

[KKKS14] Petteri Kaski, Mikko Koivisto, Janne H. Korhonen, and Igor S. Sergeev. Fast monotone summation over disjoint sets. Information Processing Letters, 114(5):264– 267, May 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013003062.

Kolpakov:2010:MRA

[KKO10] Roman Kolpakov, Gregory Kucherov, and Pascal Ochem. On maximal repetitions of arbitrary exponent. Information Processing Letters, 110(7):252–256, March 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kubica:2013:LTA

[KKR^{+}13] M. Kubica, T. Kulczyński, J. Radoszewski, W. Rytter, and T. Waleń. A linear time algorithm for consecutive permutation pattern matching. Information Processing Letters, 113(12): 430–433, June 30, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013000926.

Karandikar:2015:ISC

[KKS15] P. Karandikar, M. Kufleitner, and Ph. Schnoebelen. On the index of Simon’s congruence for piecewise testability. Information Processing Letters, 115(4):515– 519, April 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002580.

Kochol:2011:MRC

[KKSŠ11] Martin Kochol, Nad’a Krivoňáková, Silvia Smejová, and Katarína Šranková. Matrix reduction in a combinatorial computation. Information Processing Letters, 111(4):164–168, January 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kasperski:2013:AMM

[KKZ13] Adam Kasperski, Adam Kurpisz, and Pawel Zieliński. Approximating the minmax (regret) selecting items problem. Information Processing Letters, 113(1–2): 23–29, January 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019012002815.

Koehler:2010:AAB

[KL10] Henning Koehler and Sebastian Link. Armstrong axioms and Boyce–Codd– Heath Normal Form under bag semantics. Information Processing Letters, 110 (16):717–724, July 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kowaluk:2018:USE

[KL18] Miroslaw Kowaluk and Andrzej Lingas. Are unique subgraphs not easier to find? Information Processing Letters, 134(??):57– 61, June 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300395.

Klein:2013:CBH

[Kle13] Shmuel T. Klein. On the connection between Hamming codes, Heapsort and other methods. Information Processing Letters, 113(17):617–620, August 30, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001555.

Kim:2018:NCP

[KLMP18] Minki Kim, Bernard Lidický, Tomás Masarík, and Florian Pfender. Notes on complexity of packing coloring. Information Processing Letters, 137(??):6–10, September 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300929.

Kozma:2010:CTG

[KLS10] Gady Kozma, Zvi Lotker, and Gideon Stupp. On the connectivity threshold for general uniform metric spaces. Information Processing Letters, 110(10): 356–359, April 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Karpinski:2013:OCP

[KLS13] Marek Karpinski, Andrzej Lingas, and Dzmitry Sledneu. Optimal cuts and partitions in tree metrics in polynomial time. Information Processing Letters, 113(12): 447–451, June 30, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013000860.

Keller:2010:DAS

[KM10] Nathan Keller and Stephen D.

Miller. Distinguishing attacks on stream ciphers based on arrays of pseudorandom words. Information Processing Letters, 110(4): 129–132, January 16, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Katz:2011:SBR

[KM11] Matthew J. Katz and Gila Morgenstern. Settling the bound on the rectilinear link radius of a simple rectilinear polygon. Information Processing Letters, 111(3): 103–106, January 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kumamoto:2012:ODE

[KM12] Masao Kumamoto and Eiji Miyano. Optimal distortion embedding of complete binary trees into lines. Information Processing Letters, 112(10):365–370, May 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000440.

Kolliopoulos:2013:VCM

[KM13] Stavros G. Kolliopoulos and Yannis Moysoglou. The 2-valued case of makespan minimization with assignment constraints. Information Processing Letters, 113(1–2):39–43, January 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901200261X.

Klavzar:2014:DNE

[KM14] Sandi Klavzar and Meijie Ma. The domination number of exchanged hypercubes. Information Processing Letters, 114(4):159– 162, April 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013003086.

Kohler:2016:LTA

[KM16a] Ekkehard Köhler and Lalla Mouatadid. A linear time algorithm to compute a maximum weighted independent set on cocomparability graphs. Information Processing Letters, 116(6):391– 395, June 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015002124.

Kundu:2016:LTA

[KM16b] Sukhamay Kundu and Subhashis Majumder. A linear time algorithm for optimal k-hop dominating set of a tree. Information Processing Letters, 116(2):197–202, February 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001362.

Kazemi:2018:AAC

[KMK18] Mohammad Reza Kazemi, Ali Mohades, and Payam Khanteimouri. Approximation algorithms for color spanning diameter. Information Processing Letters, 135(??):53–56, July 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019018300164.

Kovacs:2015:OSA

[KMMN15] Annamária Kovács, Ulrich Meyer, Gabriel Moruz, and Andrei Negoescu. The optimal structure of algorithms for α-paging. Information Processing Letters, 115(12):932–938, December 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001337.

Kosinski:2018:CT

[KMN18] Karol Kosiński, Robert Mercas, and Dirk Nowotka. Corrigendum to “A note on Thue games” [Inf. Process. Lett. 118 (2017) 75– 77]. Information Processing Letters, 130(??):63–65, February 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301825.

See [MN17].

Kovalyov:2019:CPF

[KMS19] Mikhail Y. Kovalyov, Gur Mosheiov, and Dmitrij Sesok. Comments on “Proportionate flowshops with general position dependent processing times” [inf. process. lett. 111 (2011) 174– 177] and “Minimizing total load on a proportionate flowshop with positiondependent processing times and job-rejection” [inf. process. lett. 132 (2018) 39– 43]. Information Processing Letters, 147(??):1–2, July 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300456.

See [Mos11b, FM18].

Kantor:2013:ASD

[KN13a] Alexander Kantor and Kobbi Nissim. Attacks on statistical databases: the highly noisy case. Information Processing Letters, 113(12):409–413, June 30, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013000823.

Kapoor:2013:MFW

[KN13b] Kalpesh Kapoor and Himadri Nayak. On multiset of factors of a word. Information Processing Letters, 113(3):78–80, February 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012002955.

Krithika:2013:ADC

[KN13c] R. Krithika and N. S. Narayanaswamy. Another disjoint compression algorithm for odd cycle transversal. Information Processing Letters, 113(22– 24):849–851, November/ December 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013002238.

Kiyomi:2016:FCG

[KO16] Masashi Kiyomi and Yota Otachi. Finding a chain graph in a bipartite permutation graph. Information Processing Letters, 116(9):569–573, September 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300485.

Kobayashi:2015:CPD

[Kob15] Yasuaki Kobayashi. Computing the pathwidth of directed graphs with small vertex cover. Information Processing Letters, 115(2):310–312, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002178.

Kochol:2012:NEL

[Koc12]  Martin Kochol.  Non-

extendible latin parallelepipeds. Information Processing Letters, 112(24):942–943, December 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002335.

Kochol:2018:TCC

[Koc18] Martin Kochol. Three colorability characterized by shrinking of locally connected subgraphs into triangles. Information Processing Letters, 135(??):33– 35, July 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300450.

Kokosinski:2017:GPI

[Kok17] Zbigniew Kokosiński. On generation of permutations of m out of n items. Information Processing Letters, 124(??):1–5, August 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017300509.

Korman:2012:MIA

[Kor12]  Matias Korman.  Mini-

mizing interference in ad hoc networks with bounded communication radius. Information Processing Letters, 112(19):748–752, October 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901200172X.

Kosolobov:2016:CRG

[Kos16] Dmitry Kosolobov. Computing runs on a general alphabet. Information Processing Letters, 116(3):241– 244, March 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015002100.

Kosolobov:2017:TLB

[Kos17] Dmitry Kosolobov. Tight lower bounds for the longest common extension problem. Information Processing Letters, 125(??):26–29, September 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300893.

Kostolanyi:2018:DWA

[Kos18] Peter Kostolányi. On deterministic weighted automata. Information Processing Letters, 140(??):37–41, December 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301704.

Kotrbcik:2012:NDC

[Kot12] Michal Kotrbcík. A note on disjoint cycles. Information Processing Letters, 112(4):135–137, February 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011003097.

Kolaitis:2012:DCC

[KP12] Phokion G. Kolaitis and Enela Pema. A dichotomy in the complexity of consistent query answering for queries with two atoms. Information Processing Letters, 112(3):77–85, January 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011002997.

Keil:2013:CMO

[KP13] J. Mark Keil and D. Pradhan. Computing a minimum outer-connected dominating set for the class of chordal graphs. Information Processing Letters, 113(14–16):552–561, July/ August 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901300135X.

Kociumaka:2014:FDF

[KP14a] Tomasz Kociumaka and Marcin Pilipczuk. Faster deterministic feedback vertex set. Information Processing Letters, 114(10):556– 560, October 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014000969.

Krasnoshchekov:2014:OHS

[KP14b] Dmitry Krasnoshchekov and Valentin Polishchuk. Orderk α-hulls and α-shapes. Information Processing Letters, 114(1–2):76–83, January/February 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013002123.

Korman:2018:LSC

[KPR18] Matias Korman, Sheung-Hung Poon, and Marcel Roeloffzen. Line segment covering of cells in arrangements. Information Processing Letters, 129(??):25– 30, January 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301515.

Koukouvinos:2011:ACV

[KPSZ11] Christos Koukouvinos, Veronika Pillwein, Dimitris E. Simos, and Zafeirakis Zafeirakopoulos. On the average complexity for the verification of compatible sequences. Information Processing Letters, 111 (17):825–830, September 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001438.

Kirsten:2011:RSR

[KQ11] D. Kirsten and K. Quaas.

Recognizability of the support of recognizable series over the semiring of the integers is undecidable. Information Processing Letters, 111(10):500–502, April 30, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Katajainen:2010:CDS

[KR10] Jyrki Katajainen and S. Srinivasa Rao. A compact data structure for representing a dynamic multiset. Information Processing Letters, 110(23):1061–1066, November 15, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Khan:2014:MNR

[KR14] Arindam Khan and Prasad Raghavendra. On mimicking networks representing minimum terminal cuts. Information Processing Letters, 114(7):365–371, July 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014000350.

Kol:2016:BQL

[KR16] Gillat Kol and Ran Raz. Bounds on 2-query Locally Testable Codes with affine tests. Information Processing Letters, 116(8):521–525, August 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016300321.

Krajicek:2012:NSA

[Kra12] Jan Krajícek. A note on SAT algorithms and proof complexity. Information Processing Letters, 112(12): 490–493, June 30, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019012000774.

Kostylev:2016:SAN

[KRV16] Egor V. Kostylev, Juan L.

Reutter, and Domagoj Vrgoc. Static analysis of navigational XPath over graph databases. Information Processing Letters, 116(7):467– 474, July 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300291.

Krzywkowski:2013:THM

[Krz13] Marcin Krzywkowski. Trees having many minimal dominating sets. Information Processing Letters, 113(8): 276–279, April 30, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019013000380.

Kulik:2010:TNE

[KS10] Ariel Kulik and Hadas Shachnai. There is no EP-TAS for two-dimensional knapsack. Information Processing Letters, 110(16): 707–710, July 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Katrenic:2011:IAB

[KS11] Ján Katrenič and Ingo Schiermeyer. Improved approximation bounds for the minimum rainbow subgraph problem. Information Processing Letters, 111(3): 110–114, January 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kochol:2012:BTG

[KS12] Martin Kochol and Riste Skrekovski. Brooks’ Theorem for generalized dart graphs. Information Processing Letters, 112(5):200– 204, February 28, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011003140.

Kauers:2018:SPS

[KS18] Manuel Kauers and Martina Seidl. Short proofs for some symmetric quantified Boolean formulas. Information Processing Letters, 140(??):1–3, December 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901830156X.

Kosolobov:2019:CLT

[KS19] Dmitry Kosolobov and Arseny M. Shur. Comparison of LZ77-type parsings. Information Processing Letters, 141(??):25–29, January 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301832.

Karaahmetoglu:2013:NMD

[KSBT13] Osman Karaahmetoglu, Muharrem Tolga Sakalli, Ercan Bulus, and Ion Tutanescu. A new method to determine algebraic expression of power mapping based Sboxes. Information Processing Letters, 113(7):229–235, April 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000070.

Keszocze:2018:CEM

[KSD18] Oliver Keszocze, Mathias Soeken, and Rolf Drechsler. The complexity of error metrics. Information Processing Letters, 139(??):??, November 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018301352.

Kshemkalyani:2011:RDC

[Ksh11] Ajay D. Kshemkalyani. Repeated detection of conjunctive predicates in distributed executions. Information Processing Letters, 111(9):447–452, April 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Khennoufa:2013:ECT

[KSK13] Riadh Khennoufa, Hamida Seba, and Hamamache Kheddouci. Edge coloring total k-labeling of generalized Petersen graphs. Information Processing Letters, 113(13):489–494, July 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001026.

Kaiser:2015:GGH

[KSV15] Tomás Kaiser, Maria Saumell, and Nico Van Cleemput. 10gabriel graphs are Hamiltonian. Information Processing Letters, 115(11):877– 881, November 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019015000903.

Krumke:2011:MCF

[KT11] Sven O. Krumke and Clemens Thielen. Minimum cost flows with minimum quantities. Information Processing Letters, 111 (11):533–537, May 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [KT12].

Krumke:2012:EMC

[KT12] Sven O. Krumke and Clemens Thielen. Erratum to “Minimum cost flows with minimum quantities” [Information Processing Letters 111 (11) (2011) 533–537]. Information Processing Letters, 112(13): 523–524, July 15, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019012000841.

See [KT11].

Kobayashi:2016:FFP

[KT16] Yasuaki Kobayashi and Hisao Tamaki. A faster fixed parameter algorithm for two-layer crossing minimization. Information Processing Letters, 116(9):547– 549, September 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019016300606.

Khoshkhah:2018:FSS

[KT18] Kaveh Khoshkhah and Dirk Oliver Theis. Fooling sets and the Spanning Tree polytope. Information Processing Letters, 132(??):11– 13, April 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017301011.

Kawamura:2017:MSN

[KTUY17] Akitoshi Kawamura, Yuichi Tatsu, Yushi Uno, and Masahide Yamato. Morpion Solitaire 5D: a new upper bound of 121 on the maximum score. Information Processing Letters, 121(??): 6–10, May 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017300030.

Kucherov:2013:CSA

[KTV13] Gregory Kucherov, Lilla Tóthmérész, and Stéphane Vialette. On the combinatorics of suffix arrays. Information Processing Letters, 113(22– 24):915–920, November/ December 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013002469.

Kundu:2017:RLS

[Kun17]  Abhisek Kundu.  Re-

laxed leverage sampling for low-rank matrix completion. Information Processing Letters, 124(??):6– 9, August 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300510.

Kundu:2018:GLT

[Kun18] Sukhamay Kundu. A generalized linear time algorithm for an optimal k-distance dominating set of a weighted tree. Information Processing Letters, 130(??):58–62, February 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301874.

Kutzkov:2012:EET

[Kut12] Konstantin Kutzkov. An exact exponential time algorithm for counting bipartite cliques. Information Processing Letters, 112(13): 535–539, July 15, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019012001032.

Kenkre:2010:AAB

[KV10] Sreyash Kenkre and Sundar Vishwanathan. Approximation algorithms for the Bipartite Multicut problem. Information Processing Letters, 110(8–9):282– 287, April 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kitaev:2016:MSW

[KV16] Sergey Kitaev and Vincent Vajnovszki. Mahonian STAT on words. Information Processing Letters, 116(2):157–162, February 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001611.

Kishore:2018:MCR

[KVK18] Ravi Kishore, Chiranjeevi Vanarasa, and Srinathan Kannan. On minimal connectivity requirements for secure message transmission in directed networks. Information Processing Letters, 131(??):1– 6, March 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301886.

Kuo:2016:SDD

[KWH16] Wen-Chung Kuo, Chun-Cheng Wang, and Hong-Ching Hou. Signed digit data hiding scheme. Information Processing Letters, 116(2):183–191, February 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001404.

Kuo:2012:SNP

[KY12] Wen-Hung Kuo and Dar-Li Yang. A short note on “Proportionate flowshops with general position-dependent processing times”. Information Processing Letters, 112(12):479–480, June 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000622.

See [Mos11b].

Ke:2013:NBS

[KYC13] Pinhui Ke, Wanghong Yu, and Zuling Chang. A note on binary sequence pairs with two-level correlation. Information Processing Letters, 113(19– 21):811–814, September/ October 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002111.

Ke:2012:NCQ

[KZ12] Pinhui Ke and Shengyuan Zhang. New classes of quaternary cyclotomic sequence of length 2p^{m} with high linear complexity. Information Processing Letters, 112(16):646–650, August 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901200141X.

Kolbe:2010:RND

[KZP10] Dashiell Kolbe, Qiang Zhu, and Sakti Pramanik. Reducing non-determinism of k-NN searching in nonordered discrete data spaces. Information Processing Letters, 110(10):420–423, April 30, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lev-Ari:2017:COS

[LABKS17] Kfir Lev-Ari, Edward Bortnikov, Idit Keidar, and Alexander Shraer. Composing ordered sequential consistency. Information Processing Letters, 123(??):47– 50, July 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017300480.

Lagerkvist:2014:WBB

[Lag14] Victor Lagerkvist. Weak bases of Boolean co-clones. Information Processing Letters, 114(9):462–468, September 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014000544.

Lai:2016:CAS

[Lai16a] Cheng-Nan Lai. On the construction of all shortest node-disjoint paths in star networks. Information Processing Letters, 116(4):299– 303, April 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001878.

Laihonen:2016:MDR

[Lai16b] Tero Laihonen. The metric dimension for resolving several objects. Information Processing Letters, 116(11):694–700, November 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300801.

Lampis:2011:KOV

[Lam11] Michael Lampis. A kernel of order 2k − c log k for vertex cover. Information Processing Letters, 111(23–24):1089–1091, December 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011002389.

Lange:2011:HEP

[Lan11] Martin Lange. P -hardness of the emptiness problem for visibly pushdown languages. Information Processing Letters, 111(7):338–341, March 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lasota:2017:EAC

[Las17] Slawomir Lasota. Equivariant algorithms for constraint satisfaction problems over coset templates. Information Processing Letters, 118(??):44–52, February 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016301363.

Lauria:2018:CET

[Lau18a] Massimo Lauria. Cliques enumeration and tree-like resolution proofs. Information Processing Letters, 135 (??):62–67, July 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019018300486.

Lauria:2018:NAK

[Lau18b] Massimo Lauria. A note about k-DNF resolution. Information Processing Letters, 137(??):33–39, September 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300942.

Lazic:2010:RPB

[Laz10] Ranko Lazić. The reachability problem for branching vector addition systems requires doubly-exponential space. Information Processing Letters, 110(17): 740–745, August 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lu:2013:CSA

[LC13] Linzhen Lu and Shaozhen Chen. A compress slide attack on the full GOST block cipher. Information Processing Letters, 113(17):634– 639, August 30, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013001543.

Lin:2017:LTA

[LC17] Min-Sheng Lin and Chien-Min Chen. Linear-time algorithms for counting independent sets in bipartite permutation graphs. Information Processing Letters, 122(??): 1–7, June 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017300182.

Lee:2017:DCN

[LCC17] Ming Che Lee, Sheng Yu Chiu, and Jia Wei Chang. A deep convolutional neural network based Chinese menu recognition app. Information Processing Letters, 128(??):14–20, December 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017301321.

Liu:2014:HNG

[LDW14] Xiaoping Liu, Zhilan Dang, and Baoyindureng Wu. The hub number, girth and Mycielski graphs. Information Processing Letters, 114(10): 561–563, October 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019014000957.

LeRoux:2019:MDI

[Le 19] Stéphane Le Roux. Memoryless determinacy of infinite parity games: Another simple proof. Information Processing Letters, 143(??): 8–13, March 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018302096.

Lee:2010:EAF

[Lee10] Jae Moon Lee. An efficient algorithm to find k-nearest neighbors in flocking behavior. Information Processing Letters, 110(14–15):576– 579, July 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:2017:AHM

[Lee17]  Euiwoong Lee.  APX-

hardness of maximizing Nash social welfare with indivisible items. Information Processing Letters, 122(??): 17–20, June 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300212.

Leitert:2017:CDC

[Lei17] Arne Leitert. 3-colouring for dually chordal graphs and generalisations. Information Processing Letters, 128(??): 21–26, December 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019017301345.

Leiva:2018:RTS

[Lei18] Luis A. Leiva. Responsive text summarization. Information Processing Letters, 130(??):52–57, February 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301850.

Luccio:2010:LBR

[LEP10] Fabrizio Luccio, Antonio Mesa Enriquez, and Linda Pagli. Lower bounds on the rotation distance of binary trees. Information Processing Letters, 110(21): 934–938, October 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:2016:LCR

[LFW^{+}16] Yu Liu, Kai Fu, Wei Wang, Ling Sun, and Meiqin Wang. Linear cryptanalysis of reduced-round SPECK. Information Processing Letters, 116(3):259– 266, March 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001891.

Lin:2018:IFA

[LFW^{+}18] Mugang Lin, Qilong Feng, Jianxin Wang, Jianer Chen, Bin Fu, and Wenjun Li. An improved FPT algorithm for Almost Forest Deletion problem. Information Processing Letters, 136(??):30– 36, August 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300772.

Liu:2017:SCE

[LFXH17] Yong-Jin Liu, Dian Fan, Chunxu Xu, and Ying He. Space complexity of exact discrete geodesic algorithms on regular triangulations. Information Processing Letters, 124(??):10– 14, August 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300625.

Li:2014:SIS

[LFZJ14] Yangyang Li, Shixia Feng, Xiangrong Zhang, and Licheng Jiao. SAR image segmentation based on quantum-inspired multiobjective evolutionary clustering algorithm. Information Processing Letters, 114(6): 287–293, June 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013003190.

Lucas:2018:ULM

[LG18] Salvador Lucas and Raúl Gutiérrez. Use of logical models for proving infeasibility in term rewriting. Information Processing Letters, 136(??):90–95, August 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300826.

Louza:2017:OSS

[LGT17] Felipe A. Louza, Simon Gog, and Guilherme P. Telles. Optimal suffix sorting and LCP array construction for constant alphabets. Information Processing Letters, 118(??):30–34, February 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016301375.

Lin:2010:EAC

[LH10] Chin-Yang Lin and Ting-Wei Hou. An efficient approach to cyclic reference counting based on a coarsegrained search. Information Processing Letters, 111(1): 1–10, December 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:2011:TCT

[LH11] Bertrand M. T. Lin and F. J. Hwang. Total completion time minimization in a 2-stage differentiation flowshop with fixed sequences per job type. Information Processing Letters, 111(5):

208–212, February 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:2011:ICA

[LHH11] Han-Yu Lin, Chien-Lung Hsu, and Shih-Kun Huang. Improved convertible authenticated encryption scheme with provable security. Information Processing Letters, 111(13):661–666, July 1, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011000846.

Lan:2016:CPI

[LHW^{+}16] Yan Lan, Xin Han, Zongtao Wu, He Guo, and Xin Chen. Complexity of problem TF2|ν = 1, c = 2|C_{max}. Information Processing Letters, 116(1):65– 69, January 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001581.

Li:2010:FTR

[Li10] Ruijuan Li. A Fan-type result on k-ordered graphs. Information Processing Letters, 110(16):651–654, July 31, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2012:ASW

[Li12]  Jiyou Li.  On the av-

erage sensitivity of the weighted sum function. Information Processing Letters, 112(4):143–148, February 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011003036.

Li:2015:APD

[Li15] Mengjun Li. Automatic proving or disproving equality loop invariants based on finite difference techniques. Information Processing Letters, 115(4):468– 474, April 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014002567.

Libura:2010:NRT

[Lib10] Marek Libura. A note on robustness tolerances for combinatorial optimization problems. Information Processing Letters, 110 (16):725–729, July 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:2010:TFG

[Liu10]  Shuli Liu.  On tough-

ness and fractional (g, f, n)- critical graphs. Information Processing Letters, 110(10): 378–382, April 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:2015:RWA

[Liu15]  Yu-Liang Liu.  Routing

and wavelength assignment for exchanged hypercubes in linear array optical networks. Information Processing Letters, 115(2):203–208, February 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001884.

Luo:2015:SVM

[LJ15] Wenchang Luo and Min Ji.

Scheduling a variable maintenance and linear deteriorating jobs on a single machine. Information Processing Letters, 115(1):33– 39, January 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001732.

Lin:2018:CPD

[LJL18] Shangwei Lin, Ya’nan Jin, and Chunfang Li. Cartesian product digraphs with optimal restricted arc connectivity. Information Processing Letters, 134(??):72– 75, June 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300437.

Li:2019:KRA

[LJP19] Rongjia Li, Chenhui Jin, and Hongchen Pan. Key recovery attacks on reducedround Joltik-BC in the single-key setting. Information Processing Letters, 151 (??):Article 105834, November 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901930105X.

Lee:2014:SPB

[LK14] Jooyoung Lee and Daesung Kwon. Security of permutation-based compression function lp231. Information Processing Letters, 114(7):372–381, July 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019014000325.

Lee:2012:HEO

[LKC^{+}12] Jaeheung Lee, Seokhyun Kim, Yookun Cho, Yoojin Chung, and Yongsu Park. HORSIC: an efficient onetime signature scheme for wireless sensor networks. Information Processing Letters, 112(20):783–787, October 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001950.

Lien:2015:DNG

[LKF15] Min-Yun Lien, Jyhmin Kuo, and Hung-Lin Fu. On the decycling number of generalized Kautz digraphs. Information Processing Letters, 115(2):209–211, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001896.

Lemire:2017:SVF

[LKR17] Daniel Lemire, Nathan Kurz, and Christoph Rupp. Stream VByte: Faster byteoriented integer compression. Information Processing Letters, 130(??):1– 6, February 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301679.

Lai:2010:SMS

[LL10a] Peng-Jen Lai and Wen-Chiung Lee. Single-machine scheduling with a nonlinear deterioration function. Information Processing Letters, 110(11):455–459, May 16, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:2010:TMT

[LL10b] Lan Lin and Yixun Lin. Two models of two-dimensional bandwidth problems. Information Processing Letters, 110(11):469–473, May 16, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:2010:MMD

[LL10c] Di Liu and Jing Li. Manyto-many n-disjoint path covers in n-dimensional hypercubes. Information Processing Letters, 110(14–15):580– 584, July 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:2010:MCB

[LL10d] Tsung-Hao Liu and Hsueh-I Lu. Minimum cycle bases of weighted outerplanar graphs. Information Processing Letters, 110(21): 970–974, October 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:2013:MSC

[LL13] Lan Lin and Yixun Lin. Machine scheduling with contiguous processing constraints. Information Processing Letters, 113(8):280– 284, April 30, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013000495.

Levcopoulos:2014:NQM

[LL14] Christos Levcopoulos and Andrzej Lingas. A note on a QPTAS for maximum weight triangulation of planar point sets. Information Processing Letters, 114(8): 414–416, August 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014000398.

Li:2016:SWS

[LL16] Jiyou Li and Chu Luo.

The simplified weighted sum function and its average sensitivity. Information Processing Letters, 116(5):331– 336, May 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901600003X.

Li:2017:IFP

[LL17] M. Li and Q. Liu. Inexact feasibility pump for mixed integer nonlinear programming. Information Processing Letters, 118(??):110– 116, February 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019016301508.

Li:2011:SBP

[LlChL11] Yin Li, Gong liang Chen, and Jian hua Li. Speedup of bit-parallel Karatsuba multiplier in GF(2^{m}) generated by trinomials. Information Processing Letters, 111(8): 390–394, March 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2010:ETI

[LlCyChL10] Yin Li, Gong liang Chen, Yi yang Chen, and Jian hua Li. An extension of TYT inversion algorithm in polynomial basis. Information Processing Letters, 110(8– 9):300–303, April 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luo:2010:PAE

[LLG10] Yiyuan Luo, Xuejia Lai, and Zheng Gong. Pseudorandomness analysis of the (extended) Lai–Massey scheme. Information Processing Letters, 111(2):90–96, December 31, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [LM91].

Li:2014:AAR

[LLG14] Weidong Li, Jianping Li, and Li Guan. Approximation algorithms for the ring loading problem with penalty cost. Information Processing Letters, 114(1–2):56–59, January/ February 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002202.

Lee:2010:SRT

[LLK10] Byoung-Hoon Lee, Sung-Hwa Lim, and Jai-Hoon Kim. Scalable real-time monitoring system for ubiquitous smart space. Information Processing Letters, 110(8–9):294–299, April 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lai:2011:MOL

[LLL^{+}11] Hong-Jian Lai, Hao Li, Ping Li, Yanting Liang, and Senmei Yao. Mod (2p + 1)- orientations in line graphs. Information Processing Letters, 111(23–24):1085–1088, December 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011002407.

Li:2018:CGG

[LLL18a] Xiaomin Li, Lan Lei, and Hong-Jian Lai. The connectivity of generalized graph products. Information Processing Letters, 136(??):37– 40, August 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300759.

Lichen:2018:AAC

[LLL18b] Junran Lichen, Jianping Li, and Ko-Wei Lih. Approximation algorithms for constructing specific subgraphs with minimum number of length-bounded stock pieces. Information Processing Letters, 137(??):11–16, September 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300930.

Lai:2018:EQK

[LLP^{+}18] Hong Lai, Mingxing Luo, Josef Pieprzyk, Zhiguo Qu, and Mehmet A. Orgun. Efficient quantum key distribution using Fibonaccinumber coding with a biased basis choice. Information Processing Letters, 134(??): 24–30, June 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300267.

LeMerrer:2014:HTF

[LLT14] Erwan Le Merrer, Nicolas Le Scouarnec, and Gilles Trédan. Heuristical top-k: fast estimation of centralities in complex networks. Information Processing Letters, 114(8):432–436, August 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014000490.

Lin:2015:FOC

[LLW15] Chiou-Jiun Lin, Jia-Jie Liu, and Yue-Li Wang. Finding outer-connected dominating sets in interval graphs. Information Processing Letters, 115(12):917–922, December 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001301.

Liu:2013:DMS

[LLWH13] Yiguang Liu, Stan Z. Li, Wei Wu, and Ronggang Huang. Dynamics of a mean-shift-like algorithm and its applications on clustering. Information Processing Letters, 113(1–2):8– 16, January 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002827.

Lai:1991:PNB

[LM91] Xuejia Lai and James L.

Massey. A proposal for a new block encryption standard. In Advances in cryptology—EUROCRYPT ’90 (Aarhus, 1990), volume 473 of Lecture Notes in Computer Science, pages 389–404. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/0473/04730389.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/0473/04730389.

pdf. See [LLG10] for proofs of requirements on the number of rounds.

Lucarelli:2011:IAA

[LM11] Giorgio Lucarelli and Ioannis Milis. Improved approximation algorithms for the Max Edge-Coloring problem. Information Processing Letters, 111(16):819– 823, August 30, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011001530.

Lu:2015:RCN

[LM15] Zaiping Lu and Yingbin Ma. Rainbow 2connection numbers of Cayley graphs. Information Processing Letters, 115(4): 486–491, April 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014002774.

Liotta:2016:VDI

[LM16] Giuseppe Liotta and Fabrizio Montecchiani. Lvisibility drawings of ICplanar graphs. Information Processing Letters, 116(3): 217–222, March 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901500201X.

Liu:2016:PBG

[LMCG16] Lingfeng Liu, Suoxia Miao, Mengfan Cheng, and Xiaojing Gao. A pseudorandom bit generator based on new multi-delayed Chebyshev map. Information Processing Letters, 116(11):674– 681, November 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019016300953.

Lokshtanov:2013:IFP

[LMS13] Daniel Lokshtanov, Neeldhara Misra, and Saket Saurabh. Imbalance is fixed parameter tractable. Information Processing Letters, 113(19–21):714–718, September/October 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0020019013001774.

Lin:2014:FAS

[LMS14] Min Chih Lin, Michel J. Mizrahi, and Jayme L. Szwarcfiter. Fast algorithms for some dominating induced matching problems. Information Processing Letters, 114(10):524–528, October 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901400091X.

Lee:2016:SPO

[LMSN16] Mun-Kyu Lee, Pierre Michaud, Jeong Seop Sim, and Dae-Hun Nyang. A simple proof of optimality for the MIN cache replacement policy. Information Processing Letters, 116(2):168–170, February 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001593.

Lambert:2015:CHS

[LMU15] Thomas Lambert, Loris Marchal, and Bora Uçar. Comments on the hierarchically structured bin packing problem. Information Processing Letters, 115(2):306– 309, February 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014002166.

Lin:2017:AWN

[LMV17] Min Chih Lin, Julián Mestre, and Saveliy Vasiliev. Approximating weighted neighborhood independent sets. Information Processing Letters, 130(??):11–15, February 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301709.

Le:2011:GCS

[LN11] Van Bang Le and Ngoc Tuy Nguyen. A good characterization of squares of strongly chordal split graphs. Information Processing Letters, 111(3):120–123, January 1, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:2011:LCP

[LNP11] Taehyung Lee, Joong Chae Na, and Kunsoo Park. Online construction of parameterized suffix trees for large alphabets. Information Processing Letters, 111(5): 201–207, February 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lofgren:2014:CMC

[Lof14] Peter Lofgren. On the complexity of the Monte Carlo method for incremental PageRank. Information Processing Letters, 114(3): 104–106, March 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019013002743.

Lee:2010:PRL

[LP10] Jooyoung Lee and Je Hong Park. Preimage resistance of LPmkr with r = m −

1. Information Processing Letters, 110(14–15):602– 608, July 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lozin:2013:BPS

[LP13a] Vadim Lozin and Christopher Purcell. Boundary properties of the satisfiability problems. Information Processing Letters, 113(9): 313–317, May 15, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019013000525.

Luccio:2013:CRN

[LP13b] Fabrizio Luccio and Linda Pagli. Chain rotations: a new look at tree distance. Information Processing Letters, 113(7):201–204, April 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000276.

Lima:2010:PKE

[LPdS10] J. B. Lima, D. Panario, and R. M. Campello de Souza. Public-key encryption based on Chebyshev polynomials over GF(q). Information Processing Letters, 111(2): 51–56, December 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lingas:2014:CNC

[LPWZ14] Andrzej Lingas, Leonidas Palios, Agnieszka Wasylewicz, and Pawel Zyliński. Corrigendum to “Note on covering monotone orthogonal polygons” [Inf. Process. Lett. 104(6) (2007) 220– 227]. Information Processing Letters, 114(11):646– 654, November 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014000593.

See [LWŻ07].

Li:2017:ADS

[LQL^{+}17] Kai Li, Daowen Qiu, Lvzhou Li, Shenggen Zheng, and Zhenbang Rong. Application of distributed semi-quantum computing model in phase estimation. Information Processing Letters, 120(??):23–29, April 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016301831.

Lal:2010:RCA

[LR10] Akash Lal and G. Ramalingam. Reference count analysis with shallow aliasing. Information Processing Letters, 111(2):57–63, December 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:2018:NOI

[LR18] Yunwen Liu and Vincent Rijmen. New observations on invariant subspace attack. Information Processing Letters, 138(??):22–26, October 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018301121.

Lalla-Ruiz:2016:HBR

[LREIMBMV16] Eduardo Lalla-Ruiz, Christopher Expósito-Izquierdo, Belén Melián-Batista, and J. Marcos Moreno-Vega. A Hybrid Biased Random Key Genetic Algorithm for the Quadratic Assignment Problem. Information Processing Letters, 116(8):513– 520, August 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300254.

Li:2011:FC

[LRS11] Xiaozhou Li, Atri Rudra, and Ram Swaminathan. Flexible coloring. Information Processing Letters, 111 (11):538–540, May 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lohrey:2019:SOT

[LRS19] Markus Lohrey, Carl Philipp Reh, and Kurt Sieber. Sizeoptimal top dag compression. Information Processing Letters, 147(??):27–31, July 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901930050X.

Lima:2017:DM

[LRSS17] Carlos V. G. C. Lima, Dieter Rautenbach, Uéverton S. Souza, and Jayme L. Szwarcfiter. Decycling with a matching. Information Processing Letters, 124(??):26– 29, August 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017300522.

Libkin:2010:DPM

[LS10] Leonid Libkin and Cristina Sirangelo. Disjoint pattern matching and implication in strings. Information Processing Letters, 110(4): 143–147, January 16, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liang:2011:AAC

[LS11] Zuosong Liang and Erfang Shan. Approximation algorithms for clique-transversal sets and clique-independent sets in cubic graphs. Information Processing Letters, 111(23–24):1104–1107, December 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011002481.

Lin:2014:CMI

[LS14] Min-Sheng Lin and Sheng-Huang Su. Counting maximal independent sets in directed path graphs. Information Processing Letters, 114(10):568–572, October 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014000982.

Liang:2015:CTS

[LS15] Zuosong Liang and Erfang Shan. The cliquetransversal set problem in claw-free graphs with degree at most 4. Information Processing Letters, 115(2):331– 335, February 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014002233.

Liang:2017:CTS

[LSK17] Zuosong Liang, Erfang Shan, and Liying Kang. The clique-transversal set problem in claw, K_{4}-free planar graphs. Information Processing Letters, 118(??):64– 68, February 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301429.

Li:2011:DFA

[LSLY11] Ruilin Li, Bing Sun, Chao Li, and Jianxiong You. Differential Fault Analysis on SMS4 using a single fault. Information Processing Letters, 111(4):156–163, January 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:2015:FAC

[LSS15] Min Chih Lin, Francisco J.

Soulignac, and Jayme L.

Szwarcfiter. A faster algorithm for the cluster editing problem on proper interval graphs. Information Processing Letters, 115(12):913–916, December 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001313.

Lou:2011:BTO

[LST11] Tiancheng Lou, Xiaoming Sun, and Christophe Tartary. Bounds and tradeoffs for Double-Base Number Systems. Information Processing Letters, 111(10): 488–493, April 30, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:2017:ISI

[LSW17] Euiwoong Lee, Melanie Schmidt, and John Wright. Improved and simplified inapproximability for kmeans. Information Processing Letters, 120(??):40– 43, April 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301739.

Li:2015:NLC

[LSZX15] Zepeng Li, Zehui Shao, Enqiang Zhu, and Jin Xu. A note on local coloring of graphs. Information Processing Letters, 115(2):302– 305, February 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014002154.

Lamberger:2013:MNC

[LT13a] Mario Lamberger and Elmar Teufl. Memoryless nearcollisions, revisited. Information Processing Letters, 113(3):60–66, February 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012003018.

Lin:2013:CTR

[LT13b] Min-Sheng Lin and Chao-Chun Ting. Computing K-terminal reliability of dtrapezoid graphs. Information Processing Letters, 113 (19–21):734–738, September/October 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001956.

Liu:2013:CGV

[LT13c] Yong-Jin Liu and Kai Tang.

The complexity of geodesic Voronoi diagrams on triangulated 2-manifold surfaces. Information Processing Letters, 113(4):132–136, February 28, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012003201.

Lin:2015:CKT

[LT15a] Min-Sheng Lin and Chao-Chun Ting. Computing the K-terminal reliability of directed path graphs. Information Processing Letters, 115(10):773–778, October 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015000757.

Lin:2015:PTA

[LT15b] Min-Sheng Lin and Chao-Chun Ting. A polynomialtime algorithm for computing K-terminal residual reliability of d-trapezoid graphs. Information Processing Letters, 115(2):371– 376, February 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019014002282.

Lauria:2018:SCP

[LT18] Massimo Lauria and Neil Thapen. On semantic cutting planes with very small coefficients. Information Processing Letters, 136(??): 70–75, August 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019018300875.

Li:2019:DGN

[LTK^{+}19] Xiaoyan Li, Yuan-Hsiang Teng, Tzu-Liang Kung, Qi Chen, and Cheng-Kuan Lin. The diagnosability and 1-good-neighbor conditional diagnosability of hypercubes with missing links and broken-down nodes. Information Processing Letters, 146(??):20– 26, June 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300158.

Li:2011:PBO

[LTWS11] M. S. Li, W. J. Tang, Q. H.

Wu, and J. R. Saunders. Paired-bacteria optimiser — a simple and fast algorithm. Information Processing Letters, 111(16):809–813, August 30, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001396.

Lu:2015:PLE

[Lu15]  Daniel L. Lu.  Planar

lower envelope of monotone polygonal chains. Information Processing Letters, 115(12):962–964, December 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001179.

Lu:2019:CVT

[Lü19] Huazhong Lü. On the conjecture of vertex-transitivity of DCell. Information Processing Letters, 142(??):80– 83, February 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018302114.

Lucas:2010:IKS

[Luc10] Joan M. Lucas. An improved kernel size for rotation distance in binary trees. Information Processing Letters, 110(12–13):481– 484, June 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lucas:2015:CCS

[Luc15] Salvador Lucas. Completeness of context-sensitive rewriting. Information Processing Letters, 115(2):87– 92, February 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014001410.

Lutz:2014:FPT

[Lut14] Jack H. Lutz. The frequent paucity of trivial strings. Information Processing Letters, 114(11):643–645, November 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901400101X.

Lin:2015:AAC

[LV15] Min Chih Lin and Saveliy Vasiliev. Approximation algorithms for clique transversals on some graph classes. Information Processing Letters, 115(9):667– 670, September 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015000630.

Lu:2019:EDH

[LW19] Huazhong Lü and Tingzeng Wu. Edge-disjoint Hamiltonian cycles of balanced hypercubes. Information Processing Letters, 144(??):25– 30, April 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018302369.

Liu:2010:NDC

[LWL10] Feng Liu, ChuanKun Wu, and XiJun Lin. A new definition of the contrast of visual cryptography scheme. Information Processing Letters, 110(7):241–246, March 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2011:PTC

[LWL11] Jing Li, Shiying Wang, and Di Liu. Pancyclicity of ternary n-cube networks under the conditional fault model. Information Processing Letters, 111(8):370–374, March 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2010:GCP

[LWS10] Hui Li, Chuan-Kun Wu, and Jun Sun. A general compiler for password-authenticated group key exchange protocol. Information Processing Letters, 110(4):160–167, January 16, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lu:2018:WLS

[LWXS18] Kejing Lu, Hongya Wang, Yingyuan Xiao, and Hui Song. Why locality sensitive hashing works: A practical perspective. Information Processing Letters, 136(??): 49–58, August 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019018300796.

Liu:2014:NBP

[LWXZ14] Mingjie Liu, Xiaoyun Wang, Guangwu Xu, and Xuexin Zheng. A note on BDD problems with λ_{2}-gap. Information Processing Letters, 114(1–2):9–12, January/February 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013002615.

Lingas:2007:NCM

[LWŻ07] Andrzej Lingas, Agnieszka Wasylewicz, and Paweł Żyliński. Note on covering monotone orthogonal polygons with star-shaped polygons. Information Processing Letters, 104(6):220– 227, December 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See corrigendum [LPWZ14].

Li:2013:EFT

[LX13a] Xiang-Jun Li and Jun-Ming Xu. Edge-fault tolerance of hypercube-like networks. Information Processing Letters, 113(19– 21):760–763, September/ October 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001993.

Li:2013:GMF

[LX13b] Xiang-Jun Li and Jun-Ming Xu. Generalized measures of fault tolerance in exchanged hypercubes. Information Processing Letters, 113(14–16):533–537, July/ August 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901300118X.

Liang:2013:MFA

[LX13c] Hao Liang and Jun-Ming Xu. Minimum feedback arc set of m-free digraphs. Information Processing Letters, 113(8):260–264, April 30, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013000379.

Li:2017:CTE

[LX17] Zepeng Li and Jin Xu. A characterization of trees with equal independent domination and secure domination numbers. Information Processing Letters, 119(??):14–18, March 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0020019016301685.

Lin:2019:VKE

[LX19] Weibo Lin and Mingyu Xiao. A (3 + ϵ)k-vertex kernel for edge-disjoint triangle packing. Information Processing Letters, 142(??):20– 26, February 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301911.

Li:2012:IAA

[LXDX12] Yu Li, Dachuan Xu, Donglei Du, and Naihua Xiu. Improved approximation algorithms for the robust fault-tolerant facility location problem. Information Processing Letters, 112(10): 361–364, May 31, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019012000452.

Lu:2014:CRU

[LXJ^{+}14] Jianfeng Lu, Dewu Xu, Lei Jin, Jianmin Han, and Hao Peng. On the complexity of role updating feasibility problem in RBAC. Information Processing Letters, 114(11):597–602, November 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014001112.

Lai:2012:RHB

[LXLY12] Hong Lai, Jinghua Xiao, Lixiang Li, and Yixian Yang. Recursive hiding of biometrics-based secret sharing scheme using adversary structure. Information Processing Letters, 112 (17–18):683–687, September 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012001573.

Lv:2019:IZN

[LXT19] Xuanjiao Lv, Lin Xiao, and Zhiguo Tan. Improved Zhang neural network with finite-time convergence for time-varying linear system of equations solving. Information Processing Letters, 147(??):88– 93, July 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019300699.

Liu:2014:FOS

[LXX14] Jinfei Liu, Li Xiong, and Xiaofeng Xu. Faster outputsensitive skyline computation algorithm. Information Processing Letters, 114(12):710–713, December 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001288.

Li:2011:OSU

[LY11] Wenhua Li and Jinjiang Yuan. Online scheduling on unbounded parallelbatch machines to minimize maximum flow-time. Information Processing Letters, 111(18):907–911, September 30, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001736.

Liu:2019:HPN

[LY19] Ziwen Liu and Chao Yang.

Hanano Puzzle is NPhard. Information Processing Letters, 145(??):6– 10, May 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300055.

Li:2011:UPB

[LYF11] Shisheng Li, Jinjiang Yuan, and Baoqiang Fan. Unbounded parallel-batch scheduling with family jobs and delivery coordination. Information Processing Letters, 111(12):575–582, June 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011000779.

Lin:2010:IST

[LYHC10] Jia-Cian Lin, Jinn-Shyong Yang, Chiun-Chieh Hsu, and Jou-Ming Chang. Independent spanning trees vs. edge-disjoint spanning trees in locally twisted cubes. Information Processing Letters, 110(10):414–419, April 30, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lu:2014:DAN

[LYHH14] Jiqiang Lu, Wun-She Yap, Matt Henricksen, and Swee-Huay Heng. Differential attack on nine rounds of the SEED block cipher. Information Processing Letters, 114(3):116–123, March 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901300272X.

Li:2010:TAM

[LZ10a] Xiuying Li and Zhao Zhang. Two algorithms for minimum 2-connected r-hop dominating set. Information Processing Letters, 110(22): 986–991, October 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:2010:PCC

[LZ10b] Hong Liu and Daming Zhu.

Parameterized complexity of control problems in Maximin election. Information Processing Letters, 110(10): 383–388, April 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:2012:EPS

[LZ12] Qiping Lin and Fangguo Zhang. Efficient precomputation schemes of kP + IQ. Information Processing Letters, 112(11):462– 466, June 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000786.

Li:2014:SBS

[LZ14] Song-Song Li and Yu-Zhong Zhang. Serial batch scheduling on uniform parallel machines to minimize total completion time. Information Processing Letters, 114(12):692–695, December 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001239.

Ladinek:2015:PCD

[LZ15] Irena Hrastnik Ladinek and Janez Zerovnik. Perfect codes in direct graph bundles. Information Processing Letters, 115(9):707–711, September 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015000496.

Liu:2010:DNC

[LZCM10] Juan Liu, Xindong Zhang, Xing Chen, and Jixiang Meng. On domination number of Cartesian product of directed cycles. Information Processing Letters, 110(5): 171–173, February 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:2012:SMS

[LZCX12] Ming Liu, Feifeng Zheng, Chengbin Chu, and Yinfeng Xu. Single-machine scheduling with past-sequence-dependent delivery times and release times. Information Processing Letters, 112(21):835–838, November 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012001901.

Li:2010:PES

[LZJX10] C. H. Li, X. F. Zhang, H. Jin, and W. Xiang. Epassport EAC scheme based on Identity-Based Cryptography. Information Processing Letters, 111(1):26– 30, December 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2012:OSE

[LZLY12] Wenjie Li, Zhenkun Zhang, Hailing Liu, and Jinjiang Yuan. Online scheduling of equal-length jobs with incompatible families on multiple batch machines to maximize the weighted number of early jobs. Information Processing Letters, 112(12):503–508, June 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901200083X.

Li:2017:NCL

[LZSX17] Zepeng Li, Enqiang Zhu, Zehui Shao, and Jin Xu. NPcompleteness of local colorings of graphs. Information Processing Letters, 130(??): 25–29, February 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019017301692.

Li:2012:OAS

[LZY12] Wenhua Li, Zhenkun Zhang, and Sufang Yang. Online algorithms for scheduling unit length jobs on parallelbatch machines with lookahead. Information Processing Letters, 112(7):292–297, March 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012000191.

M:2015:AAC

[M.15] Jagadish M. An approximation algorithm for the cutting-sticks problem. Information Processing Letters, 115(2):170–174, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014001823.

Machacek:2019:PWU

[Mac19] John Machacek. Partial words with a unique position starting a square. Information Processing Letters, 145(??):44–47, May 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0020019019300183.

Mannaa:2010:CEP

[Man10] Bassel Mannaa. Cluster editing problem for points on the real line: a polynomial time algorithm. Information Processing Letters, 110(21):961–965, October 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Manurangsi:2019:NDV

[Man19]  Pasin Manurangsi.  A

note on degree vs gap of Min-Rep Label Cover and improved inapproximability for connectivity problems. Information Processing Letters, 145(??):24–29, May 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300043.

Martin:2011:LLD

[Mar11] Barnaby Martin. Low-level dichotomy for quantified constraint satisfaction problems. Information Processing Letters, 111(20):999– 1003, October 31, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901100202X.

Mashhadi:2017:NMS

[Mas17] Samaneh Mashhadi. New multi-stage secret sharing in the standard model. Information Processing Letters, 127(??):43–48, November 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301242.

Massberg:2015:GHC

[Maß15] Jens Maßberg. Generalized Huffman coding for binary trees with choosable edge lengths. Information Processing Letters, 115(4):502– 506, April 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002634.

Matsuki:2012:ACC

[Mat12] Norichika Matsuki. An analytic criterion for CSAT. Information Processing Letters, 112(5):164–165, February 28, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011003152.

Matsubara:2015:EAA

[Mat15] Shunichi Matsubara. An enumeration algorithm for all integers nonrepresentable by some positive integers. Information Processing Letters, 115(2):280–284, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002117.

Matsumoto:2019:DBG

[Mat19] Naoki Matsumoto. The difference between game chromatic number and chromatic number of graphs. Information Processing Letters, 151(??):Article 105835, November 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019301061.

Moezkarimi:2014:PGF

[MB14] Zahra Moezkarimi and Alireza Bagheri. A PTAS for geometric 2-FTP. Information Processing Letters, 114(12):670–675, December 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001409.

Ma:2019:MPI

[MBW19] Gang Ma, Qiuju Bian, and Jianfeng Wang. The maximum PI index of bicyclic graphs with even number of edges. Information Processing Letters, 146(??):13– 16, June 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019300249.

Merouane:2015:SDG

[MC15] Houcine Boumediene Merouane and Mustapha Chellali. On secure domination in graphs. Information Processing Letters, 115(10): 786–790, October 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019015000836.

Malekesmaeili:2012:TBL

[MCS12] Mehrnoush Malekesmaeili, Cedric Chauve, and Tamon Stephen. A tight bound on the length of odd cycles in the incompatibility graph of a non-C1P matrix. Information Processing Letters, 112(20):799–803, October 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012001925.

Maretic:2014:LCU

[MDB14] Grgur Petric Maretić, Mohammad Torabi Dashti, and David Basin. LTL is closed under topological closure. Information Processing Letters, 114(8):408–413, August 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014000386.

Meir:2012:RMP

[Mei12] Or Meir. On the rectangle method in proofs of robustness of tensor products. Information Processing Letters, 112(6):257–260, March 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011003115.

Meir:2018:DSU

[Mei18] Or Meir. The direct sum of universal relations. Information Processing Letters, 136(??):105–111, August 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300899.

Mennink:2012:IFH

[Men12] Bart Mennink. Increasing the flexibility of the herding attack. Information Processing Letters, 112(3):98– 105, January 31, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011002857.

Meshram:2015:EIB

[Mes15] Chandrashekhar Meshram. An efficient ID-based cryptographic encryption based on discrete logarithm problem and integer factorization problem. Information Processing Letters, 115(2):351–358, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002221.

Maiseli:2016:RED

[MG16] Baraka Jacob Maiseli and Huijun Gao. Robust edge detector based on anisotropic diffusion-driven process. Information Processing Letters, 116(5):373– 378, May 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015002148.

Martinez-Garcia:2011:FRB

[MGNAB11]  Carles  Martínez-García,

Guillermo Navarro-Arribas, and Joan Borrell. Fuzzy Role-Based Access Control. Information Processing Letters, 111(10):483–487, April 30, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mihaljevic:2012:ISR

[MGPI12] Miodrag J. Mihaljević, Sugata Gangopadhyay, Goutam Paul, and Hideki Imai. Internal state recovery of keystream generator LILI-128 based on a novel weakness of the employed Boolean function. Information Processing Letters, 112(21):805–810, November 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002013.

Malekimajd:2011:POS

[MHFMSa11] M. Malekimajd, M. R.

Hoseiny-Farahabady, A. Movaghar, and H. Sarbazi-azad. Pancyclicity of OTIS (swapped) networks based on properties of the factor graph. Information Processing Letters, 111(23–24):1114–1119, December 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011002201.

Morinaga:2018:CBN

[MHIU18] Atsushi Morinaga, Kenji Hara, Kohei Inoue, and Kiichi Urahama. Classification between natural and graphics images based on generalized Gaussian distributions. Information Processing Letters, 138(??):27– 30, October 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018301236.

Magdon-Ismail:2017:NHI

[MI17] Malik Magdon-Ismail. NPhardness and inapproximability of sparse PCA. Information Processing Letters, 126(??):35–38, October 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901730100X.

Milshtein:2015:NBC

[Mil15] Moshe Milshtein. A new binary code of length 16 and minimum distance 3. Information Processing Letters, 115(12):975–976, December 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001143.

Milovanov:2019:PCC

[Mil19] Alexey Milovanov. #Pcompleteness of counting roots of a sparse polynomial. Information Processing Letters, 142(??):77–79, February 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018302011.

Mota:2016:PSM

[MIM16] Alexandre Mota, Juliano Iyoda, and Heitor Maranhão. Program synthesis by model finding. Information Processing Letters, 116(11): 701–705, November 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019016300813.

Minh:2011:RLS

[Min11] Ha Quang Minh. The regularized least squares algorithm and the problem of learning halfspaces. Information Processing Letters, 111(8):395–401, March 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Misra:2015:MAN

[Mis15] Jayadev Misra. Mapping among the nodes of infinite trees: a variation of König’s infinity lemma. Information Processing Letters, 115(5): 548–549, May 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015000149.

Miyazaki:2014:ACO

[Miy14] Shuichi Miyazaki. On the advice complexity of online bipartite matching and online stable marriage. Information Processing Letters, 114(12):714–717, December 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001276.

Mahmood:2011:ISI

[MK11] Sajjad Mahmood and Azhar Khan. An industrial study on the importance of software component documentation: a system integrator’s perspective. Information Processing Letters, 111(12): 583–590, June 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011000755.

Mikulski:2014:FHD

[MK14] Lukasz Mikulski and Maciej Koutny. Folded Hasse diagrams of combined traces. Information Processing Letters, 114(4):208– 216, April 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002858.

Manak:2016:EET

[MK16] Martin Manak and Ivana Kolingerova. Extension of the edge tracing algorithm to disconnected Voronoi skeletons. Information Processing Letters, 116(2):85– 92, February 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901500174X.

Matsumoto:2011:AAD

[MKI11] Yusuke Matsumoto, Naoyuki Kamiyama, and Keiko Imai. An approximation algorithm dependent on edgecoloring number for minimum maximal matching problem. Information Processing Letters, 111(10): 465–468, April 30, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Martin:2010:SCE

[ML10a] Cristian Martín and Mikel Larrea. A simple and communication-efficient Omega algorithm in the crashrecovery model. Information Processing Letters, 110 (3):83–87, January 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Miao:2010:SOC

[ML10b] Huifang Miao and Guoping Lin. Strong orientations of complete k-partite graphs achieving the strong diameter. Information Processing Letters, 110(6):206–210, February 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Min:2011:OSO

[MLW11] Xiao Min, Jing Liu, and Yuqing Wang. Optimal semi-online algorithms for scheduling problems with reassignment on two identical machines. Information Processing Letters, 111 (9):423–428, April 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mor:2012:BSI

[MM12] Baruch Mor and Gur Mosheiov. Batch scheduling of identical jobs on parallel identical machines. Information Processing Letters, 112(20):762–766, October 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001718.

Meshram:2013:IBC

[MM13] Chandrashekhar Meshram and Suchitra A. Meshram. An identity-based cryptographic model for discrete logarithm and integer factoring based cryptosystem.

Information Processing Letters, 113(10–11):375–380, May/June 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000550.

Mavronicolas:2015:CPE

[MM15a] Marios Mavronicolas and Burkhard Monien. The complexity of pure equilibria in mix-weighted congestion games on parallel links. Information Processing Letters, 115(12):927–931, December 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001349.

Mor:2015:NMM

[MM15b] Baruch Mor and Gur Mosheiov. A note: Minimizing maximum earliness on a proportionate flowshop. Information Processing Letters, 115(2):253–255, February 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002002.

Madireddy:2019:AHG

[MM19a] Raghunath Reddy Madireddy and Apurva Mudgal. Approximability and hardness of geometric hitting set with axis-parallel rectangles. Information Processing Letters, 141(??):9–15, January 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018301819.

Madireddy:2019:NHG

[MM19b] Raghunath Reddy Madireddy and Apurva Mudgal. NPhardness of geometric set cover and hitting set with rectangles containing a common point. Information Processing Letters, 141(??): 1–8, January 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018301807.

Maiseli:2018:PMM

[MMKK18] Baraka Maiseli, Hubert Msuya, Suzan Kessy, and Michael Kisangiri. Perona– Malik model with selfadjusting shape-defining constant. Information Processing Letters, 137(??):26– 32, September 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019018300966.

Malvestuto:2011:CSP

[MMM11] Francesco M. Malvestuto, Mauro Mezzini, and Marina Moscarini. Computing simple-path convex hulls in hypergraphs. Information Processing Letters, 111(5): 231–234, February 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mancini:2017:MME

[MMM^{+}17] Toni Mancini, Federico Mari, Annalisa Massini, Igor Melatti, Ivano Salvo, and Enrico Tronci. On minimising the maximum expected verification time. Information Processing Letters, 122 (??):8–16, June 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019017300170.

Magirius:2015:CPL

[MMP15] Marco Magirius, Martin Mundhenk, and Raphaela Palenta. The complexity of primal logic with disjunction. Information Processing Letters, 115(5):536– 542, May 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000125.

Marenco:2015:TAN

[MMS15] Javier Marenco, Marcelo Mydlarz, and Daniel Severín. Topological additive numbering of directed acyclic graphs. Information Processing Letters, 115(2):199– 202, February 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014001872.

Meshram:2012:IBC

[MMZ12] Chandrashekhar Meshram, Suchitra A. Meshram, and Mingwu Zhang. An IDbased cryptographic mechanisms based on GDLP and IFP. Information Processing Letters, 112(19):753– 758, October 15, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901200169X.

Mukhopadhyay:2014:EMP

[MN14] Debapriyay Mukhopadhyay and Subhas C. Nandy. Efficient multiple-precision integer division algorithm. Information Processing Letters, 114(3):152–157, March 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019013002627. This

paper provides a correction to the algorithm presented in [HZSL05], and also supplies a complicated correctness proof.

Momtazi:2015:WWC

[MN15] Saeedeh Momtazi and Felix Naumann. Who wants a computer to be a millionaire? Information Processing Letters, 115(4):481– 485, April 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002750.

Mitzenmacher:2016:HPS

[MN16] Michael Mitzenmacher and Vikram Nathan. Hardness of peeling with stashes. Information Processing Letters, 116(11):682–688, November 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300783.

Mercas:2017:NTG

[MN17] Robert Mercas and Dirk Nowotka. A note on Thue games. Information Processing Letters, 118(??):75– 77, February 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301454.

See corrigendum [KMN18].

Makinen:2019:APB

[MN19] Veli Mäkinen and Tuukka Norri. Applying the Positional Burrows–Wheeler Transform to all-pairs Hamming distance. Information Processing Letters, 146(??): 17–19, June 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300262.

Minier:2012:RKI

[MNP12] Marine Minier and María Naya-Plasencia. A related key impossible differential attack against 22 rounds of the lightweight block cipher LBlock. Information Processing Letters, 112(16):624–629, August 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001238.

Mansour:2011:LFG

[MNV11] Toufik Mansour, Ghalib Nassar, and Vincent Vajnovszki. Loop-free Gray code algorithm for the erestricted growth functions. Information Processing Letters, 111(11):541–544, May 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mathieu:2012:LBR

[MO12] Claire Mathieu and Olga Ohrimenko. Lower bounds for randomized algorithms for online chain partitioning. Information Processing Letters, 112(17–18):663–666, September 30, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019012001391.

Mathew:2015:HPB

[MÖ15] K. Ashik Mathew and Patric R. J. Östergård. On hypercube packings, blocking sets and a covering problem. Information Processing Letters, 115(2):141–145, February 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001677.

Mollard:2018:EPC

[Mol18] Michel Mollard. The existence of perfect codes in a family of generalized Fibonacci cubes. Information Processing Letters, 140(??): ii, December 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301571.

Momke:2015:IAA

[Möm15] Tobias Mömke. An improved approximation algorithm for the traveling salesman problem with relaxed triangle inequality. Information Processing Letters, 115(11):866–871, November 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901500099X.

Montanaro:2010:NQQ

[Mon10] Ashley Montanaro. Nonadaptive quantum query complexity. Information Processing Letters, 110(24): 1110–1113, November 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Montanaro:2012:QQC

[Mon12] Ashley Montanaro. The quantum query complexity of learning multilinear polynomials. Information Processing Letters, 112(11): 438–442, June 15, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019012000701.

Monnot:2015:NTS

[Mon15] Jerome Monnot. A note on the traveling salesman reoptimization problem under vertex insertion. Information Processing Letters, 115(3):435–438, March 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002361.

Moody:2011:UIQ

[Moo11] Dustin Moody. Using 5isogenies to quintuple points on elliptic curves. Information Processing Letters, 111 (7):314–317, March 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Morizumi:2011:IAA

[Mor11] Hiroki Morizumi. Improved approximation algorithms for minimum AND-circuits problem via k-set cover. Information Processing Letters, 111(5):218–221, February 1, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Morsy:2016:EHT

[Mor16]  Ehab Morsy.  An ex-

tension of Hall’s theorem for partitioned bipartite graphs. Information Processing Letters, 116(11):706–709, November 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300898.

Moser:2011:ZOS

[Mos11a] Philippe Moser. A zero-one SUBEXP-dimension law for BPP. Information Processing Letters, 111(9):429–432, April 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mosheiov:2011:PFG

[Mos11b] Gur Mosheiov. Proportionate flowshops with general position-dependent processing times. Information Processing Letters, 111 (4):174–177, January 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See note [KY12] and comments [KMS19].

Mosca:2013:MWI

[Mos13a] Raffaele Mosca. Maximum weight independent sets in (P_{6},co-banner)-free graphs. Information Processing Letters, 113(3):89–93, February 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002967.

Moszkowski:2013:IBC

[Mos13b] Ben Moszkowski. Interconnections between classes of sequentially compositional temporal formulas. Information Processing Letters, 113(9):350–353, May 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000501.

Molinero:2016:CE

[MOS16] Xavier Molinero, Martin Olsen, and Maria Serna. On the complexity of exchanging. Information Processing Letters, 116(6):437– 441, June 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016000065.

Mousavi:2019:TTE

[Mou19] Ramin Mousavi. Thin trees in 8-edge-connected planar graphs. Information Processing Letters, 143(??):51– 55, March 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018302321.

Moulton:2017:CTA

[MOW17] Vincent Moulton, James Oldman, and Taoyang Wu. A cubic-time algorithm for computing the trinet distance between level-1 networks. Information Processing Letters, 123(??):36– 41, July 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300388.

Maitra:2013:ESK

[MP13] Arpita Maitra and Goutam Paul. Eavesdropping in semiquantum key distribution protocol. Information Processing Letters, 113(12): 418–422, June 30, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019013000859.

Mesnard:2015:SOF

[MP15] Fred Mesnard and Étienne Payet. A second-order formulation of non-termination. Information Processing Letters, 115(11):882–885, November 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015000897.

Mostefaoui:2019:CTC

[MPRC19] Achour Mostéfaoui, Matthieu Perrin, Michel Raynal, and Jiannong Cao. Crashtolerant causal broadcast in O ( n ) messages. Information Processing Letters, 151(??):Article 105837, November 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019301206.

Menc:2017:TSO

[MPU17] Artur Menc, Dominik Pajak, and Przemyslaw Uznański. Time and space optimality of rotor-router graph exploration. Information Processing Letters, 127(??):17–20, November 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301199.

Moosa:2010:IPB

[MR10] Tanaeem M. Moosa and M. Sohel Rahman. Indexing permutations for binary strings. Information Processing Letters, 110(18– 19):795–798, September 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Montassier:2010:DSG

[MRZ10] Mickael Montassier, André Raspaud, and Xuding Zhu. Decomposition of sparse graphs into two forests, one having bounded maximum degree. Information Processing Letters, 110 (20):913–916, September 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mathew:2012:CPB

[MS12] Sunil Mathew and M. S.

Sunitha. A characterization of partial blocks in weighted graphs. Information Processing Letters, 112 (17–18):706–710, September 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001664.

Morawiecki:2013:SBP

[MS13] Pawel Morawiecki and Marian Srebrny. A SAT-based preimage analysis of reduced Keccak hash functions. Information Processing Letters, 113(10–11):392–397, May/June 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000811.

Mkrtchyan:2015:ALS

[MS15] Vahan Mkrtchyan and K. Subramani. On the approximability of the Largest Sphere Rule Ensemble Classification problem. Information Processing Letters, 115(11):833–837, November 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001088.

Mahajan:2016:LOR

[MS16] Meena Mahajan and Anil Shukla. Level-ordered Qresolution and tree-like Qresolution are incomparable. Information Processing Letters, 116(3):256–258, March 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015002112.

Maneth:2018:BMT

[MS18a] S. Maneth and H. Seidl. Balancedness of MSO transductions in polynomial time. Information Processing Letters, 133(??):26–32, May 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300115.

Marcilon:2018:ITH

[MS18b] Thiago Marcilon and Rudini Sampaio. The P_{3} infection time is W [1]- hard parameterized by the treewidth. Information Processing Letters, 132(??):55– 61, April 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901730220X.

Moser:2018:LDD

[MS18c] Philippe Moser and Frank Stephan. Limit-depth and DNR degrees. Information Processing Letters, 135(??): 36–40, July 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300449.

Meyer:2019:ERV

[MS19] Maxime Meyer and Ben Smyth. Exploiting re-voting in the Helios election system. Information Processing Letters, 143(??):14–19, March 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901830214X.

Matsui:2014:FPF

[MSM14] Tomomi Matsui, Noriyoshi Sukegawa, and Atsushi Miyauchi. Fractional programming formulation for the vertex coloring problem. Information Processing Letters, 114(12):706–709, December 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014001240.

Mahajan:2016:VVM

[MST16] Meena Mahajan, Nitin Saurabh, and Sébastien Tavenas. VNP = VP in the multilinear world. Information Processing Letters, 116(2):179–182, February 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001416.

Muller:2014:UAO

[MSV14] Richard Müller, Christian Stahl, and Walter Vogler. Undecidability of accordance for open systems with unbounded message queues. Information Processing Letters, 114(12):663–669, December 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001483.

Maheshwari:2011:LIN

[MSZ11] Anil Maheshwari, Michiel Smid, and Norbert Zeh. Low-interference networks in metric spaces of bounded doubling dimension. Information Processing Letters, 111(23–24):1120–1123, December 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011002626.

Mubayi:2010:FBS

[MT10] Dhruv Mubayi and György Turán. Finding bipartite subgraphs efficiently. Information Processing Letters, 110(5):174–177, February 1, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Meir:2018:CAP

[MT18] Or Meir and Avishay Tal. The choice and agreement problems of a random function. Information Processing Letters, 133(??):16–20, May 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017302211.

Melkman:2010:DSA

[MTA10] Avraham A. Melkman, Takeyuki Tamura, and Tatsuya Akutsu. Determining a singleton attractor of an AND/OR Boolean network in O(1.587^{n}) time. Information Processing Letters, 110(14–15):565–569, July 1, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moll:2012:CHW

[MTT12] Lukas Moll, Siamak Tazari, and Marc Thurley. Computing hypergraph width measures exactly. Information Processing Letters, 112(6):238–242, March 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011003243.

Mapa:2015:MAS

[MU15] Sílvia Maria Santana Mapa and Sebastián Urrutia. On the maximum acyclic subgraph problem under disjunctive constraints. Information Processing Letters, 115(2):119–124, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001513.

Mansour:2013:EGR

[MV13] Toufik Mansour and Vincent Vajnovszki. Efficient generation of restricted growth words. Information Processing Letters, 113(17):613–616, August 30, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001427.

Ma:2012:PAE

[MWZ12] Beibei Ma, Baoyindureng Wu, and Wanping Zhang. Proximity and average eccentricity of a graph. Information Processing Letters, 112(10):392–395, May 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000427.

Ma:2013:OSS

[MY13] Ran Ma and Jinjiang Yuan. Online scheduling on a single machine with rejection under an agreeable condition to minimize the total completion time plus the total rejection cost. Information Processing Letters, 113(17):593–598, August 30, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001403.

Ma:2018:NST

[MY18] Fei Ma and Bing Yao. The number of spanning trees of a class of self-similar fractal models. Information Processing Letters, 136(??):64– 69, August 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901830084X.

See corrected proof [EAE21].

Ma:2011:SCE

[MZ11] Meijie Ma and Liying Zhu. The super connectivity of exchanged hypercubes. Information Processing Letters, 111(8):360–364, March 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Meng:2015:CRS

[MZ15] Xianmeng Meng and Xuexin Zheng. Cryptanalysis of RSA with a small parameter revisited. Information Processing Letters, 115(11):858–862, November 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901500109X.

Miao:2011:BPB

[MZC11] Cuixia Miao, Yuzhong Zhang, and Zhigang Cao. Bounded parallel-batch scheduling on single and multi machines for deteriorating jobs. Information Processing Letters, 111(16):798–803, August 30, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001529.

Ma:2014:CRI

[MZQL14] Ying Ma, Shunzhi Zhu, Ke Qin, and Guangchun Luo. Combining the requirement information for software defect estimation in design time. Information Processing Letters, 114(9):469– 474, September 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014000556.

Neves:2019:ONB

[NA19] Samuel Neves and Filipe Araujo. An observation on NORX, BLAKE2, and ChaCha. Information Processing Letters, 149(??):1–5, September 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019300821.

Nofal:2019:CSI

[NAD19] Samer Nofal, Katie Atkinson, and Paul E. Dunne. On checking skeptical and ideal admissibility in abstract argumentation frameworks. Information Processing Letters, 148(??):7– 12, August 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019300705.

Nanongkai:2013:SFS

[Nan13] Danupon Nanongkai. Simple FPTAS for the subsetsums ratio problem. Information Processing Letters, 113(19–21):750–753, September/October 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0020019013001981.

Nandi:2015:CDF

[Nan15] Mrinal Nandi. On commutativity of Discrete Fourier Transform. Information Processing Letters, 115(10): 779–785, October 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015000769.

Nakade:2012:EIE

[NB12] Apurv Nakade and Somenath Biswas. Effect of increasing the energy gap between the two lowest energy states on the mixing time of the Metropolis algorithm. Information Processing Letters, 112(23):922–927, December 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002311.

Niehues:2018:FMR

[NCP18] L. Boppre Niehues, R. Custódio, and D. Panario. Fast modular reduction and squaring in GF(2^{m}). Information Processing Letters, 132(??): 33–38, April 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017302168.

Nederlof:2017:SNM

[Ned17] Jesper Nederlof. A short note on Merlin–Arthur protocols for subset sum. Information Processing Letters, 118(??):15–16, February 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016301296.

Ning:2015:CEC

[NFW15] Wantao Ning, Xiaoli Feng, and Ling Wang. The connectivity of exchanged crossed cube. Information Processing Letters, 115(2):394–396, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014002324.

Nash:2010:OSA

[NG10] Nicholas Nash and David Gregg. An output sensitive algorithm for computing a maximum independent set of a circle graph. Information Processing Letters, 110 (16):630–634, July 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Neggazi:2015:NSS

[NHK15] Brahim Neggazi, Mohammed Haddad, and Hamamache Kheddouci. A new selfstabilizing algorithm for maximal p-star decomposition of general graphs. Information Processing Letters, 115(11):892–898, November 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000873.

Nicholson:2019:REA

[Nic19] Patrick K. Nicholson. Revisiting explicit adaptive twoprobe schemes. Information Processing Letters, 143(??): 1–3, March 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018302035.

Nakashima:2015:CLT

[NII^{+}15] Yuto Nakashima, Tomohiro I., Shunsuke Inenaga, Hideo Bannai, and Masayuki Takeda. Constructing LZ78 tries and position heaps in linear time for large alphabets. Information Processing Letters, 115(9):655–659, September 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000629.

Ning:2013:FTD

[Nin13] Bo Ning. Fan-type degree condition restricted to triples of induced subgraphs ensuring Hamiltonicity. Information Processing Letters, 113(19– 21):823–826, September/ October 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013002032.

Ning:2015:NCM

[Nin15] Bo Ning. Notes on a conjecture of Manoussakis concerning Hamilton cycles in digraphs. Information Processing Letters, 115(2):221– 224, February 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019014001914.

Ning:2016:SCE

[Nin16] Wantao Ning. The super connectivity of exchanged crossed cube. Information Processing Letters, 116(2): 80–84, February 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019015001751.

Nisan:1992:PGS

[Nis92]  N. Nisan.  Pseudoran-

dom generators for spacebounded computation. Combinatorica, 12(4):449–461, December 1992. CODEN COMBDI. ISSN 0209-9683 (print), 1439-6912 (electronic). See comments [DPS11].

Ni:2015:OOM

[NLXX15] Guanqun Ni, Li Luo, Yinfeng Xu, and Jiuping Xu. Optimal online markdown and markup pricing policies with demand uncertainty. Information Processing Letters, 115(11):804– 811, November 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015001027.

Nie:2014:IMC

[NLZX14] Lei Nie, Jingjie Liu, Haicang Zhang, and Zhiwei Xu. On the inapproximability of minimizing cascading failures under the deterministic threshold model. Information Processing Letters, 114(1–2):1–4, January/ February 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002640.

Nandy:2010:RLE

[NMB10] Subhas C. Nandy, Krishnendu Mukhopadhyaya, and Bhargab B. Bhattacharya. Recognition of largest empty orthoconvex polygon in a point set. Information Processing Letters, 110(17):746–752, August 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nakprasit:2017:SEV

[NN17] Keaitsuda Maneeruk Nakprasit and Kittikorn Nakprasit. The strong equitable vertex 2-arboricity of complete bipartite and tripartite graphs. Information Processing Letters, 117(??):40– 44, January 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016301272.

Nose:2011:SWA

[Nos11] Peter Nose. Security weaknesses of authenticated key agreement protocols. Information Processing Letters, 111(14):687–696, July 31, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011001074.

Nose:2014:SWS

[Nos14]  Peter Nose.  Security

weaknesses of a signature scheme and authenticated key agreement protocols. Information Processing Letters, 114(3):107– 115, March 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002731.

Nandy:2017:FAM

[NPR17] Subhas C. Nandy, Supantha Pandit, and Sasanka Roy. Faster approximation for maximum independent set on unit disk graph. Information Processing Letters, 127(??):58–61, November 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301291.

Nicolae:2017:PMM

[NR17] Marius Nicolae and Sanguthevar Rajasekaran. On pattern matching with k mismatches and few don’t cares. Information Processing Letters, 118(??):78–82, February 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301442.

Narayanaswamy:2011:DSB

[NS11] N. S. Narayanaswamy and C. R. Subramanian. Dominating set based exact algorithms for 3-coloring. Information Processing Letters, 111(6):251–255, February 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Norouzi:2014:SEC

[NT14] Nima Norouzi and Yung H. Tsin. A simple 3-edge connected component algorithm revisited. Information Processing Letters, 114(1–2):50–55, January/ February 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013002470.

Nguyen:2016:EFV

[NTD16] Minh Tan Nguyen, Quang Hai Truong, and Tran Khanh Dang. Enhance fuzzy vault security using nonrandom chaff point generator. Information Processing Letters, 116(1):53–64, January 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001568.

Nutov:2018:IAA

[Nut18] Zeev Nutov. Improved approximation algorithms for k-connected m-dominating set problems. Information Processing Letters, 140(??): 25–29, December 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019018301601.

Nanasi:2015:SAH

[NVB15] Michal Nánási, Tomás Vinar, and Brona Brejová. Sequence annotation with HMMs: New problems and their complexity. Information Processing Letters, 115(6–8):635–639, June/August 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901500040X.

Xie:2014:NBP

[nXlCL14] Xiao ning Xie, Gong liang Chen, and Yin Li. Novel bit-parallel multiplier for GF(2^{m}) defined by allone polynomial using generalized Karatsuba algorithm. Information Processing Letters, 114(3):140– 146, March 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002664.

Ni:2019:CAO

[NZX19] Guanqun Ni, Feifeng Zheng, and Yinfeng Xu. Competitive analysis of online revenue management with hierarchical resources. Information Processing Letters, 142(??):41–45, February 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018302023.

Oanea:2011:BAN

[Oan11] Olivia Oanea. Boundedness of adaptive nets is decidable. Information Processing Letters, 111(15):738–743, August 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011001244.

Olsen:2012:NTN

[OBT12] Martin Olsen, Lars Bækgaard, and Torben Tambo. On non-trivial Nash stable partitions in additive hedonic games with symmetric 0/1-utilities. Information Processing Letters, 112(23):903–907, December 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012002438.

Ochem:2017:SNC

[Och17] Pascal Ochem. 2-subcoloring is NP-complete for planar comparability graphs. Information Processing Letters, 128(??):46–48, December 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301461.

Ohlebusch:2010:EAA

[OG10] Enno Ohlebusch and Simon Gog. Efficient algorithms for the all-pairs suffix-prefix problem and the all-pairs substring-prefix problem. Information Processing Letters, 110(3):123–128, January 1, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ortin:2011:UIT

[OG11] Francisco Ortin and Miguel García. Union and intersection types to support both dynamic and static typing. Information Processing Letters, 111(6):278–286, February 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Oh:2013:IBA

[OKM13] Jinsoo Oh, Jin Kim, and Byung-Ro Moon. On the inequivalence of bilinear algorithms for 3 × 3 matrix multiplication. Information Processing Letters, 113(17):640–645, August 30, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001531.

Ochem:2017:OEC

[OM17] Pascal Ochem and Nazanin Movarraei. Oriented, 2-edge-colored, and 2vertex-colored homomorphisms. Information Processing Letters, 123(??):42– 46, July 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017300364.

Orumiehchiha:2014:PAN

[OPS14] Mohammad Ali Orumiehchiha, Josef Pieprzyk, and Ron Steinfeld. Practical attack on NLM–MAC scheme. Information Processing Letters, 114(10):547–550, October 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014000696.

Oron:2011:SBM

[Oro11]  Daniel Oron.  Schedul-

ing a batching machine with convex resource consumption functions. Information Processing Letters, 111(19):962–967, October 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001918.

Ou:2016:FAS

[OZL16] Jinwen Ou, Xueling Zhong, and Chung-Lun Li. Faster algorithms for single machine scheduling with release dates and rejection. Information Processing Letters, 116(8):503–507, August 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901630014X.

Pallo:2010:RRK

[Pal10] Jean Marcel Pallo. Rootrestricted Kleenean rotations. Information Processing Letters, 111(1):46– 50, December 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Park:2011:CGR

[Par11] Moonju Park. Comments on “Generalized rate monotonic schedulability bounds using relative period ratios”. Information Processing Letters, 111(7):334–337, March 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [WLLS08].

Parys:2016:WCP

[Par16] Pawel Parys. Weak containment for partial words is coNP-complete. Information Processing Letters, 116(2):125–127, February 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001672.

Pasalic:2015:CNN

[Pas15a] E. Pasalic. Corrigendum to “A note on nonexistence of vectorial bent functions with binomial trace representation in the PS^{−} class” [Information Processing Letters 115 (2) (2015) 139–140]. Information Processing Letters, 115(4):520, April 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002762.

See [Pas15b].

Pasalic:2015:NNV

[Pas15b] E. Pasalic. A note on nonexistence of vectorial bent functions with binomial trace representation in the PS^{−} class. Information Processing Letters, 115:139– 140, February 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014001550.

See corrigendum [Pas15a].

Papadopoulos:2018:FAG

[PC18] Kleitos Papadopoulos and Demetres Christofides. A fast algorithm for the gas station problem. Information Processing Letters, 131(??):55–59, March 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002001901730203X.

Pai:2019:IDC

[PC19] Kung-Jui Pai and Jou-Ming Chang. Improving the diameters of completely independent spanning trees in locally twisted cubes. Information Processing Letters, 141(??):22–24, January 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301844.

Park:2013:PMT

[PCH13] Sun-Mi Park, Ku-Young Chang, and Dowon Hong. Parallel GF(3^{m}) multiplier for trinomials. Information Processing Letters, 113(4):111–115, February 28, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012003055.

Park:2010:VSA

[PCK10] Jonghun Park, Byung-Cheon Choi, and Kwanho Kim. A vector space approach to tag cloud similarity ranking. Information Processing Letters, 110(12– 13):489–496, June 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pai:2016:VTF

[PCY16] Kung-Jui Pai, Jou-Ming Chang, and Jinn-Shyong Yang. Vertex-transitivity on folded crossed cubes. Information Processing Letters, 116(11):689–693, November 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016300795.

Panday:2018:FWT

[PdAL18] Deepak Panday, Renato Cordeiro de Amorim, and Peter Lane. Feature weighting as a tool for unsupervised feature selection. Information Processing Letters, 129(??):44– 52, January 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017301618.

Pearce:2016:SEA

[Pea16] David J. Pearce. A spaceefficient algorithm for finding strongly connected components. Information Processing Letters, 116(1):47– 52, January 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001532.

Pelc:2010:FTS

[Pel10] Andrzej Pelc. Fault-tolerant strategies in the Iterated Prisoner’s Dilemma. Information Processing Letters, 110(10):389–395, April 30, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Peltomaki:2015:CRS

[Pel15] Jarkko Peltomäki. Characterization of repetitions in Sturmian words: a new proof. Information Processing Letters, 115(11):886– 891, November 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015000885.

Pelc:2018:RTP

[Pel18] Andrzej Pelc. Reaching a target in the plane with no information. Information Processing Letters, 140(??): 8–12, December 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019018300863.

Perez:2017:FIC

[Pér17] Guillermo A. Pérez. The fixed initial credit problem for partial-observation energy games is Ack-complete. Information Processing Letters, 118(??):91–99, February 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016301466.

Pettie:2012:SRM

[Pet12] S. Pettie. A simple reduction from maximum weight matching to maximum cardinality matching. Information Processing Letters, 112(23):893–898, December 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012002293.

Panda:2012:LDC

[PG12] B. S. Panda and Preeti Goel. L(2, 1)-labeling of dually chordal graphs and strongly orderable graphs. Information Processing Letters, 112(13):552–556, July 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001056.

Pasalic:2019:DMS

[PGZB19] E. Pasalic, S. Gangopadhyay, W.-G. Zhang, and S. Bajrić. Design methods for semi-bent functions. Information Processing Letters, 143(??):61– 70, March 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018302345.

Puente:2011:NRL

[PH11] Víctor Lavín Puente and Montserrat Hermo. Negative results on learning multivalued dependencies with queries. Information Processing Letters, 111(19):968–972, October 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001992.

Park:2013:SST

[PI13] Jung-Heum Park and Insung Ihm. Single-source three-disjoint path covers in cubes of connected graphs. Information Processing Letters, 113(14–16):527–532, July/August 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001294.

Park:2019:LTA

[PI19] Jung-Heum Park and Insung Ihm. A linear-time algorithm for finding a one-to-many 3-disjoint path cover in the cube of a connected graph. Information Processing Letters, 142(??):57– 63, February 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018302047.

Piestrak:2015:NRA

[Pie15] Stanislaw J. Piestrak. A note on RNS architectures for the implementation of the diagonal function. Information Processing Letters, 115(4):453–457, April 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014002671.

Pang:2013:IMA

[PLPW13] Liaojun Pang, Huixian Li, Qingqi Pei, and Yumin Wang. Improvement on Meshram et al.’s IDbased cryptographic mechanism. Information Processing Letters, 113(19– 21):789–792, September/ October 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002081.

Podolskii:2012:ELB

[Pod12] Vladimir V. Podolskii. Exponential lower bound for bounded depth circuits with few threshold gates. Information Processing Letters, 112(7):267–271, March 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011003504.

Polak:2018:WIH

[Pol18] Adam Polak. Why is it hard to beat O(n^{2}) for longest common weakly increasing subsequence? Information Processing Letters, 132(??): 1–5, April 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017302016.

Panda:2010:LCS

[PP10] B. S. Panda and D. Pradhan. Locally connected spanning trees in cographs, complements of bipartite graphs and doubly chordal graphs. Information Processing Letters, 110(23): 1067–1073, November 15, 2010. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Panda:2013:LTA

[PP13] B. S. Panda and S. Paul. A linear time algorithm for liar’s domination problem in proper interval graphs. Information Processing Letters, 113(19– 21):815–822, September/ October 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002019.

Park:2014:EUB

[PP14] Moonju Park and Xuefeng Piao. Enhanced utilization bound of rate-monotonic scheduling in Controller Area Networks. Information Processing Letters, 114(11):603–606, November 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001100.

Pei:2018:SII

[PPH18] Li-Dan Pei, Xiang-Feng Pan, and Fu-Tao Hu. Some improved inequalities related to Vizing’s conjecture. Information Processing Letters, 135(??):85–91, July 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300565.

Peng:2017:ODF

[PPN^{+}17] Weiguang Peng, NingNing Peng, KengMeng Ng, Kazuyuki Tanaka, and Yue Yang. Optimal depth-first algorithms and equilibria of independent distributions on multi-branching trees. Information Processing Letters, 125(??):41–45, September 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300881.

Ponti:2017:ILT

[PR17] Moacir Ponti and Mateus Riva. An incremental lineartime learning algorithm for the Optimum-Path Forest classifier. Information Processing Letters, 126(??):1– 6, October 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901730090X.

Pradhan:2012:AAT

[Pra12] D. Pradhan. Algorithmic aspects of k-tuple total domination in graphs. Information Processing Letters, 112(21):816–822, November 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012001986.

Pontarelli:2014:IPI

[PRM14] Salvatore Pontarelli, Pedro Reviriego, and Michael Mitzenmacher. Improving the performance of Invertible Bloom Lookup Tables. Information Processing Letters, 114(4):185–191, April 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013002950.

Pontarelli:2016:ICB

[PRM16] Salvatore Pontarelli, Pedro Reviriego, and Juan Antonio Maestro. Improving counting Bloom filter performance with fingerprints. Information Processing Letters, 116(4):304– 309, April 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001866.

Paulusma:2019:PCK

[PS19a] Daniël Paulusma and Stefan Szeider. On the parameterized complexity of (k, s)- SAT. Information Processing Letters, 143(??):34–36, March 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901830228X.

Polak:2019:NLB

[PS19b] Sven C. Polak and Alexander Schrijver. New lower bound on the Shannon capacity of C_{7} from circular graphs. Information Processing Letters, 143(??):37– 40, March 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018302291.

Policriti:2011:CEA

[PT11] Alberto Policriti and Alexandru I. Tomescu. Counting extensional acyclic digraphs. Information Processing Letters, 111(16):787–791, August 30, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001426.

Pagh:2012:CTC

[PT12] Rasmus Pagh and Charalampos E. Tsourakakis. Colorful triangle counting and a MapReduce implementation. Information Processing Letters, 112(7):277–281, March 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011003358.

Pudlak:2012:LBS

[Pud12] Pavel Pudlák. A lower bound on the size of resolution proofs of the Ramsey theorem. Information Processing Letters, 112 (14–15):610–611, August 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012001275.

Puente:2011:LSQ

[Pue11] Víctor Lavín Puente. Learning a subclass of k-quasi-Horn formulas with membership queries. Information Processing Letters, 111 (11):550–555, May 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Puleo:2016:CDM

[Pul16] Gregory J. Puleo. Complexity of a disjoint matching problem on bipartite graphs. Information Processing Letters, 116(10):649–652, October 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016300771.

Pang:2015:TSD

[PWC^{+}15] Chaoyi Pang, Junhu Wang, Yu Cheng, Haolan Zhang, and Tongliang Li. Topological sorts on DAGs. Information Processing Letters, 115(2):298–301, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002142.

Peng:2010:EIA

[PYHA10] Yung-Hsing Peng, Chang-Biau Yang, Kuo-Si Huang, and Hsing-Yen Ann. Efficient indexing algorithms for one-dimensional discretelyscaled strings. Information Processing Letters, 110 (16):730–734, July 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Peng:2011:IOD

[PYTH11] Yung-Hsing Peng, Chang-Biau Yang, Chiou-Ting Tseng, and Chiou-Yi Hor. The indexing for onedimensional proportionallyscaled strings. Information Processing Letters, 111(7): 318–322, March 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Peng:2013:FGP

[PYW^{+}13] Xi Peng, Zhang Yi, Xiao-Yong Wei, De-Zhong Peng, and Yong-Sheng Sang. Freegram phrase identification for modeling Chinese text. Information Processing Letters, 113(4):137–144, February 28, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901200302X.

Park:2016:TCB

[PYYC16] Jong-Min Park, Hyung-Jeong Yang, Jae-Dong Yang, and Dong-Hoon Choi. A technique for the conceptbased detection of functional modules in an interaction network. Information Processing Letters, 116(10): 611–617, October 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019016300631.

Pasalic:2012:MOB

[PZ12] E. Pasalic and W. G. Zhang. On multiple output bent functions. Information Processing Letters, 112(21):811–815, November 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001974.

Qin:2016:VTQ

[QD16] Huawang Qin and Yuewei Dai. Verifiable (t, n) threshold quantum secret sharing using d-dimensional Bell state. Information Processing Letters, 116(5):351– 355, May 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016000077.

Qi:2017:BPO

[Qi17] Hao Qi. Bounds on partial online list colouring. Information Processing Letters, 127(??):23–26, November 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301102.

Queyroi:2015:SDC

[QK15] François Queyroi and Sergey Kirgizov. Suppression distance computation for hierarchical clusterings. Information Processing Letters, 115(9):689–693, September 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000678.

Qian:2010:NIM

[QX10] Haifeng Qian and Shouhuai Xu. Non-interactive multisignatures in the plain public-key model with efficient verification. Information Processing Letters, 111(2):82–89, December 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Qin:2016:STI

[QYWX16] Zhen Qin, Chen Yuan, Yilei Wang, and Hu Xiong. On the security of two identitybased signature schemes based on pairings. Information Processing Letters, 116 (6):416–418, June 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019016300096.

Rackham:2010:CCC

[Rac10] Tom Rackham. The complexity of changing colourings with bounded maximum degree. Information Processing Letters, 110(17): 735–739, August 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rajasingh:2012:MWH

[RARM12] Indra Rajasingh, Micheal Arockiaraj, Bharati Rajan, and Paul Manuel. Minimum wirelength of hypercubes into n-dimensional grid networks. Information Processing Letters, 112(14–15):583– 586, August 15, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901200110X.

Rangasamy:2013:IFS

[RAT13] Parvathi Rangasamy, Muhammad Akram, and S. Thilagavathi. Intuitionistic fuzzy shortest hyperpath in a network. Information Processing Letters, 113(17):599– 603, August 30, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013001385.

Ren:2013:CTS

[RDX13] Jianfeng Ren, Donglei Du, and Dachuan Xu. The complexity of two supply chain scheduling problems. Information Processing Letters, 113(17):609–612, August 30, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001397.

Roh:2010:BSW

[RH10] Dongyoung Roh and Sang Geun Hahn. On the bit security of the weak Diffie– Hellman problem. Information Processing Letters, 110 (18–19):799–802, September 15, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rial:2017:IFO

[Ria17] Alfredo Rial. Issuer-free oblivious transfer with access control revisited. Information Processing Letters, 126(??):7–11, October 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017300923.

Rial:2019:CAS

[Ria19] Alfredo Rial. A conditional access system with revocation for mobile pay-TV systems revisited. Information Processing Letters, 147(??): 6–9, July 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901930047X.

Roh:2010:LVV

[RJS^{+}10] Hyun-Gul Roh, Myeongjae Jeon, Euiseong Seo, Jinsoo Kim, and Joonwon Lee. Log’ version vector: Logging version vectors concisely in dynamic replication. Information Processing Letters, 110(14–15):614– 620, July 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ristov:2015:USS

[RK15] Strahil Ristov and Damir Korencić. Using static suffix array in dynamic application: Case of text compression by longest first substitution. Information Processing Letters, 115(2):175–181, February 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001835.

Reis:2018:AET

[RM18] Jacinto Reis and Alexandre Mota. Aiding exploratory testing with pruned GUI models. Information Processing Letters, 133(??):49– 55, May 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300176.

Roman:2011:NCR

[Rom11] Adam Roman. The NPcompleteness of the Road Coloring Problem. Information Processing Letters, 111 (7):342–347, March 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reviriego:2019:CCF

[RPL19] Pedro Reviriego, Salvatore Pontarelli, and Gil Levy. CuCoTrack: Cuckoo filter based connection tracking. Information Processing Letters, 147(??):55–60, July 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019300651.

Roayaei:2016:FAM

[RR16] Mehdy Roayaei and Mohammadreza Razzazi. An FPT-algorithm for modifying a graph of bounded treewidth to decrease the size of its dominating set using minimum modification. Information Processing Letters, 116(9):590–594, September 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016300448.

Rajasingh:2012:EHN

[RRR12] Indra Rajasingh, Bharati Rajan, and R. Sundara Rajan. Embedding of hypercubes into necklace, windmill and snake graphs. Information Processing Letters, 112(12):509–515, June 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012000749.

Rampersad:2010:DPF

[RS10] Narad Rampersad and Jeffrey Shallit. Detecting patterns in finite regular and context-free languages. Information Processing Letters, 110(3):108–112, January 1, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reidenbach:2012:MHA

[RS12a] Daniel Reidenbach and Markus L. Schmid. On multi-head automata with restricted nondeterminism. Information Processing Letters, 112(14–15):572–577, August 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001111.

Rueda:2012:PAP

[RS12b] Sonia L. Rueda and Juana Sendra. On the performance of the approximate parametrization algorithm for curves. Information Processing Letters, 112(5):172– 178, February 28, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011003127.

Roetteler:2015:NQR

[RS15] Martin Roetteler and Rainer Steinwandt. A note on quantum related-key attacks. Information Processing Letters, 115(1):40– 44, January 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014001719.

Rinemberg:2019:EDS

[RS19] Martín Rinemberg and Francisco J. Soulignac. The eternal dominating set problem for interval graphs. Information Processing Letters, 146(??):27– 29, June 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019300213.

Rampersad:2011:ISB

[RSwW11] Narad Rampersad, Jeffrey Shallit, and Ming wei Wang. Inverse star, borders, and palstars. Information Processing Letters, 111(9):420– 422, April 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rizzi:2013:RUR

[RT13] Romeo Rizzi and Alexandru I. Tomescu. Ranking, unranking and random generation of extensional acyclic digraphs. Information Processing Letters, 113(5–6):183–187, March 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000161.

Rudow:2017:DLM

[Rud17] M. Rudow. Discrete logarithm and minimum circuit size. Information Processing Letters, 128(??):1–4, December 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301278.

Rusu:2018:SSP

[Rus18]  Irena Rusu.  Sorting

signed permutations by reversals using link-cut trees. Information Processing Letters, 132(??):44–48, April 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017302193.

Radonjic:2010:ISD

[RV10] Aleksandar Radonjic and Vladimir Vujicic. Integer SEC–DED codes for low power communications. Information Processing Letters, 110(12–13):518–520, June 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See comments [BB11].

Rotter:2013:DAR

[RV13] Daniel Rotter and Jens Vygen. d-dimensional arrangement revisited. Information Processing Letters, 113(13): 498–505, July 15, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013001142.

Radonjic:2017:ICC

[RV17] Aleksandar Radonjic and Vladimir Vujicic. Integer codes correcting single errors and burst asymmetric errors within a byte. Information Processing Letters, 121(??):45–50, May 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300169.

Rodrigues:2017:NCC

[RX17] Félix Carvalho Rodrigues and Eduardo Candido Xavier. Non-cooperative capacitated facility location games. Information Processing Letters, 117(??):45–53, January 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301284.

Roussel:2010:TCP

[RZ10] Nicolas Roussel and Xuding Zhu. Total coloring of planar graphs of maximum degree eight. Information Processing Letters, 110(8–9):321– 324, April 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rovan:2014:MTC

[RZ14] Branislav Rovan and Marek Zeman. Modeling time criticality of information. Information Processing Letters, 114(3):147–151, March 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019013002652.

Rzazewski:2014:EAG

[Rza14] Pawel Rzazewski. Exact algorithm for graph homomorphism and locally injective graph homomorphism. Information Processing Letters, 114(7):387– 391, July 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014000362.

Sasikala:2017:NAF

[SaBG17] S. Sasikala, S. Appavu alias Balamurugan, and S. Geetha. A novel adaptive feature selector for supervised classification. Information Processing Letters, 117(??):25–34, January 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301119.

Salmela:2012:ACB

[Sal12] Leena Salmela. Average complexity of backward qgram string matching algorithms. Information Processing Letters, 112(11): 433–437, June 15, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019012000592.

Sandeep:2011:PCG

[San11] R. B. Sandeep. Perfectly colorable graphs. Information Processing Letters, 111(19):960–961, October 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001876.

Sarkar:2011:TES

[Sar11] Palash Sarkar. Tweakable enciphering schemes using only the encryption function of a block cipher. Information Processing Letters, 111(19):945–955, October 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001852.

Sarkar:2014:PEK

[Sar14] Santanu Sarkar. Proving empirical key-correlations in RC4. Information Processing Letters, 114(5):234–238, May 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014000039.

Saxena:2010:FFC

[Sax10] Sanjeev Saxena. On finding fundamental cut sets. Information Processing Letters, 110(4):168–170, January 16, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Suttichaya:2013:SLP

[SB13] Vasin Suttichaya and Pattarasinee Bhattarakosol. Solving the learning parity with noise’s open question. Information Processing Letters, 113(14–16):562– 566, July/August 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013001336.

Sadeghi:2019:SAS

[SB19] Sadegh Sadeghi and Nasour Bagheri. Security analysis of SIMECK block cipher against related-key impossible differential. Information Processing Letters, 147(??): 14–21, July 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300468.

Sergey:2012:CBT

[SC12] Ilya Sergey and Dave Clarke. A correspondence between type checking via reduction and type checking via evaluation. Information Processing Letters, 112(1–2):13–20, January 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011002791.

Steyaert:2013:BFI

[SCB13] Bart Steyaert, Dieter Claeys, and Herwig Bruneel. The burst factor and its impact on buffer performance. Information Processing Letters, 113(18): 681–684, September 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019013001762.

Schmid:2013:NCM

[Sch13a]  Markus L. Schmid.  A

note on the complexity of matching patterns with variables. Information Processing Letters, 113(19– 21):729–733, September/ October 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001786.

Schmidt:2013:STV

[Sch13b] Jens M. Schmidt. A simple test on 2-vertex- and 2-edge-connectivity. Information Processing Letters, 113(7):241–244, April 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000288.

Schewe:2014:DSS

[Sch14] Sven Schewe. Distributed synthesis is simply undecidable. Information Processing Letters, 114(4):203– 207, April 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002925.

Schaudt:2015:DMI

[Sch15] Oliver Schaudt. On disjoint maximal independent sets in graphs. Information Processing Letters, 115(1):23– 27, January 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001720.

Soraluze:2011:CEF

[SCL^{+}11] Iratxe Soraluze, Roberto Cortiñas, Alberto Lafuente, Mikel Larrea, and Felix Freiling. Communicationefficient failure detection and consensus in omission environments. Information Processing Letters, 111(6): 262–268, February 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sokolov:2013:LLR

[SD13] A. V. Sokolov and A. V. Drac. The linked list representation of n LIFOstacks and/or FIFO-queues in the single-level memory. Information Processing Letters, 113(19– 21):832–835, September/ October 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901300210X.

Sanyal:2019:SSA

[SD19] Deepayan Sanyal and Swagatam Das. On semisupervised active clustering of stable instances with oracles. Information Processing Letters, 151(??): Article 105833, November 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019301048.

Shakiba:2014:CCI

[SDM14] Mohsen Shakiba, Mohammad Dakhilalian, and Hamid Mala. On computational complexity of impossible differential cryptanalysis. Information Processing Letters, 114(5):252–255, May 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013003116.

Schubert:2014:NSR

[SeF14] Aleksy Schubert and Ken etsu Fujita. A note on subject reduction in (→, ∃)- Curry with respect to complete developments. Information Processing Letters, 114(1–2):72–75, January/ February 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002160.

Service:2010:NFL

[Ser10] Travis C. Service. A No Free Lunch theorem for multiobjective optimization. Information Processing Letters, 110(21):917–923, October 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Su:2012:IIN

[SF12] Chen Su and Haining Fan. Impact of Intel’s new instruction sets on software implementation of GF(2)[x] multiplication. Information Processing Letters, 112(12): 497–502, June 30, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019012000804.

Sun:2013:ISO

[SF13] Hongyang Sun and Rui Fan. Improved semi-online makespan scheduling with a reordering buffer. Information Processing Letters, 113(12):434–439, June 30, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013000847.

Supeesun:2017:LNS

[SF17] Adisak Supeesun and Jittat Fakcharoenphol. Learning network structures from contagion. Information Processing Letters, 121(??):11– 16, May 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901730011X.

Sanyal:2013:CTR

[SGM13] Biswajit Sanyal, Prosenjit Gupta, and Subhashis Majumder. Colored top-K range-aggregate queries. Information Processing Letters, 113(19– 21):777–784, September/ October 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002020.

Sawada:2017:FLF

[SH17] Joe Sawada and Patrick Hartman. Finding the largest fixed-density necklace and Lyndon word. Information Processing Letters, 125(??):15–19, September 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300698.

Shafiei:2011:PPO

[SHFMSA11] T. Shafiei, M. R. Hoseiny-Farahabady, A. Movaghar, and H. Sarbazi-Azad. On pancyclicity properties of OTIS-mesh. Information Processing Letters, 111(8):

353–359, March 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shi:2012:SSA

[Shi12]  Zhengnan Shi.  A self-

stabilizing algorithm to maximal 2-packing with improved complexity. Information Processing Letters, 112(13):525–531, July 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012000968.

Shin:2013:NMS

[Shi13] Chan-Su Shin. A note on minimum-sum coverage by aligned disks. Information Processing Letters, 113 (22–24):871–875, November/December 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019013002251.

Shin:2015:TED

[Shi15] Kilho Shin. Tree Edit Distance and Maximum Agreement Subtree. Information Processing Letters, 115(1): 69–73, January 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019014001768.

Shitov:2018:CGC

[Shi18] Yaroslav Shitov. On the complexity of graph coloring with additional local conditions. Information Processing Letters, 135(??):92– 94, July 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300644.

Shparlinski:2013:CNE

[Shp13] Igor E. Shparlinski. Correcting noisy exponentiation black-boxes modulo a prime. Information Processing Letters, 113(12):414– 417, June 30, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000999.

Shraibman:2018:NMC

[Shr18] Adi Shraibman. A note on multiparty communication complexity and the Hales–Jewett theorem. Information Processing Letters, 139(??):39–43, November 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301467.

Shraibman:2019:CBL

[Shr19] Adi Shraibman. The corruption bound, log-rank, and communication complexity. Information Processing Letters, 141(??):16– 21, January 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301820.

Sattari:2018:UBM

[SI18] Sattar Sattari and Mohammad Izadi. Upper bounds for minimum dilation triangulation in two special cases. Information Processing Letters, 133(??):33– 38, May 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300103.

Siebertz:2019:GDB

[Sie19] Sebastian Siebertz. Greedy domination on biclique-free graphs. Information Processing Letters, 145(??):64– 67, May 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300080.

Simon:2016:ECA

[Sim16] Hans Ulrich Simon. Efficient computation of approximate isomorphisms between Boolean functions. Information Processing Letters, 116(3):237– 240, March 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001933.

Skowronek-Kaziow:2012:MVC

[SK12] Joanna Skowronek-Kaziów.

Multiplicative vertex-colouring weightings of graphs. Information Processing Letters, 112(5):191–194, February 28, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011003139.

Szalachowski:2010:CCG

[SKK10] P. Szalachowski, B. Ksiezopolski, and Z. Kotulski. CMAC, CCM and GCM/GMAC: Advanced modes of operation of symmetric block ciphers in wireless sensor networks. Information Processing Letters, 110(7):247–251, March 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Saso:2011:PPI

[SKN11] Takashi Saso, Kojiro Kobayashi, and Atsuyoshi Nakamura. On the possible patterns of inputs for block sorting in the Burrows–Wheeler transformation. Information Processing Letters, 111(12): 595–599, June 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019011000731.

Stambouli:2019:DFA

[SL19] Abdelouadoud Stambouli and Luigi Logrippo. Data flow analysis from capability lists, with application to RBAC. Information Processing Letters, 141(??):30– 40, January 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018301790.

Saettler:2015:ADT

[SLC15] Aline Saettler, Eduardo Laber, and Ferdinando Cicalese. Approximating decision trees with value dependent testing costs. Information Processing Letters, 115(6–8):594–599, June/ August 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901500023X.

Saettler:2017:DTC

[SLdAMP17] Aline Saettler, Eduardo Laber, and Felipe de A. Mello Pereira. Decision tree classification with bounded number of errors. Information Processing Letters, 127(??):27–31, November 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301205.

Sun:2019:SCR

[SLF19] Lei Sun, Jian Liu, and Fang-Wei Fu. Secondary constructions of RSBFs with good cryptographic properties. Information Processing Letters, 147(??):44– 48, July 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300547.

Sliva:2012:ALG

[Slí12] Radek Slíva. Antimagic labeling graphs with a regular dominating subgraph. Information Processing Letters, 112(21):844– 847, November 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019012002116.

Sousa:2013:GDC

[SLL13] L. S. Sousa, J. C. B. Leite, and Orlando Loques. Green data centers: Using hierarchies for scalable energy efficiency in large Web clusters. Information Processing Letters, 113(14–16):507– 515, July/August 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001901300121X.

Suksompong:2016:ELW

[SLS16] Warut Suksompong, Charles E. Leiserson, and Tao B. Schardl. On the efficiency of localized work stealing. Information Processing Letters, 116(2):100–106, February 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001726.

Sarkar:2010:CRM

[SM10a] Santanu Sarkar and Subhamoy Maitra. Cryptanalysis of RSA with more than one decryption exponent. Information Processing Letters, 110(8–9):336– 340, April 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sarkar:2010:CRT

[SM10b] Santanu Sarkar and Subhamoy Maitra. Cryptanalysis of RSA with two decryption exponents. Information Processing Letters, 110(5): 178–181, February 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Small:2013:NEC

[SM13] Lucy Small and Oliver Mason. Nash Equilibria for competitive information diffusion on trees. Information Processing Letters, 113(7):217–219, April 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000173.

Sherkhonov:2017:CAC

[SM17] Evgeny Sherkhonov and Maarten Marx. Containment of acyclic conjunctive queries with negated atoms or arithmetic comparisons. Information Processing Letters, 120(??):30– 39, April 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301867.

Sabir:2019:PRR

[SM19] Eminjan Sabir and Jixiang Meng. Parallel routing in regular networks with faults. Information Processing Letters, 142(??):84–89, February 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018302138.

Shen:2010:DEN

[SML^{+}10] Yang Shen, LiZhuang Ma, Hai Liu, Yanxia Bao, and Zhihua Chen. Detecting and extracting natural snow from videos. Information Processing Letters, 110(24): 1124–1130, November 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sakai:2010:WNO

[SO10] Masahiko Sakai and Mizuhito Ogawa. Weakly-non-overlapping non-collapsing shallow term rewriting systems are confluent. Information Processing Letters, 110 (18–19):810–814, September 15, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Siala:2019:CSE

[SO19] Mohamed Siala and Barry O’Sullivan. Combinatorial search from an energy perspective. Information Processing Letters, 148(??):23– 27, August 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019300791.

Sorenson:2010:RST

[Sor10] Jonathan P. Sorenson. A randomized sublinear time parallel GCD algorithm for the EREW PRAM. Information Processing Letters, 110(5):198–201, February 1, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Seo:2018:FBM

[SP18] Jungjoo Seo and Kunsoo Park. Fast batch modular exponentiation with common-multiplicand multiplication. Information Processing Letters, 129(??): 5–10, January 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019017301527.

Simoes:2013:UHC

[SPdR13] D. R. Simões, J. Portugheis, and V. C. da Rocha, Jr. Universal homophonic coding scheme using differential encoding and interleaving. Information Processing Letters, 113(17):628–633, August 30, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901300152X.

Shi:2017:GSD

[SQLS17] Minjia Shi, Liqin Qian, Yan Liu, and Patrick Solé. Good self-dual generalized quasicyclic codes exist. Information Processing Letters, 118(??):21–24, February 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016301326.

Seidl:2011:ECD

[SR11] Helmut Seidl and Andreas Reuß. Extending H_{1}-clauses with disequalities. Information Processing Letters, 111(20):1007–1013, October 31, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011002031.

Schweitzer:2010:CFH

[SS10a] Pascal Schweitzer and Patrick Schweitzer. Connecting face hitting sets in planar graphs. Information Processing Letters, 111(1): 11–15, December 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Simon:2010:OIH

[SS10b] Hans Ulrich Simon and Balázs Szörényi. Oneinclusion hypergraph density revisited. Information Processing Letters, 110(8– 9):341–344, April 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schaudt:2012:CCD

[SS12] Oliver Schaudt and Rainer Schrader. The complexity of connected dominating sets and total dominating sets with specified induced subgraphs. Information Processing Letters, 112(24):953–957, December 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002517.

Sanders:2016:SGS

[SS16] Peter Sanders and Christian Schulz. Scalable generation of scale-free graphs. Information Processing Letters, 116(7):489– 491, July 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300102.

Sadhya:2017:PRE

[SS17] Debanjan Sadhya and Sanjay Kumar Singh. Privacy risks ensuing from crossmatching among databases: a case study for soft biometrics. Information Processing Letters, 128(??):38–45, December 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301436.

Sakharov:2018:VAS

[SS18] Alexander Sakharov and Timothy Sakharov. The Viterbi algorithm for subsets of stochastic contextfree languages. Information Processing Letters, 135(??): 68–72, July 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300589.

Schlipf:2019:SCE

[SS19a] Lena Schlipf and Jens M. Schmidt. Simple computation of st-edge- and stnumberings from ear decompositions. Information Processing Letters, 145(??):58– 63, May 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901930016X.

Sheng:2019:ILK

[SS19b] Bin Sheng and Yuefang Sun. An improved linear kernel for the cycle contraction problem. Information Processing Letters, 149(??):14– 18, September 2019. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019019300912.

Su:2012:FHG

[SSK12] Hsun Su, Yuan-Kang Shih, and Shin-Shin Kao. On the 1-fault hamiltonicity for graphs satisfying Ore’s theorem. Information Processing Letters, 112(21):839–843, November 15, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019012002025.

Son:2017:TMS

[SSKA17] Wanbin Son, Fabian Stehn, Christian Knauer, and Hee-Kap Ahn. Top-k Manhattan spatial skyline queries. Information Processing Letters, 123(??):27– 35, July 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901730039X.

Schmidt-Schauss:2010:CMS

[SSS10] Manfred Schmidt-Schauß and David Sabel. Closures of may-, should- and mustconvergences for contextual equivalence. Information Processing Letters, 110(6): 232–235, February 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schmidt-Schauss:2012:FET

[SSS12] Manfred Schmidt-Schauß and Georg Schnitger. Fast equality test for straight-line compressed strings. Information Processing Letters, 112(8–9):341–345, April 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000257.

Scott:2015:AAH

[SSS15] Allan Scott, Venkatesh Srinivasan, and Ulrike Stege. k-Attribute-Anonymity is hard even for k = 2. Information Processing Letters, 115(2):368–370, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002269.

Schmidt-Schauss:2011:CAS

[SSSM11] Manfred Schmidt-Schauß, David Sabel, and Elena Machkasova. Counterexamples to applicative simulation and extensionality in non-deterministic call-by-need lambda-calculi with letrec. Information Processing Letters, 111(14):711– 716, July 31, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001177.

Shiau:2015:ICC

[SSW15] Alexander Chane Shiau, Tzong-Huei Shiau, and Yue-Li Wang. Incidence coloring of Cartesian product graphs. Information Processing Letters, 115(10):765–768, October 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000721.

See corrigendum [SSW16].

Shiau:2016:CIC

[SSW16] Alexander Chane Shiau, Tzong-Huei Shiau, and Yue-Li Wang. Corrigendum to “Incidence coloring of Cartesian product graphs” [Inf. Process. Lett. 115 (2015) 765–768]. Information Processing Letters, 116 (2):216, February 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015001428.

See [SSW15].

Song:2016:ATM

[SSZW16] Yalin Song, Yaoru Sun, Hong Zhang, and Fang Wang. Activity testing model for automatic correction of hand pointing. Information Processing Letters, 116(11):653–659, November 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016300928.

Smal:2018:PPI

[ST18] Alexander V. Smal and Navid Talebanfard. Prediction from partial information and hindsight, an alternative proof. Information Processing Letters, 136(??): 102–104, August 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019018300917.

Santiesteban-Toca:2015:NMC

[STAR15] Cosme E. Santiesteban-Toca and Jesús S. Aguilar-Ruiz. A new multiple classifier system for the prediction of protein’s contacts map. Information Processing Letters, 115(12):983–990, December 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001040.

Soeken:2018:PMD

[STMD18] Mathias Soeken, Eleonora Testa, Alan Mishchenko, and Giovanni De Micheli. Pairs of majority-decomposing functions. Information Processing Letters, 139(??):30– 34, November 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019018301510.

Storch:2016:BBC

[Sto16] Tobias Storch. Black-box complexity: Advantages of memory usage. Information Processing Letters, 116(6): 428–432, June 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019016000119.

Struth:2016:EPK

[Str16] Georg Struth. On the expressive power of Kleene algebra with domain. Information Processing Letters, 116(4):284–288, April 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002001901500191X.

Shrestha:2012:BCB

[STU12] Anish Man Singh Shrestha, Satoshi Tayu, and Shuichi Ueno. Bandwidth of convex bipartite graphs and related graphs. Information Processing Letters, 112(11): 411–417, June 15, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019012000610.

Shieh:2012:IMI

[STY12] Min-Zheng Shieh, Shi-Chun Tsai, and Ming-Chuan Yang. On the inapproximability of maximum intersection problems. Information Processing Letters, 112(19):723–727, October 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012001652.

Sukhamay:2019:RBO

[Suk19] Kundu Sukhamay. Relationship between optimal kdistance dominating sets in a weighted graph and its spanning trees. Information Processing Letters, 147(??): 3–5, July 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019300250.

Sung:2011:DCE

[Sun11] Jaechul Sung. Differential cryptanalysis of eightround SEED. Information Processing Letters, 111(10): 474–478, April 30, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sun:2016:NEB

[Sun16] Shuliang Sun. A novel edge based image steganography with 2^{k} correction and Huffman encoding. Information Processing Letters, 116(2): 93–99, February 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015001738.

Sun:2019:NCS

[Sun19]  Hongyang Sun.  Non-

clairvoyant scheduling with conflicts for unit-size jobs. Information Processing Letters, 144(??):1–8, April 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019018302400.

Suzuki:2018:NDF

[Suz18] Toshio Suzuki. Non-depth-first search against independent distributions on an AND–OR tree. Information Processing Letters, 139(??): 8–12, November 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019018301388.

Surinx:2019:MPR

[SV19] Dimitri Surinx and Jan Van den Bussche. A monotone preservation result for Boolean queries expressed as a containment of conjunctive queries. Information Processing Letters, 150(??): 1–5, October 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019300936.

Sviridenko:2012:NKR

[Svi12]  Maxim Sviridenko.  A

note on the Kenyon–Remila strip-packing algorithm. Information Processing Letters, 112(1–2):10–12, January 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011002742.

Srivathsan:2012:APS

[SW12] B. Srivathsan and Igor Walukiewicz. An alternate proof of Statman’s finite completeness theorem. Information Processing Letters, 112(14–15):612–616, August 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012001263.

Schmitt:2013:EIS

[SW13] Manuel Schmitt and Rolf Wanka. Exploiting independent subformulas: a faster approximation scheme for #k-SAT. Information Processing Letters, 113(9):337– 344, May 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000677.

Stokes:2018:CGW

[SW18] Jonathan Stokes and Steven Weber. Common greedy wiring and rewiring heuristics do not guarantee maximum assortative graphs of given degree. Information Processing Letters, 139(??): 49–52, November 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019018301509.

Srimani:2019:SSA

[SW19] Pradip K. Srimani and James Z. Wang. Selfstabilizing algorithm for two disjoint minimal dominating sets. Information Processing Letters, 147(??):38– 43, July 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300572.

Shen:2018:NCZ

[SWF18] Lin-Zhi Shen, Jie-Jing Wen, and Fang-Wei Fu. A new class of zero-difference balanced functions. Information Processing Letters, 136(??):9–11, August 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300668.

Song:2015:CDM

[SWLX15] Yafei Song, Xiaodan Wang, Lei Lei, and Yaqiong Xing. Credibility decay model in temporal evidence combination. Information Processing Letters, 115(2):248–252, February 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014001999.

Sun:2016:ZNN

[SWW16] Jia Sun, Shiheng Wang, and Ke Wang. Zhang neural networks for a set of linear matrix inequalities with timevarying coefficient matrix. Information Processing Letters, 116(10):603–610, October 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901630062X.

Sparl:2012:LTA

[SWZ12] Petra Sparl, Rafal Witkowski, and Janez Zerovnik. A linear time algorithm for 7 − [3] coloring triangle-free hexagonal graphs. Information Processing Letters, 112(14–15):567–571, August 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000579.

Shao:2015:SAS

[SY15a] Zhi-Yi Shao and Bo Yang.

On security against the server in designated tester public key encryption with keyword search. Information Processing Letters, 115(12):957–961, December 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001283.

Song:2015:FLA

[SY15b] Yinglei Song and Menghong Yu. On finding the longest antisymmetric path in directed acyclic graphs. Information Processing Letters, 115(2):377–381, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002294.

Sun:2018:GOF

[SZ18] Zhiren Sun and Sizhong Zhou. A generalization of orthogonal factorizations in digraphs. Information Processing Letters, 132(??):49– 54, April 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901730217X.

Su:2017:GBC

[SZC^{+}17] Sen Su, Sen Zhao, Xiang Cheng, Rong Bi, Xin Cao, and Jie Wang. Group-based collective keyword querying in road networks. Information Processing Letters, 118(??):83–90, February 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016301491.

Szymanska:2012:CDH

[Szy12]  Edyta Szymańska.  H-

colorings of dense hypergraphs. Information Processing Letters, 112 (23):899–902, December 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901200230X.

Takazawa:2016:AAM

[Tak16] Kenjiro Takazawa. A 7/6approximation algorithm for the minimum 2-edge connected subgraph problem in bipartite cubic graphs. Information Processing Letters, 116(9):550– 553, September 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901630059X.

Tarlecki:2014:WMT

[Tar14] Andrzej Tarlecki. Władyslaw Marek Turski (1938–2013). Information Processing Letters, 114(8):397–398, August 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014000581.

Tatti:2019:SPE

[Tat19]  Nikolaj Tatti.  Strongly

polynomial efficient approximation scheme for segmentation. Information Processing Letters, 142(??):1– 8, February 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301856.

Tang:2011:EKA

[TC11] Qiang Tang and Liqun Chen. Extended KCI attack against two-party key establishment protocols. Information Processing Letters, 111(15):744–747, August 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001293.

Tsai:2017:DEE

[TC17] Chang-Hsiung Tsai and Jheng-Cheng Chen. Diagnosable evaluation of enhanced optical transpose interconnection system networks. Information Processing Letters, 118(??):105– 109, February 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019016301478.

Tao:2010:OSO

[TCXT10] Jiping Tao, Zhijun Chao, Yugeng Xi, and Ye Tao. An optimal semi-online algorithm for a single machine scheduling problem with bounded processing time. Information Processing Letters, 110(8–9):325– 330, April 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tiplea:2014:NSC

[TD14] Ferucio Laurentiu Tiplea and Constantin Catalin Dragan. A necessary and sufficient condition for the asymptotic idealness of the GRS threshold secret sharing scheme. Information Processing Letters, 114(6): 299–303, June 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014000179.

Tell:2019:PPP

[Tel19] Roei Tell. Proving that prBPP = prP is as hard as proving that “almost NP” is not contained in P /poly. Information Processing Letters, 152(??):Article 105841, December 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019301243.

Tian:2011:LAS

[TFY11] Ji Tian, Ruyan Fu, and Jinjiang Yuan. An on-line algorithm for the single machine unbounded parallelbatching scheduling with large delivery times. Information Processing Letters, 111(21–22):1048–1053, November 15, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011002249.

Thomas:2012:APL

[Tho12] Michael Thomas. On the applicability of Post’s lattice. Information Processing Letters, 112(10):386– 391, May 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012000439.

Thomasian:2018:VSM

[Tho18] Alexander Thomasian. Vacationing server model for M/G/1 queues for rebuild processing in RAID5 and threshold scheduling for readers and writers. Information Processing Letters, 135(??):41–46, July 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300401.

Takehara:2012:CPS

[THS12] Reiko Takehara, Masahiro Hachimori, and Maiko Shigeno. A comment on pure-strategy Nash equilibria in competitive diffusion games. Information Processing Letters, 112(3):59–60, January 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011002869.

Tian:2015:IBP

[Tia15] Miaomiao Tian. Identitybased proxy re-signatures from lattices. Information Processing Letters, 115(4): 462–467, April 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901400266X.

Tille:2019:GRS

[Til19] Yves Tillé. A general result for selecting balanced unequal probability samples from a stream. Information Processing Letters, 152(??):Article 105840, December 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019301231.

Tiwary:2014:LCS

[Tiw14]  Hans Raj Tiwary.  On

the largest convex subsets in Minkowski sums. Information Processing Letters, 114(8):405–407, August 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014000374.

Tan:2014:OSS

[TJ14] Xuehou Tan and Bo Jiang.

Optimum sweeps of simple polygons with two guards. Information Processing Letters, 114(3):130– 136, March 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002706.

Tan:2018:IAC

[TJ18] Xuehou Tan and Bo Jiang. An improved algorithm for computing a shortest watchman route for lines. Information Processing Letters, 131(??):51–54, March 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017302053.

Thakkar:2015:NDT

[TK15] Jay Thakkar and Aditya Kanade. Non-deterministic transducer models of retransmission protocols over noisy channels. Information Processing Letters, 115(9):694–698, September 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015000691.

Tian:2012:PSO

[TL12] Fang Tian and Zi-Long Liu. Probabilistic single obnoxious facility location with fixed budget. Information Processing Letters, 112(5):195–199, February 28, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011002365.

Ta:2016:EAC

[TLL16] Toan Thang Ta, Cheng-Yao Lin, and Chin Lung Lu. An efficient algorithm for computing nonoverlapping inversion and transposition distance. Information Processing Letters, 116(12):744–749, December 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301089.

Ta:2018:TSC

[TLL18] Toan Thang Ta, Cheng-Yao Lin, and Chin Lung Lu. Two-string consensus problem under non-overlapping inversion and transposition distance. Information Processing Letters, 130(??):46– 51, February 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301849.

Tax:2018:IPM

[TLS^{+}18] Niek Tax, Xixi Lu, Natalia Sidorova, Dirk Fahland, and Wil M. P. van der Aalst. The imprecisions of precision measures in process mining. Information Processing Letters, 135(??):1– 8, July 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300280.

Tan:2016:CRD

[TLWZ16] Haisheng Tan, Hongyu Liang, Rui Wang, and Jipeng Zhou. Computing Roman domatic number of graphs. Information Processing Letters, 116(9):554– 559, September 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019016300588.

Taheri:2011:HTB

[TNN11] Hoda Taheri, Peyman Neamatollahi, and Mahmoud Naghibzadeh. A hybrid token-based distributed mutual exclusion algorithm using wraparound twodimensional array logical topology. Information Processing Letters, 111 (17):841–847, September 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011001554.

Tan:2016:BIB

[TPL16] Chik How Tan, Theo Fanuela Prabowo, and Duc-Phong Le. Breaking an ID-based encryption based on discrete logarithm and factorization problems. Information Processing Letters, 116(2):116– 119, February 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015001696.

Tezcan:2016:IID

[TS16] Cihangir Tezcan and Ali Aydin Selçuk. Improved improbable differential attacks on ISO standard CLEFIA: Expansion technique revisited. Information Processing Letters, 116(2):136–143, February 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001659.

Tsai:2011:NOR

[Tsa11] Ping-Ying Tsai. A note on an optimal result on faulttolerant cycle-embedding in alternating group graphs. Information Processing Letters, 111(8):375–378, March 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsai:2015:PDA

[Tsa15] Chang-Hsiung Tsai. The pessimistic diagnosability of alternating group graphs under the PMC model. Information Processing Letters, 115(2):151–154, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901400177X.

Tsur:2013:TDR

[Tsu13] Dekel Tsur. Top-k document retrieval in optimal space. Information Processing Letters, 113(12):440– 443, June 30, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013000896.

Tsur:2018:SDS

[Tsu18]  Dekel Tsur.  Succinct

data structures for nearest colored node in a tree. Information Processing Letters, 132(??):6–10, April 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017301710.

Tsur:2019:EEN

[Tsu19a] Dekel Tsur. The effective entropy of next/previous larger/smaller value queries. Information Processing Letters, 145(??):39–43, May 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0020019019300195.

Tsur:2019:FPA

[Tsu19b] Dekel Tsur. Faster parameterized algorithm for pumpkin vertex deletion set. Information Processing Letters, 147(??):74–76, July 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300663.

Tu:2015:FPA

[Tu15]  Jianhua Tu.  A fixed-

parameter algorithm for the vertex cover P_{3} problem. Information Processing Letters, 115(2):96–99, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001446.

Turetsky:2012:TSW

[Tur12] Daniel Turetsky. A K-trivial set which is not jump traceable at certain orders. Information Processing Letters, 112(13):544–547, July 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901200107X.

Turau:2013:SSA

[Tur13] Volker Turau. Self-stabilizing algorithms for efficient sets of graphs and trees. Information Processing Letters, 113(19–21):771–776, September/October 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002001901300197X.

Tutu:2013:CSI

[Tut13]  Ionut Tutu.  Comor-

phisms of structured institutions. Information Processing Letters, 113(22– 24):894–900, November/ December 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002408.

Tseng:2015:BUC

[TVB15] Lewis Tseng, Nitin Vaidya, and Vartika Bhandari. Broadcast using certified propagation algorithm in presence of Byzantine faults. Information Processing Letters, 115(4):512–514, April 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019014002609.

Tan:2017:NAN

[TWW17] Lijun Tan, Shiheng Wang, and Ke Wang. A new adaptive network-based fuzzy inference system with adaptive adjustment rules for stock market volatility forecasting. Information Processing Letters, 127(??):32– 36, November 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019017301217.

Tiwary:2017:ECC

[TWZ17] Hans Raj Tiwary, Stefan Weltge, and Rico Zenklusen. Extension complexities of Cartesian products involving a pyramid. Information Processing Letters, 128(??): 11–13, December 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019017301333.

Thomasian:2011:CDP

[TX11] Alexander Thomasian and Jun Xu. X-code double parity array operation with two disk failures. Information Processing Letters, 111 (12):568–574, June 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tang:2011:FPC

[TXQ11] Chunming Tang, Maozhi Xu, and Yanfeng Qi. Faster pairing computation on genus 2 hyperelliptic curves. Information Processing Letters, 111(10):494–499, April 30, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tu:2013:VCI

[TY13] Jianhua Tu and Fengmei Yang. The vertex cover P_{3} problem in cubic graphs. Information Processing Letters, 113(13):481–485, July 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001130.

Tan:2016:CCA

[TY16a] Syh-Yuan Tan and Wun-She Yap. Cryptanalysis of a CP– ABE scheme with policy in normal forms. Information Processing Letters, 116(7): 492–495, July 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019016300126.

Tolba:2016:GMA

[TY16b] Mohamed Tolba and Amr M. Youssef. Generalized MitM attacks on full TWINE. Information Processing Letters, 116(2):128–135, February 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001660.

Tyszka:2013:CCF

[Tys13] Apoloniusz Tyszka. Conjecturally computable functions which unconditionally do not have any finitefold Diophantine representation. Information Processing Letters, 113(19– 21):719–722, September/ October 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001932.

Tu:2011:FAA

[TZ11] Jianhua Tu and Wenli Zhou.

A factor 2 approximation algorithm for the vertex cover P_{3} problem. Information Processing Letters, 111(14): 683–686, July 31, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011001153.

Tan:2016:PQT

[TZF16] Xiaoqing Tan, Xiaoqian Zhang, and Junbin Fang. Perfect quantum teleportation by four-particle cluster state. Information Processing Letters, 116(5):347– 350, May 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016000089.

Vadlamudi:2013:IBS

[VAC13] S. G. Vadlamudi, Sandip Aine, and P. P. Chakrabarti. Incremental Beam search. Information Processing Letters, 113(22–24):888–893, November/December 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0020019013002391.

Vagvolgyi:2018:IRT

[Vág18] Sándor Vágvölgyi. Intersection of the reflexive transitive closures of two rewrite relations induced by term rewriting systems. Information Processing Letters, 134 (??):47–51, June 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019018300346.

Vajnovszki:2018:ESL

[Vaj18] Vincent Vajnovszki. The equidistribution of some length-three vincular patterns on S_{n}(132). Information Processing Letters, 130(??):40–45, February 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017301837.

Vallee:2010:NEC

[Val10] Thierry Vallée. Normal Eulerian clique-covering and Hamiltonicity. Information Processing Letters, 110 (16):697–701, July 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Valmari:2012:FBP

[Val12] Antti Valmari. Fast brief practical DFA minimization. Information Processing Letters, 112(6):213–217, March 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011003267.

Vaya:2013:RCL

[Vay13] Shailesh Vaya. Round complexity of leader election and gossiping in bidirectional radio networks. Information Processing Letters, 113(9): 307–312, May 15, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901300046X.

vanBreugel:2012:BPC

[vB12] Franck van Breugel. On behavioural pseudometrics and closure ordinals. Information Processing Letters, 112(19):715–718, October 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001706.

Veanes:2015:STA

[VB15] Margus Veanes and Nikolaj Bjørner. Symbolic tree automata. Information Processing Letters, 115(3):418– 424, March 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002555.

Vianna:2019:SMB

[VB19] Henrique Damasceno Vianna and Jorge Luis Victória Barbosa. A scalable model for building context-aware applications for noncommunicable diseases prevention. Information Processing Letters, 148(??):1– 6, August 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019300675.

vanBevern:2015:APC

[vBCH^{+}15] René van Bevern, Jiehua Chen, Falk Hüffner, Stefan Kratsch, Nimrod Talmon, and Gerhard J. Woeginger. Approximability and parameterized complexity of multicover by c-intervals. Information Processing Letters, 115(10):744–749, October 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000423.

vonderBruggen:2017:ESF

[vdBCDH17] Georg von der Brüggen, Jian-Jia Chen, Robert I. Davis, and Wen-Hung Huang. Exact speedup factors for linear-time schedulability tests for fixedpriority preemptive and non-preemptive scheduling. Information Processing Letters, 117(??):1–5, January 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301090.

vanEe:2017:SNB

[vE17] Martijn van Ee. Some notes on bounded starwidth graphs. Information Processing Letters, 125(??):9– 14, September 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019017300704.

vanGlabbeek:2018:VES

[vG18] Rob J. van Glabbeek. On the validity of encodings of the synchronous in the asynchronous π-calculus. Information Processing Letters, 137(??):17–25, September 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300954.

vanGlabbeek:2011:APP

[vGGS11] Rob J. van Glabbeek, Ursula Goltz, and Jens-Wolfhard Schicke. Abstract processes of place/transition systems. Information Processing Letters, 111(13): 626–633, July 1, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019011000767.

vanGarderen:2012:UPS

[vGLM12] Mereke van Garderen, Giuseppe Liotta, and Henk Meijer. Universal point sets for 2-coloured trees. Information Processing Letters, 112(8–9):346–350, April 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012000269.

Viderman:2013:LDC

[Vid13] Michael Viderman. LP decoding of codes with expansion parameter above 2/3. Information Processing Letters, 113(7):225–228, April 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000240.

vanIersel:2013:QKC

[vIL13] Leo van Iersel and Simone Linz. A quadratic kernel for computing the hybridization number of multiple trees. Information Processing Letters, 113(9):318– 323, May 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000562.

Virza:2011:SVB

[Vir11] Madars Virza. Sensitivity versus block sensitivity of Boolean functions. Information Processing Letters, 111 (9):433–435, April 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanIersel:2010:LTP

[vISS10] Leo van Iersel, Charles Semple, and Mike Steel. Locating a tree in a phylogenetic network. Information Processing Letters, 110(23):1037–1043, November 15, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Venkatakrishnan:2018:IUB

[VK18] Y. B. Venkatakrishnan and B. Krishnakumari. An improved upper bound of edgevertex domination number of a tree. Information Processing Letters, 134(??):14– 17, June 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300279.

Vecchio:2015:GTD

[VLV15] Massimo Vecchio and Roberto López-Valcarce. A greedy topology design to accelerate consensus in broadcast wireless sensor networks. Information Processing Letters, 115(3):408– 413, March 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014002592.

Viard:2018:EMC

[VML18] Tiphaine Viard, Clémence Magnien, and Matthieu Latapy. Enumerating maximal cliques in link streams with durations. Information Processing Letters, 133(??): 44–48, May 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300152.

Vollala:2017:EEM

[VN17] Satyanarayana Vollala and Ramasubramanian N. Energy efficient modular exponentiation for publickey cryptography based on bit forwarding techniques. Information Processing Letters, 119(??):25– 38, March 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301715.

Vidali:2010:CVB

[VNP10] Janoš Vidali, Peter Nose, and Enes Pašalić. Collisions for variants of the BLAKE hash function. Information Processing Letters, 110(14– 15):585–590, July 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Voudouris:2019:NEP

[Vou19] Alexandros A. Voudouris. A note on the efficiency of position mechanisms with budget constraints. Information Processing Letters, 143(??): 28–33, March 2019. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019018302278.

Vrgoc:2015:UVA

[Vrg15] Domagoj Vrgoc. Using variable automata for querying data graphs. Information Processing Letters, 115(3): 425–430, March 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014002506.

Visconti:2018:IUB

[VSP18] Andrea Visconti, Chiara Valentina Schiavo, and René Peralta. Improved upper bounds for the expected circuit complexity of dense systems of linear equations over GF(2). Information Processing Letters, 137(??):1–5, September 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300905.

Vajnovszki:2011:RCP

[VV11] V. Vajnovszki and R. Vernay. Restricted compositions and permutations: From old to new Gray codes. Information Processing Letters, 111(13):650–655, July 1, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011000913.

Vygen:2011:FAO

[Vyg11] Jens Vygen. Faster algorithm for optimum Steiner trees. Information Processing Letters, 111(21– 22):1075–1079, November 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011002341.

vanZuylen:2016:TUB

[vZBSY16] Anke van Zuylen, James Bieron, Frans Schalekamp, and Gexin Yu. A tight upper bound on the number of cyclically adjacent transpositions to sort a permutation. Information Processing Letters, 116(11):718– 722, November 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019016300904.

vonzurGathen:2015:CGG

[vzGS15] Joachim von zur Gathen and Igor E. Shparlinski. Circulant graphs and GCD and LCM of subsets. Information Processing Letters, 115(2):134–138, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001549.

Walczak:2010:SRS

[Wal10] Bartosz Walczak. A simple representation of subwords of the Fibonacci word. Information Processing Letters, 110(21):956–960, October 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wawrzyniak:2014:SAL

[Waw14] Wojciech Wawrzyniak. A strengthened analysis of a local algorithm for the minimum dominating set problem in planar graphs. Information Processing Letters, 114(3):94–98, March 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0020019013002767.

Wang:2012:PCE

[WB12] Qingju Wang and Andrey Bogdanov. The provable constructive effect of diffusion switching mechanism in CLEFIA-type block ciphers. Information Processing Letters, 112(11):427– 432, June 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012000555.

Wang:2012:OSO

[WBC^{+}12] Yuxin Wang, Attila Benko, Xin Chen, György Dósa, He Guo, Xin Han, and Cecilia Sik Lanyi. Online scheduling with one rearrangement at the end: Revisited. Information Processing Letters, 112(16): 641–645, August 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002001901200138X.

Wu:2015:INE

[WCJ15] Yong Wu, T. C. E. Cheng, and Min Ji. Inefficiency of the Nash equilibrium for selfish machine covering on two hierarchical uniform machines. Information Processing Letters, 115(11):838–844, November 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001015.

Wu:2011:MBR

[WCW11] Chih-Jen Wu, Yu-Wei Chen, and Yue-Li Wang. The minimum bandwidth required at each time slot of the fast broadcasting scheme. Information Processing Letters, 111(20):1014–1018, October 31, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011002055.

Wang:2012:TOA

[WCZZ12] GuiPing Wang, ShuYu Chen, XiaoQin Zhang, and Zhen Zhou. Topological ordering algorithm for LDAG. Information Processing Letters, 112 (21):823–828, November 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012002098.

Wang:2011:ECG

[WD11] Yongge Wang and Yvo Desmedt. Edge-colored graphs with applications to homogeneous faults. Information Processing Letters, 111(13):634–641, July 1, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011000809.

Wu:2016:CDA

[WDH16] Yahui Wu, Su Deng, and Hongbin Huang. Capacity and delay analysis in delay tolerant network with multiple communities. Information Processing Letters, 116(4):310– 315, April 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001854.

Wan:2013:INE

[WDT13] Long Wan, Xiaofang Deng, and Zhiyi Tan. Inefficiency of Nash equilibria with parallel processing policy. Information Processing Letters, 113(13):465–469, July 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000835.

Wang:2016:GNC

[WH16] Shiying Wang and Weiping Han. The g-good-neighbor conditional diagnosability of n-dimensional hypercubes under the MM^{∗} model. Information Processing Letters, 116(9):574–577, September 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300473.

Widel:2017:CHS

[Wid17] Wojciech Widel. Cliqueheavy subgraphs and pancyclicity of 2-connected graphs. Information Processing Letters, 117(??):6– 9, January 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301120.

Wijsen:2010:RCC

[Wij10] Jef Wijsen. A remark on the complexity of consistent conjunctive query answering under primary key violations. Information Processing Letters, 110(21): 950–955, October 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Winzen:2013:DRQ

[Win13] Carola Winzen. Directionreversing quasi-random rumor spreading with restarts. Information Processing Letters, 113(22–24):921–926, November/December 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0020019013002433.

Witt:2014:FLT

[Wit14] Carsten Witt. Fitness levels with tail bounds for the analysis of randomized search heuristics. Information Processing Letters, 114(1–2):38–41, January/ February 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013002500.

Wu:2010:ECS

[WJS10] Zong-Da Wu, Tao Jiang, and Wu-Jie Su. Efficient computation of shortest absent words in a genomic sequence. Information Processing Letters, 110(14–15): 596–601, July 1, 2010. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:2011:TDH

[WL11a] Bo-Yi Wang and Hsueh-I Lu. Two-dimensional homing sort. Information Processing Letters, 111(21– 22):1067–1071, November 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011002213.

Wang:2011:GPC

[WL11b] Feng Wang and Wensong Lin. Group path covering and distance two labeling of graphs. Information Processing Letters, 111(13): 621–625, July 1, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011000792.

Wang:2012:GPC

[WL12] Feng Wang and Wensong Lin. Group path covering and L(j, k)-labelings of diameter two graphs. Information Processing Letters, 112(4):124–128, February 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011003073.

Wei:2008:GRM

[WLLS08] Hsin-Wen Wei, Kwei-Jay Lin, Wan-Chen Lu, and Wei-Kuan Shih. Generalized rate monotonic schedulability bounds using relative period ratios. Information Processing Letters, 107(5):142–148, August 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Par11].

Wu:2011:TCP

[WLW11] Qian Wu, Qiuli Lu, and Yingqian Wang. (Δ + 1)- total-colorability of plane graphs of maximum degree Δ ≥ 6 with neither chordal 5-cycle nor chordal 6-cycle. Information Processing Letters, 111(15):767–772, August 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011001268.

Wang:2010:ECS

[WMLN10] Shuaiqiang Wang, Jun Ma, Jiming Liu, and Xiaofei Niu. Evolving choice structures for genetic programming. Information Processing Letters, 110(20):871– 876, September 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:2010:IKP

[WNFC10] Jianxin Wang, Dan Ning, Qilong Feng, and Jianer Chen. An improved kernelization for P_{2}-packing. Information Processing Letters, 110(5):188–192, February 1, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Weimann:2011:NED

[WP11] Oren Weimann and David Peleg. A note on exact distance labeling. Information Processing Letters, 111(14): 671–673, July 31, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011001062.

Wang:2010:MGS

[WS10] Xiaorong Wang and Hongbo Shi. A modified Guruswami– Sudan algorithm for decoding Reed–Solomon codes. Information Processing Letters, 110(22):992–997, October 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:2013:PPC

[WS13] Gaoli Wang and Yanzhao Shen. Preimage and pseudocollision attacks on stepreduced SM3 hash function. Information Processing Letters, 113(8):301–306, April 30, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000513.

Wang:2016:SSC

[WS16] Chao Wang and René Sitters. On some special cases of the restricted assignment problem. Information Processing Letters, 116(11):723–728, November 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300916.

Wang:2012:SCK

[WSW12] Hechao Wang, Erfang Shan, and Wei Wang. On the super connectivity of Kronecker products of graphs. Information Processing Letters, 112(10):402–405, May 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000348.

Wang:2013:NMC

[WT13] Qichun Wang and Chik How Tan. A new method to construct Boolean functions with good cryptographic properties. Information Processing Letters, 113(14–16):567–571, July/ August 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001348.

Wu:2013:IVC

[Wu13] Yunjian Wu. An improvement on Vizing’s conjecture. Information Processing Letters, 113(3):87–88, February 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002979.

Wu:2014:AMM

[Wu14] Bang Ye Wu. On approximating metric 1-median in sublinear time. Information Processing Letters, 114(4): 163–166, April 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013003074.

Wang:2011:TCP

[WW11] Bing Wang and Jian-Liang Wu. Total coloring of planar graphs with maximum degree 7. Information Processing Letters, 111(20):1019– 1021, October 31, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011002043.

Watel:2015:FAP

[WWBB15] Dimitri Watel, Marc-Antoine Weisser, Cédric Bentz, and Dominique Barth. An FPT algorithm in polynomial space for the Directed Steiner Tree problem with Limited number of Diffusing nodes. Information Processing Letters, 115(2):275–279, February 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901400204X.

Wen:2014:MZC

[WWBC14] Long Wen, Meiqin Wang, Andrey Bogdanov, and Huaifeng Chen. Multidimensional zero-correlation attacks on lightweight block cipher HIGHT: Improved cryptanalysis of an ISO standard. Information Processing Letters, 114(6):322– 330, June 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014000088.

Wang:2014:LAG

[WWLC14] Huijuan Wang, Jianliang Wu, Bin Liu, and Hongyu Chen. On the linear arboricity of graphs embeddable in surfaces. Information Processing Letters, 114(9):475– 479, September 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014000568.

Wang:2012:FTA

[WWS12] Gaocai Wang, Guojun Wang, and Zhiguang Shan. Fault tolerance analysis of mesh networks with uniform versus nonuniform node failure probability. Information Processing Letters, 112(10): 396–401, May 31, 2012. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019012000373.

Wang:2018:SCI

[WWW18] Yiqiao Wang, Ping Wang, and Weifan Wang. Strong chromatic index of K_{4}- minor free graphs. Information Processing Letters, 129(??):53–56, January 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301631.

Wang:2013:DPS

[WWWZ13] Lei Wang, Xiaodong Wang, Yingjie Wu, and Daxin Zhu. A dynamic programming solution to a generalized LCS problem. Information Processing Letters, 113 (19–21):723–728, September/October 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001944.

Wang:2011:CIB

[WWYY11] Xu An Wang, Jian Weng, Xiaoyuan Yang, and Yanjiang Yang. Cryptanalysis of an identity based broadcast encryption scheme without random oracles. Information Processing Letters, 111(10):461–464, April 30, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:2015:NPS

[WWZ15] Yanning Wang, Yanyan Wang, and Xuguang Zhang. Note on property M(k) of some complete multipartite graphs. Information Processing Letters, 115(2):228– 231, February 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001901400194X.

Wang:2013:PGC

[WX13] Yingqian Wang and Jinghan Xu. Planar graphs with cycles of length neither 4 nor 6 are (2, 0, 0)-colorable. Information Processing Letters, 113(18):659–663, September 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001580.

Wu:2019:ELC

[WXCK19] Chenhuang Wu, Chunxiang Xu, Zhixiong Chen, and Pinhui Ke. On error linear complexity of new generalized cyclotomic binary sequences of period p^{2}. Information Processing Letters, 144(??):9–15, April 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018302473.

Wang:2012:BFN

[WXZ^{+}12] Jian Wang, Xirong Xu, Dejun Zhu, Liqing Gao, and Jun-Ming Xu. On the bounds of feedback numbers of (n, k)-star graphs. Information Processing Letters, 112(12):473–478, June 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012000828.

Wang:2017:RRC

[WYLP17] Junhu Wang, Jeffrey Xu Yu, Jixue Liu, and Chaoyi Pang. A revised result on chasing tree patterns under schema graphs. Information Processing Letters, 119(??):19– 24, March 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016301697.

Wang:2014:CGR

[WYW14] Zongyue Wang, Hongbo Yu, and Xiaoyun Wang. Cryptanalysis of GOST R hash function. Information Processing Letters, 114(12):655–662, December 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014001458.

Wang:2012:NCG

[WZL12] Shiying Wang, Lei Zhang, and Shangwei Lin. A neighborhood condition for graphs to be maximally krestricted edge connected. Information Processing Letters, 112(3):95–97, January 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011002833.

Wang:2016:FPA

[WZQH16] Rui Wang, Yongquan Zhou, Shilei Qiao, and Kang Huang. Flower pollination algorithm with bee pollinator for cluster analysis. Information Processing Letters, 116(1):1–14, January 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001507.

Wang:2018:CSA

[WZS^{+}18] Qiang Wang, Hao Zhang, Jianfei Sun, Hu Xiong, and Zhiguang Qin. Comments on “A secure anti-collusion data sharing scheme for dynamic groups in the cloud”. Information Processing Letters, 140(??):30–33, December 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300851.

See [ZJ16].

Xavier:2012:NMS

[Xav12] Eduardo C. Xavier. A note on a Maximum k-Subset Intersection problem. Information Processing Letters, 112(12):471–472, June 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012000750.

Xu:2011:STI

[XCYZ11] Haifeng Xu, Lin Chen, Deshi Ye, and Guochuan Zhang. Scheduling on two identical machines with a speed-up resource. Information Processing Letters, 111(17):831–835, September 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011001566.

Xu:2011:TUB

[XGX11] XueLin Xu, ZongSheng Gao, and Ke Xu. A tighter upper bound for random MAX 2-SAT. Information Processing Letters, 111(3): 115–119, January 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xiao:2010:FMW

[Xia10] Mingyu Xiao. Finding minimum 3-way cuts in hypergraphs. Information Processing Letters, 110(14– 15):554–558, July 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xu:2015:MRS

[XL15] Yinfeng Xu and Hongmei Li. Minimax regret 1sink location problem in dynamic cycle networks. Information Processing Letters, 115(2):163–169, February 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001811.

Xiao:2019:NGN

[XLT^{+}19] Lin Xiao, Kenli Li, Zhiguo Tan, Zhijun Zhang, Bolin Liao, Ke Chen, Long Jin, and Shuai Li. Nonlinear gradient neural network for solving system of linear equations. Information Processing Letters, 142(??):35– 40, February 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301893.

Xiang:2016:EMP

[XLWZ16] Xinyin Xiang, Hui Li, Mingyu Wang, and Xingwen Zhao. Efficient multi-party concurrent signature from lattices. Information Processing Letters, 116(8):497– 502, August 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016300138.

Xu:2016:PTC

[XLZ16] Jin Xu, Zepeng Li, and Enqiang Zhu. On purely treecolorable planar graphs. Information Processing Letters, 116(8):532–536, August 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016300357.

Xu:2011:CSM

[XMLL11] Zhenpeng Xu, Chaoguang Men, Weiwei Li, and Xiang Li. Checkpoint scheduling model for optimality. Information Processing Letters, 111(19):979–984, October 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011002018.

Xiao:2012:FAE

[XN12] Mingyu Xiao and Hiroshi Nagamochi. An FPT algorithm for edge subset feedback edge set. Information Processing Letters, 112(1–2):5–9, January 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901100278X.

Xiao:2010:FSS

[XPC^{+}10] Wenjun Xiao, Behrooz Parhami, Weidong Chen, Mingxin He, and Wenhong Wei. Fully symmetric swapped networks based on bipartite cluster connectivity. Information Processing Letters, 110(6):211–215, February 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xing:2015:BLT

[XS15] Huaming Xing and Moo Young Sohn. Bounds on locating total domination number of the Cartesian product of cycles and paths. Information Processing Letters, 115(12):950–956, December 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001325.

Xie:2012:EAP

[XTTH12] Y. Xie, S. Tang, C. Tang, and X. Huang. An efficient algorithm for parameterizing HsMM with Gaussian and Gamma distributions. Information Processing Letters, 112(19):732–737, October 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001561.

Xu:2011:AEC

[Xu11] Yian Xu. On acyclic edge coloring of toroidal graphs. Information Processing Letters, 111(13):647–649, July 1, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011001037.

Xu:2012:CMC

[Xu12] Ning Xu. Complexity of minimum corridor guarding problems. Information Processing Letters, 112 (17–18):691–696, September 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012001433.

Xie:2012:ORI

[XW12] Min Xie and Libin Wang. One-round identity-based key exchange with Perfect Forward Security. Information Processing Letters, 112(14–15):587–591, August 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012001196.

Xu:2016:CCP

[XWLJ16] Jie Xu, Qiaoyan Wen, Wenmin Li, and Zhengping Jin. Circuit ciphertextpolicy attribute-based hybrid encryption with verifiable delegation in cloud computing. IEEE Transactions on Parallel and Distributed Systems, 27(1):119– 129, January 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

www.computer.org/csdl/
trans/td/2016/01/07010954-

abs.html. See comments [XWS17].

Xiong:2017:CCC

[XWS17] Hu Xiong, Qiang Wang, and Jianfei Sun. Comments on “Circuit ciphertext-policy attribute-based hybrid encryption with verifiable delegation”. Information Processing Letters, 127(??):67– 70, November 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019017301308.

See [XWLJ16].

Xie:2014:SCC

[XXZ14] Dezheng Xie, Huanhuan Xiao, and Zhihong Zhao. Star coloring of cubic graphs. Information Processing Letters, 114(12): 689–691, December 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0020019014001082.

Xu:2010:RMS

[XZ10] Qiwen Xu and Naijun Zhan. Rate monotonic scheduling re-analysed. Information Processing Letters, 110(6): 226–231, February 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xing:2017:TLD

[XZ17a] Rundan Xing and Bo Zhou. On the two largest distance eigenvalues of graph powers. Information Processing Letters, 119(??):39–43, March 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016301727.

Xu:2017:LTA

[XZ17b] Zhaowei Xu and Wenhui Zhang. Linear templates of ACTL formulas with an application to SAT-based verification. Information Processing Letters, 127(??):6– 16, November 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019017301096.

Xu:2018:SKD

[XZ18] Chao Xu and Qian Zhang. The shortest kinship description problem. Information Processing Letters, 138(??):57–60, October 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301303.

Xie:2016:IAS

[XZBX16] Fang Xie, Yuzhong Zhang, Qingguo Bai, and Zhe Xu. Inefficiency analysis of the scheduling game on limited identical machines with activation costs. Information Processing Letters, 116(4): 316–320, April 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015001842.

Xu:2015:MAC

[XZW15] Chuandong Xu, Shenggui Zhang, and Yi Wang. On the maximum arc-chromatic number of digraphs with bounded outdegrees or indegrees. Information Processing Letters, 115(12):939– 944, December 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015001350.

Y:2013:LCE

[YA13] Asnath Victy Phamila Y.

and R. Amutha. Low complexity energy efficient very low bit-rate image compression scheme for wireless sensor network. Information Processing Letters, 113(18):672–676, September 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001750.

Yamakami:2014:CSQ

[Yam14] Tomoyuki Yamakami. Constantspace quantum interactive proofs against multiple provers. Information Processing Letters, 114(11):611–619, November 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001070.

Yamamoto:2019:FAF

[Yam19] Hiroaki Yamamoto. A faster algorithm for finding shortest substring matches of a regular expression. Information Processing Letters, 143(??):56–60, March 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019018302333.

Yang:2014:TOK

[Yan14] Yongjie Yang. Towards optimal kernel for edge-disjoint triangle packing. Information Processing Letters, 114 (7):344–348, July 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014000210.

Yao:2011:WHM

[YBC11] Erlin Yao, Yungang Bao, and Mingyu Chen. What Hill–Marty model learn from and break through Amdahl’s law? Information Processing Letters, 111 (23–24):1092–1095, December 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901100250X.

Yahiaoui:2013:SSA

[YBHK13] Saïd Yahiaoui, Yacine Belhoul, Mohammed Haddad, and Hamamache Kheddouci. Self-stabilizing algorithms for minimal global powerful alliance sets in graphs. Information Processing Letters, 113(10–11):365–370, May/ June 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000720.

Yu:2010:ETD

[YBM^{+}10] Cui Yu, James Bailey, Julian Montefusco, Rui Zhang, and Jiling Zhong. Enhancing the B^{+}-tree by dynamic node popularity caching. Information Processing Letters, 110(7):268–273, March 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yazdi:2015:AAD

[YBMK15] Ehsan Najafi Yazdi, Alireza Bagheri, Zahra Moezkarimi, and Hamidreza Keshavarz. An O(1)-approximation algorithm for the 2-dimensional geometric freeze-tag problem. Information Processing Letters, 115(6–8):618–622, June/August 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015000344.

Yan:2011:AAF

[YC11] Li Yan and Marek Chrobak. Approximation algorithms for the Fault-Tolerant Facility Placement problem. Information Processing Letters, 111(11):545–549, May 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yu:2019:NGE

[YC19] Unjong Yu and Jeong-Ok Choi. A note on general epidemic region for infinite regular graphs. Information Processing Letters, 143(??): 41–46, March 2019. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019018302308.

Yi:2011:IGB

[YCL11] Chenfu Yi, Yuhuan Chen, and Zhongliang Lu. Improved gradient-based neural networks for online solution of Lyapunov matrix equation. Information Processing Letters, 111(16):780–786, August 30, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001384.

Yang:2010:EMT

[YDT10] Xiaofan Yang, Qiang Dong, and Yuan Yan Tang. Embedding meshes/tori in faulty crossed cubes. Information Processing Letters, 110(14–15):559–564, July 1, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:2014:VDP

[YeCM14] Fang Yang, Xiang en Chen, and Chunyan Ma. On the vertex-distinguishing proper edge coloring of composition of complete graph and star. Information Processing Letters, 114(4):217–221, April 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901300269X.

Yang:2013:PPP

[YEMR13] Qingxuan Yang, John Ellis, Khalegh Mamakani, and Frank Ruskey. In-place permuting and perfect shuffling using involutions. Information Processing Letters, 113(10–11):386–391, May/ June 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013000719.

Yusuf:2019:TAB

[YEMR19] Marwa Yusuf, Ahmed El-Mahdy, and Erven Rohou. Towards automatic binary runtime loop deparallelization using onstack replacement. Information Processing Letters, 145 (??):53–57, May 2019. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019019300171.

Yuan:2012:NEC

[YGK12] Chen Yuan, Qian Guo, and Haibin Kan. A novel elementary construction of matching vectors. Information Processing Letters, 112(12):494–496, June 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000762.

Yuan:2019:DCF

[YH19] Yuan Yuan and Rong-Xia Hao. A degree condition for fractional [a, b]-covered graphs. Information Processing Letters, 143(??):20– 23, March 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018302151.

Yu:2012:IMB

[YHLC12] Shiwei Yu, T.-Z. Huang, Xiaoyun Liu, and Wufan Chen. Information measures based on fractional calculus. Information Processing Letters, 112(23):916–921, December 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012002463.

Yamanaka:2010:CEP

[YiN10] Katsuhisa Yamanaka and Shin ichi Nakano. A compact encoding of plane triangulations with efficient query supports. Information Processing Letters, 110 (18–19):803–809, September 15, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jiang:2016:EFS

[yJxW16] Sheng yi Jiang and Lian xi Wang. Efficient feature selection based on correlation measure between continuous and discrete features. Information Processing Letters, 116(2):203–215, February 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001271.

Yuan:2011:HRS

[YK11] Chen Yuan and Haibin Kan. Holographic reduction for some counting problems. Information Processing Letters, 111(15):761–766, August 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011001402.

Yamakami:2013:DPR

[YK13] Tomoyuki Yamakami and Yuichi Kato. The dissecting power of regular languages. Information Processing Letters, 113(4):116–122, February 28, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012003109.

Yuan:2015:RAJ

[YK15] Chen Yuan and Haibin Kan. A refined analysis on the jump number problem of interval orders. Information Processing Letters, 115(11):797–800, November 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015000988.

Yoshinaka:2012:CLS

[YKD^{+}12] Ryo Yoshinaka, Jun Kawahara, Shuhei Denzumi, Hiroki Arimura, and Shin ichi Minato. Counterexamples to the long-standing conjecture on the complexity of BDD binary operations. Information Processing Letters, 112(16):636–640, August 31, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001305.

Yang:2011:NZF

[YL11a] Fan Yang and Xiangwen Li.

Nowhere-zero 3-flows in dihedral Cayley graphs. Information Processing Letters, 111(9):416–419, April 1, 2011. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Yum:2011:ACO

[YL11b] Dae Hyun Yum and Pil Joong Lee. On the average cost of order-preserving encryption based on hypergeometric distribution. Information Processing Letters, 111(19):956–959, October 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001906.

Yang:2016:SAA

[YL16a] Zheng Yang and Shuangqing Li. On security analysis of an after-the-fact leakage resilient key exchange protocol. Information Processing Letters, 116(1):33– 40, January 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001490.

Yuan:2016:SCT

[YL16b] Jun Yuan and Aixia Liu. Sufficient conditions for triangle-free graphs to be super k-restricted edgeconnected. Information Processing Letters, 116(2): 163–167, February 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002001901500160X.

Yu:2017:LTA

[YL17] Ting Yu and Mengchi Liu.

A linear time algorithm for maximal clique enumeration in large sparse graphs. Information Processing Letters, 125(??):35–40, September 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017300911.

Yang:2010:KCF

[YLG10] Weihua Yang, Hengzhe Li, and Xiaofeng Guo. A kind of conditional fault tolerance of (n, k)-star graphs. Information Processing Letters, 110(22):1007–1011, October 31, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:2016:SIF

[YLGC16] Bin Yang, Jie Luo, Ling Guo, and Fang Cheng. Simultaneous image fusion and demosaicing via compressive sensing. Information Processing Letters, 116 (7):447–454, July 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019016300242.

Yang:2016:CLG

[YLLL16] Liang Yang, Bing Liu, Hongfei Lin, and Yuan Lin. Combining local and global information for product feature extraction in opinion documents. Information Processing Letters, 116(10): 623–627, October 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019016300576.

Yang:2010:CCC

[YLM10] Weihua Yang, Hengzhe Li, and Jixiang Meng. Conditional connectivity of Cayley graphs generated by transposition trees. Information Processing Letters, 110(23):1027–1030, November 15, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:2018:DLA

[YLZZ18a] Mengduo Yang, Fanzhang Li, Li Zhang, and Zhao Zhang. Deep learning algorithm with visual impression. Information Processing Letters, 136(??):1– 4, August 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300577.

Yang:2018:LGI

[YLZZ18b] Mengduo Yang, Fanzhang Li, Li Zhang, and Zhao Zhang. Lie group impression for deep learning. Information Processing Letters, 136(??):12–16, August 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300619.

Yu:2017:RPU

[YMJ17] Dongjin Yu, Yunlei Mu, and Yike Jin. Rating prediction using review texts with underlying sentiments. Information Processing Letters, 117(??):10–18, January 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019016301107.

Yu:2014:SPE

[YMSA14] Yong Yu, Yi Mu, Willy Susilo, and Man Ho Au. Security pitfalls of an efficient threshold proxy signature scheme for mobile agents. Information Processing Letters, 114(1–2):5–8, January/ February 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002639.

Yamashita:2014:FSS

[YNH^{+}14] Kazuya Yamashita, Yasuaki Nishitani, Sadaki Hirose, Satoshi Okawa, and Nobuyasu Osato. The firing squad synchronization problem with sub-generals. Information Processing Letters, 114(1–2):60–65, January/February 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019013002196.

Yu:2010:PAP

[YSGY10] Lingfei Yu, Kun She, Haigang Gong, and Changyuan Yu. Price of anarchy in parallel processing. Information Processing Letters, 110(8–9):288–293, April 1, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yonta:2010:RAH

[YTN10] Paulin Melatagia Yonta, Maurice Tchuente, and René Ndoundam. Routing automorphisms of the hypercube. Information Processing Letters, 110(20): 854–860, September 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:2015:PIC

[YTYZ15] Hao Yang, Keming Tang, Jianjiang Yu, and Licai Zhu.

A practical information coverage approach in wireless sensor network. Information Processing Letters, 115 (1):6–10, January 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019014001434.

Yuster:2011:SCE

[Yus11] Raphael Yuster. A shortest cycle for each vertex of a graph. Information Processing Letters, 111(21– 22):1057–1061, November 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011002195.

Yu:2012:NOO

[YW12] Sheng Yu and Prudence W. H. Wong. A note on “An optimal online algorithm for single machine scheduling to minimize total general completion time”. Information Processing Letters, 112(1–2):55–58, January 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011002638.

Yuan:2014:CSS

[YW14] Yumin Yuan and Chenhui Wang. Certificateless signature scheme with security enhanced in the standard model. Information Processing Letters, 114(9):492–499, September 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014000635.

Yan:2011:PAB

[YWC11] Guofeng Yan, Jianxin Wang, and Shuhong Chen. Performance analysis for (X , S)- bottleneck cell in large-scale wireless networks. Information Processing Letters, 111(6):269–277, February 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yao:2014:NHS

[YWcW^{+}14] Xin-Wei Yao, Wan-Liang Wang, Teng cao Wu, Xiao min Yao, and Shuang-Hua Yang. A Novel Hybrid Slot Allocation Mechanism for 802.11e EDCA Protocol. Information Processing Letters, 114(1–2):25–30, January/February 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019013002585.

Yang:2015:SPB

[YWL15] Yuxing Yang, Shiying Wang, and Jing Li. Subnetwork preclusion for bubblesort networks. Information Processing Letters, 115(11):817–821, November 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015001076.

Yang:2014:CIA

[YWWW14] Hedong Yang, Lijie Wen, Jianmin Wang, and Raymond K. Wong. CPL+: an improved approach for evaluating the local completeness of event logs. Information Processing Letters, 114(11):607–610, November 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014001094.

Yang:2012:OBL

[YWY12] Xiaofan Yang, Lei Wang, and Luxing Yang. Optimal broadcasting for locally twisted cubes. Information Processing Letters, 112(4):129–134, February 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901100305X.

Yu:2013:OIR

[YXW13] Haiyan Yu, Yinfeng Xu, and Tengyu Wu. Online inventory replenishment scheduling of temporary orders. Information Processing Letters, 113(5–6):188– 192, March 15, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019013000021.

Yang:2011:CDH

[YYDL11] Erjie Yang, Xiaofan Yang, Qiang Dong, and Jing Li. Conditional diagnosability of hypermeshes under the comparison model. Information Processing Letters, 111(4):188–193, January 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yenigun:2017:CCE

[YYK17] Hüsnü Yenigün, Nina Yevtushenko, and Natalia Kushik. The complexity of checking the existence and derivation of adaptive synchronizing experiments for deterministic FSMs. Information Processing Letters, 127(??):49–53, November 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301230.

Yang:2016:NGP

[YYR16] Chao Yang, Bing Yao, and Han Ren. A note on graph proper total colorings with many distinguishing constraints. Information Processing Letters, 116(6):396– 400, June 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019015002082.

Yu:2012:RWA

[YYYZ12] Cui Yu, Xiaofan Yang, LuXing Yang, and Jing Zhang. Routing and wavelength assignment for 3-ary n-cube in array-based optical network. Information Processing Letters, 112(6):252–256, March 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011003206.

Yu:2013:RWAa

[YYZH13] Cui Yu, Xiaofan Yang, Jing Zhang, and Li He. Routing and wavelength assignment for 3-ary ncube communication patterns in linear array optical networks for n communication rounds. Information Processing Letters, 113(18):677–680, September 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019013001749.

Yang:2019:FTP

[YZ19] Yuxing Yang and Lingling Zhang. Faulttolerant-prescribed Hamiltonian laceability of balanced hypercubes. Information Processing Letters, 145 (??):11–15, May 2019. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019019300031.

Yu:2014:MMS

[YZH14] Xianyu Yu, Yulin Zhang, and Kai Huang. Multimachine scheduling with general position-based deterioration to minimize total load revisited. Information Processing Letters, 114(8): 399–404, August 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019014000337.

Yan:2018:DSC

[YZRC18] Jin Yan, Shaohua Zhang, Yanyan Ren, and Junqing Cai. Degree sum conditions on two disjoint cycles in graphs. Information Processing Letters, 138(??):1– 6, October 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301091.

Yang:2017:SMC

[YZZ17] Weihua Yang, Shuli Zhao, and Shurong Zhang. Strong Menger connectivity with conditional faults of folded hypercubes. Information Processing Letters, 125(??): 30–34, September 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901730087X.

Yang:2012:ORA

[YZZX12] Xingyu Yang, Weiguo Zhang, Yong Zhang, and Weijun Xu. Optimal randomized algorithm for a generalized ski-rental with interest rate. Information Processing Letters, 112(13):548–551, July 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001081.

Zhang:2017:APM

[ZA17] Peng Zhang and Mikhail J.

Atallah. On approximate pattern matching with thresholds. Information Processing Letters, 123(??): 21–26, July 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017300376.

Zalinescu:2011:SSC

[Zal11] Eugen Zalinescu. Shorter strings containing all kelement permutations. Information Processing Letters, 111(12):605–608, June 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011000810.

Zhang:2010:SCP

[ZC10] Guo-Qiang Zhang and Licong Cui. A set coverage problem. Information Processing Letters, 110(4): 158–159, January 16, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2012:CNM

[ZC12] Zongyang Zhang and Zhenfu Cao. Concurrent nonmalleable statistically hiding commitment. Information Processing Letters, 112(11):443–448, June 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012000713.

Zhang:2011:UIP

[ZCC^{+}11] Yong Zhang, Zhuo Chang, Francis Y. L. Chin, Hing-Fung Ting, and Yung H. Tsin. Uniformly inserting points on square grid. Information Processing Letters, 111(16):773–779, August 30, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001219.

Zhi:2018:TSI

[ZCC18] Ruicong Zhi, Lianyu Cao, and Gang Cao. Translation and scale invariants of Krawtchouk moments. Information Processing Letters, 130(??):30–35, February 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301667.

Zhang:2012:OCC

[ZCT^{+}12] Yong Zhang, Francis Y. L.

Chin, Hing-Fung Ting, Joseph Wun-Tat Chan, Xin Han, and Ka-Cheong Lam. Online call control in cellular networks revisited. Information Processing Letters, 112(1–2):21–25, January 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011002766.

Zhu:2018:BFC

[ZD18] Chuangying Zhu and Junping Du. Background feature clustering and its application to social text. Information Processing Letters, 136(??):44–48, August 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300784.

Zhang:2017:SFA

[ZDQ^{+}17] Yunong Zhang, Yaqiong Ding, Binbin Qiu, Yinyan Zhang, and Xiaodong Li. Signum-function array activated ZNN with easier circuit implementation and finite-time convergence for linear systems solving. Information Processing Letters, 124(??):30–34, August 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017300674.

Zehavi:2016:RAL

[Zeh16] Meirav Zehavi. A randomized algorithm for long directed cycle. Information Processing Letters, 116(6): 419–422, June 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019016300114.

Zelke:2011:IMM

[Zel11] Mariano Zelke. Intractability of min- and max-cut in streaming graphs. Information Processing Letters, 111(3):145–150, January 1, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhou:2010:SCS

[ZF10] Jin-Xin Zhou and Yan-Quan Feng. Super-connected but not super edge-connected graphs. Information Processing Letters, 111(1):22– 25, December 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhou:2011:SCN

[ZFJZ11] Wujun Zhou, Jianxi Fan, Xiaohua Jia, and Shukui Zhang. The spined cube: a new hypercube variant with smaller diameter. Information Processing Letters, 111 (12):561–567, June 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zivkovic:2010:IUS

[ŽGH10] Aleš Živkovič, Uroš Goljat, and Marjan Heričko. Improving the usability of the source code quality index with interchangeable metrics sets. Information Processing Letters, 110(6):236– 240, February 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2011:FMI

[ZGL11] Kejia Zhang, Hong Gao, and Jianzhong Li. Finding multiple induced disjoint paths in general graphs. Information Processing Letters, 111(20):1022–1026, October 31, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011002080.

Zhu:2012:ISA

[ZGLY12] Xiaomin Zhu, Hao Guo, Shaoshuai Liang, and Xiaoling Yang. An improved security-aware packet scheduling algorithm in real-time wireless networks. Information Processing Letters, 112(7):282–288, March 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011003498.

Zhang:2010:ILB

[Zha10a] Chi Zhang. An improved lower bound on query complexity for quantum PAC learning. Information Processing Letters, 111(1):40– 45, December 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2010:RAC

[Zha10b] Haihui Zhang. RETRACTED ARTICLE: On 3-choosability of planar graphs with neither adjacent triangles nor 5-, 6- and 9-cycles. Information Processing Letters, 110(24):1084–1087, November 30, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See corrigendum [Zha13].

Zhang:2011:PSR

[Zha11] Liyu Zhang. Proof system representations of degrees of disjoint NP-pairs. Information Processing Letters, 111 (7):348–351, March 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2013:CSC

[Zha13] Haihui Zhang. Corrigendum to “On 3-choosability of planar graphs with neither adjacent triangles nor 5-, 6- and 9-cycles” [Information Processing Letters 110 (24) (2010) 1084– 1087]. Information Processing Letters, 113(9):354–356, May 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012003092.

See [Zha10b].

Zhang:2018:LMR

[Zha18] Hailiang Zhang. On the largest matching roots of graphs with cut edges. Information Processing Letters, 136(??):41–43, August 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019018300310.

Zhou:2015:NLB

[Zho15] Xue Zhou. A new lower bound for the number of perfect matchings of line graph. Information Processing Letters, 115(2):269–274, February 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014002038.

Zhou:2016:TDR

[Zho16] Gelin Zhou. Two-dimensional range successor in optimal time and almost linear space. Information Processing Letters, 116(2):171–174, February 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901500157X.

Zhu:2012:TSS

[Zhu12] Hongli Zhu. A two stage scheduling with transportation and batching. Information Processing Letters, 112(19):728–731, October 15, 2012. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001640.

Zhang:2013:IST

[ZHX13] Yan-Hong Zhang, Wei Hao, and Tao Xiang. Independent spanning trees in crossed cubes. Information Processing Letters, 113 (18):653–658, September 15, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001610.

See comment [CWYP14].

Zhou:2010:EDG

[ZHXS10] Zhenghua Zhou, Zhi Hu, Maozhi Xu, and Wangan Song. Efficient 3dimensional GLV method for faster point multiplication on some GLS elliptic curves. Information Processing Letters, 110 (22):1003–1006, October 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhou:2011:TCG

[ZJ11] Sizhong Zhou and Jiashang Jiang. Toughness and (a, b, k)-critical graphs. Information Processing Letters, 111(9):403–407, April 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhu:2016:SAC

[ZJ16] Zhongma Zhu and Rui Jiang. A secure anticollusion data sharing scheme for dynamic groups in the cloud. IEEE Transactions on Parallel and Distributed Systems, 27(1):40–50, January 2016. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        www.computer.org/csdl/
                                        trans/td/2016/01/07004069-

abs.html. See comments [WZS^{+}18].

Zemke:2010:GAG

[ZJN10] Andrew Zemke, Sandra James, and Darren A. Narayan. Greedy algorithms for generalized k-rankings of paths. Information Processing Letters, 110(22): 979–985, October 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2010:TVS

[ZKXY10] Yunong Zhang, Zhende Ke, Peng Xu, and Chenfu Yi.

Time-varying square roots finding via Zhang dynamics versus gradient dynamics and the former’s link and new explanation to Newton– Raphson iteration. Information Processing Letters, 110(24):1103–1109, November 30, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2011:ECC

[ZL11] Xia Zhang and Guizhen Liu. On edge covering colorings of graphs. Information Processing Letters, 111(20):1004–1006, October 31, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011002067.

Zhang:2012:ECP

[ZL12] Xin Zhang and Guizhen Liu. On edge colorings of 1-planar graphs without adjacent triangles. Information Processing Letters, 112(4):138–142, February 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011003024.

Zhu:2017:SWR

[ZL17] Chun Jiang Zhu and Kam-Yiu Lam. Source-wise round-trip spanners. Information Processing Letters, 124(??):42–45, August 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017300686.

Zhu:2018:DIR

[ZL18] Chun Jiang Zhu and Kam-Yiu Lam. Deterministic improved round-trip spanners. Information Processing Letters, 129(??):57–60, January 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019017301643.

Zhang:2011:APF

[ZLB11] Jinshan Zhang, Heng Liang, and Fengshan Bai. Approximating partition functions of the two-state spin system. Information Processing Letters, 111(14):702– 710, July 31, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011001189.

Zhang:2010:DNC

[ZLCM10] Xindong Zhang, Juan Liu, Xing Chen, and Jixiang Meng. Domination number of Cartesian products of directed cycles. Information Processing Letters, 111(1): 36–39, December 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2018:DRD

[ZLJS18] Xiujun Zhang, Zepeng Li, Huiqin Jiang, and Zehui Shao. Double Roman domination in trees. Information Processing Letters, 134(??): 31–34, June 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300139.

Zhang:2018:LPE

[ZLLZ18] Xin Zhang, Jingfen Lan, Bi Li, and Qiang Zhu. Light paths and edges in families of outer-1-planar graphs. Information Processing Letters, 136(??):83– 89, August 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018300887.

Zhang:2017:RDH

[ZLM17] Bo Zhang, Qianmu Li, and Yuanyuan Ma. Research on dynamic heuristic scanning technique and the application of the malicious code detection model. Information Processing Letters, 117(??):19–24, January 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019016300989.

Zhu:2019:VDU

[ZLNB19] Chun Jiang Zhu, Kam-Yiu Lam, Joseph Kee Yin Ng, and Jinbo Bi. On the VC-dimension of unique round-trip shortest path systems. Information Processing Letters, 145(??):1– 5, May 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002001901930002X.

Zhu:2015:TCT

[ZLS^{+}15] Enqiang Zhu, Zepeng Li, Zehui Shao, Jin Xu, and Chanjuan Liu. Tree-core and tree-coritivity of graphs. Information Processing Letters, 115(10):754–759, October 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015000447.

Zhang:2017:PMP

[ZLS^{+}17] Zhao Zhang, Xiaoting Li, Yishuo Shi, Hongmei Nie, and Yuqing Zhu. PTAS for minimum k-path vertex cover in ball graph. Information Processing Letters, 119(??):9–13, March 2017.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019016301673.

Zhu:2016:ACT

[ZLSX16] Enqiang Zhu, Zepeng Li, Zehui Shao, and Jin Xu. Acyclically 4-colorable triangulations. Information Processing Letters, 116(6): 401–408, June 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015002227.

Zhang:2013:LMR

[ZLY13] Hailiang Zhang, Rongfei Lin, and Guanglong Yu. The largest matching root of unicyclic graphs. Information Processing Letters, 113 (19–21):804–806, September/October 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013001920.

Zhang:2019:SML

[ZLZX19] E Zhang, Ming Liu, Feifeng Zheng, and Yinfeng Xu. Single machine lot scheduling to minimize the total weighted (discounted) completion time. Information Processing Letters, 142(??): 46–51, February 2019. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001901830187X.

Zhu:2019:WBL

[ZMZ19] Changming Zhu, Chengjiu Mei, and Rigui Zhou. Weight-based label-unknown multi-view data set generation approach. Information Processing Letters, 146(??): 1–12, June 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019019300237.

Zhao:2015:MWS

[ZpH15] Shuhuan Zhao and Zheng ping Hu. A modular weighted sparse representation based on Fisher discriminant and sparse residual for face recognition with occlusion. Information Processing Letters, 115(9):677– 683, September 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019015000642.

Zhang:2011:SRP

[ZPX11] Guiqing Zhang, Chung Keung Poon, and Yinfeng Xu. The ski-rental problem with multiple discount options. Information Processing Letters, 111 (18):903–906, September 30, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001839.

Zhang:2015:IMZ

[ZQJ^{+}15] Yunong Zhang, Binbin Qiu, Long Jin, Dongsheng Guo, and Zhi Yang. Infinitely many Zhang functions resulting in various ZNN models for time-varying matrix inversion with link to Drazin inverse. Information Processing Letters, 115(9):703– 706, September 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019015000460.

Zhu:2018:EPW

[ZS18] Enqiang Zhu and Zehui Shao. Extremal problems on weak Roman domination number. Information Processing Letters, 138(??):7– 11, October 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019018301224.

Shi:2011:OCB

[zSfNbL11] Hai zhong Shi, Pan feng Niu, and Jian bo Lu. One conjecture of bubblesort graphs. Information Processing Letters, 111 (18):926–929, September 30, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011001700.

Zhang:2016:PHG

[ZSLW16] Defu Zhang, Leyuan Shi, Stephen C. H. Leung, and Tao Wu. A priority heuristic for the guillotine rectangular packing problem. Information Processing Letters, 116(1):15–21, January 2016. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019015001519.

Zawidzki:2013:SPM

[ZST13] Michal Zawidzki, Renate A. Schmidt, and Dmitry Tishkovsky. Satisfiability problem for modal logic with global counting operators coded in binary is NExpTimecomplete. Information Processing Letters, 113(1–2): 34–38, January 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019012002645.

Zhou:2013:TCF

[ZSY13] Sizhong Zhou, Zhiren Sun, and Hui Ye. A toughness condition for fractional (k, m)-deleted graphs. Information Processing Letters, 113(8):255–259, April 30, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013000392.

Zunic:2018:SDO

[Zun18] Anastazia Zunić. Shape diameter for object analysis. Information Processing Letters, 136(??):76–79, August 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901830067X.

Zhang:2014:POS

[ZV14] Xiaowang Zhang and Jan Van den Bussche. On the primitivity of operators in SPARQL. Information Processing Letters, 114(9):480– 485, September 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901400057X.

Zhang:2010:LTM

[ZvdV10] Xiandong Zhang and Steef van de Velde. On-line twomachine job shop scheduling with time lags. Information Processing Letters, 110(12– 13):510–513, June 15, 2010.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2010:TCP

[ZW10] Jingwen Zhang and Yingqian Wang. (Δ + 1)-total-colorability of plane graphs with maximum degree Δ at least 6 and without adjacent short cycles. Information Processing Letters, 110 (18–19):830–834, September 15, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2011:ECP

[ZW11] Xin Zhang and Jian-Liang Wu. On edge colorings of 1planar graphs. Information Processing Letters, 111(3): 124–128, January 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2014:DAE

[ZW14] Lei Zhang and Wenling Wu. Differential analysis of the Extended Generalized Feistel Networks. Information Processing Letters, 114(12):723–727, December 2014. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014001318.

Zhang:2012:TAL

[ZWWL12] Liting Zhang, Wenling Wu, Peng Wang, and Bo Liang. TrCBC: Another look at CBC-MAC. Information Processing Letters, 112(7):302–307, March 31, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012000221.

Zhou:2010:CDA

[ZX10] Shuming Zhou and Wenjun Xiao. Conditional diagnosability of alternating group networks. Information Processing Letters, 110(10): 403–409, April 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhou:2011:CFT

[ZX11] Shuming Zhou and Jun-Ming Xu. Conditional fault tolerance of arrangement graphs. Information Processing Letters, 111 (21–22):1037–1043, November 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011002092.

Zong:2011:HNH

[ZXJ^{+}11] Yu Zong, Guandong Xu, Ping Jin, Yanchun Zhang, and Enhong Chen. HC AB: a new heuristic clustering algorithm based on Approximate Backbone. Information Processing Letters, 111(17):857–863, September 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011001542.

Zhou:2019:DCF

[ZXS19] Sizhong Zhou, Yang Xu, and Zhiren Sun. Degree conditions for fractional (a, b, k)- critical covered graphs. Information Processing Letters, 152(??):Article 105838, December 2019. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019019301218.

Zhang:2010:KCF

[ZXY10] Zhao Zhang, Wei Xiong, and Weihua Yang. A kind of conditional fault tolerance of alternating group graphs. Information Processing Letters, 110(22):998–1002, October 31, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2011:OAG

[ZXZL11] Wenming Zhang, Yinfeng Xu, Feifeng Zheng, and Ming Liu. Online algorithms for the general k-search problem. Information Processing Letters, 111(14): 678–682, July 31, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019011001141.

Zhang:2012:MST

[ZY12] Tongquan Zhang and Ying Yin. The minimum spanning tree problem with nonterminal set. Information Processing Letters, 112 (17–18):688–690, September 30, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019012001639.

Zhou:2017:LRC

[ZY17] Yanwei Zhou and Bo Yang.

Leakage-resilient CCA2secure certificateless publickey encryption scheme without bilinear pairing. Information Processing Letters, 130(??):16–24, February 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019017301680.

Zhou:2013:ELO

[ZYC13] Xiangqian Zhou, Bing Yao, and Xiang’en Chen. Every lobster is odd-elegant.

Information Processing Letters, 113(1–2):30–33, January 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019012002797.

Zhang:2015:GSD

[ZYL15] Hailiang Zhang, Guanglong Yu, and Shanlin Li. Graphs with six distinct matching roots. Information Processing Letters, 115(5):521– 526, May 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019014002270.

Zhao:2011:TBR

[ZZ11] Chunyan Zhao and Zhiming Zheng. Threshold behaviors of a random constraint satisfaction problem with exact phase transitions. Information Processing Letters, 111(20):985–988, October 31, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0020019011001980.

Zunic:2013:SEB

[ZZ13] Dragisa Zunić and Jovisa Zunić. Shape ellipticity based on the first Hu moment invariant. Information Processing Letters, 113 (19–21):807–810, September/October 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019013002093.

Zhang:2014:TSS

[ZZ14] Zan-Bo Zhang and Xiaoyan Zhang. Triangle strings: Structures for augmentation of vertex-disjoint triangle sets. Information Processing Letters, 114(8):450– 456, August 2014. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019014000520.

Zhang:2018:CNB

[ZZ18] Yuzhuo Zhang and Xia Zhang. On f-colorings of nearly bipartite graphs. Information Processing Letters, 134(??):67–71, June 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0020019018300334.

Zhao:2019:IAR

[ZZ19] Chunhuan Zhao and Zhongxiang Zheng. Improved analysis of the reduction from BDD to uSVP. Information Processing Letters, 148(??): 28–32, August 2019. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001901930081X.

Zhang:2010:FAM

[ZZH10a] Meng Zhang, Yi Zhang, and Liang Hu. A faster algorithm for matching a set of patterns with variable length don’t cares. Information Processing Letters, 110(6):216–220, February 15, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2010:PMW

[ZZH10b] Meng Zhang, Yi Zhang, and Liang Hu. Pattern matching with wildcards using words of shorter length. Information Processing Letters, 110(24):1099–1102, November 30, 2010. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2016:CRA

[ZZH16] Meng Zhang, Yi Zhang, and Chen Hou. Compact representations of automata for regular expression matching. Information Processing Letters, 116(12):750– 756, December 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019016301016.

Zhu:2011:IAF

[ZZJ11] Jianghan Zhu, Xiaomin Zhu, and Jianqing Jiang. Improving adaptivity and fairness of processing realtime tasks with QoS requirements on clusters through dynamic scheduling. Information Processing Letters, 111(12):609–613, June 15, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0020019011000834.

Zhang:2011:RRM

[ZZXL11] Yong Zhang, Weiguo Zhang, Weijun Xu, and Hongyi Li. A risk-reward model for the on-line leasing of depreciable equipment. Information Processing Letters, 111(6): 256–261, February 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2015:GSE

[ZZZ15] Yan Zhang, Zhaohui Zhu, and Jinjin Zhang. On the greatest solution of equations in CLL_{R}. Information Processing Letters, 115(11):903–907, November 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002001901500085X.